### v1

# Attention as an Adaptive Filter

Abstract
We introduce Adaptive Filter Attention (AFA), a novel attention mechanism that incorporates a learnable dynamics model directly into the computation of attention weights. Rather than comparing queries and keys directly, we model the input sequence as discrete observations of a linear stochastic differential equation (SDE). By imposing a linear dynamics model with simultaneously diagonalizable state matrices and noise covariances, we can make use of a closed-form solution to the differential Lyapunov equation to efficiently propagate pairwise uncertainties through the dynamics. Attention naturally arises as the maximum likelihood solution for this linear SDE, with attention weights corresponding to robust residual-based reweightings of the propagated pairwise precisions. Imposing an additional constraint on the state matrix’s eigenvalues leads to a simplified variant with the same computational and memory complexity as standard attention. In the limit of vanishing dynamics and process noise, and using a small-angle approximation, we recover ordinary dot-product attention.
1 Introduction
Self-attention has become the dominant paradigm for sequence modeling due to its parallelism and scalability. However, this parallelism comes at the cost of losing the recursive structure that RNNs naturally impose, which provides an implicit temporal regularization that encourages consistency over time. State space models (SSMs) offer a complementary approach: they retain temporal structure through shared linear dynamics while supporting parallel training via convolutional implementations. However, most SSMs assume deterministic dynamics, limiting their ability to capture stochasticity and uncertainty.
Our goal is to bridge these two paradigms. We view self-attention as a form of filtering: it selectively averages across inputs, denoising by pooling information from similar states. However, unlike Kalman Filtering, standard attention does not explicitly propagate states through time. This motivates incorporating explicit temporal structure—through known or learned dynamics—into attention, restoring the inductive bias that standard attention loses by forgoing recurrence.
This technical report is structured as follows:
Section 2: Preliminaries
We introduce the model setup and review standard results for propagating covariances through linear SDEs. We then discuss the batch maximum likelihood estimator and its connection to the Kalman Filter. Next, we derive a closed-form solution of the differential Lyapunov equation for diagonalizable systems with simultaneously diagonalizable process and measurement noise covariances. Finally, we discuss an adaptively reweighted MLE that reweights propagated covariances using the pair-wise residuals, and its relationship to the iteratively reweighted least squares (IRLS) algorithm and a robustified form of the Kalman Filter.
Section 3: Adaptive Filter Attention
We generalize the robustified MLE to a tensor form of attention, and briefly analyze its time and memory complexity. We show that the algorithm can be simplified to a form with the same time and memory complexity as standard attention, through the use of convolutional kernels for the (diagonalized) matrix exponential and pair-wise propagated covariances, and by imposing additional structure on the linear SDE. We show that this formulation reduces to a complex-valued variant of standard attention in the case of a skew-symmetric state matrix and normalized inputs. We then describe an implementation of this algorithm, and an unrolled, recursive approximation for inference. We describe basic experiments on simple simulated linear dynamical systems.
Section 4: Radial-Tangential Model
We generalize the SDE model from Section 3, which assumed isotropic noise, to allow for separate noise covariances in the tangential and radial directions. We show that the propagated covariances in this case can be written in closed-form for the case of diagonal plus rank-1 covariances that co-rotate with the dynamics. We derive a form of attention that solves the MLE while respecting the geometric constraints imposed by the assumed dynamics model, and compare it to the standard Transformer architecture.
Section 5: Related Work
We overview related work, including classical methods for filtering and adaptive filtering; the use of neural networks for modeling and filtering dynamical systems; neural state space models; connections between attention and structured state space models; related architectures; and various other topics.
Section 6: Discussion and Conclusion
Appendix A: Background
We briefly overview basic background on linear systems, stochastic processes, the Kalman Filter, adaptive filtering, the attention mechanism, and neural state space models.
Appendix B: Additional Material
We cover some additional topics, including a separable approximation of the propagated precision kernel, discussion of normal plus rank-1 noise covariance, and the expressivity of our structured SDE model.
Note: This technical report describes ongoing work. Evaluation on real-world datasets is in progress and will be reported in future work.
2 Preliminaries
We review standard results on vector SDEs and linear Gaussian filtering. The reader familiar with these topics may skip to Section 3. Additional background material can be found in Appendix A.
2.1 Model Setup and Definitions
Consider a linear time-invariant (LTI) Itô stochastic differential equation (SDE):
$d\mathbf{x}(t)=\mathbf{A}\mathbf{x}(t)dt+\mathbf{B}d\mathbf{w}(t),\quad\mathbf{z}(t_{k})=\mathbf{C}\mathbf{x}(t_{k})+\mathbf{v}(t_{k}),$
where $\mathbf{x}(t)\in\mathbb{R}^{d_{x}}$ is the latent state, $\mathbf{z}_{k}\in\mathbb{R}^{d_{z}}$ is the $k$th measurement, $\mathbf{A}$ is the state matrix, $\mathbf{B}$ is the input matrix, and $\mathbf{C}$ is the output matrix, $d\mathbf{w}(t)\sim\mathcal{N}(\mathbf{0},\mathbf{I}dt)$ is a standard Wiener process (Brownian motion), $\mathbf{Q}:=\mathbf{B}\mathbf{B}^{\top}$ is the continuous-time process noise covariance, and $\mathbf{v}(t_{k})$ is zero-mean Gaussian measurement noise with covariance $\mathbf{R}$, i.e. $\mathbf{v}(t_{k})\sim\mathcal{N}(\mathbf{0},\mathbf{R})$.
With initial condition $\mathbf{x}(t_{i})$, the state at a later time $t_{j}>t_{i}$ can be written explicitly using the matrix exponential and an Itô integral:
$\mathbf{x}(t_{j})=e^{\mathbf{A}(t_{j}-t_{i})}\mathbf{x}(t_{i})+\int_{t_{i}}^{t_{j}}e^{\mathbf{A}(t_{j}-s)}\mathbf{B}d\mathbf{w}(s),$
Letting $\tau=s-t_{i}$, this is:
$\mathbf{x}(t_{j})=e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{\mathbf{A}(\Delta t_{ij}-\tau)}\mathbf{B}d\mathbf{w}(\tau),$
where $\Delta t_{ij}=t_{j}-t_{i}$.
Given $\mathbf{x}(t_{j})$, we can estimate $\mathbf{x}(t_{i})$ using the matrix exponential:
$\mathbf{\hat{x}}_{i}(t_{j}):=e^{-\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{j}).$
Likewise, if we measure $\mathbf{z}(t_{j})$, and the measurement matrix $\mathbf{C}$ is square and invertible, we can estimate the latent state $\mathbf{x}(t_{i})$, or equivalently $\mathbf{z}^{C}(t_{i}):=\mathbf{C}\mathbf{x}(t_{i})$, by computing a “pulled-back” or "pulled-forward" measurement:
$\mathbf{\hat{z}}_{i}(t_{j}):=\mathbf{C}e^{-\mathbf{A}\Delta t_{ij}}\mathbf{C}^{-1}\mathbf{z}(t_{j}).$
2.2 Propagation of Uncertainty
Recall that the state at time $t_{j}$ can be written as:
$\mathbf{x}(t_{j})=e^{\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}\,d\mathbf{w}(\tau)\right).$
The covariance of the integral term is:
$\mathbf{V}_{B}(\Delta t_{ij}):=\mathbb{E}\Bigg{[}\left(\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau_{1}}\mathbf{B}\,d\mathbf{w}(\tau_{1})\right)\left(\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau_{2}}\mathbf{B}\,d\mathbf{w}(\tau_{2})\right)^{\top}\Bigg{]}$
$\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{Q}e^{-\mathbf{A}^{\top}\tau}\,d\tau,\quad\mathbf{Q}=\mathbf{B}\mathbf{B}^{\top}$
We will refer to this as the backward-propagated state covariance. As the system evolves, the state distribution evolves as:
$\mathbf{x}(t_{j})\sim\mathcal{N}\left(e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i}),\mathbf{V}_{F}(\Delta t_{ij})\right)$
where the forward-propagated state covariance is:
$\mathbf{V}_{F}(\Delta t_{ij})=e^{\mathbf{A}\Delta t_{ij}}\mathbf{V}_{B}(\Delta t_{ij})e^{\mathbf{A}^{\top}\Delta t_{ij}}$
$e^{\mathbf{A}\Delta t_{ij}}\left[\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{Q}e^{-\mathbf{A}^{\top}\tau}\,d\tau\right]e^{\mathbf{A}^{\top}\Delta t_{ij}}.$
$\int_{0}^{\Delta t_{ij}}e^{\mathbf{A}(\Delta t_{ij}-\tau)}\mathbf{Q}e^{\mathbf{A}^{\top}(\Delta t_{ij}-\tau)}\,d\tau.$
Using the change of variables $s=\Delta t_{ij}-\tau\implies ds=-d\tau$ and flipping the limits, we get:
$\mathbf{V}_{F}(\Delta t_{ij})=\int_{0}^{\Delta t_{ij}}e^{\mathbf{A}s}\mathbf{Q}e^{\mathbf{A}^{\top}s}\,ds.$
which is a solution of the differential Lyapunov equation:
$\frac{d}{ds}\mathbf{V}_{F}(s)=\mathbf{A}\mathbf{V}_{F}(s)+\mathbf{V}_{F}(s)\mathbf{A}^{\top}+\mathbf{Q},$
with initial condition $\mathbf{V}_{F}(0)=\mathbf{0}$.
Recall that the estimate of the state is:
$\mathbf{\hat{x}}_{i}(t_{j}):=e^{\mathbf{-A}\Delta t_{ij}}\mathbf{x}(t_{j})=\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}d\mathbf{w}(\tau).$
Hence, $\mathbf{\hat{x}}_{i}(t_{j})\sim\mathcal{N}\left(\mathbf{x}(t_{i}),\mathbf{V}_{B}(\Delta t_{ij})\right)$, where: $\mathbf{V}_{B}(\Delta t_{ij})$ quantifies the uncertainty of the estimate $\mathbf{\hat{x}}_{i}(t_{j})$ of the state $\mathbf{x}(t_{i})$ at an earlier time $t_{i}$, given the state $\mathbf{x}(t_{j})$ at a later time $t_{j}$.
Conversely, if we observe the state at an earlier time $t_{j}$ and wish to quantify our uncertainty about an estimate of its state at a later time $t_{i}$ (i.e. the causal case), we can switch indices $i$ and $j$ and solve for $\mathbf{x}(t_{j})$:
$\mathbf{x}(t_{i})=e^{\mathbf{A}(t_{i}-t_{j})}\mathbf{x}(t_{j})+\int_{t_{j}}^{t_{i}}e^{\mathbf{A}(t_{i}-\tau)}\mathbf{B}d\mathbf{w}(\tau)$
$e^{-\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{j})+\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}d\mathbf{w}(\tau)\right).$
$\mathbf{x}(t_{j})=e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i})-\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}d\mathbf{w}(\tau)$
So:
$\mathbf{x}(t_{j})\sim\mathcal{N}\left(e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i}),\mathbf{V}_{B}(-\Delta t_{ij})\right)$
(Note that, according to our definition, $\Delta t_{ij}<0$ in the causal case, so $\Delta t_{ij}=|\Delta t_{ij}|$.)
Hence, in an LTI with Gaussian noise, the covariance of the state propagated forward in time under $\mathbf{A}$ is equal to the covariance of the state propagated backward in time under $\mathbf{A}$.
We again form an estimate:
$\mathbf{\hat{x}}_{i}(t_{j})=e^{-\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{j})=\mathbf{x}(t_{i})-e^{-\mathbf{A}\Delta t_{ij}}\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}d\mathbf{w}(\tau)$
which has distribution: $\mathbf{\hat{x}}_{i}(t_{j})\sim\mathcal{N}\left(\mathbf{x}(t_{i}),\mathbf{V}_{F}(|\Delta t_{ij}|)\right)$,
where $\mathbf{V}_{F}(|\Delta t_{ij}|)=e^{-\mathbf{A}\Delta t_{ij}}\mathbf{V}_{B}(|\Delta t_{ij}|)e^{-\mathbf{A}^{\top}\Delta t_{ij}}=e^{\mathbf{A}|\Delta t_{ij}|}\mathbf{V}_{B}(|\Delta t_{ij}|)e^{\mathbf{A}^{\top}|\Delta t_{ij}|}$.
So, in general, $\mathbf{\hat{x}}_{i}(t_{j})\sim\mathcal{N}\left(\mathbf{x}(t_{i}),\mathbf{V}(\Delta t_{ij})\right)$, where the propagated state covariance is:
[…]
Recall that the measurement at time $t_{j}$ is:
$\mathbf{z}(t_{j})=\mathbf{C}e^{\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}d\mathbf{w}(\tau)\right)+\mathbf{v}(t_{j}).$
The measurement process is distributed as:
$\mathbf{z}(t_{j})\sim\mathcal{N}\left(\mathbf{C}e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i}),\mathbf{C}\mathbf{V}_{F}(\Delta t_{ij})\mathbf{C}^{\top}+\mathbf{R}\right)$
Recall that our "pulled-forward" estimated measurement is defined as:
$\mathbf{\hat{z}}_{i}(t_{j}):=\mathbf{M}(\Delta t_{ij})\mathbf{z}(t_{j})\quad\text{where}\quad\mathbf{M}(\Delta t):=\mathbf{C}e^{\mathbf{-A}\Delta t}\mathbf{C}^{-1}.$
In terms of the "pulled-forward" state estimate, this is:
$\mathbf{\hat{z}}_{i}(t_{j})=\mathbf{C}e^{\mathbf{-A}\Delta t_{ij}}\mathbf{x}(t_{j})+\mathbf{C}e^{\mathbf{-A}\Delta t_{ij}}\mathbf{C}^{-1}\mathbf{v}(t_{j})$
$\mathbf{C}\mathbf{\hat{x}}_{i}(t_{j})+\mathbf{M}(\Delta t_{ij})\mathbf{v}(t_{j})$
Hence,
$\mathbf{\hat{z}}_{i}(t_{j})\sim\mathcal{N}\left(\mathbf{z}^{C}(t_{i}),\mathbf{V}^{C}(\Delta t_{ij})\right)$
where the propagated measurement covariance is:
$\mathbf{V}^{C}(\Delta t_{ij})=\mathbf{C}\mathbf{V}(\Delta t_{ij})\mathbf{C}^{\top}+\mathbf{M}(\Delta t_{ij})\mathbf{R}\mathbf{M}^{\top}(\Delta t_{ij})$
We can also define a propagated state precision matrix $\mathbf{P}(\Delta t_{ij})=\big{(}\mathbf{V}(\Delta t_{ij})\big{)}^{-1}$ and a propagated measurement precision matrix $\mathbf{P}^{C}(\Delta t_{ij})=\big{(}\mathbf{V}^{C}(\Delta t_{ij})\big{)}^{-1}$. Note that $\mathbf{V}(\Delta t)$ and $\mathbf{V}^{C}(\Delta t)$ and their inverses are positive definite (PD) and symmetric for all $\Delta t$, provided that $\mathbf{Q},\mathbf{R}$ are PD and symmetric.
2.3 Batch Maximum Likelihood Trajectory Estimation
We wish to estimate the trajectory of the system from $m$ measurements taken at times $\{t_{1},...,t_{m}\}$, which we denote $\{\mathbf{z}(t_{1}),...,\mathbf{z}(t_{m})\}$. We can do so by solving a separate maximum likelihood estimation (MLE) problem for each $\mathbf{x}_{i}$.
Let $\mathbf{x}_{i}=\mathbf{x}(t_{i})$, $\mathbf{\hat{x}}_{ij}=\mathbf{\hat{x}}_{i}(t_{j})$ and $\mathbf{V}_{ij}=\mathbf{V}(\Delta t_{ij})$. The log-likelihood is:
$\ell=-\frac{d_{x}m}{2}\log(2\pi)-\frac{1}{2}\sum_{j=1}^{m}\log|\det(\mathbf{V}_{ij})|-\frac{1}{2}\sum_{j=1}^{m}\left(\mathbf{\hat{x}}_{ij}-\mathbf{x}_{i}\right)^{\top}\mathbf{V}_{ij}^{-1}\left(\mathbf{\hat{x}}_{ij}-\mathbf{x}_{i}\right)$
Taking the derivative with respect to $\mathbf{x}_{i}$:
$\frac{\partial\ell}{\partial\mathbf{x}_{i}}=\sum_{j}\left(\mathbf{x}_{i}^{\top}\left(\mathbf{V}_{ij}^{-1}+\mathbf{V}_{ij}^{-\top}\right)-2\left(\mathbf{V}_{ij}^{-1}\mathbf{\hat{x}}_{ij}\right)^{\top}\right)=\mathbf{0}.$
Defining a precision matrix $\mathbf{P}_{ij}=\mathbf{V}_{ij}^{-1}$, the MLE solution for $\mathbf{x}(t_{i})$ is a precision-weighted estimator:
$\mathbf{\bar{x}}_{i}:=\left(\sum_{j}\mathbf{P}_{ij}\right)^{-1}\sum_{j}\mathbf{P}_{ij}\mathbf{\hat{x}}_{ij}$
Letting $\mathbf{z}_{i}=\mathbf{z}(t_{i})$, $\mathbf{\hat{z}}_{ij}=\mathbf{\hat{z}}_{i}(t_{j})$ and $\mathbf{V}^{C}_{ij}=\mathbf{V}^{C}(\Delta t_{ij})$, and repeating the above with the PDF for $\mathbf{\hat{z}}_{ij}$, the MLE solution for $\mathbf{z}^{C}(t_{i})$ is:
$\mathbf{\bar{z}}_{i}:=\left(\sum_{j}\mathbf{P}_{ij}^{C}\right)^{-1}\sum_{j}\mathbf{P}_{ij}^{C}\mathbf{\hat{z}}_{ij},\quad\text{where}\quad\mathbf{P}_{ij}^{C}=\mathbf{V}_{ij}^{C-1}.$
For the causal case:
$\mathbf{\bar{z}}_{i}:=\left(\sum_{j\leq i}\mathbf{P}_{ij}^{C}\right)^{-1}\sum_{j\leq i}\mathbf{P}_{ij}^{C}\mathbf{\hat{z}}_{ij}$
Repeating this computation for all $t_{i}$ produces an estimate of the full trajectory.
After estimating $\mathbf{\bar{z}}_{i}$, we can predict the next measurement $\mathbf{z}(t_{i+1})$ using the system dynamics:
$\mathbf{z}(t_{i+1})=\mathbf{C}e^{\mathbf{A}\Delta t}\mathbf{C}^{-1}\mathbf{z}^{C}(t_{i})+\mathbf{C}\int_{0}^{\Delta t}e^{-\mathbf{A}\tau}\mathbf{B}d\mathbf{w}(\tau)+\mathbf{v}(t_{i+1}),$
where $\Delta t=t_{i+1}-t_{i}$. Since the noise terms are zero mean and independent, the minimum mean squared error (MMSE) prediction given $\mathbf{\bar{z}}_{i}$ is:
$\mathbf{z}^{\mathrm{pred}}(t_{i+1}):=\mathbb{E}[\mathbf{z}(t_{i+1})\mid\mathbf{\bar{z}}_{i}]=\mathbf{C}e^{\mathbf{A}\Delta t}\mathbf{C}^{-1}\mathbf{\bar{z}}_{i}.$
which matches the standard approach in Kalman Filtering of propagating the current estimate forward through the deterministic part of the dynamics.
2.4 Recursive (Kalman Filter) Formulation
We can express the batch MLE for a trajectory in a recursive (information-form) manner, while accounting for the dynamics of the system. Consider estimating the state $\mathbf{z}^{C}(t_{i})$ given all measurements $\{\mathbf{z}(t_{j})\}_{j\leq i}$.
Prediction (Propagate Prior Through Dynamics).
Suppose we have already incorporated measurements up to time $t_{i-1}$. Let the posterior at $t_{i-1}$ be $\mathbf{\bar{z}}_{i-1}^{+}$ with covariance $\bm{\Lambda}_{\mathbf{V},i-1}^{+}$ and precision $\bm{\Lambda}_{\mathbf{P},i-1}^{+}=(\bm{\Lambda}_{\mathbf{V},i-1}^{+})^{-1}$. Propagating through the linear dynamics:
$\mathbf{z}_{i}=\mathbf{A}\mathbf{z}_{i-1}+\mathbf{q}_{i-1},\quad\mathbf{q}_{i-1}\sim\mathcal{N}(0,\mathbf{Q}),$
the prior for time $t_{i}$ becomes:
$\mathbf{\bar{z}}_{i}^{-}=\mathbf{A}\mathbf{\bar{z}}_{i-1}^{+},\quad\bm{\Lambda}_{\mathbf{V},i}^{-}=\mathbf{A}\bm{\Lambda}_{\mathbf{V},i-1}^{+}\mathbf{A}^{\top}+\mathbf{Q},\quad\bm{\Lambda}_{\mathbf{P},i}^{-}=(\bm{\Lambda}_{\mathbf{V},i}^{-})^{-1},\quad\bm{\eta}_{i}^{-}=\bm{\Lambda}_{\mathbf{P},i}^{-}\,\mathbf{\bar{z}}_{i}^{-}$
Update (Incorporate Measurement).
Now include the measurement at $t_{i}$ with measurement model:
$\mathbf{z}_{i}=\mathbf{C}\mathbf{z}^{-}_{i}+\mathbf{r}_{i},\quad\mathbf{r}_{i}\sim\mathcal{N}(0,\mathbf{R}),$
and define the information-form measurement contribution:
$\mathbf{P}^{C}_{ii}=\mathbf{C}^{\top}\mathbf{R}^{-1}\mathbf{C},\qquad\mathbf{b}_{i}=\mathbf{C}^{\top}\mathbf{R}^{-1}\mathbf{z}_{i}.$
The posterior precision and information vector are then:
$\bm{\Lambda}_{\mathbf{P},i}^{+}=\bm{\Lambda}_{\mathbf{P},i}^{-}+\mathbf{P}^{C}_{ii},\qquad\bm{\eta}_{i}^{+}=\bm{\eta}_{i}^{-}+\mathbf{b}_{i},$
and the posterior mean (MAP estimate) is:
$\mathbf{\bar{z}}_{i}^{+}=(\bm{\Lambda}_{\mathbf{P},i}^{+})^{-1}\bm{\eta}_{i}^{+}.$
Equivalently, in the standard Kalman Filter (covariance) form:
$\mathbf{K}_{i}=\bm{\Lambda}_{\mathbf{V},i}^{-}\mathbf{C}^{\top}(\mathbf{C}\bm{\Lambda}_{\mathbf{V},i}^{-}\mathbf{C}^{\top}+\mathbf{R})^{-1},\quad\mathbf{\bar{z}}_{i}^{+}=\mathbf{\bar{z}}_{i}^{-}+\mathbf{K}_{i}(\mathbf{z}_{i}-\mathbf{C}\mathbf{\bar{z}}_{i}^{-}),\quad\bm{\Lambda}_{\mathbf{V},i}^{+}=(\mathbf{I}-\mathbf{K}_{i}\mathbf{C})\bm{\Lambda}_{\mathbf{V},i}^{-}.$
2.5 Closed-form Precision Matrix Calculation for Diagonalizable Matrices
The estimation procedure above requires computing $\mathbf{P}^{C}(\Delta t_{ij})$, which is generally expensive, as it involves integration and matrix inversion. However, if we assume that $\mathbf{A}$ is diagonalizable, that is we can write it in the form $\mathbf{A}=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1}\in\mathbb{R}^{d\times d}$, where $\bm{\Lambda}\in\mathbb{C}^{d\times d}$ is a diagonal matrix and $\mathbf{S}\in\mathbb{C}^{d\times d}$ an invertible matrix, the backward-propagated covariance becomes:
$\mathbf{V}_{B}(\Delta t)=\int_{0}^{|\Delta t|}e^{\mathbf{-A}\tau}\mathbf{Q}e^{\mathbf{-A}^{\dagger}\tau}d\tau$
$\mathbf{S}\left(\int_{0}^{|\Delta t|}e^{-\bm{\Lambda}\tau}(\mathbf{S}^{-1}\mathbf{Q}\mathbf{S}^{-\dagger})e^{-\bm{\Lambda}^{\dagger}\tau}d\tau\right)\mathbf{S}^{\dagger}$
Let $\mathbf{G}:=\mathbf{S}^{-1}\mathbf{Q}\mathbf{S}^{-\dagger}$ and $\bm{\Lambda}_{\mathbf{V}B}(\Delta t):=\int_{0}^{|\Delta t|}e^{-\bm{\Lambda}\tau}(\mathbf{S}^{-1}\mathbf{Q}\mathbf{S}^{-\dagger})e^{-\bm{\Lambda}^{\dagger}\tau}d\tau$.
Note that $(e^{-\bm{\Lambda}t}\mathbf{G}e^{-\bm{\Lambda}^{\dagger}t})_{kl}=\mathbf{G}_{kl}e^{-\lambda_{k}t}e^{-\bar{\lambda}_{l}t}=\mathbf{G}_{kl}e^{-(\lambda_{k}+\bar{\lambda}_{l})t}$. Hence, the $kl$th component of $\bm{\Lambda}_{\mathbf{V}B}(\Delta t)$ is:
$\bm{\Lambda}_{\mathbf{V}B}(\Delta t)_{kl}=\mathbf{G}_{kl}\int_{0}^{|\Delta t|}e^{-(\lambda_{k}+\bar{\lambda}_{l})\tau}d\tau$
$\begin{cases}\mathbf{G}_{kl}\,\dfrac{1-e^{-(\lambda_{k}+\bar{\lambda}_{l})|\Delta t|}}{\lambda_{k}+\bar{\lambda}_{l}},&\lambda_{k}+\bar{\lambda}_{l}\neq 0\\[5.69054pt] \mathbf{G}_{kl}\,|\Delta t|,&\lambda_{k}+\bar{\lambda}_{l}=0\end{cases}$
If we assume that $\mathbf{Q}=\mathbf{S}\bm{\Lambda}_{Q}\mathbf{S}^{\dagger}$, i.e. that the process noise acts independently along the decoupled modes of the system, then $\mathbf{G}=\bm{\Lambda}_{Q}$ and $\bm{\Lambda}_{\mathbf{V}B}$ becomes diagonal:
$\bm{\Lambda}_{\mathbf{V}B}(|\Delta t_{ij}|)=\bm{\Lambda}_{Q}\frac{e^{-2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}-\mathbf{I}}{-2\mathrm{Re}(\bm{\Lambda})}$
where division is taken element-wise (ignoring, for now, the case $\mathrm{Re}(\lambda_{k})=0$).
For the causal case, the forward-propagated covariance is:
$\mathbf{V}_{F}(|\Delta t_{ij}|)=e^{\mathbf{A}|\Delta t_{ij}|}\mathbf{V}_{B}(|\Delta t_{ij}|)e^{\mathbf{A}^{\top}|\Delta t_{ij}|}$
$\mathbf{S}e^{\bm{\Lambda}|\Delta t_{ij}|}\bm{\Lambda}_{\mathbf{V}B}(|\Delta t_{ij}|)e^{\bm{\Lambda}^{\dagger}|\Delta t_{ij}|}\mathbf{S}^{\dagger}$
$\mathbf{S}e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}\left(\bm{\Lambda}_{Q}\frac{e^{-2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}-\mathbf{I}}{-2\mathrm{Re}(\bm{\Lambda})}\right)\mathbf{S}^{\dagger}$
$\mathbf{S}\bm{\Lambda}_{\mathbf{V}}(|\Delta t_{ij}|)\mathbf{S}^{\dagger}$
where
$\bm{\Lambda}_{\mathbf{V}}(|\Delta t_{ij}|)=\bm{\Lambda}_{Q}\frac{\mathbf{I}-e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}}{-2\mathrm{Re}(\bm{\Lambda})}$
$\bm{\Lambda}_{\mathbf{V}}(|\Delta t_{ij}|)$ is positive semi-definite provided $\bm{\Lambda}_{Q}\geq\mathbf{0}$ (which is true by definition). (Here, the inequalities are applied element-wise.)
The propagated precision matrix $\mathbf{P}(t)$ can be easily obtained by inverting $\bm{\Lambda}_{\mathbf{V}}(\Delta t)$ element-wise: $\mathbf{P}(\Delta t)=\mathbf{S}^{-\dagger}\bm{\Lambda}_{\mathbf{P}}(\Delta t)\mathbf{S}^{-1}$, where $\bm{\Lambda}_{\mathbf{P}}(\Delta t)=\bm{\Lambda}_{\mathbf{V}}(\Delta t)^{-1}$.
To include the measurement model, recall that the propagated measurement covariance is:
$\mathbf{V}_{C}(\Delta t_{ij})=\mathbf{C}\mathbf{V}(\Delta t_{ij})\mathbf{C}^{\dagger}+\mathbf{M}(\Delta t_{ij})\mathbf{R}\mathbf{M}^{\dagger}(\Delta t_{ij}),\quad\text{where}\quad\mathbf{M}(\Delta t_{ij})=\mathbf{C}e^{\mathbf{-A}\Delta t_{ij}}\mathbf{C}^{-1}.$
If $\mathbf{C}=\mathbf{S}\bm{\Lambda}_{C}\mathbf{S}^{-1}$, this becomes: $\mathbf{M}(\Delta t_{ij})=\mathbf{S}\bm{\Lambda}_{C}e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{C}^{-1}\mathbf{S}=\mathbf{S}e^{-\bm{\Lambda}\Delta t_{ij}}\mathbf{S}^{-1}$. If, in addition, $\mathbf{R}=\mathbf{S}\bm{\Lambda}_{R}\mathbf{S}^{\dagger}$, then:
$\mathbf{V}^{C}(\Delta t_{ij})=\mathbf{S}\bm{\Lambda}_{C}\bm{\Lambda}_{\mathbf{V}}(\Delta t_{ij})\bm{\Lambda}_{C}^{\dagger}\mathbf{S}^{\dagger}+\mathbf{S}e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{R}e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}\mathbf{S}^{\dagger}$
$\mathbf{S}\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})\mathbf{S}^{\dagger}$
where:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\bm{\Lambda}_{C}^{2}\bm{\Lambda}_{\mathbf{V}}(\Delta t_{ij})+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R}$
$\bm{\Lambda}_{C}^{2}\bm{\Lambda}_{Q}\frac{\mathbf{I}-e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}}{-2\mathrm{Re}(\bm{\Lambda})}+e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}\bm{\Lambda}_{R}$
(Stable dynamics attenuate the effect of measurement noise.)
Finally, $\mathbf{P}^{C}(\Delta t_{ij})=\mathbf{S}^{-\dagger}\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{ij})\mathbf{S}^{-1}$, where $\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{ij})=\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})^{-1}$.
Letting $\mathbf{z}_{\mathbf{s}_{i}}=\mathbf{S}^{-1}\mathbf{z}_{i}$, and $\mathbf{\hat{z}}_{\mathbf{s}ij}:=e^{-\bm{\Lambda}\Delta t_{ij}}\mathbf{z}_{\mathbf{s}j}$, the MLE is now tractable:
$\mathbf{\bar{z}}_{i}=\mathbf{S}\left[\left(\sum_{j\leq i}\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{ij})\right)^{-1}\sum_{j\leq i}\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{ij})\mathbf{\hat{z}}_{\mathbf{s}ij}\right]\mathbf{S}^{-1}.$
For the non-causal case,
$\bm{\Lambda}_{\mathbf{V}}(\Delta t)=\bm{\Lambda}_{\mathbf{V}B}(\Delta t)=\bm{\Lambda}_{Q}\frac{e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t}-\mathbf{I}}{-2\mathrm{Re}(\bm{\Lambda})}$
So,
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\bm{\Lambda}_{C}^{2}\bm{\Lambda}_{\mathbf{V}}(\Delta t_{ij})+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R}$
$\bm{\Lambda}_{C}^{2}\bm{\Lambda}_{Q}\frac{e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}-\mathbf{I}}{-2\mathrm{Re}(\bm{\Lambda})}+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R}$
In summary, the general expression for the propagated measurement covariance, including both the causal and non-causal cases, and all values of $\operatorname{Re}(\lambda)$ is:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\text{sign}(\Delta t_{ij})\bm{\Lambda}_{C}^{2}\bm{\Lambda}_{Q}\mathcal{K}(\bm{\Lambda},\Delta t_{ij})+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R},$
$\mathcal{K}(\bm{\Lambda},\Delta t):=\operatorname{diag}\!\big{(}\varphi(\lambda_{k},\Delta t)\big{)}_{k=1}^{d},\qquad\varphi(\lambda,\Delta t):=\begin{cases}\dfrac{e^{-2\operatorname{Re}(\lambda)\,\Delta t}-1}{-2\operatorname{Re}(\lambda)},&\operatorname{Re}(\lambda)\neq 0,\\[6.0pt] \Delta t,&\operatorname{Re}(\lambda)=0,\end{cases}$
This is bounded for all $\Delta t_{ij}$ in the causal case if and only if every component of $\mathrm{Re}(\bm{\Lambda})$ is negative.
In particular, consider $\mathbf{A}=\alpha\mathbf{I}+\bm{\Omega}\in\mathbb{R}^{d\times d}$ where $\alpha\in\mathbb{R}$ and $\bm{\Omega}\in\mathbb{R}^{d\times d}$ is diagonalizable over $\mathbb{C}$, and has strictly imaginary eigenvalues, i.e. $\bm{\Omega}=\mathbf{S}\bm{\Lambda}_{\Omega}\mathbf{S}^{-1}$, where $\mathbf{S}\in\mathbb{C}^{d\times d}$, $\bm{\Lambda}_{\Omega}=\text{diag}(\lambda_{\Omega,1},\dots\lambda_{\Omega,d})$, with $\lambda_{\Omega,k}\in i\mathbb{R}$. If we also assume that the process and measurement noise are isotropic $\bm{\Lambda}_{Q}=\sigma^{2}\mathbf{I}$, $\bm{\Lambda}_{R}=\eta^{2}\mathbf{I}$, and $\bm{\Lambda}_{C}=\lambda_{C}\mathbf{I}$, then we have:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\sigma_{V}^{2}(\Delta t_{ij})\mathbf{I},\quad\text{where}\quad\sigma_{V}^{2}(\Delta t_{ij}):=\text{sign}(\Delta t_{ij})\lambda_{C}^{2}\sigma^{2}\frac{e^{-2\alpha\Delta t_{ij}}-1}{-2\alpha}+e^{-2\alpha\Delta t_{ij}}\eta^{2}$
We can also generalize the above to the case where continuous-time measurements are incorporated. In this case, the covariance evolution is governed by the differential Riccati equation, which modifies the differential Lyapunov equation to include the measurement update:
$\mathbf{\dot{V}}_{F}(t)=\mathbf{A}\mathbf{V}_{F}(t)+\mathbf{V}_{F}(t)\mathbf{A}^{\top}+\mathbf{Q}-\mathbf{V}_{F}(t)\mathbf{C}^{\top}\mathbf{R}^{-1}\mathbf{C}\mathbf{V}_{F}(t)$
Again assuming simultaneously diagonalizble $\mathbf{A},\mathbf{C},\mathbf{Q},\mathbf{R}$, the matrix equations decouple into separate scalar Riccati differential equations for each diagonal element, with $k$th diagonal term given by:
$\dot{\lambda}_{vk}(t)=2\mathrm{Re}(\lambda_{i})\lambda_{vk}(t)+\lambda_{\Omega k}-\frac{\lambda_{Ck}^{2}}{\lambda_{\Gamma k}}\lambda_{vk}(t)^{2}$
Letting $\alpha_{k}=\mathrm{Re}(\lambda_{k})$, $\beta_{k}=\sqrt{\alpha_{k}^{2}+\frac{\lambda_{Qk}}{\lambda_{Rk}}\lambda_{Ck}^{2}}$, $\gamma_{k}=\frac{1}{\beta_{k}}\log\left(\frac{\beta_{k}-\alpha_{k}}{\beta_{k}+\alpha_{k}}\right)$, the $k$th element of $\bm{\Lambda}_{\mathbf{V}}$ becomes:
$\lambda_{vk}(\Delta t_{ij})=\frac{\lambda_{Rk}}{\lambda_{Ck}^{2}}\left(\beta_{k}\tanh\left(\frac{\beta_{k}}{2}\left(\Delta t_{ij}-\gamma_{k}\right)\right)+\alpha_{k}\right)$
2.6 Robust Maximum Likelihood Trajectory Estimation
Maximum Likelihood Estimation (MLE) of a system’s trajectory relies on accurate precision matrices, which represent the uncertainty of our estimates. If our dynamic model, defined by $\{\mathbf{S},\mathbf{S}^{-1},\bm{\Lambda},\bm{\Lambda}_{C},\bm{\Lambda}_{Q},\bm{\Lambda}_{R}\}$, does not perfectly match reality, these theoretical precision matrices may misrepresent the true uncertainty, leading to suboptimal state estimates.
One approach to mitigate model mis-specification is to adaptively adjust the precision matrix $\mathbf{P}^{C}(\Delta t_{ij})$ using observed residuals. Since $\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\left(\mathbf{z}^{C}_{i},\mathbf{V}^{C}_{ij}\right)$, it follows that the pair-wise residual $\mathbf{r}_{ij}:=\mathbf{z}_{i}^{C}-\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big{(}0,\mathbf{V}^{C}_{ij}\big{)}$. We generally can’t observe $\mathbf{z}_{i}^{C}$, but we can also define a measurement residual, $\mathbf{r_{z}}_{ij}:=\mathbf{z}_{i}-\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big{(}0,\mathbf{V}^{C}_{ij}+\mathbf{R})$. Note that these pair-wise residuals are distinct from the total innovation $\mathbf{y}_{ij}:=\mathbf{z}_{i}^{C}-\mathbf{\bar{z}}_{ij}$. Since $\mathrm{Cov}\!\left(\mathbf{y}_{i}\right)=\left(\sum_{j}\mathbf{P}^{C}_{ij}\right)^{-1}\preceq\big{(}\mathbf{P}^{C}_{ij}\big{)}^{-1},\,\forall j$, the combined estimator’s innovation covariance is no larger than the residual covariance of any single predictor.
2.6.1 Variational Formulation of Robust Estimation
Suppose we have vector-valued measurements:
$\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big{(}\bm{\mu}_{i},\mathbf{\Sigma}_{ij}\big{)},$
where $\bm{\mu}_{i}\in\mathbb{R}^{d}$ is the output of the true latent state at time $i$ (which we distinguish from $\mathbf{z}^{C}_{i}$, which denotes the output under the latent state generated by our LTI model), and $\mathbf{\Sigma}_{ij}$ is the unknown measurement covariance. Our objective is to maximize the marginal log-likelihood:
$\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i})\;=\;\log\int p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{ij})\,p(\mathbf{\Sigma}_{ij})\,d\mathbf{\Sigma}_{ij}.$
i.e. a Gaussian scale mixture. Introducing a variational distribution $q(\mathbf{\Sigma}_{ij})$ over the unknown covariance and applying Jensen’s inequality yields the Evidence Lower Bound (ELBO):
$\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i})\geq\mathcal{L}(q,\bm{\mu}_{i}):=\mathbb{E}_{q(\mathbf{\Sigma}_{ij})}\big{[}\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{ij})\big{]}-\mathrm{KL}\big{(}q(\mathbf{\Sigma}_{ij})\,\|\,p(\mathbf{\Sigma}_{ij})\big{)}.$
Assuming $\mathbf{\Sigma}_{ij}$ is diagonal, the expected log-likelihood term becomes:
$\mathbb{E}_{q}\big{[}\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{ij})\big{]}=-\frac{1}{2}\mathbb{E}_{q}[\log\det\mathbf{\Sigma}_{ij}]-\frac{1}{2}(\mathbf{\hat{z}}_{ij}-\bm{\mu}_{i})^{\top}\mathbb{E}_{q}[\mathbf{\Sigma}_{ij}^{-1}](\mathbf{\hat{z}}_{ij}-\bm{\mu}_{i})+\text{const}.$
Define the expected precision: $\mathbf{W}_{ij}:=\mathbb{E}_{q}\big{[}\mathbf{\Sigma}_{ij}^{-1}\big{]}.$
Then, the expected log-likelihood is proportional to a weighted quadratic form:
$\mathbb{E}_{q}\big{[}\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{ij})\big{]}\;\propto\;-\frac{1}{2}(\mathbf{\hat{z}}_{ij}-\bm{\mu}_{i})^{\top}\mathbf{W}_{ij}(\mathbf{\hat{z}}_{ij}-\bm{\mu}_{i}).$
We model the unknown measurement precision as a scaled version of the dynamics-based precision:
$\mathbf{\Sigma}_{ij}^{-1}=\tau_{ij}\,\mathbf{P}^{C}_{ij},$
where $\mathbf{P}^{C}_{ij}$ is the precision predicted by the LTI model and $\tau_{ij}>0$ is a latent scalar that modulates the overall confidence in this observation.
Conditional on $\tau_{ij}$, the Gaussian likelihood for the measurement $\mathbf{\hat{z}}_{ij}$ is:
$p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\tau_{ij})=\mathcal{N}\big{(}\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},(\tau_{ij}\mathbf{P}^{C}_{ij})^{-1}\big{)}.$
The optimal variational distribution over $\tau_{ij}$ is proportional to the prior times the likelihood:
$q^{*}(\tau_{ij})\propto p(\tau_{ij})\,p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\tau_{ij}).$
To enable Bayesian inference, we place a conjugate prior on $\tau_{ij}$. Using a Gamma prior:
$\tau_{ij}\sim\mathrm{Gamma}(\alpha,\beta).$
Since the prior is Gamma and the likelihood is Gaussian with a precision multiplier, the posterior is again a Gamma distribution:
$q^{*}(\tau_{ij})=\mathrm{Gamma}\!\Big{(}\alpha+\frac{d}{2},\;\beta+\frac{1}{2}\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}\Big{)},\quad\mathbf{r}_{ij}:=\mathbf{\hat{z}}_{ij}-\bm{\mu}_{i}.$
The expected precision under the posterior is:
$\mathbf{W}_{ij}:=\mathbb{E}_{q^{*}}[\mathbf{\Sigma}_{ij}^{-1}]=\mathbb{E}_{q^{*}}[\tau_{ij}]\,\mathbf{P}^{C}_{ij}=w_{ij}\,\mathbf{P}^{C}_{ij},$
where:
$w_{ij}=\frac{\alpha+d/2}{\beta+\frac{1}{2}\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}}.$
The corresponding (marginal) likelihood $p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i})$ is a Student’s $t$. This is equivalent to an M-estimator with loss
$\rho(\mathbf{r}_{ij})\;=\;\Big{(}\alpha+\tfrac{d}{2}\Big{)}\,\log\!\Big{(}\beta+\tfrac{1}{2}\,\mathbf{r}_{ij}^{\top}\mathbf{P}_{ij}\,\mathbf{r}_{ij}\Big{)},\qquad\mathbf{r}_{ij}:=\bm{\mu}_{i}-\mathbf{\hat{z}}_{ij}.$
Dividing numerator and denominator by $\beta$ and setting the numerator to $1$ (since it is canceled by the normalization), there is only one independent constant, $\nu=2\beta$, which controls the residual scaling:
$w_{ij}=\frac{1}{1+\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}/\nu}.$
The term $\mathbf{r}^{\top}\mathbf{V}^{-1}\mathbf{r}$ is known as the Mahalanobis distance, which quantifies the difference between a vector $\mathbf{r}$ and a distribution with covariance $\mathbf{V}$. Here, $\mathbf{P}^{C}_{ij}$ encodes the directional confidence from the LTI model, while $w_{ij}$ adaptively rescales the precision based on the observed residual $\mathbf{r}_{ij}$, down-weighting measurements with larger than expected deviations.
The metric $1/(1+\mathbf{r}^{\top}\mathbf{P}\mathbf{r}/\nu)$ behaves similarly to the unnormalized softmax: if one residual is much smaller than the others, it dominates, while if all residuals are similar and not too large, the output approximates a uniform weighting. The key difference arises when all residuals are small relative to $\nu$: unlike softmax, which picks a dominant term whenever relative differences are large, this similarity score reflects uncertainty and maintains a roughly uniform attention distribution. In this sense, $\nu$ acts as a kind of noise floor.
Using the fact that $\log(1+x)\approx x\quad\text{for }x\ll 1$, we can expand the rational form of $w_{ij}$:
$w_{ij}=\frac{1}{1+\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}/\nu}\approx\exp(-\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}/\nu).$
which results in a more aggressive down-weighting of large residuals (and $\nu$ becomes a temperature parameter).
2.6.2 Iteratively-Reweighted Least Squares
As we noted above, we cannot observe the true latent state $\bm{\mu}_{i}$, so we cannot compute the residual $\mathbf{r}_{ij}:=\mathbf{\hat{z}}_{ij}-\bm{\mu}_{i}$. However, we can use our current best estimate of $\bm{\mu}_{i}$, which we denote $\mathbf{\hat{z}}_{ii}$, and a corresponding residual: $\mathbf{\hat{r}}_{ij}:=\mathbf{\hat{z}}_{ij}-\mathbf{\hat{z}}_{ii}$. This leads naturally to an iteratively-reweighted least squares (IRLS)-style update for the latent state (at iteration $k$):
$\mathbf{\hat{z}}^{(k+1)}_{ii}=\Big{(}\sum_{j}w_{ij}^{(k)}\mathbf{P}^{C(k)}_{ij}\Big{)}^{-1}\sum_{j}w_{ij}^{(k)}\mathbf{P}^{C(k)}_{ij}\mathbf{\hat{z}}_{ij}^{(k)},$
Recall that (by construction) $\mathbf{P}_{ij}^{C(k)}=\mathbf{S}^{-\dagger}\bm{\lambda}_{\mathbf{P}ij}^{C}\mathbf{S}^{-1}$. Let $\mathbf{z}_{\mathbf{s}i}:=\mathbf{S}^{-1}\mathbf{z}_{i}$ and $\mathbf{\hat{z}}_{\mathbf{s}ij}=\mathbf{S}^{-1}\mathbf{\hat{z}}_{ij}=e^{-\bm{\Lambda}\Delta t_{ij}}$, define a corresponding residual $\mathbf{r_{zs}}_{ij}:=\mathbf{z}_{\mathbf{s}i}-\mathbf{\hat{z}}_{\mathbf{s}ij}$, and let $\bm{\lambda}_{\mathbf{P}ij}^{C}$ be the diagonal of $\bm{\lambda}_{\mathbf{P}ij}^{C}$. We have an iterative update rule for the state estimate in the eigenbasis:
$\mathbf{\hat{z}}_{\mathbf{s}ii}^{(k+1)}=\mathbf{\hat{z}}_{\mathbf{s}i}^{\text{MLE}(k)},\quad\text{where}\quad\mathbf{\hat{z}}_{\mathbf{s}i}^{\text{MLE}(k)}=\left(\sum_{j}w_{ij}^{(k)}\bm{\lambda}_{\mathbf{P}ij}^{C}\right)^{-1}\odot\sum_{j}w_{ij}^{(k)}\bm{\lambda}_{\mathbf{P}ij}^{C}\odot\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)},$
$w_{ij}^{(k)}=\frac{1}{1+(\bm{\lambda}_{\mathbf{P}ij}^{C})^{\top}\|\mathbf{r}^{(k)}_{\mathbf{s}ij}\|^{2}/\nu},\quad\text{where}\quad\mathbf{r}_{\mathbf{s}ij}^{(k)}:=\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ii}-\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ij}$.
and initial condition $\mathbf{\hat{z}}_{\mathbf{s}ii}^{(1)}=\mathbf{z}_{\mathbf{s}i}$. (Alternatively, if we only have a single iteration, we can replace $\bm{\lambda}_{\mathbf{P}ij}^{C}$ with $\bm{\lambda}_{\mathbf{P}ij}^{C}+\bm{\lambda}_{R}$ (the covariance of $\mathbf{z}_{\mathbf{s}i}$).) To stabilize the iteration, we can instead use an affine combination of the current and previous estimate:
$\mathbf{\hat{z}}^{(k+1)}_{\mathbf{s}ii}=(1-\delta)\mathbf{\hat{z}}_{\mathbf{s}ii}^{(k)}+\delta\mathbf{\hat{z}}_{\mathbf{s}i}^{\text{MLE}(k)},$
This IRLS iteration, with precision matrices derived from an LTI model, performs a batched maximum likelihood update of latent states using pairwise uncertainty propagation. We will refer to it as Kalman-IRLS, as it is closely related to adaptive Kalman Filtering, although unlike the latter, it does not rely on recursive updates. It can be implemented with a multi-layer neural network with learnable parameters $\mathbf{S},\bm{\Lambda},\bm{\Lambda}_{Q},\bm{\Lambda}_{R}$ shared across layers. If $\mathbf{\hat{z}}^{(k)}_{ii}\rightarrow\mathbf{z}^{C}$ in expectation with increasing $k$, we can take $\mathbf{\bar{z}}_{\mathbf{s}}\approx\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ii}$ at the last layer, and then predict the measurement at the next time step in the eigenbasis using the (diagonalized) matrix exponential:
$\mathbf{z}^{\text{pred}}_{\mathbf{s}i}=\bm{\Lambda}_{C}e^{\bm{\Lambda}(t_{i+1}-t_{i})}\bm{\Lambda}_{C}^{-1}\mathbf{\bar{z}}_{\mathbf{s}i}=e^{\bm{\Lambda}(t_{i+1}-t_{i})}\mathbf{\bar{z}}_{\mathbf{s}i}$
Finally, we compute the output prediction: $\mathbf{z}^{\text{pred}}_{i}=\mathbf{S}\mathbf{z}^{\text{pred}}_{\mathbf{s}i}$, which we then compare to the actual measurement at the next time step $\mathbf{z}_{i+1}$. We can then define a loss, e.g. $\mathcal{L}=\sum_{i}\|\mathbf{z}^{\text{pred}}_{i}-\mathbf{z}_{i+1}\|_{2}^{2}$ and train our model parameters $\mathbf{S},\bm{\Lambda},\bm{\Lambda}_{Q},\bm{\Lambda}_{R}$ using backpropagation.
2.6.3 Iteratively-Reweighted Kalman Filter (IRKF)
We now describe a way to incorporate state-dependent weighting from the full batch MLE while retaining the recursive structure of the Kalman Filter. The key idea is to use the innovation residual at each step to reweight the update, and iterate within each time step to approximate the effect of all pairwise comparisons. This iteration again takes place in the eigenbasis.
Prediction Step (Include Process Noise).
At time $t_{i}$, suppose we have the posterior from the previous time step $\mathbf{\bar{z}}_{\mathbf{s},i-1}^{+}$ with covariance $\bm{\Lambda}_{\mathbf{V},i-1}^{+}$ and precision $\bm{\Lambda}_{\mathbf{P},i-1}^{+}=(\bm{\Lambda}_{\mathbf{V},i-1}^{+})^{-1}$. Propagating through the linear dynamics with process noise $\bm{\Lambda}_{Q}$, the prior for the current time step becomes:
$\mathbf{\bar{z}}_{\mathbf{s}i}^{-}=\bm{\Lambda}\mathbf{\bar{z}}_{\mathbf{s},i-1}^{+},\quad\bm{\Lambda}_{\mathbf{V},i}^{-}=\bm{\Lambda}\bm{\Lambda}_{\mathbf{V},i-1}^{+}\bm{\Lambda}^{\dagger}+\bm{\Lambda}_{Q}=\|\bm{\Lambda}\|^{2}\bm{\Lambda}_{\mathbf{V},i-1}^{+}+\bm{\Lambda}_{Q},\quad\bm{\Lambda}_{\mathbf{P},i}^{-}=(\bm{\Lambda}_{\mathbf{V},i}^{-})^{-1}.$
The propagated precision for a single time step is constant: $\bm{\Lambda}_{\mathbf{P}}^{C}:=\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{i,i-1})$.
Weighted Update (Measurement + IRKF Iteration).
Within each iteration $k$ at time $t_{i}$, compute the innovation:
$\mathbf{y}_{\mathbf{s}i}^{(k)}=\mathbf{z}_{\mathbf{s}i}-\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}.$
Define the nominal posterior precision from the measurement model:
$\bm{\Lambda}_{\mathbf{P},i}^{+(k)}=\bm{\Lambda}_{\mathbf{P},i}^{-}+\bm{\Lambda}_{\mathbf{P}}^{C},$
and compute a residual-based robust weight:
$w_{i}^{(k)}=\frac{1}{1+(\mathbf{y}_{\mathbf{s}i}^{(k)})^{\dagger}\bm{\Lambda}_{\mathbf{P},i}^{+(k)}\mathbf{y}_{\mathbf{s}i}^{(k)}/\nu}.$
The weighted posterior precision is then:
$\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)}=\bm{\Lambda}_{\mathbf{P},i}^{-}+w_{i}^{(k)}\,\bm{\Lambda}_{\mathbf{P}}^{C},$
Notice that the posterior information vector can be written as:
$\bm{\tilde{\eta}}_{\mathbf{s}i}^{+}=\bm{\eta}_{\mathbf{s}i}^{-}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{z}_{\mathbf{s}i}=\bm{\Lambda}_{\mathbf{P},i}^{-}\mathbf{\bar{z}}_{\mathbf{s}i}^{-}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{z}_{\mathbf{s}i}$
$\bm{\Lambda}_{\mathbf{P},i}^{-}\mathbf{\bar{z}}_{\mathbf{s}i}^{-}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{y}_{\mathbf{s}i}^{(k)}.$
So the total posterior mean in the eigenbasis becomes:
$\mathbf{\bar{z}}_{\mathbf{s}i}^{+(k)}=(\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)})^{-1}\bm{\tilde{\eta}}_{\mathbf{s}i}^{+}$
$(\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)})^{-1}\Big{[}\bm{\Lambda}_{\mathbf{P},i}^{+(k)}\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{y}_{\mathbf{s}i}^{(k)}\Big{]}.$
$(\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)})^{-1}\Big{[}\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)}\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{y}_{\mathbf{s}i}^{(k)}\Big{]}=\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}+w_{i}^{(k)}(\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)})^{-1}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{y}_{\mathbf{s}i}^{(k)}.$
(the robust weight only scales the measurement contribution, not the prior).
Alternatively, we can use a damped update:
$\mathbf{\bar{z}}_{\mathbf{s}i}^{+(k)}=\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}+\delta\,w_{i}^{(k)}(\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)})^{-1}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{y}_{\mathbf{s}i}^{(k)}.$
where $\delta\in(0,1]$. Set the prior for the next inner iteration as: $\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k+1)}=\mathbf{\bar{z}}_{\mathbf{s}i}^{+(k)}$.
Iterate until convergence to produce the final posterior $\mathbf{\bar{z}}_{\mathbf{s}i}^{+}$ and $\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+}$ for the current time step.
While the batch MLE uses the global information provided by pair-wise residuals, the innovation-based weight $w_{i}$ in the recursive update depends only on the current residual $r_{i}=\mathbf{z}_{i}-\mathbf{\bar{z}}_{i}^{-}$. However, when the dynamics are slow (small eigenvalues), $\Delta t$ is sufficiently small, and the residuals are not too large, $\mathbf{\bar{z}}_{\mathbf{s}i}^{-}$ effectively encodes past measurements. In this case, with enough inner iterations, IRKF should approximate the batch IRLS solution, while retaining the recursive structure of the Kalman Filter. The number of pair-wise comparisons for a sequence of length $m$ is then reduced from $\mathcal{O}(m^{2})$ to $\mathcal{O}(Km)$, where $K\ll m$ is the number of inner iterations.
3 Adaptive Filter Attention
3.1 Generalizing the Adaptive Filter to a Tensor Form of Attention
We now generalize the batch MLE form of the above adaptive filter using learned, complex-valued query, key, value, and output matrices: $\mathbf{W_{q}}\in\mathbb{C}^{d_{k}\times d_{e}}$, $\mathbf{W_{k}}\in\mathbb{C}^{d_{k}\times d_{e}}$, $\mathbf{W_{v}}\in\mathbb{C}^{d_{v}\times d_{e}}$, $\mathbf{W_{p}}\in\mathbb{C}^{d_{e}\times d_{v}}$, where $d_{e}$ is the input embedding dimension $d_{k}$ is the query/key embedding dimension, and $d_{v}$ is the value embedding dimension.
In this formulation, $\mathbf{W_{q}}$, $\mathbf{W_{k}}$, $\mathbf{W_{v}}$ replace $\mathbf{S}^{-1}$ and $\mathbf{W_{p}}$ replaces $\mathbf{S}$. When $\mathbf{W_{q}}=\mathbf{W_{k}}=\mathbf{W_{v}}=\mathbf{W_{p}}^{-1}$, the attention layer corresponds to an adaptive MLE. By allowing these parameter matrices to vary freely, the network can learn approximate inverses as needed, enabling attention to implement an MLE while retaining the flexibility to represent richer transformations.
We can define separate key/query and value dynamics parameters: $\bm{\lambda}_{\mathbf{qk}},\bm{\lambda}_{Q\mathbf{qk}},\bm{\lambda}_{R\mathbf{qk}},\bm{\lambda}_{C\mathbf{qk}}\in\mathbb{R}^{d_{k}}$; $\bm{\lambda}_{\mathbf{v}},\bm{\lambda}_{Q\mathbf{v}},\bm{\lambda}_{R\mathbf{v}},\bm{\lambda}_{C\mathbf{v}}\in\mathbb{R}^{d_{v}}$. If we set $d_{k}=d_{v}$, then we can use a shared set of parameters for the key and value dynamics: $\bm{\lambda},\bm{\lambda}_{Q},\bm{\lambda}_{R},\bm{\lambda}_{C}\in\mathbb{R}^{d_{v}}$. Note that we can absorb $\bm{\lambda}_{C}^{2}$ into $\bm{\lambda}_{Q}$, but we keep it here for clarity.
The inputs to the network are: $\quad\mathbf{Z}=\left[\mathbf{z}_{1},\dots,\mathbf{z}_{m}\right]\in\mathbb{C}^{d_{e}\times m}$.
We compute the following:
Query, key, and value matrices:
$\mathbf{Z_{q}}=\mathbf{W_{q}}\mathbf{Z}\in\mathbb{C}^{d_{k}\times m},\quad\mathbf{Z_{k}}=\mathbf{W_{k}}\mathbf{Z}\in\mathbb{C}^{d_{k}\times m},\quad\mathbf{Z_{v}}=\mathbf{W_{v}}\mathbf{Z}\in\mathbb{C}^{d_{v}\times m}$
Exponential tensors: $\mathcal{E}_{\mathbf{qk}}\in\mathbb{C}^{d_{k}\times m\times m},\quad\mathcal{E}_{\mathbf{v}}\in\mathbb{C}^{d_{v}\times m\times m}$:
$\mathcal{E}_{\mathbf{qk}}[k,i,j]=e^{\lambda_{\mathbf{qk}k}(t_{i}-t_{j})},\quad\mathcal{E}_{\mathbf{v}}[k,i,j]=e^{\lambda_{\mathbf{v}k}(t_{i}-t_{j})}$
"Pulled-forward" estimates: $\mathcal{\hat{Z}}_{\mathbf{k}}\in\mathbb{C}^{d_{k}\times m\times m},\quad\mathcal{\hat{Z}}_{\mathbf{v}}\in\mathbb{C}^{d_{v}\times m\times m}$:
$\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]=\mathcal{E}_{\mathbf{qk}}[k,i,j]\cdot\mathbf{Z_{k}}[k,j],\quad\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]=\mathcal{E}_{\mathbf{v}}[k,i,j]\cdot\mathbf{Z_{v}}[k,j],$
The residuals: $\mathcal{R}\in\mathbb{C}^{d_{k}\times m\times m}$:
$\mathcal{R}_{\mathbf{qk}}[k,i,j]=\mathbf{Z_{q}}[k,i]-\hat{\mathcal{Z}}_{\mathbf{k}}[k,i,j]$
The norms of the residuals:
$\|\mathcal{R}_{\mathbf{qk}}[k,i,j]\|^{2}=\mathrm{Re}^{2}\left(\mathcal{R}_{\mathbf{qk}}[k,i,j]\right)+\mathrm{Im}^{2}\left(\mathcal{R}_{\mathbf{qk}}[k,i,j]\right)$
The key and value precision tensors: $\mathcal{P}^{C}_{\mathbf{qk}}\in\mathbb{R}^{d_{k}\times m\times m},\quad\mathcal{P}^{C}_{\mathbf{v}}\in\mathbb{R}^{d_{v}\times m\times m}$:
$\mathcal{P}^{C}_{\mathbf{qk}}[:,i,j]=\left(\bm{\lambda}_{C\mathbf{qk}}^{2}\odot\bm{\lambda}_{Q\mathbf{qk}}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})|t_{j}-t_{i}|}}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})}+\bm{\lambda}_{R\mathbf{qk}}\odot e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})|t_{j}-t_{i}|}\right)^{-1}$
$\mathcal{P}^{C}_{\mathbf{v}}[:,i,j]=\left(\bm{\lambda}_{C\mathbf{v}}^{2}\odot\bm{\lambda}_{Q\mathbf{v}}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})|t_{j}-t_{i}|}}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})}+\bm{\lambda}_{R\mathbf{v}}\odot e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})|t_{j}-t_{i}|}\right)^{-1}$
We apply a causal attention mask to $\mathcal{P}^{C}_{\mathbf{qk}},\mathcal{P}^{C}_{\mathbf{v}}$, masking out all $j>i$.
Compute the weights: $\mathbf{W}\in\mathbb{R}^{m\times m}$:
$\mathbf{W}[i,j]=\left(1+\sum_{k}\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]\cdot\|\mathcal{R}_{\mathbf{qk}}[k,i,j]\|^{2}/\nu\right)^{-1}$
Here, $\nu$ is a learnable scalar. The value parameters $\mathbf{W_{v}},\bm{\lambda}_{\mathbf{v}},\bm{\lambda}_{Q\mathbf{v}},\bm{\lambda}_{R\mathbf{v}}$ are responsible for the prior, while the query and key parameters $\mathbf{W_{q}},\mathbf{W_{k}},\bm{\lambda}_{\mathbf{qk}},\bm{\lambda}_{Q\mathbf{qk}},\bm{\lambda}_{R\mathbf{qk}}$ are responsible for the adaptive reweighting.
Compute the unnormalized attention scores: $\tilde{\mathcal{A}}\in\mathbb{R}^{d_{v}\times m\times m}$:
$\tilde{\mathcal{A}}[k,i,j]=\mathbf{W}[i,j]\cdot\mathcal{P}_{\mathbf{v}}^{C}[k,i,j]$
Normalize to obtain the attention tensor: $\mathcal{A}\in\mathbb{R}^{d_{v}\times m\times m}$:
$\mathcal{A}[k,i,j]=\tilde{\mathcal{A}}[k,i,j]\big{/}\sum_{j^{\prime}\leq i}\tilde{\mathcal{A}}[k,i,j^{\prime}]$
Compute the estimate: $\mathbf{\bar{Z}}\in\mathbb{C}^{d_{v}\times m}$:
$\mathbf{\bar{Z}}[k,i]=\sum_{{j\leq i}}\mathcal{A}[k,i,j]\cdot\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$
Optionally, add a residual connection:
$\mathbf{\bar{Z}}[k,i]\leftarrow(\mathbf{1}-\bm{\delta})\odot\mathbf{Z}_{\mathbf{v}}[k,i]+\bm{\delta}\odot\mathbf{\bar{Z}}[k,i],\quad\bm{\delta}\in(0,1]^{d_{v}}$
After $n$ layers with shared weights, compute the prediction $\mathbf{Z}^{\text{pred}}\in\mathbb{C}^{d_{e}\times m}$ using a one-step forward transition:
$\mathbf{Z}_{\mathbf{v}}^{\text{pred}}[k,i]=\mathcal{E}_{\mathbf{v}}^{\text{pred}}[k,i+1,i]\cdot\mathbf{\bar{Z}}[k,i]$
Map back to the original basis:
$\mathbf{Z}^{\text{pred}}[k,i]=\mathbf{W_{p}}\mathbf{Z}_{\mathbf{v}}^{\text{pred}}[k,i]$
This adaptive filter generalizes attention by introducing a more structured similarity measure that integrates uncertainty and time decay. If $\bm{\lambda}_{\mathbf{v}},\bm{\lambda}_{\mathbf{k}},\bm{\lambda}_{Q\mathbf{v}},\bm{\lambda}_{Q\mathbf{v}}\rightarrow\mathbf{0}$, the relative temporal decay on the attention weights disappears and we are left with a tensor generalization of standard attention.
By using the same (diagonal) matrix exponential to filter and to predict the next step, we enforce self-consistency in the model. One issue is that it may be preferable to do multiple steps of iterative refinement of the estimated trajectory before prediction. One way to address this is to give the model extra flexibility by outputting a gated interpolation between estimate and prediction. That is:
$\mathbf{\hat{z}}^{k+1}_{ii}=\bigl{(}1-\mathbf{g}^{(k)}\bigr{)}\,\mathbf{e}_{i}^{(k)}+\mathbf{g}^{(k)}\,\mathbf{p}_{i}^{(k)}.$
where $\mathbf{e}_{i}^{(k)}$ is the estimate, $\mathbf{p}_{i}^{(k)}$ is the prediction, and $\mathbf{g}^{(k)}\in(0,1]^{d_{v}}$ is a learnable gate. We can also make $\mathbf{g}^{(k)}$ a function of local residuals or features, for example:
$\mathbf{g}^{(k)}=\sigma\!\left(\mathbf{a}^{\top}\phi\!\left(\|\mathbf{r}^{(k)}\|^{2}\right)+\mathbf{b}\right),$
so the layer decides (data-dependently) whether to predict or to keep refining the estimate.
3.2 Complexity
Vanilla attention with input length $m$ and embedding dimension $d$ has time complexity $\mathcal{O}(m^{2}d)$ to compute attention scores and apply them across the sequence, plus $\mathcal{O}(md^{2})$ to transform the inputs into query, key, and value vectors and to transform the outputs through a linear layer. The storage of the query, key, and value matrices requires $\mathcal{O}(md)$ memory, and the attention matrix requires $\mathcal{O}(m^{2})$ memory.
The above tensor form of attention involves element-wise multiplication of $d\times m\times m$ tensors, resulting in time complexity $\mathcal{O}(m^{2}d+md^{2})$, which is the same as vanilla attention, but slower in practice due to less-optimized element-wise operations. The time complexity to propagate the precision matrices $\mathbf{P}_{ij}^{C}$ through the differential Lyapunov equation in the full (non-diagonalizable case) is $\mathcal{O}(d^{3})$, plus another $\mathcal{O}(d^{3})$ to invert each of these matrices, for each of the $\mathcal{O}(m^{2})$ time differences. Hence, the total time complexity is $\mathcal{O}(m^{2}d^{3})$. In contrast, for systems that admit a closed-form solution, the propagation and matrix inversions become $\mathcal{O}(d)$, so the total time complexity is $\mathcal{O}(m^{2}d)$. If the measurements are sampled at equal time intervals, the time complexity reduces to $\mathcal{O}(md)$.
Hence, the main bottleneck is not run-time, but memory, as we need to store all $m^{2}$ residuals, pulled-forward estimates, and the attention tensor, which are all tensors of size $\mathcal{O}(m^{2}d)$. In the following sections, we examine how to simplify the algorithm to avoid this bottleneck.
3.3 Convolutional Representation
If the time steps are equal, i.e. $\Delta t=t_{i+1}-t_{i}$ is fixed for all $i$, we can store all $ij$ exponentials in a kernel of size 2(m1)d2(m-1)d, or (m1)d(m-1)d in the causal case, rather than $m^{2}d$. In particular, we compute: $\mathcal{K}_{\mathbf{v}}[\tau]=e^{\bm{\lambda}_{\mathbf{v}}\Delta t\tau}$ and $\mathcal{K}_{\mathbf{qk}}[\tau]=e^{\bm{\lambda}_{\mathbf{qk}}\Delta t\tau}$, for $\tau\in[m-1,-(m-1)]$:
$\mathcal{K}_{\mathbf{v}}=\left[e^{\bm{\lambda}_{\mathbf{v}}(m-1)\Delta t},e^{\bm{\lambda}_{\mathbf{v}}(m-2)\Delta t},\dots,e^{\bm{\lambda}_{\mathbf{v}}\Delta t},e^{\mathbf{0}},e^{-\bm{\lambda}_{\mathbf{v}}\Delta t},\dots,e^{-\bm{\lambda}_{\mathbf{v}}(m-1)\Delta t},e^{-\bm{\lambda}_{\mathbf{v}}m\Delta t}\right],$
or in the causal case, $\tau\in[m-1,0]$:
$\mathcal{K}_{\mathbf{v}}=\left[e^{\bm{\lambda}_{\mathbf{v}}(m-1)\Delta t},e^{\bm{\lambda}_{\mathbf{v}}(m-2)\Delta t},\dots,e^{\bm{\lambda}_{\mathbf{v}}\Delta t},1,\dots,1\right],$
(where we pad the kernel with $m$ 1s for easier indexing.) We define $\mathcal{K}_{\mathbf{qk}}$ similarly. We can then efficiently compute the exponentials using a Toeplitz matrix to index from $\mathcal{K}_{\mathbf{v}}$ and $\mathcal{K}_{\mathbf{qk}}$. In particular, $e^{\bm{\lambda}_{\mathbf{v}k}(t_{i}-t_{j})}=\mathcal{K}_{\mathbf{v}}[m-(i-j)]$. The estimate tensors are then: $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]=\mathcal{K}_{\mathbf{v}}[m-(i-j)]\cdot\mathbf{Z}_{\mathbf{v}}[k,j]$ and $\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]=\mathcal{K}_{\mathbf{qk}}[m-(i-j)]\cdot\mathbf{Z}_{\mathbf{k}}[k,j]$.
Similarly, the precision tensor is Toeplitz, so we can again represent it with a kernel of size $md$:
$\mathcal{K}_{\mathbf{qk}}^{V}[\tau]=\bm{\lambda}_{C\mathbf{qk}}^{2}\odot\bm{\lambda}_{Q\mathbf{qk}}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})\Delta t\tau}}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})}+\bm{\lambda}_{R\mathbf{qk}}\odot e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})\Delta t\tau}$
which we can express in terms of $\mathcal{K}_{\mathbf{qk}}[\tau]$:
$\mathcal{K}_{\mathbf{qk}}^{V}[\tau]=\bm{\lambda}_{C\mathbf{qk}}^{2}\odot\bm{\lambda}_{Q\mathbf{qk}}\odot\frac{\mathbf{1}-\mathcal{K}^{\dagger}_{\mathbf{k}}[\tau]\mathcal{K}_{\mathbf{qk}}[\tau]}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})}+\bm{\lambda}_{R\mathbf{qk}}\odot\big{(}\mathcal{K}^{\dagger}_{\mathbf{k}}[\tau]\mathcal{K}_{\mathbf{qk}}[\tau]\big{)}$
The full precision tensor is then:
$\mathcal{P}_{\mathbf{k}}^{C}[k,i,j]=\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|],$
where $\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]=\mathcal{K}_{\mathbf{qk}}^{V}[k,|i-j|]^{-1}$, and similarly for $\mathcal{K}_{\mathbf{v}}^{P}[\tau]$ and $\mathcal{P}_{\mathbf{v}}^{C}$.
However, we must still store the full tensors of pulled-forward estimates $\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]$ and $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$, which are of size $d_{k}\times m\times m$ and $d_{v}\times m\times m$, respectively.
3.4 Broadcasting the Estimates
To avoid explicitly constructing the pulled-forward estimates $\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]$ and $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$, we can define separate arrays for the $i$ and $j$ components:
$\bm{\Phi}_{\mathbf{v}}[k,i]:=e^{\bm{\lambda}_{\mathbf{v}k}t_{i}},\quad\mathbf{\hat{Z}}_{\mathbf{v}}[k,j]:=e^{-\bm{\lambda}_{\mathbf{v}k}t_{j}}\mathbf{Z_{v}}[k,j].$
We can then factorize $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$:
$\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]=\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\mathbf{\hat{Z}}_{\mathbf{v}}[k,j],$
And likewise for $\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]$. This requires $2md$ computations rather than $md$, but reduces the memory requirement from $m^{2}d$ to $2md$.
Since $\bm{\Phi}_{\mathbf{v}}[k,i]$ does not depend on $j$, we can pull it outside the sum:
$\mathbf{\bar{Z}_{v}}[k,i]:=\sum_{j\leq i}\mathcal{A}[k,i,j]\cdot\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]=\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\sum_{j\leq i}\mathcal{A}[k,i,j]\cdot\mathbf{\hat{Z}}_{\mathbf{v}}[k,j]$
So we can avoid forming the full $d_{v}\times m\times m$ tensor $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$.
3.5 Factorizing the Residuals
Recall that:
$\mathcal{R}_{\mathbf{qk}}[k,i,j]=\mathbf{Z_{q}}[k,i]-\hat{\mathcal{Z}}_{\mathbf{k}}[k,i,j]$
Letting:
$\bm{\Phi}_{\mathbf{qk}}[k,i]=e^{\bm{\lambda}_{\mathbf{qk}}t_{i}},\quad\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]=e^{-\bm{\lambda}_{\mathbf{qk}}t_{j}}\mathbf{Z}_{\mathbf{k}}[k,j],\quad\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]=e^{-\bm{\lambda}_{\mathbf{qk}}t_{i}}\mathbf{Z}_{\mathbf{q}}[k,i],$
[…]
where:
[…]
We need to compute the robust weights:
[…]
$\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\left(\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]-\mathbf{\bar{\hat{Z}}}_{\mathbf{k}}[k,j]\right)\cdot\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]\cdot\left(\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]-\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)$
(where the overbar denotes element-wise complex conjugation). We can avoid constructing $\mathcal$$\hat${R}} by expanding the product, and $\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]$ by using the kernel $\mathcal{K}_{\mathbf{qk}}^{P}$:
$\mathbf{D}[i,j]=\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\|^{2}+\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\Big{[}\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\|^{2}$
$-2\mathrm{Re}\left(\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big{]}$
But computing the product $\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]$ still requires constructing a $d_{k}\times m\times m$ tensor.
3.6 Isotropic Decay and Noise
If $\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]$ loses its $k$-dependence, each of the terms in $\mathbf{D}[i,j]$ simplifies to a matrix multiplication. In particular, let us assume that the real part of the eigenvalues is a shared scalar:
$\lambda_{k}=\alpha+i\omega_{k},\quad\alpha,\omega_{k}\in\mathbb{R}$
This corresponds to a system with state matrix $\mathbf{A}=\alpha\mathbf{I}+\bm{\Omega}$, where $\bm{\Omega}\in\mathbb{R}^{d\times d}$ is diagonalizable over $\mathbb{C}$, and has strictly imaginary eigenvalues, i.e. $\bm{\Omega}=\mathbf{S}\bm{\Lambda}_{\Omega}\mathbf{S}^{-1}$, where $\mathbf{S}\in\mathbb{C}^{d\times d}$, $\bm{\Lambda}_{\Omega}=\text{diag}(\lambda_{\Omega,1},\dots\lambda_{\Omega,d})$, with $\lambda_{\Omega,k}\in i\mathbb{R}$. We also assume that the noise is isotropic, i.e. that the noise parameters are scalar multiples of identity: $\bm{\Lambda}_{C}=\lambda_{C}\mathbf{I},\bm{\Lambda}_{Q}=\sigma^{2}\mathbf{I},\bm{\Lambda}_{R}=\eta^{2}\mathbf{I}$.
Note that we could define separate query/key and value dynamic parameters $\alpha$ and $\bm{\omega}$, but here we use shared parameters for simplicity.
Our covariance kernel becomes:
$\mathcal{K}^{V}(\tau)=\sigma_{V}^{2}(\tau)\mathbf{I}_{d},\quad\text{where:}\quad\sigma_{V}^{2}(\tau)=\lambda_{C}^{2}\sigma^{2}\frac{1-e^{-2\alpha\Delta t\tau}}{-2\alpha}+\eta^{2}e^{-2\alpha\Delta t\tau}$
and the $ij$th component of the precision matrix becomes:
$\mathbf{P}_{ij}^{C}=\frac{1}{\sigma_{V}^{2}(|i-j|)}.$
We can then pull $\mathbf{P}^{C}_{\mathbf{k}}\in\mathbb{R}^{m\times m}$ outside the sum:
$\mathbf{D}[i,j]=\mathbf{P}^{C}_{\mathbf{k}}[i,j]\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\Big{[}\|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\|^{2}+\|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\|^{2}-2\mathrm{Re}\left(\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big{]}$
We can now express each of these three sums as matrix products:
$\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\|^{2}=\sum_{k}\big{(}\|\bm{\Phi}_{\mathbf{qk}}\|^{2}\odot\|\mathbf{\hat{Z}}_{\mathbf{q}}\|^{2}\big{)}\in\mathbb{R}^{m},$
$\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\|^{2}=\big{(}\|\bm{\Phi}_{\mathbf{qk}}\|^{2}\big{)}^{\top}\|\mathbf{\hat{Z}}_{\mathbf{k}}\|^{2}\in\mathbb{R}^{m\times m},$
$\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]=\big{(}\|\bm{\Phi}_{\mathbf{qk}}\|^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big{)}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\in\mathbb{C}^{m\times m}$
We then sum terms, where the first term is broadcast from an $m\times 1$ vector to an $m\times m$ matrix:
$\|\mathbf{R}_{\mathbf{qk}}\|^{2}:=\mathbf{1}_{m}\big{(}\big{(}\|\bm{\Phi}_{\mathbf{qk}}\|^{2}\odot\|\mathbf{\hat{Z}}_{\mathbf{q}}\|^{2}\big{)}^{\top}\mathbf{1}_{d}\big{)}^{\top}+\big{(}\|\bm{\Phi}_{\mathbf{qk}}\|^{2}\big{)}^{\top}\|\mathbf{\hat{Z}}_{\mathbf{k}}\|^{2}-2\mathrm{Re}\Big{(}\big{(}\|\bm{\Phi}_{\mathbf{qk}}\|^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big{)}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\Big{)}$
We can factor out $e^{\alpha t_{i}}$. Let us define: Let $\mathbf{E_{qk}}_{ij}=e^{\alpha(t_{i}-t_{j})}$:
$\mathbf{E_{qk}}[i,j]=e^{\alpha(t_{i}-t_{j})},\quad\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}[k,i]:=e^{i\omega_{k}t_{i}},\quad\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}[k,i]:=e^{-i\omega_{k}t_{i}}$
And:
$\mathbf{\tilde{Z}_{q}}[k,i]:=e^{-i\omega_{k}t_{i}}\mathbf{Z_{q}}[k,i],\quad\mathbf{\tilde{Z}_{k}}[k,j]:=e^{-i\omega_{k}t_{j}}\mathbf{Z_{k}}[k,j].$
that is,
$e^{\alpha t_{i}}\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}[k,i]=\bm{\Phi}_{\mathbf{qk}}[k,i],\quad e^{-\alpha t_{i}}\mathbf{\tilde{Z}_{q}}[k,i]:=\mathbf{\hat{Z}_{q}}[k,i],\quad e^{-\alpha t_{j}}\mathbf{\tilde{Z}_{k}}[k,j]:=\mathbf{\hat{Z}_{k}}[k,j].$
Or in matrix form:
$\mathbf{\tilde{Z}_{q}}:=\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}\odot\mathbf{Z_{q}},\quad\mathbf{\tilde{Z}_{k}}:=\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}\odot\mathbf{Z_{k}}$
So:
$\mathcal{R}_{\mathbf{qk}}[k,i,j]=e^{i\omega_{k}t_{i}}\mathbf{\tilde{Z}_{q}}[k,i]-e^{\alpha(t_{i}-t_{j})}\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}[k,i]\cdot\mathbf{\tilde{Z}_{k}}[k,j]$
Or in matrix form:
$\mathcal{R}_{\mathbf{qk}}=\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}\odot\mathbf{\tilde{Z}_{q}}-\mathbf{E_{qk}}\odot(\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+})^{\dagger}\mathbf{\tilde{Z}_{k}}$
Hence,
$\|\mathbf{R}_{\mathbf{qk}}\|^{2}=e^{2\alpha t_{i}}\sum_{k}\Big{[}e^{-2\alpha t_{i}}\|\mathbf{Z_{q}}[k,i]\|^{2}+e^{-2\alpha t_{j}}\|\mathbf{Z_{k}}[k,j]\|^{2}-2\mathrm{Re}\left(e^{-\alpha t_{i}}e^{-\alpha t_{j}}\mathbf{\bar{\tilde{Z}}}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)\Big{]}$
$\sum_{k}\|\mathbf{Z_{q}}[k,i]\|^{2}+e^{2\alpha(t_{i}-t_{j})}\sum_{k}\|\mathbf{Z_{k}}[k,j]\|^{2}-2e^{\alpha(t_{i}-t_{j})}\sum_{k}\mathrm{Re}\left(\mathbf{\bar{\tilde{Z}}}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)$
Let $\mathbf{1}_{m}\in\mathbb{R}^{m\times 1}$ and $\mathbf{1}_{d_{k}}\in\mathbb{R}^{d_{k}\times 1}$ be column vectors of $1$s. In matrix form:
$\|\mathbf{R}_{\mathbf{qk}}\|^{2}=\mathbf{1}_{m}\big{(}\mathbf{1}_{d_{k}}^{\top}|\mathbf{Z_{q}}|^{2}\big{)}+\mathbf{E_{qk}}^{2}\odot\big{(}\mathbf{1}_{d_{k}}^{\top}|\mathbf{Z_{k}}|^{2}\big{)}^{\top}\mathbf{1}_{m}^{\top}-2\mathbf{E_{qk}}\odot\mathrm{Re}\big{(}\mathbf{\tilde{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{k}}\big{)}$
(where the modulus is taken element-wise). So we need three matrix products and two element-wise multiplications, as opposed to dot-product attention’s one matrix product.
We can then compute the un-normalized attention matrix:
$\mathbf{\tilde{A}}=\frac{\mathbf{P}^{C}_{\mathbf{v}}}{1+\mathbf{P}^{C}_{\mathbf{k}}\odot\|\mathbf{R}_{\mathbf{qk}}\|^{2}/\nu}$
(where division is again taken element-wise.) The above formulation reduces the required memory to $\mathcal{O}(m^{2})$.
Define a score matrix: $\mathbf{L}=\log(\mathbf{P_{v}}^{C})-\log\left(1+\mathbf{P_{k}}^{C}\odot\|\mathbf{R}_{\mathbf{qk}}\|^{2}/\nu\right)$,
and a causal mask $\mathbf{M}_{\text{causal}}$.
We can then express the row-normalization using softmax:
$\mathbf{A}[i,j]=e^{\mathbf{L}[i,j]\odot\mathbf{M}_{\text{causal}}}\big{/}\sum_{j}e^{\mathbf{L}[i,j]\odot\mathbf{M}_{\text{causal}}}$.
If the query-key and value dynamics parameters are shared, $\mathbf{P}^{C}_{\mathbf{k}}=\mathbf{P}^{C}_{\mathbf{v}}:=\mathbf{P}^{C}$, so:
$\mathbf{L}=-\log\left(\mathbf{V}^{C}+\|\mathbf{R}_{\mathbf{qk}}\|^{2}/\nu\right)$
The estimate is:
$\mathbf{\bar{Z}_{v}}[k,i]=\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\sum_{j}\mathbf{A}[i,j]\cdot\mathbf{\hat{Z}}_{\mathbf{v}}[k,j]$
Define:
$\mathbf{E_{v}}[i,j]=e^{\alpha(t_{i}-t_{j})},\quad\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}[k,i]:=e^{i\omega_{k}t_{i}},\quad\bm{\tilde{\Phi}}_{\mathbf{v}}^{-}[k,i]:=e^{-i\omega_{k}t_{i}}$
And define:
$\mathbf{\tilde{Z}_{v}}:=\bm{\tilde{\Phi}}_{\mathbf{v}}^{-}\odot\mathbf{Z_{v}}.$
That is,
$e^{\alpha t_{i}}\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}[k,i]=\bm{\Phi}_{\mathbf{v}}[k,i],\quad e^{-\alpha t_{i}}\mathbf{\tilde{Z}_{v}}[k,i]:=\mathbf{\hat{Z}_{v}}[k,i].$
Let $\mathbf{\hat{A}}=\mathbf{A}\odot\mathbf{E_{v}}$. Then,
$\mathbf{\bar{Z}_{v}}[k,i]=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}[k,i]\odot\sum_{j}\mathbf{\hat{A}}[i,j]\cdot\mathbf{\tilde{Z}}_{\mathbf{v}}[k,j]$
which, in matrix form, is:
$\mathbf{\bar{Z}_{v}}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}\odot(\mathbf{\tilde{Z}}_{\mathbf{v}}\mathbf{\hat{A}}^{\dagger})$
Or, in a form more typical for attention:
$\mathbf{\bar{Z}_{v}}^{\dagger}=(\bm{\tilde{\Phi}}_{\mathbf{v}}^{+})^{\dagger}\odot(\mathbf{\hat{A}}\mathbf{\tilde{Z}}_{\mathbf{v}}^{\dagger})$
Finally, the prediction is:
$\mathbf{Z}_{\mathbf{v}}^{\text{pred}}=(e^{\alpha\Delta t}e^{\bm{\lambda}_{\Omega}\Delta t})\odot\mathbf{\bar{Z}_{v}}$
This factorization, which separates scalar decay/growth (magnitude) from oscillatory components (direction), reduces the memory requirement of the algorithm from $\mathcal{O}(m^{2}d)$ to $\mathcal{O}(m^{2}+md)$, as in vanilla attention, while maintaining the $\mathcal{O}(m^{2}d)$ compute time. This comes at the price of losing the second-order information provided by the full tensors $\mathcal{P}^{C}_{\mathbf{qk}}$ and $\mathcal{P}^{C}_{\mathbf{v}}$. Whereas $\mathcal{P}^{C}_{\mathbf{qk}},\mathcal{P}^{C}_{\mathbf{v}}$ represent the confidence of the $k$-th feature’s projection from time $t_{j}$ to time $t_{i}$, $\mathbf{P}^{C}_{\mathbf{k}},\mathbf{P}^{C}_{\mathbf{v}}$ represent measurements of overall confidence between query and key.
The assumption that the real part of the eigenvalues is a shared scalar serves two purposes: (a) it allows the cross term to be represented as a matrix multiplication, which reduces the memory requirement, and (b) it improves the exploding/vanishing gradient problem for $\alpha<0$, because rather than computing $e^{\alpha t_{i}}$ and $e^{-\alpha t_{j}}($ separately, we compute $e^{\alpha(t_{i}-t_{j})}$, and $t_{i}>t_{j}$ for causal systems.
An alternative approach to simplify the tensor precision-weighted average into a tractable form with $\mathcal{O}(m^{2}+md)$ memory is to factorize the precision kernel into a product of factors depending separately on indices $i$ and $j$. We briefly explore this approach in the Appendix.
3.7 Unitary Dynamics
If we further constrain our state matrix $\mathbf{A}$ to have purely imaginary eigenvalues, i.e. $\alpha=0$, then $e^{\mathbf{\Lambda}}$ is orthogonal, and our dynamical system is unitary (norm preserving) in the eigenbasis.
Using L’Hôpital’s rule:
$\lim_{\mathrm{Re}(\lambda_{k})\to 0}\frac{1-e^{2\mathrm{Re}(\lambda_{k})\Delta t\tau}}{-2\mathrm{Re}(\lambda_{k})}=\Delta t\,\tau$
Our covariance kernel becomes:
$\mathcal{K}^{V}(\tau)=\sigma_{V}^{2}(\tau)\mathbf{I}_{d},\quad\text{where:}\quad\sigma_{V}^{2}(\tau):=\lambda_{C}^{2}\sigma^{2}\Delta t\tau+\eta^{2}$
And the $ij$th component of the precision matrix becomes:
$\mathbf{P}_{ij}=\frac{1}{\sigma_{V}^{2}(|i-j|)},$
Since the eigenvalues are purely imaginary, $\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|=\|e^{\bm{\lambda}t_{i}}\|=1$, and: $\|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\|=\|\mathbf{Z}_{\mathbf{k}}[k,j]\|$, $\|\mathbf{\hat{Z}}_{\mathbf{q}}[k,j]\|=\|\mathbf{Z}_{\mathbf{q}}[k,j]\|$. Hence:
$\|\mathbf{R_{qk}}[i,j]\|^{2}=\sum_{k}\Big{(}\|\mathbf{Z}_{\mathbf{q}}[k,i]\|^{2}+\|\mathbf{Z}_{\mathbf{k}}[k,j]\|^{2}-2\mathrm{Re}\big{(}\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big{)}\Big{)}$
If, in addition, $\mathbf{S}$ is unitary (in which case $\mathbf{A}$ is skew-symmetric), then $\|\mathbf{Z}_{\mathbf{q}}[k,i]\|^{2}=\|\mathbf{Z}[k,i]\|^{2}$, $\|\mathbf{Z}_{\mathbf{v}}[k,i]\|^{2}=\|\mathbf{Z}[k,i]\|^{2}$, so if the inputs have been normalized to unit length, then: $\sum_{k}\|\mathbf{Z}_{\mathbf{q}}[k,i]\|^{2}=\sum_{k}\|\mathbf{Z}_{\mathbf{k}}[k,j]\|^{2}=1$. Hence,
$\|\mathbf{R_{qk}}\|^{2}=2\Big{(}1-\mathrm{Re}\big{(}\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big{)}\Big{)}$
The score matrix is:
$\mathbf{L}=-\log\Big(\nu\mathbf{V}^{C}+2-2\mathrm{Re}\big{(}\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big{)}\Big{missing})$
$\log(\nu\mathbf{V}^{C}+2)-\log\left(1-\frac{2\mathrm{Re}\big{(}\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big{)}}{\nu\mathbf{V}^{C}+2}\right)$
Expanding around $\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}=\mathbf{0}$, this is approximately:
$\mathbf{L}\approx-\log(\nu\mathbf{V}^{C}+2)+\frac{2\mathrm{Re}\big{(}\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big{)}}{\nu\mathbf{V}^{C}+2}$
If we set $\sigma^{2}\rightarrow 0$ (zero process noise), then $\mathbf{V}^{C}=\eta^{2}\mathbf{1}$ becomes a simple scalar bias. Letting $\tau=2/(\nu\eta^{2}+2)$, the unnormalized softmax becomes:
$\frac{\tau}{2}\cdot\exp(\tau\mathrm{Re}\big{(}\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}))$
The constant term $\frac{\tau}{2}$ outside the exponential cancels in the normalization. The imaginary part of the matrix exponential becomes a multiplicative relative positional encoding, and we recover a complex-valued form of ordinary dot-product attention.
3.8 Implementation
3.8.1 Adaptive Filter Attention
We now present an implementation of the above algorithm, for the case of isotropic noise and decay.
We use scalar noise parameters $\sigma^{2},\eta^{2}$ and dynamics parameters $\alpha,\bm{\omega}$, where the real part of the eigenvalues is set to $\alpha^{2}$ to ensure stability of our LTI model. Note that the eigenvalues of $\bm{\Omega}$ appear in complex conjugate pairs, as a necessary condition for $\mathbf{A}$ to be real, so we can write the vector of eigenvalues as:
$\bm{\lambda}_{\Omega}=\big{\{}i\omega_{1},-i\omega_{1},\dots,i\omega_{d/2},-i\omega_{d/2}\big{\}},\,\omega_{k}\in\mathbb{R}$
We switch from the rational weight $\tfrac{1}{1+r_{ij}^{2}/\nu}$ to the equivalent exponential–log form $\exp\!\big{(}-\tau_{s}\log(1+r_{ij}^{2}/\nu)\big{)}$, which both improves numerical stability and introduces a temperature parameter $\tau_{s}$ that controls the sharpness of the weighting; this reduces exactly to the rational form when $\tau_{s}=1$. We drop $\lambda_{C}^{2}$ since we can absorb it into $\sigma^{2}$, and drop the Mahalanobis scale $\nu$, since we can absorb it into $\sigma,\eta,\tau_{s},\mathbf{W}_{Q},\mathbf{W}_{K}$.
Note that $\bm{\Omega}$ is similar to a block diagonal matrix of $2\times 2$ rotation matrices, i.e. there exists some $\mathbf{P}\in\mathbb{R}^{d\times d}$ such that:
[…]
However, to retain our fully diagonalized representation, we must work with complex numbers. We define the following linear layer variant to reproduce the effect of complex multiplication:
Complex Linear Layer. We denote a complex linear layer by $\mathcal{L}(\cdot)$. Instead of operating directly on complex vectors, we represent inputs and outputs by stacking their real and imaginary parts. For an input $\mathbf{x}=\left[\mathbf{x}_{r},\mathbf{x}_{i}\right]^{\top}$ with $\mathbf{x}_{r},\mathbf{x}_{i}\in\mathbb{R}^{d}$, the layer computes:
$\mathcal{L}(\mathbf{x})=\begin{bmatrix}\mathbf{W}_{r}&-\mathbf{W}_{i}\\[4.0pt] \mathbf{W}_{i}&\mathbf{W}_{r}\end{bmatrix}\mathbf{x}+\begin{bmatrix}\mathbf{b}_{r}\\[4.0pt] \mathbf{b}_{i}\end{bmatrix}$,
where $\mathbf{W}_{r},\mathbf{W}_{i}\in\mathbb{R}^{d_{e}\times d}$ are the real and imaginary weight matrices, and $\mathbf{b}_{r},\mathbf{b}_{i}\in\mathbb{R}^{d_{e}}$ are real-valued biases (where $d_{e}$ is the embedding dimension). This construction is equivalent to multiplication by a complex weight matrix $\mathbf{W}_{r}+i\mathbf{W}_{i}$ with bias $\mathbf{b}_{r}+i\mathbf{b}_{i}$, while remaining entirely in the real domain.
We initialize the complex linear layer weights isotropically in the complex plane as:
$\mathbf{W}_{ij}:=\mathbf{W}_{r,ij}+i\mathbf{W}_{i,ij}=\alpha_{ij}\begin{bmatrix}\cos(\phi_{ij})\\ \sin(\phi_{ij})\end{bmatrix},\quad\alpha_{ij}\sim\sqrt{\tfrac{2}{d_{\text{in}}+d_{\text{{out}}}}}\,\mathcal{N}(0,1),\;\;\phi_{ij}\sim\mathcal{U}(0,2\pi).$
Since the inputs to the network layer are real-valued, we can concatenate a vector of zeros to the inputs to represent the imaginary part. In general, the outputs of the network layer may be complex-valued. We pass the real part of the output to the next layer and discard the imaginary part. Alternatively, we can design the input and output projections to map between real and complex inputs/outputs. We pursue the former option here, for simplicity.
Another option to enforce real inputs is to impose a penalty on $\|\operatorname{Im}(\mathbf{Z_{p}})\|$ or on $\|\mathbf{W_{v}}\mathbf{W_{p}}-\mathbf{I}^{d_{v}\times d_{v}}\|$. For simplicity, we do not impose penalties and allow these matrices to freely vary.
If our data has a known time scale, we can set $\Delta t$ as a known constant, otherwise we can treat it as a learned parameter.
In our implementation, queries and keys are computed from the current values rather than directly from the raw inputs. This design enables a natural recursive approximation of full batch attention: as the value estimates are iteratively updated at each step, the queries and keys reflect the latest state, allowing each iteration to propagate residual information through the trajectory. This is optional.
For clarity, we present the single-head version. A multi-head extension may be obtained in the usual way by applying the procedure in parallel with separate projections and concatenating the outputs, with each head having its own set of complex linear layers $\mathcal{L}_{Q},\mathcal{L}_{K},\mathcal{L}_{V},\mathcal{L}_{P}$ and dynamics parameters $\alpha,\bm{\omega},\sigma,\eta$.
For simplicity, we use shared dynamic parameters $\alpha,\bm{\omega}$ for the queries/keys and values, so $\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}$ and $\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{-}$. We define these as separate variables however, since in general they can be different.
Input: Input sequence $\mathbf{Z}_{r}\in\mathbb{R}^{d\times m}$
Complex linear layers: $\mathcal{L}_{Q},\mathcal{L}_{K},\mathcal{L}_{V},\mathcal{L}_{P}$; dynamics/noise parameters: $\alpha,\bm{\omega},\sigma,\eta$; time step $\Delta t$; softmax temperature $\tau_{s}$; residual weight $\delta$.
Constants: Causal mask $\mathbf{M}_{\text{causal}}\in\{0,1\}^{m\times m}$; stability constant $\epsilon$.
(Define $t_{i}:=\Delta t\cdot i,\quad t_{j}:=\Delta t\cdot j\$.) Complex-valued Projections:
Lift real inputs to complex by concatenating zeros: $\mathbf{Z}=\left[\mathbf{Z}_{r},\mathbf{0}\right]^{\top}\in\mathbb{C}^{2\times d\times m}$
$\big{(}\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big{)}\leftarrow\mathcal{L}_{V}\big{(}\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z})\big{)}.$
$\big{(}\operatorname{Re}(\mathbf{Z_{q}}),\operatorname{Im}(\mathbf{Z_{q}})\big{)}\leftarrow\mathcal{L}_{Q}\big{(}\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big{)}$,
$\big{(}\operatorname{Re}(\mathbf{Z_{k}}),\operatorname{Im}(\mathbf{Z_{k}})\big{)}\leftarrow\mathcal{L}_{K}\big{(}\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big{)}$.
Exponential kernels:
$\mathcal{K}_{\alpha}[\tau]=\mathbf{1}_{\{\tau>0\}}\odot e^{-\alpha^{2}\Delta t\tau}+\mathbf{1}_{\{\tau\leq 0\}},\quad\mathcal{K}_{\bm{\omega}}[\tau]=\mathbf{1}_{\{\tau>0\}}\odot e^{\bm{\omega}\Delta t\tau}+\mathbf{1}_{\{\tau\leq 0\}}$
Decay kernel: $\kappa[\tau]=\frac{1-\mathcal{K}^{2}_{\alpha}[\tau]+\epsilon}{2\alpha^{2}+\epsilon}\mathbf{1}_{\{\alpha^{2}>\epsilon\}}+\tau\mathbf{1}_{\{\alpha^{2}\leq\epsilon\}}$
Covariance kernel: $\sigma^{2}_{V}(\tau)=\sigma^{2}\kappa[|\tau|]+\eta^{2}\mathcal{K}^{2}_{\alpha}[|\tau|],\quad\mathbf{V}^{C}[i,j]=\sigma^{2}_{V}(i-j)$
Query/Key dynamics:
$\mathbf{E_{qk}}[i,j]=\mathcal{K}_{\alpha}[i-j]=\mathcal{K}_{\alpha}[i]\mathcal{K}_{\alpha}[j]^{-\top},\quad\bm{\tilde{\Phi}}^{+}_{\mathbf{qk}}[k,i]=e^{i\bm{\omega}_{k}t_{i}},\quad\bm{\tilde{\Phi}}^{-}_{\mathbf{qk}}[k,i]=e^{-i\bm{\omega}_{k}t_{i}}$
$\mathbf{\tilde{Z}_{q}}[k,i]=\bm{\tilde{\Phi}}^{-}_{\mathbf{qk}}\odot\mathbf{Z_{q}}[k,i],\quad\mathbf{\tilde{Z}_{k}}[k,j]=\bm{\tilde{\Phi}}^{-}_{\mathbf{qk}}\odot\mathbf{Z_{k}}[k,j]$
Value dynamics:
$\mathbf{E_{v}}[i,j]=\mathcal{K}_{\alpha}[i-j]=\mathcal{K}_{\alpha}[i]\mathcal{K}_{\alpha}[j]^{-\top},\quad\bm{\tilde{\Phi}}^{+}_{\mathbf{v}}[k,i]=e^{i\bm{\omega}_{k}t_{i}},\quad\bm{\tilde{\Phi}}^{-}_{\mathbf{v}}[k,i]=e^{-i\bm{\omega}_{k}t_{i}}$
$\mathbf{\tilde{Z}_{v}}[k,i]=\bm{\tilde{\Phi}}_{\mathbf{v}}^{-}\odot\mathbf{Z_{v}}[k,i]$
Squared Residuals:
$\|\mathbf{R}_{\mathbf{qk}}\|^{2}=\mathbf{1}_{m}(\mathbf{1}_{d}^{\top}\|\mathbf{Z_{q}}\|^{2})+\mathbf{E_{qk}}^{2}\odot(\mathbf{1}_{d}^{\top}\|\mathbf{Z_{k}}\|^{2})^{\top}\mathbf{1}_{m}^{\top}-2\mathbf{E_{qk}}\odot\mathrm{Re}(\mathbf{\tilde{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{k}})$
Attention matrix:
$\mathbf{L}=-\log\left(\mathbf{V}^{C}+\|\mathbf{R}_{\mathbf{qk}}\|^{2}+\epsilon\right)\odot\mathbf{M}_{\text{causal}},$
$\mathbf{A}[i,j]=\text{softmax}_{j}\big{(}\tau_{s}\mathbf{L}[i,j]\big{)},\quad\mathbf{\hat{A}}=\mathbf{E_{v}}\odot\mathbf{A}$
Output with value dynamics: $\mathbf{\bar{Z}_{v}}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}\odot(\mathbf{\tilde{Z}}_{\mathbf{v}}\mathbf{\hat{A}}^{\dagger})$
Residual connection: $\mathbf{\bar{Z}}=(1-\text{sigmoid}(\delta))\mathbf{Z_{v}}+\text{sigmoid}(\delta)\mathbf{\bar{Z}_{v}}$
Prediction: $\mathbf{Z}_{\mathbf{v}}^{\text{pred}}=(e^{\alpha\Delta t}e^{\bm{\lambda}_{\Omega}\Delta t})\odot\mathbf{\bar{Z}_{v}}$
Output Projection: $\big{(}\operatorname{Re}(\mathbf{Z_{p}}),\operatorname{Im}(\mathbf{Z_{p}})\big{)}\leftarrow\mathcal{L}_{P}\big{(}\operatorname{Re}(\mathbf{Z}_{\mathbf{v}}^{\text{pred}}),\operatorname{Im}(\mathbf{Z}_{\mathbf{v}}^{\text{pred}})\big{)}$
Return: $\operatorname{Re}(\mathbf{Z_{p}})$ ;
  // We discard $\operatorname{Im}(\mathbf{Z_{p}})$.
Algorithm 1 Adaptive Filter Attention (Single Head)
This defines a single neural network layer, which we can train with backpropagation, as usual.
In summary, this formulation modifies standard attention by including:
  1. 1.
    Complex-valued linear layers;
  2. 2.
    Complex-valued multiplicative positional encodings with learned parameters;
  3. 3.
    Scalar decay constants (one per head) with pair-wise uncertainty propagation via the closed-form solution of the differential Lyapunov equation;
  4. 4.
    A similarity function based on the Mahalanobis-distance;
  5. 5.
    A shared linear dynamics model for state estimation and prediction, to enforce model self-consistency.
A multi-head implementation of Algorithm 1 is available at the following: Github Repository. (This repository also contains an implementation of the tensor form discussed in Section 3.1.)
3.8.2 Unrolled Adaptive Filter Attention
At inference time, we can approximately unroll our learned model into a reweighted Kalman Filter. The outer iteration propagates latent states through the sequence for fast, online inference, while the inner iterations at each step approximate the effect of full pairwise attention without incurring $\mathcal{O}(m^{2})$ computations. The recursive precision update acts as an implicit normalization of the "attention" weights.
Input: Prior state $\mathbf{\bar{z}}_{0}^{-}$, measurements $\{\mathbf{\hat{z}}_{\mathbf{s}ii}\}_{i=1}^{m}$ Parameters: Complex linear layers $\mathcal{L}_{Q},\mathcal{L}_{K},\mathcal{L}_{V},\mathcal{L}_{P}$; dynamics/noise parameters: $\alpha,\bm{\omega},\sigma,\eta$; timestep $\Delta t$; softmax temperature $\tau_{s}$, residual weight $\delta$. Constants: Causal mask $\mathbf{M}_{\text{causal}}\in\{0,1\}^{m\times m}$, stability constant $\epsilon$.
Complex-valued Projections:
Lift real inputs to complex by concatenating zeros: $\mathbf{z}=\left[\mathbf{z}_{r},\mathbf{0}\right]^{\top}$
$\big{(}\operatorname{Re}(\mathbf{z_{v}}),\operatorname{Im}(\mathbf{z_{v}})\big{)}\leftarrow\mathcal{L}_{V}\big{(}\operatorname{Re}(\mathbf{z}),\operatorname{Im}(\mathbf{z})\big{)}.$
State transition matrix:
$\Phi_{\alpha}:=e^{-\alpha^{2}\Delta t},\quad\bm{\Phi}_{\bm{\omega}}:=e^{\bm{\omega}\Delta t},\quad\bm{\Phi}:=\Phi_{\alpha}\bm{\Phi}_{\bm{\omega}}$
Innovation covariance: $\sigma^{2}_{V}=\sigma^{2}\kappa+\eta^{2}\Phi^{2}_{\alpha},\quad\kappa=\frac{1-\Phi^{2}_{\alpha}+\epsilon}{2\alpha^{2}+\epsilon}\mathbf{1}_{\{\alpha^{2}>\epsilon\}}+\Delta t\mathbf{1}_{\{\alpha^{2}\leq\epsilon\}}$
Initial conditions: $\mathbf{\bar{z}}^{+}_{\mathbf{v},0}=\mathbf{z_{v}},\quad p_{0}^{-(0)}=1/\sigma_{V}^{2}$
for $i\leftarrow 1$ to $m$ do
   Initialize prior mean: $\mathbf{\bar{z}}_{\mathbf{v}i}^{-(0)}=\bm{\Phi}\mathbf{\bar{z}}^{+}_{\mathbf{v},i-1}$
 $p_{i}^{-(0)}=1/(\Phi_{\alpha}^{2}/p_{i-1}^{+}+\sigma^{2})$
    // Propagated precision
 for $k\leftarrow 0$ to K1K-1
   // Inner iteration within time step
   do
    Recompute queries and keys:
    $\big{(}\operatorname{Re}(\mathbf{\bar{z}}_{\mathbf{q}}^{-(k)}),\operatorname{Im}(\mathbf{\bar{z}}_{\mathbf{q}}^{-(k)})\big{)}\leftarrow\mathcal{L}_{Q}\big{(}\operatorname{Re}(\mathbf{\bar{z}}_{\mathbf{v}}^{-(k)}),\operatorname{Im}(\mathbf{\bar{z}}_{\mathbf{v}}^{-(k)})\big{)}$
    $\big{(}\operatorname{Re}(\mathbf{\bar{z}}_{\mathbf{k}}^{-(k)}),\operatorname{Im}(\mathbf{\bar{z}}_{\mathbf{k}}^{-(k)})\big{)}\leftarrow\mathcal{L}_{K}\big{(}\operatorname{Re}(\mathbf{\bar{z}}_{\mathbf{v}}^{-(k)}),\operatorname{Im}(\mathbf{\bar{z}}_{\mathbf{v}}^{-(k)})\big{)}$
    Compute innovations: $\mathbf{y}_{\mathbf{qk}i}^{(k)}=\mathbf{z}_{\mathbf{q}i}-\mathbf{\bar{z}}_{\mathbf{k}i}^{-(k)},\quad\mathbf{y}_{\mathbf{v}i}^{(k)}=\mathbf{z}_{\mathbf{v}i}-\mathbf{\bar{z}}_{\mathbf{v}i}^{-(k)}$
    Nominal posterior precision (before robust weighting): $p_{i}^{+(k)}=p_{i}^{-(k)}+1/\sigma^{2}_{V}$
    "Attention" weight: $w_{i}^{(k)}=-\log\left(1+p_{i}^{+(k)}\mathbf{y}_{\mathbf{v}i}^{(k)\dagger}\mathbf{y}_{\mathbf{v}i}^{(k)}+\epsilon\right)$
    Weighted posterior precision: $\tilde{p}_{i}^{+(k)}=p_{i}^{-(k)}+\exp(\tau_{s}w_{i}^{(k)})/\sigma_{V}^{2}$
    Posterior mean update:
$\mathbf{\bar{z}}_{\mathbf{v}i}^{+(k)}=\mathbf{\bar{z}}_{\mathbf{v}i}^{-(k)}+\text{sigmoid}(\delta)\,w_{i}^{(k)}\mathbf{y}_{\mathbf{v}i}^{(k)}/(\tilde{p}_{i}^{+(k)}\sigma_{V}^{2})$
Set prior for next inner iteration: $\mathbf{\bar{z}}_{\mathbf{v}i}^{-(k+1)}=\mathbf{\bar{z}}_{\mathbf{v}i}^{+(k)},\quad p_{i}^{-(k+1)}=\tilde{p}_{i}^{+(k)}$
 Final posterior for this step: $\mathbf{\bar{z}}_{\mathbf{v}i}^{+}=\mathbf{\bar{z}}_{\mathbf{v}i}^{+(K)}$, $p_{i}^{+}=p_{i}^{+(K)}$
 Propagate mean and precision to next time step:
 $\mathbf{\bar{z}}_{\mathbf{v},i+1}^{-}=\bm{\Phi}\mathbf{\bar{z}}_{\mathbf{v}i}^{+}$,  $p_{i+1}^{-}=1/(\Phi_{\alpha}^{2}/p_{i}^{+}+\sigma^{2})$
Map back to original basis: $\big{(}\operatorname{Re}(\mathbf{z_{p}}),\operatorname{Im}(\mathbf{z_{p}})\big{)}\leftarrow\mathcal{L}_{P}\big{(}\operatorname{Re}(\mathbf{z}_{\mathbf{v}i}^{+}),\operatorname{Im}(\mathbf{z}_{\mathbf{v}i}^{+})\big{)}$
Return: $\operatorname{Re}(\mathbf{z_{p}})$
Algorithm 2 Unrolled Adaptive Filter Attention (Inference)
3.9 Basic Experiments
We tested our model (Algorithm 1) on a diagonalizable 2D linear system with zero-mean Gaussian process and measurement noise. For reproducibility, we used a fixed state matrix $\mathbf{A}=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1}$, where:
$\bm{\Lambda}=\begin{bmatrix}-0.1-i&0\\ 0&-0.1+i\end{bmatrix},\quad\mathbf{S}=\begin{bmatrix}1-i&1+i\\ 1&1\end{bmatrix},\quad\mathbf{S}^{-1}=\frac{1}{2}\begin{bmatrix}i&1-i\\ -i&1+i\end{bmatrix}$
We simulated this linear system starting from a random initial condition, over 1,000 time steps. We use simple Euler updates, with additive zero-mean Gaussian process noise scaled by the square root of the time step injected between time steps:
$\mathbf{x}_{t+\Delta t}=\mathbf{A}\mathbf{x}_{t}\,\Delta t+\mathbf{x}_{t}+\sqrt{\Delta t}\,\bm{\epsilon}_{t},\quad\bm{\epsilon}_{t}\sim\mathcal{N}(0,\sigma^{2})$
We sampled measurements at regular intervals (we chose every 10 time steps), added zero-mean Gaussian measurement noise, and then projected the data into an embedding dimension of $d_{e}=128$ (an effective embedding dimension of 256 since the real/imag stacking doubles the input dimension). We trained the model on one-step prediction, using only the noisy measurements, with the model having no access to the ground-truth trajectory. After learning, we projected the model’s output back into 2D to assess its performance.
The ground-truth and measured trajectories and the model predictions with a single layer of AFA after training are shown in Fig. 1.
Figure 1: Filter performance on different 2D systems: ground-truth trajectory (black), measured (blue), and predicted (red). (a) system with only measurement noise ($\sigma^{2}=0.0,\,\eta^{2}=1.0$). (b) system with both process noise and measurement noise ($\sigma^{2}=0.3,\,\eta^{2}=0.5$). (c) higher noise ($\sigma^{2}=0.5,\,\eta^{2}=2.0$).
To check that the model has learned the right dynamics, we compute "pulled-forward" estimates at four time points. As training progresses, the "pulled-forward" estimates converge to point clouds near the ground truth, as shown in Fig. 2.
Figure 2: Comparison of an AFA layer’s "pulled-forward" state estimates at different stages of training. The true trajectory is shown as a solid black line, and the pulled-forward estimates as colored point clouds. (a) State estimates early in training. (b) State estimates midway through training. (c) State estimates after training is complete.
While standard attention is also able to filter this noisy trajectory, it is not able to do so effectively in a single layer, and the resulting attention matrices lack the consistent and interpretable attention matrices of AFA. We compare the attention matrices from a trained model comprised of two layers of standard attention to a trained model comprised of a single layer of AFA in Fig. 3. The AFA attention matrix displays a clear periodic structure, corresponding to pair-wise precisions propagated through the underlying oscillatory dynamics.
Figure 3: Attention matrices produced by training standard attention and AFA on a 2D LTI with process and measurement noise. (a) First layer of standard attention (b) Second layer of standard attention. (c) Single layer of AFA.
These experiments, along with animations of the predicted trajectories, pulled-forward estimates, and attention matrices during the course of training are available at the same Github link provided above.
Testing the model on real-world language and time series data will be the subject of future work.
4 Radial-Tangential Model
We were able to significantly simplify our algorithm by assuming that the dynamics decouple into separate magnitude and angular components, and that the process and measurement noise are isotropic. Here, we will slightly relax these assumptions by assuming that the process and measurement noise also decompose into separate radial and tangential components, and derive a modification of the above algorithm.
4.1 Itô SDE with Radial–Tangential Dynamics and Noise
Consider the linear stochastic differential equation:
$d\mathbf{x}(t)=\mathbf{A}\mathbf{x}(t)dt+\mathbf{B}(t)d\mathbf{w}(t),\quad\mathbf{z}(t_{k})=\mathbf{C}\mathbf{x}(t_{k})+\mathbf{v}(t_{k}),$
Let $\mathbf{A}(t)=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1}$, with $\bm{\Lambda}=\alpha\mathbf{I}+\bm{\Lambda}_{\Omega}$, where $\bm{\Lambda}_{\Omega}\in i\mathbb{R}^{d\times d}$ is diagonal, and $\mathbf{B}(t)=\mathbf{S}\bm{\Lambda}_{B}(t)\mathbf{S}^{\dagger}$, $\mathbf{C}(t)=\lambda_{C}\mathbf{I}$. We can also write the eigenvalues as $\bm{\lambda}=\alpha\mathbf{1}+\bm{\omega}$, where $\alpha\in\mathbb{R}$, $\bm{\omega}=\text{diag}(\bm{\Lambda}_{\Omega})\in i\mathbb{R}^{d}$.
We transform the state and measurements into the eigenbasis:
$\mathbf{x}_{\mathbf{s}}(t)=\mathbf{S}^{-1}\mathbf{x}(t)\in\mathbb{C}^{d},\quad\mathbf{z}_{\mathbf{s}}(t_{k})=\mathbf{S}^{-1}\mathbf{z}(t_{k})\in\mathbb{C}^{d}.$
And we decompose the state and measurement into radial and tangential components:
$\mathbf{x}_{\mathbf{s}}(t)=m_{\mathbf{s}}(t)\mathbf{u}_{\mathbf{s}}(t),\quad\mathbf{z}_{\mathbf{s}}(t)=m_{\mathbf{zs}}(t)\mathbf{u}_{\mathbf{zs}}(t)\quad\text{where:}$
$m_{\mathbf{s}}(t)=\|\mathbf{x}_{\mathbf{s}}(t)\|_{2}=\sqrt{\mathbf{x}_{\mathbf{s}}^{\dagger}(t)\mathbf{x}_{\mathbf{s}}(t)}\in\mathbb{R},\quad\mathbf{u}_{\mathbf{s}}(t)=\mathbf{x}_{\mathbf{s}}(t)/\|\mathbf{x}_{\mathbf{s}}(t)\|_{2}\in\mathbb{S}^{d-1}.$
$m_{\mathbf{zs}}(t)=\|\mathbf{z}_{\mathbf{s}}(t)\|_{2}=\sqrt{\mathbf{z}_{\mathbf{s}}^{\dagger}(t)\mathbf{z}_{\mathbf{s}}(t)}\in\mathbb{R},\quad\mathbf{u}_{\mathbf{zs}}(t)=\mathbf{z}_{\mathbf{s}}(t)/\|\mathbf{z}_{\mathbf{s}}(t)\|_{2}\in\mathbb{S}^{d-1}.$
The dynamics in the eigenbasis are:
$d\mathbf{x_{s}}(t)=\bm{\Lambda}\mathbf{x_{s}}(t)dt+\bm{\Lambda}_{B}d\mathbf{w_{s}}(t)$
We assume independent process noise acting in the radial and tangential directions in the eigenbasis:
$d\mathbf{x}_{\mathbf{s}}=(\alpha\mathbf{I}+\bm{\Lambda}_{\Omega})\mathbf{x}_{\mathbf{s}}\,dt+\sigma_{r}\mathbf{u}_{\mathbf{s}}\,dw_{r}+\sigma_{t}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,d\mathbf{w}_{t},$
where $\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}):=\mathbf{I}-\mathbf{u}_{\mathbf{s}}\mathbf{u}_{\mathbf{s}}^{\dagger}$ projects onto the tangent space, $\mathbf{P}_{R}(\mathbf{u}_{\mathbf{s}}):=\mathbf{u}_{\mathbf{s}}\mathbf{u}_{\mathbf{s}}^{\dagger}$ projects into the radial direction, and $w_{r},\mathbf{w}_{t}$ are independent Wiener processes.
Hence, the process noise covariance is:
$\bm{\Lambda}_{Q}(t)=\bm{\Lambda}_{B}(t)\bm{\Lambda}_{B}(t)^{\dagger}=\sigma_{r}^{2}\mathbf{u_{s}}(t)\mathbf{u_{s}}(t)^{\dagger}+\sigma_{t}^{2}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}).$
Applying Itô’s lemma for $m_{\mathbf{s}}=\sqrt{s}$,
$dm_{\mathbf{s}}=f^{\prime}(s)\,ds+\tfrac{1}{2}f^{\prime\prime}(s)\,d\langle s\rangle=\frac{1}{2m_{\mathbf{s}}}\,ds-\frac{1}{8m_{\mathbf{s}}^{3}}\,d\langle s\rangle.$
Let $s=\mathbf{x}_{\mathbf{s}}^{\dagger}\mathbf{x}_{\mathbf{s}}$. Then:
$dm_{\mathbf{s}}=\frac{1}{2m_{\mathbf{s}}}d(\mathbf{x}_{\mathbf{s}}^{\dagger}\mathbf{x}_{\mathbf{s}})=\frac{1}{2m_{\mathbf{s}}}\big{(}d\mathbf{x}_{\mathbf{s}}^{\dagger}\mathbf{x}_{\mathbf{s}}+\mathbf{x}_{\mathbf{s}}^{\dagger}d\mathbf{x}_{\mathbf{s}}+d\mathbf{x}_{\mathbf{s}}^{\dagger}d\mathbf{x}_{\mathbf{s}}\big{)}.$
Substituting the SDE,
$\mathbf{x}_{\mathbf{s}}^{\dagger}d\mathbf{x}_{\mathbf{s}}=\mathbf{x}_{\mathbf{s}}^{\dagger}\Big{(}(\alpha\mathbf{I}+\bm{\Lambda}_{\Omega})\mathbf{x}_{\mathbf{s}}\,dt+\sigma_{r}\mathbf{u}_{\mathbf{s}}\,dw_{r}+\sigma_{t}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,d\mathbf{w}_{t}\Big{)}$
$\mathbf{x}_{\mathbf{s}}^{\dagger}(\alpha\mathbf{I}+\bm{\Lambda}_{\Omega})\mathbf{x}_{\mathbf{s}}\,dt+\sigma_{r}m_{\mathbf{s}}\,dw_{r}$
so:
$\mathbf{x}_{\mathbf{s}}^{\dagger}d\mathbf{x}_{\mathbf{s}}+d\mathbf{x}_{\mathbf{s}}d\mathbf{x}_{\mathbf{s}}^{\dagger}=2\alpha m_{\mathbf{s}}^{2}dt+2\sigma_{r}m_{\mathbf{s}}\,dw_{r}$
The quadratic term is: $d\mathbf{x}_{\mathbf{s}}^{\dagger}d\mathbf{x}_{\mathbf{s}}=\sigma_{r}^{2}dt+\sigma_{t}^{2}\mathbf{u}_{\mathbf{s}}^{\dagger}\mathbf{P}_{T}\mathbf{u}_{\mathbf{s}}dt=\sigma_{r}^{2}dt,$ since $\mathbf{u}_{\mathbf{s}}^{\dagger}\mathbf{P}_{T}\mathbf{u}_{\mathbf{s}}=0$.
Altogether, we have:
$dm_{\mathbf{s}}=\frac{1}{2m_{\mathbf{s}}}\big{(}\alpha m_{\mathbf{s}}^{2}dt+\sigma_{r}m_{\mathbf{s}}dw_{r}+\sigma_{r}^{2}dt\big{)}-\frac{1}{8m_{\mathbf{s}}^{3}}(2\sigma_{r}m_{\mathbf{s}})^{2}dt$
$\alpha m_{\mathbf{s}}\,dt+\sigma_{r}\,dw_{r}+\frac{1}{2m_{\mathbf{s}}}\sigma_{r}^{2}dt-\frac{1}{2m_{\mathbf{s}}}\sigma_{r}^{2}dt$
Hence the Itô correction cancels, and the radial SDE simplifies to:
$dm_{\mathbf{s}}=\alpha m_{\mathbf{s}}\,dt+\sigma_{r}\,dw_{r}.$
For the unit direction $\mathbf{u}_{\mathbf{s}}=\mathbf{x}_{\mathbf{s}}/m_{\mathbf{s}}$, the projected Itô SDE is:
$d\mathbf{u}_{\mathbf{s}}=\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,d\mathbf{x}_{\mathbf{s}}/m_{\mathbf{s}}.$
Substituting $d\mathbf{x}_{\mathbf{s}}$ and using the fact that $\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\mathbf{u}_{\mathbf{s}}=\mathbf{0}$ gives:
$d\mathbf{u}_{\mathbf{s}}=\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,\bm{\Lambda}_{\Omega}\mathbf{u}_{\mathbf{s}}\,dt+\frac{\sigma_{t}}{m_{\mathbf{s}}}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,d\mathbf{w}_{t}.$
Finally, we assume that measurement noise in the eigenbasis also has a radial–tangential decomposition:
$\mathbf{z}_{\mathbf{s}}(t_{k})=\lambda_{C}\mathbf{x}_{\mathbf{s}}(t_{k})+v_{r}^{(k)}\mathbf{u}_{\mathbf{s}}(t_{k})+\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))\mathbf{v}_{t}^{(k)}$
$v_{r}^{(k)}\sim\mathcal{N}(0,\eta_{r}^{2}),\quad\mathbf{v}_{t}^{(k)}\sim\mathcal{N}(\mathbf{0},\eta_{t}^{2}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))).$
So, the measurement noise covariance is:
$\bm{\Lambda}_{R}(t)=\eta_{r}^{2}\mathbf{u_{s}}(t)\mathbf{u_{s}}(t)^{\dagger}+\eta_{t}^{2}\big{(}\mathbf{I}-\mathbf{u_{s}}(t)\mathbf{u_{s}}(t)^{\dagger}\big{)}.$
The radial measurement is obtained by projecting $\mathbf{z}_{\mathbf{s}}(t_{k})$ onto the state direction $\mathbf{u}_{\mathbf{s}}(t_{k})$:
$m_{\mathbf{zs}}(t_{k})$ $=\mathbf{u}_{\mathbf{s}}^{\dagger}(t_{k})\Big{(}\lambda_{C}\mathbf{x}_{\mathbf{s}}(t_{k})+v_{r}^{(k)}\mathbf{u}_{\mathbf{s}}(t_{k})+\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))\mathbf{v}_{t}^{(k)}\Big{)}$
$=\lambda_{C}\mathbf{u}_{\mathbf{s}}^{\dagger}\mathbf{x}_{\mathbf{s}}+v_{r}^{(k)}\mathbf{u}_{\mathbf{s}}^{\dagger}\mathbf{u}_{\mathbf{s}}+\mathbf{u}_{\mathbf{s}}^{\dagger}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\mathbf{v}_{t}^{(k)}$
$=\lambda_{C}m_{\mathbf{s}}(t_{k})+v_{r}^{(k)}.$
Hence, the radial measurement is 1D Gaussian:
$m_{\mathbf{zs}}(t_{k})\sim\mathcal{N}\big{(}\lambda_{C}m_{\mathbf{s}}(t_{k}),\eta_{r}^{2}\big{)}.$
The tangential (unit-direction) measurement is:
$\mathbf{u}_{\mathbf{zs}}=\exp_{\mathbf{u}_{\mathbf{s}}}\!\Big{(}\mathbf{v}_{t}/m_{\mathbf{s}}\Big{)}$
where $\exp_{\mathbf{u}}$ is the spherical exponential map, which maps a vector in the tangent space of the sphere at $\mathbf{u}\in\mathbb{S}^{d-1}$ to a point on the sphere along the geodesic defined by that tangent vector $\mathbf{v}\in T_{\mathbf{u}}\mathbb{S}^{d-1}$, i.e. $\mathbf{u}^{\top}\mathbf{v}=0$. The spherical exponential map is defined as:
$\exp_{\mathbf{u}}:T_{\mathbf{u}}\mathbb{S}^{d-1}\to\mathbb{S}^{d-1},\quad\exp_{\mathbf{u}}(\mathbf{v})=\cos\big(\|\mathbf{v}\|\big{missing})\,\mathbf{u}+\sin\big(\|\mathbf{v}\|\big{missing})\,\frac{\mathbf{v}}{\|\mathbf{v}\|},$
For small tangential noise ($\eta_{t}\ll m_{\mathbf{s}}$), we can linearize on the tangent plane:
$\mathbf{u}_{\mathbf{zs}}(t_{k})\approx\mathbf{u}_{\mathbf{s}}(t_{k})+\frac{1}{m_{\mathbf{s}}(t_{k})}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))\mathbf{v}_{t}^{(k)}.$
(The $1/m_{\mathbf{s}}$ factor comes from projecting the noise onto the unit sphere (radial normalization), and $\mathbf{P}_{T}\mathbf{v}_{t}^{(k)}$ ensures the noise is tangential.) Hence, the tangential distribution is approximately Gaussian:
$\mathbf{u}_{\mathbf{zs}}(t_{k})\sim\mathcal{N}\Big{(}\mathbf{u}_{\mathbf{s}}(t_{k}),\frac{\eta_{t}^{2}}{m_{\mathbf{s}}^{2}(t_{k})}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))\Big{)}.$
In summary, our radial-tangential SDE (RT-SDE) model is:
$d\mathbf{x}_{\mathbf{s}}$ $=(\alpha\mathbf{I}+\bm{\Lambda}_{\Omega})\mathbf{x}_{\mathbf{s}}\,dt+\sigma_{r}\mathbf{u}_{\mathbf{s}}\,dw_{r}+\sigma_{t}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,d\mathbf{w}_{t},$
$\mathbf{z}_{\mathbf{s}}(t_{k})$ $\approx\lambda_{C}\mathbf{x}_{\mathbf{s}}(t_{k})+v_{r}^{(k)}\mathbf{u}_{\mathbf{s}}(t_{k})+\frac{1}{m_{\mathbf{s}}(t_{k})}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))\mathbf{v}_{t}^{(k)},$
$v_{r}^{(k)}$ $\sim\mathcal{N}(0,\eta_{r}^{2}),\quad\mathbf{v}_{t}^{(k)}\sim\mathcal{N}(\mathbf{0},\eta_{t}^{2}\mathbf{I}).$
Or, in the decoupled radial-tangential coordinates:
$dm_{\mathbf{s}}$ $=\alpha m_{\mathbf{s}}\,dt+\sigma_{r}\,dw_{r},$
$d\mathbf{u}_{\mathbf{s}}$ $=\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,\bm{\Lambda}_{\Omega}\mathbf{u}_{\mathbf{s}}\,dt+\frac{\sigma_{t}}{m_{\mathbf{s}}}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,d\mathbf{w}_{t},$
$m_{\mathbf{zs}}(t_{k})$ $=\lambda_{C}\,m_{\mathbf{s}}(t_{k})+v_{r}^{(k)},$ $v_{r}^{(k)}$ $\sim\mathcal{N}(0,\eta_{r}^{2}),$
$\mathbf{u}_{\mathbf{zs}}(t_{k})$ $\approx\mathbf{u}_{\mathbf{s}}(t_{k})+\frac{1}{m_{\mathbf{s}}(t_{k})}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))\mathbf{v}_{t}^{(k)},$ $\mathbf{v}_{t}^{(k)}$ $\sim\mathcal{N}(\mathbf{0},\eta_{t}^{2}\mathbf{I}).$
Remark: Polar vs Cartesian parameterizations.
Although the radial and angular dynamics decouple, operating in polar coordinates $(r,\theta)$ makes the MLE nonlinear. Even if $\mathbf{x}(t)$ evolves linearly, the radial and angular dynamics become state-dependent, and the noise covariance depends on the current state. This makes the log-likelihood non-quadratic, so the MLE cannot be expressed in closed form and requires iterative optimization. Cartesian coordinates are therefore preferred, as they preserve linear dynamics with Gaussian noise and allow exact, closed-form propagation of covariances and MLE computation.
4.2 Propagation of Uncertainty through the RT-SDE
If $\mathbf{Q}$ is a function of $t$, the state evolution (for the causal case) is governed by:
$\mathbf{x}(t_{i})=e^{\mathbf{A}(t_{i}-t_{j})}\mathbf{x}(t_{j})+\int_{t_{j}}^{t_{i}}e^{\mathbf{A}(t_{i}-s)}\mathbf{B}(s)d\mathbf{w}(s)$
Letting $\tau=s-t_{j}$, this is:
$\mathbf{x}(t_{i})=e^{-\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{j})+\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}(t_{j}+\tau)d\mathbf{w}(\tau)\right).$
So the backward-propagated covariance becomes:
$\mathbf{V}_{B}(t_{i},t_{j})=\int_{0}^{|\Delta t_{ij}|}e^{-\mathbf{A}\tau}\mathbf{Q}(t_{j}+\tau)e^{-\mathbf{A}^{\top}\tau}\,d\tau.$
Plugging in $\mathbf{A}=\mathbf{S}(\alpha\mathbf{I}+\bm{\Lambda}_{\Omega})\mathbf{S}^{-1}$,
$\mathbf{V}_{B}(t_{i},t_{j})=\mathbf{S}\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})\mathbf{S}^{-1},\quad\text{where:}\quad\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})=\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}e^{-\bm{\Lambda}_{\Omega}\tau}\bm{\Lambda}_{Q}(t_{j}+\tau)e^{-\bm{\Lambda}_{\Omega}^{\dagger}\tau}\,d\tau$
where the process noise is:
$\bm{\Lambda}_{Q}(t)=\sigma_{r}^{2}\mathbf{u_{s}}(t)\mathbf{u_{s}}(t)^{\dagger}+\sigma_{t}^{2}\big{(}\mathbf{I}-\mathbf{u_{s}}(t)\mathbf{u_{s}}(t)^{\dagger}\big{)}.$
Note that $\mathbf{u_{s}}(t_{j}+\tau)=e^{\bm{\Lambda}_{\Omega}\tau}\mathbf{u_{s}}(t_{j})$. Hence, the rotation cancels inside the integral:
$\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})=\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}e^{-\bm{\Lambda}_{\Omega}\tau}\Big{(}\sigma_{r}^{2}\mathbf{u_{s}}(t_{j}+\tau)\mathbf{u_{s}}(t_{j}+\tau)^{\dagger}+\sigma_{t}^{2}\big{(}\mathbf{I}-\mathbf{u_{s}}(t_{j}+\tau)\mathbf{u_{s}}(t_{j}+\tau)^{\dagger}\big{)}\Big{)}e^{-\bm{\Lambda}_{\Omega}^{\dagger}\tau}\,d\tau$
$\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}\Big{(}\sigma_{r}^{2}\mathbf{u_{s}}(t_{j})\mathbf{u_{s}}(t_{j})^{\dagger}+\sigma_{t}^{2}\big{(}\mathbf{I}-\mathbf{u_{s}}(t_{j})\mathbf{u_{s}}(t_{j})^{\dagger}\big{)}\Big{)}\,d\tau$
$\bm{\Lambda}_{Q}(t_{j})\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha s}\,ds=\kappa_{B}(|\Delta t_{ij}|)\,\bm{\Lambda}_{Q}(t_{j}),$
where:
$\kappa_{B}(\Delta t)=\begin{cases}\dfrac{e^{-2\alpha\Delta t}-1}{-2\alpha},&\alpha\neq 0,\\ \Delta t,&\alpha=0.\end{cases}$
The forward-propagated covariance is then:
$\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})=e^{\bm{\Lambda}|\Delta t_{ij}|}\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})e^{\bm{\Lambda}^{\dagger}|\Delta t_{ij}|}$
$e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\kappa_{F}(|\Delta t_{ij}|)e^{-\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{\Lambda}_{Q}(t_{j})e^{-\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}$
where:
$\kappa_{F}(\Delta t)=e^{2\alpha\Delta t}\kappa_{B}(\Delta t)=\begin{cases}\dfrac{1-e^{2\alpha\Delta t}}{-2\alpha},&\alpha\neq 0,\\ \Delta t,&\alpha=0.\end{cases}$
(Note: we prefer to keep things in terms of $\kappa_{F}(\Delta t)$ rather than $\kappa_{B}(\Delta t)$, because the former is bounded for causal models, while the latter is unbounded.)
Plugging in $\bm{\Lambda}_{Q}(t_{j})$:
$\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})=\kappa_{F}(|\Delta t_{ij}|)e^{-\bm{\Lambda}_{\Omega}\Delta t_{ij}}\Big{(}\sigma_{r}^{2}\mathbf{u_{s}}(t_{j})\mathbf{u_{s}}(t_{j})^{\dagger}+\sigma_{t}^{2}\big{(}\mathbf{I}-\mathbf{u_{s}}(t_{j})\mathbf{u_{s}}(t_{j})^{\dagger}\big{)}\Big{)}e^{-\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}$
$\kappa_{F}(|\Delta t_{ij}|)\Big{(}\sigma_{r}^{2}\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}+\sigma_{t}^{2}\big{(}\mathbf{I}-\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}\big{)}\Big{)}$
$\kappa_{F}(|\Delta t_{ij}|)\Big{(}\sigma_{t}^{2}\mathbf{I}+(\sigma_{r}^{2}-\sigma_{t}^{2})\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}\Big{)}$
Finally,
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\lambda_{C}^{2}\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})+e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{R}(t_{j})e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\lambda_{C}^{2}\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})+e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{R}(t_{j})e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\lambda_{C}^{2}\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})+e^{-2\alpha\Delta t_{ij}}\bm{\Lambda}_{R}(t_{i})$
Altogether, the propagated measurement covariance is:
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})$ $=\Big{(}\lambda_{C}^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma_{t}^{2}+e^{-2\alpha\Delta t_{ij}}\eta_{t}^{2}\Big{)}\mathbf{I}$
$\quad+\Big{(}\lambda_{C}^{2}\kappa_{F}(|\Delta t_{ij}|)(\sigma_{r}^{2}-\sigma_{t}^{2})+e^{-2\alpha\Delta t_{ij}}(\eta_{r}^{2}-\eta_{t}^{2})\Big{)}\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$
$\sigma_{Vr}^{2}(|\Delta t_{ij}|)\mathbf{I}+\Big{(}\sigma_{Vr}^{2}(|\Delta t_{ij}|)-\sigma_{Vt}^{2}(|\Delta t_{ij}|)\Big{)}\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$
where:
$\sigma_{Vr}^{2}(|\Delta t_{ij}|)$ $=\lambda_{C}^{2}\,\kappa_{F}(|\Delta t_{ij}|)\,\sigma_{r}^{2}+e^{-2\alpha\Delta t_{ij}}\,\eta_{r}^{2},$
$\sigma_{Vt}^{2}(|\Delta t_{ij}|)$ $=\lambda_{C}^{2}\,\kappa_{F}(|\Delta t_{ij}|)\,\sigma_{t}^{2}+e^{-2\alpha\Delta t_{ij}}\,\eta_{t}^{2}.$
Rearranging, this is:
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\sigma_{Vt}^{2}(|\Delta t_{ij}|)(\mathbf{I}-\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger})+\sigma_{Vr}^{2}(|\Delta t_{ij}|)\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$
If $\sigma_{r}=\sigma_{t}:=\sigma$ and $\eta_{r}=\eta_{t}:=\eta$, then this simplifies to:
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\Big{(}\lambda_{C}^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma^{2}+e^{-2\alpha\Delta t_{ij}}\eta^{2}\Big{)}\mathbf{I}$
which was exactly the result for the case of isotropic noise.
Since the propagated covariance is of the form $\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=a\mathbf{I}+b\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$, where $a=\sigma_{Vt}^{2},\,b=\sigma_{Vr}^{2}-\sigma_{Vt}^{2}$, i.e. a rank-1 correction of identity, we can invert it with the Sherman-Morrison formula:
$\bm{\Lambda}_{\mathbf{P}}^{C}=\big{(}\bm{\Lambda}_{\mathbf{V}}^{C}\big{)}^{-1}=\frac{1}{a}\Big{(}\mathbf{I}+\frac{b}{a}\mathbf{u_{s}}\mathbf{u_{s}}^{\dagger}\Big{)}^{-1}=\frac{1}{a}\Big{(}\mathbf{I}-\frac{b}{a}\frac{\mathbf{u_{s}}\mathbf{u_{s}}^{\dagger}}{1+\frac{b}{a}\mathbf{u_{s}}^{\dagger}\mathbf{u_{s}}}\Big{)}$
$\frac{1}{a}\Big{(}\mathbf{I}-\mathbf{u_{s}}\mathbf{u_{s}}^{\dagger}\Big{)}+\Big{(}-\frac{b}{a(a+b)}+\frac{1}{a}\Big{)}\mathbf{u_{s}}\mathbf{u_{s}}^{\dagger},$
$\frac{1}{a}\Big{(}\mathbf{I}-\mathbf{u_{s}}\mathbf{u_{s}}^{\dagger}\Big{)}+\Big{(}\frac{1}{a+b}\Big{)}\mathbf{u_{s}}\mathbf{u_{s}}^{\dagger},$
Hence:
$\bm{\Lambda}_{\mathbf{P}}^{C}(t_{i},t_{j})=\frac{1}{\sigma_{Vt}^{2}(|\Delta t_{ij}|)}\Big{(}\mathbf{I}-\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}^{\dagger}(t_{i})\Big{)}+\frac{1}{\sigma_{Vr}^{2}(|\Delta t_{ij}|)}\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}^{\dagger}(t_{i}).$
(which we could also have obtained by just inverting each term, since they are along orthogonal subspaces).
4.3 Radial–Tangential Kalman IRLS
We now seek to estimate the latent states $\{\mathbf{z}_{i}^{C}\}_{i=1}^{m}\in\mathbb{R}^{d}$ from the observations $\{\mathbf{z}_{i}\}_{i=1}^{m}\in\mathbb{R}^{d}$, assuming the latent states were generated by an RT-SDE.
The joint log-likelihood of the measurements factors into radial and tangential parts:
$\log p\big{(}\{\mathbf{z}_{s}(t_{j})\}\mid m_{s}(\cdot),\mathbf{u}_{s}(\cdot)\big{)}=\sum_{j}\log p\!\left(m_{\mathbf{z}s}(t_{j})\mid m_{s}(t_{j})\right)\,+\,\sum_{j}\log p\!\left(\mathbf{u}_{\mathbf{z}s}(t_{j})\mid\mathbf{u}_{s}(t_{j}),m_{s}(t_{j})\right).$
$\text{where:}\quad m_{\mathbf{zs}}(t_{j})\sim\mathcal{N}\big{(}\lambda_{C}m_{\mathbf{s}}(t_{j}),\eta_{r}^{2}\big{)},\quad\mathbf{u}_{\mathbf{zs}}(t_{j})\sim\mathcal{N}\Big{(}\mathbf{u}_{\mathbf{s}}(t_{j}),\frac{\eta_{t}^{2}}{m_{\mathbf{s}}^{2}(t_{j})}\,\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{j}))\Big{)}.$
Hence, we can treat the estimation problem as two decoupled weighted least squares problems: one for magnitude and one for direction. This motivates the following iteration:
Let $\mathbf{z}_{\mathbf{s}i}:=\mathbf{z}_{\mathbf{s}}(t_{i})$. The initial condition is: $\mathbf{z}_{\mathbf{s}i}^{(0)}=\mathbf{z}_{\mathbf{s}i}$.
At iteration $k$, we compute the magnitudes and directions:
$m_{\mathbf{zs}i}^{(k)}=\|\mathbf{z}_{\mathbf{s}i}^{(k)}\|_{2},\quad\mathbf{u}_{\mathbf{zs}i}^{(k)}=\mathbf{z}_{\mathbf{s}i}^{(k)}/\|\mathbf{z}_{\mathbf{s}i}^{(k)}\|_{2}.$
And the estimates:
$\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}=e^{-\bm{\lambda}\Delta t_{ij}}\mathbf{z_{s}}_{j}^{(k)}=\Big{(}e^{-\alpha\Delta t_{ij}}m_{\mathbf{zs}j}^{(k)}\Big{)}\Big{(}e^{-\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{u}_{\mathbf{zs}j}^{(k)}\Big{)}$
We showed above that:
$\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}\sim\mathcal{N}\Big{(}\mathbf{z}_{\mathbf{s}i}^{C},\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})\Big{)}$
where:
$\begin{gathered}\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\sigma_{Vt}^{2}(|\Delta t_{ij}|)\big{(}\mathbf{I}-\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}\big{)}+\sigma_{Vr}^{2}(|\Delta t_{ij}|)\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger},\\[5.69054pt] \sigma_{Vr}^{2}(|\Delta t_{ij}|)=\lambda_{C}^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma_{r}^{2}+e^{-2\alpha\Delta t_{ij}}\eta_{r}^{2},\\[2.84526pt] \sigma_{Vt}^{2}(|\Delta t_{ij}|)=\lambda_{C}^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma_{t}^{2}+e^{-2\alpha\Delta t_{ij}}\eta_{t}^{2}.\end{gathered}$
It follows that:
$(\mathbf{u}_{\mathbf{s}i}^{(k)})^{\dagger}\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}\sim\mathcal{N}\left((\mathbf{u}_{\mathbf{s}i}^{(k)})^{\dagger}\mathbf{z}_{\mathbf{s}i}^{C},\sigma_{Vr}^{2}(|\Delta t_{ij}|)\right),\quad\mathbf{P}_{T\mathbf{zs}i}\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}\sim\mathcal{N}\left(\mathbf{P}_{T\mathbf{zs}i}\mathbf{z}_{\mathbf{s}i}^{C},\sigma_{Vt}^{2}(|\Delta t_{ij}|)\mathbf{P}_{T\mathbf{zs}i}^{(k)}\right)$
We do not have access to the true tangent projection matrix at $\mathbf{x}_{\mathbf{s}i}$, $\mathbf{P}_{T\mathbf{s}i}:=\mathbf{I}-\mathbf{u}_{\mathbf{s}i}\mathbf{u}_{\mathbf{s}i}^{\dagger}$, but we can approximate it using the tangent projection matrix at $\mathbf{z}_{\mathbf{s}i}^{(k)}/\lambda_{C}$:
$\mathbf{P}_{T\mathbf{zs}i}^{(k)}\Big{(}\mathbf{z}_{\mathbf{s}i}^{(k)}/\lambda_{C}\Big{)}:=\mathbf{I}-\frac{\mathbf{z}_{\mathbf{s}i}^{(k)}\Big{(}\mathbf{z}_{\mathbf{s}i}^{(k)}\Big{)}^{\dagger}/\lambda_{C}^{2}}{\Big{(}\mathbf{z}_{\mathbf{s}i}^{(k)}\Big{)}^{\dagger}\Big{(}\mathbf{z}_{\mathbf{s}i}^{(k)}\Big{)}/\lambda_{C}^{2}}=\mathbf{I}-\mathbf{u}_{\mathbf{zs}i}^{(k)}(\mathbf{u}_{\mathbf{zs}i}^{(k)})^{\dagger}.$
Define separate radial and tangential estimates:
$\hat{m}_{\mathbf{zs}ij}^{(k)}:=(\mathbf{u}_{\mathbf{zs}i}^{(k)})^{\dagger}\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)},\qquad\mathbf{\hat{v}}_{\mathbf{zs}ij}^{(k)}:=\mathbf{P}_{T\mathbf{zs}i}^{(k)}\mathbf{\hat{z}}_{\mathbf{zs}ij}^{(k)}.$
We assume:
$\hat{m}_{\mathbf{zs}ij}^{(k)}\overset{\text{approx}}{\sim}\mathcal{N}\left((\mathbf{u}_{\mathbf{zs}i}^{(k)})^{\dagger}\mathbf{z}_{\mathbf{s}i}^{C},\sigma_{Vr}^{2}(|\Delta t_{ij}|)\right),\quad\mathbf{\hat{v}}_{\mathbf{zs}ij}^{(k)}\overset{\text{approx}}{\sim}\mathcal{N}\left(\mathbf{P}_{T\mathbf{zs}i}^{(k)}\mathbf{z}_{\mathbf{s}i}^{C},\sigma_{Vt}^{2}(|\Delta t_{ij}|)\mathbf{P}_{T\mathbf{zs}i}^{(k)}\right)$
And define corresponding radial and tangential residuals:
$\mathbf{r}_{\mathbf{zs}ij}=\mathbf{z}_{\mathbf{s}i}^{(k)}-\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)},\quad\ell_{\mathbf{zs}ij}^{(k)}:=(\mathbf{u}_{\mathbf{zs}i}^{(k)})^{\dagger}\mathbf{r}_{\mathbf{zs}ij},\quad\mathbf{d}_{\mathbf{zs}ij}^{(k)}:=\mathbf{P}_{T\mathbf{zs}i}^{(k)}\mathbf{r}_{\mathbf{zs}ij}$
where:
$\ell_{\mathbf{zs}ij}^{(k)}\overset{\text{approx}}{\sim}\mathcal{N}\left(0,\sigma_{Vr}^{2}(|\Delta t_{ij}|)\right),\quad\mathbf{d}_{\mathbf{zs}ij}^{(k)}\overset{\text{approx}}{\sim}\mathcal{N}\left(\mathbf{0},\sigma_{Vt}^{2}(|\Delta t_{ij}|)\mathbf{P}_{T\mathbf{zs}i}^{(k)}\right)$
Defining robust weights:
$w_{r,ij}^{(k)}=\frac{1}{1+(\ell_{ij}^{(k)})^{2}/(\nu_{r}\sigma_{Vr,ij}^{2})},\quad w_{t,ij}^{(k)}=\frac{1}{1+(\mathbf{d}_{\mathbf{zs}ij}^{(k)})^{\dagger}\mathbf{d}_{\mathbf{zs}ij}^{(k)}/(\nu_{t}\sigma_{Vt,ij}^{2})}.$
the IRLS update for the magnitude is:
$\bar{m}_{\mathbf{zs}i}^{(k)}:=\frac{\sum_{j\leq i}\big{(}w_{r,ij}^{(k)}/\sigma_{Vr,ij}^{2}\big{)}\hat{m}_{\mathbf{zs}ij}^{(k)}}{\sum_{j\leq i}\big{(}w_{r,ij}^{(k)}/\sigma_{Vr,ij}^{2}\big{)}}$
And we can update the magnitude with the MLE:
$m_{\mathbf{zs}i}^{(k+1)}=\bar{m}_{\mathbf{zs}i}^{(k)}$
Letting $\mathbf{\Sigma}_{Tij}^{+(k)}=\mathbf{P}_{T\mathbf{zs}i}^{(k)}/\sigma_{Vt,ij}^{2}$, we compute a tangent-space precision-weighted average:
$\mathbf{\bar{v}}_{i}^{(k)}:=\left(\sum_{j\leq i}w_{t,ij}^{(k)}\mathbf{\Sigma}_{Tij}^{+(k)}\right)^{-1}\sum_{j\leq i}w_{t,ij}^{(k)}\mathbf{\Sigma}_{Tij}^{+(k)}\mathbf{\hat{v}}_{\mathbf{zs}ij}^{(k)}$
$\left(\sum_{j\leq i}w_{t,ij}^{(k)}/\sigma_{Vt,ij}^{2}\right)^{-1}\sum_{j\leq i}\big{(}w_{t,ij}^{(k)}/\sigma_{Vt,ij}^{2}\big{)}\mathbf{\hat{v}}_{\mathbf{zs}ij}^{(k)}$
We can then update the full state estimate with a simple first-order update:
$\mathbf{z}_{\mathbf{s}i}^{(k+1)}=m_{\mathbf{zs}i}^{(k+1)}\mathbf{u}_{\mathbf{zs}i}^{(k)}+\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}$
We can do better using the (spherical) exponential map. Let:
$\theta_{i}^{(k)}:=\big{\|}\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}\big{\|},\qquad\mathbf{a}_{i}^{(k)}:=\frac{\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}}{\|\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}\|},$
The exact geodesic step is:
$\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\begin{cases}\cos\theta_{i}^{(k)}\,\mathbf{u}_{\mathbf{zs}i}^{(k)}+\sin\theta_{i}^{(k)}\,\mathbf{a}_{i}^{(k)},&\text{if }\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}\neq 0,\\ \mathbf{u}_{\mathbf{zs}i}^{(k)},&\text{if }\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}=0.\end{cases}$
A first-order approximation is:
$\mathbf{y}_{\mathbf{zs}i}^{(k+1)}=\mathbf{u}_{\mathbf{zs}i}^{(k)}+\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}/m_{\mathbf{zs}i}^{(k)},\qquad\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\frac{\mathbf{y}_{\mathbf{zs}i}^{(k+1)}}{\|\mathbf{y}_{\mathbf{zs}i}^{(k+1)}\|},$
Or, equivalently:
$\mathbf{y}_{\mathbf{zs}i}^{(k+1)}=\mathbf{z}_{\mathbf{s}i}^{(k)}+\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)},\qquad\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\frac{\mathbf{y}_{\mathbf{zs}i}^{(k+1)}}{\|\mathbf{y}_{\mathbf{zs}i}^{(k+1)}\|},$
Having updated the magnitude and direction, set:
$\mathbf{z}_{\mathbf{s}i}^{(k+1)}=m_{\mathbf{zs}i}^{(k+1)}\,\mathbf{u}_{\mathbf{zs}i}^{(k+1)}.$
Or we can use a conservative update:
$\mathbf{z}_{\mathbf{s}i}^{(k+1)}=\gamma\,\mathbf{z}_{\mathbf{s}i}^{(k)}+(1-\gamma)\,m_{\mathbf{zs}i}^{(k+1)}\mathbf{u}_{\mathbf{zs}i}^{(k+1)}.$
4.4 Implementation
In this section, we present a vectorized implementation of the RT–Kalman IRLS update. The algorithm combines scalar precision-weighted updates for radial components with a separate precision-weighted update followed by a geodesic step for tangential components.
Input sequence: $\mathbf{Z}_{r}\in\mathbb{R}^{d\times m}$
Complex linear layers: $\mathcal{L}_{Q},\mathcal{L}_{K},\mathcal{L}_{V},\mathcal{L}_{P}$
Parameters: Dynamics and noise $\{\alpha,\bm{\omega},\sigma_{r},\sigma_{t},\eta_{r},\eta_{t}\}$; temperatures $\{\tau_{r},\tau_{t}\}$; timestep $\Delta t$; residual weight $\delta$. Define $t_{i}=\Delta t\cdot i;$$t_{j}=\Delta t\cdot j.$
Constants: stability $\epsilon$, causal mask $\mathbf{M}_{\text{causal}}$ Complex-valued projections:
Lift inputs to complex form: $\mathbf{Z}=[\mathbf{Z}_{r},0]^{\top}$
$\big{(}\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big{)}\leftarrow\mathcal{L}_{V}(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z}))$
$\big{(}\operatorname{Re}(\mathbf{Z_{q}}),\operatorname{Im}(\mathbf{Z_{q}})\big{)}\leftarrow\mathcal{L}_{Q}(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}}))$
$\big{(}\operatorname{Re}(\mathbf{Z_{k}}),\operatorname{Im}(\mathbf{Z_{k}})\big{)}\leftarrow\mathcal{L}_{K}(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}}))$
Exponential kernels:
$\mathcal{K}_{\alpha}[\tau]=\mathbf{1}_{\{\tau>0\}}\odot e^{-\alpha^{2}\Delta t\tau}+\mathbf{1}_{\{\tau\leq 0\}},\quad\mathcal{K}_{\bm{\omega}}[\tau]=\mathbf{1}_{\{\tau>0\}}\odot e^{\bm{\omega}\Delta t\tau}+\mathbf{1}_{\{\tau\leq 0\}}$
Decay kernel: $\kappa[\tau]=\frac{1-\mathcal{K}_{\alpha}^{2}[\tau]+\epsilon}{2\alpha^{2}+\epsilon}\mathbf{1}_{\{\alpha^{2}>\epsilon\}}+\tau\mathbf{1}_{\{\alpha^{2}\leq\epsilon\}}$
Covariance kernels:
$\mathbf{V}^{C}_{r}[i,j]=\kappa(|i-j|)\sigma_{r}^{2}+e^{-2\alpha^{2}\Delta t|i-j|}\eta_{r}^{2}$
$\mathbf{V}^{C}_{t}[i,j]=\kappa(|i-j|)\sigma_{t}^{2}+e^{-2\alpha^{2}\Delta t|i-j|}\eta_{t}^{2}$
Precompute kernels:
$\mathbf{E}_{\mathbf{qk}}[i,j]=e^{-\alpha^{2}(t_{i}-t_{j})},\quad\mathbf{E}_{\mathbf{v}}[i,j]=e^{-\alpha^{2}(t_{i}-t_{j})},$
$\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}[k,j]=e^{-i\omega_{k}t_{j}},\quad\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}[k,j]=e^{i\omega_{k}t_{j}},$
$\bm{\tilde{\Phi}}_{\mathbf{v}}^{-}[k,j]=e^{-i\omega_{k}t_{j}},\quad\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}[k,j]=e^{i\omega_{k}t_{j}},$
$\mathbf{\tilde{Z}_{q}}=\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}\odot\mathbf{Z_{q}},\quad\mathbf{\tilde{Z}_{k}}=\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}\odot\mathbf{Z_{k}}$
Compute normalized value directions: $\mathbf{U}_{\mathbf{v}}[:,i]=\mathbf{Z_{v}}[:,i]\big{/}\left(\|\mathbf{Z_{v}}[:,i]\|+\epsilon\right)$
Radial and tangential decompositions:
$\mathbf{\tilde{m}}_{\mathbf{q}}=\mathbf{U}_{\mathbf{v}}^{\dagger}\mathbf{\tilde{Z}_{q}},\quad\mathbf{\tilde{m}}_{\mathbf{k}}=\mathbf{U}_{\mathbf{v}}^{\dagger}\mathbf{\tilde{Z}_{k}},\quad\mathbf{\tilde{m}}_{\mathbf{v}}=\mathbf{U}_{\mathbf{v}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{v}})$
$\mathbf{\tilde{V}}_{\mathbf{q}}=\mathbf{\tilde{Z}_{q}}-\mathbf{U}_{\mathbf{v}}\mathbf{\tilde{m}}_{\mathbf{q}},\quad\mathbf{\tilde{V}}_{\mathbf{k}}=\mathbf{\tilde{Z}_{k}}-\mathbf{U}_{\mathbf{v}}\mathbf{\tilde{m}}_{\mathbf{k}},\quad\mathbf{\tilde{V}}_{\mathbf{v}}=\mathbf{\tilde{Z}_{v}}-\mathbf{U}_{\mathbf{v}}\mathbf{\tilde{m}}_{\mathbf{v}}$
Compute residuals:
Tangential: $\mathbf{R}_{\text{tan}}^{2}=\|\mathbf{\tilde{V}}_{\mathbf{q}}\|_{2}^{2}\mathbf{1}_{m}^{\top}+\mathbf{1}_{m}\|\mathbf{\tilde{V}}_{\mathbf{k}}\|_{2}^{2}{}^{\top}-2\mathrm{Re}(\mathbf{E}_{\mathbf{qk}}\odot\mathbf{\tilde{V}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{V}}_{\mathbf{k}})$
Radial: $\mathbf{R}_{\text{rad}}^{2}=(\mathbf{\tilde{m}}_{\mathbf{q}}\mathbf{\tilde{m}}_{\mathbf{q}}^{\dagger})\mathbf{1}_{m}^{\top}+\mathbf{1}_{m}(\mathbf{\tilde{m}}_{\mathbf{k}}\mathbf{\tilde{m}}_{\mathbf{k}}^{\dagger})^{\top}-2\mathrm{Re}(\mathbf{E}_{\mathbf{qk}}\odot\mathbf{\tilde{m}}_{\mathbf{q}}\mathbf{\tilde{m}}_{\mathbf{k}}^{\dagger})$
Dynamic precision kernels:
$\mathbf{P}^{C}_{\text{rad}}[i,j]=1/\sigma_{V_{r}}^{2}(|t_{i}-t_{j}|),\quad\mathbf{P}^{C}_{\text{tan}}[i,j]=1/\sigma_{V_{t}}^{2}(|t_{i}-t_{j}|)$
Robust weights and attention:
$\mathbf{\tilde{A}}_{\text{rad}}[i,j]=-\log(\mathbf{V}^{C}_{\text{rad}}[i,j]+\mathbf{R}_{\text{rad}}^{2}[i,j]+\epsilon),$
$\mathbf{\tilde{A}}_{\text{tan}}[i,j]=-\log(\mathbf{V}^{C}_{\text{tan}}[i,j]+\mathbf{R}_{\text{tan}}^{2}[i,j]+\epsilon)$
$\mathbf{A}_{\text{rad}}=\operatorname{softmax}\Big{(}\tau_{r}\,\mathbf{\tilde{A}}_{\text{rad}}\odot\mathbf{M}_{\text{causal}},\;\text{dim}=1\Big{)},$
$\mathbf{A}_{\text{tan}}=\operatorname{softmax}\Big{(}\tau_{t}\,\mathbf{\tilde{A}}_{\text{tan}}\odot\mathbf{M}_{\text{causal}},\;\text{dim}=1\Big{)}$
Apply value decay and compute outputs:
$\mathbf{\hat{A}}_{\text{rad}}=\mathbf{E}_{\mathbf{v}}\odot\mathbf{A}_{\text{rad}},\quad\mathbf{\hat{A}}_{\text{tan}}=\mathbf{E}_{\mathbf{v}}\odot\mathbf{A}_{\text{tan}}$
$\mathbf{\bar{m}}_{\mathbf{v}}=\mathbf{\hat{A}}_{\text{rad}}\,\mathbf{\tilde{m}}_{\mathbf{v}},\quad\mathbf{\bar{V}}_{\mathbf{v}}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}\odot(\mathbf{\tilde{V}}_{\mathbf{v}}\,(\mathbf{\hat{A}}_{\text{tan}})^{\dagger})$
Geodesic step: $\mathbf{\bar{U}}_{\mathbf{v}}[:,i]=\left(\mathbf{Z}_{\mathbf{v}}[:,i]+\mathbf{\bar{V}}_{\mathbf{v}}[:,i]\right)\Big{/}\left(\|\mathbf{Z}_{\mathbf{v}}[:,i]+\mathbf{\bar{V}}_{\mathbf{v}}[:,i]\|+\epsilon\right)$
Reconstruct output: $\mathbf{\bar{Z}}_{\mathbf{v}}[:,i]=\mathbf{\bar{m}}_{\mathbf{v}}[i]\cdot\mathbf{\bar{U}}_{\mathbf{v}}[:,i]$
Residual connection and prediction:
$\mathbf{\bar{Z}_{v}}\leftarrow(1-\text{sigmoid}(\delta))\mathbf{Z_{v}}+\text{sigmoid}(\delta)\mathbf{\bar{Z}_{v}},\quad\mathbf{Z}_{\mathbf{v}}^{\text{pred}}=\mathbf{E_{v}}[:,1]\odot\mathbf{\bar{Z}}_{\mathbf{v}}$
Output projection:
$\big{(}\operatorname{Re}(\mathbf{Z_{p}}),\operatorname{Im}(\mathbf{Z_{p}})\big{)}\leftarrow\mathcal{L}_{P}(\operatorname{Re}(\mathbf{Z}_{\mathbf{v}}^{\text{pred}}),\operatorname{Im}(\mathbf{Z}_{\mathbf{v}}^{\text{pred}}))$
Return: $\operatorname{Re}(\mathbf{Z_{p}})$
Algorithm 3 RT Adaptive Filter Attention
4.5 Interpreting the Transformer as an MLE on a Hypersphere
The algorithm above faithfully implements RT-MLE with geometric constraints, but is somewhat intricate. We propose the following simplifications for clarity and practical use.
Recall that, in RT-Kalman IRLS our estimated state decomposes as:
$\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}=e^{-\bm{\lambda}\Delta t_{ij}}\mathbf{z_{s}}_{j}^{(k)}=\Big{(}e^{-\alpha\Delta t_{ij}}m_{\mathbf{zs}j}^{(k)}\Big{)}\Big{(}e^{-\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{u}_{\mathbf{zs}j}^{(k)}\Big{)}$
where $\mathbf{\hat{u}}_{\mathbf{zs}ij}^{(k)}=\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}/\|\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}\|=e^{-\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{u}_{\mathbf{zs}j}^{(k)}$. Assuming that the transformed magnitude is approximately constant across time steps $e^{-\alpha\Delta t_{ij}}m_{\mathbf{zs}j}^{(k)}\approx m_{\mathbf{zs}i}^{(k)}$, gives a simple precision-weighted average of $\mathbf{\hat{u}}$:
$\mathbf{\bar{v}}_{\mathbf{s}i}^{(k)}\approx m_{\mathbf{zs}i}^{(k)}\mathbf{\bar{u}}_{\mathbf{zs}i}^{(k)},\quad\text{where:}\quad\mathbf{\bar{u}}_{\mathbf{zs}i}^{(k)}\approx\left(\sum_{j\leq i}w_{t,ij}^{(k)}/\sigma_{Vt,ij}^{2}\right)^{-1}\sum_{j\leq i}\big{(}w_{t,ij}^{(k)}/\sigma_{Vt,ij}^{2}\big{)}\mathbf{\hat{u}}_{\mathbf{zs}ij}^{(k)}$
For small angles, the attention update is largely aligned with the tangent space and any component along $u$ is partially corrected by the subsequent geodesic normalization, so we omit the explicit tangent projection for simplicity.
The geodesic update step becomes:
$\mathbf{y}_{\mathbf{zs}i}^{(k)}\approx\mathbf{u}_{\mathbf{zs}i}^{(k)}+\mathbf{\bar{u}}_{\mathbf{zs}i}^{(k)},\qquad\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\frac{\mathbf{y}_{\mathbf{zs}i}^{(k)}}{\|\mathbf{y}_{\mathbf{zs}i}^{(k)}\|},$
And the output is again:
$\mathbf{z}_{\mathbf{s}i}^{(k+1)}=m_{\mathbf{zs}i}^{(k+1)}\,\mathbf{u}_{\mathbf{zs}i}^{(k+1)}.$
This structure — Norm, Attention, Add & Norm — closely resembles a Transformer.
This suggests that the Transformer’s effectiveness may stem, at least in part, from its ability to approximate a principled filtering mechanism, with attention functioning as a generalized MLE for an SDE with unitary (or norm-preserving) linear dynamics, and normalization layers realizing approximate geodesic steps along the hypersphere.
We present the following Transformer-inspired simplification of RT-AFA:
Input: Input sequence $\mathbf{Z}_{r}\in\mathbb{R}^{d\times m}$
Parameters: AFA layer $\mathrm{AFA}_{\mathrm{dir}}$, scalar AFA layer $\mathrm{AFA}_{\mathrm{mag}}$, output complex linear layer $\mathcal{L}_{O}$, real-valued FFN layers: $\mathrm{FFN}=\mathcal{L}_{2}\circ\mathrm{ReLU}\circ\mathcal{L}_{1}$, stability const. $\epsilon$.
Magnitude and unit direction:
$\mathbf{m}=\|\mathbf{Z}_{r}\|_{\text{rowwise}}\in\mathbb{R}^{1\times m}$
$\mathbf{U}_{r}=\gamma_{1}\mathbf{Z}_{r}\big{/}(\mathbf{m}+\epsilon)+\beta_{1}\quad(\text{element-wise division})$
Lift to complex:
$\mathbf{U}=[\mathbf{U}_{r};\mathbf{0}]\in\mathbb{C}^{2\times d\times m}$
Apply AFA attention:
$\mathbf{\hat{m}}=\mathrm{AFA}_{\mathrm{mag}}(\mathbf{m})$
$\mathbf{\hat{U}}=\mathrm{AFA}_{\mathrm{dir}}(\mathbf{U})$
Geodesic step:
$\mathbf{Y}=\mathbf{U}+\mathbf{\hat{U}}$
$\mathbf{m}_{y}=\|\mathbf{Y}\|_{\text{row-wise}}$
$\mathbf{\tilde{U}}=\gamma_{2}\mathbf{Y}\big{/}(\mathbf{m}_{y}+\epsilon)+\beta_{2}\quad(\text{element-wise division})$
Combine magnitude and direction estimates:
$\mathbf{Z}=\mathbf{\hat{m}}\odot\mathbf{\tilde{U}}$
Output complex linear projection:
$(\operatorname{Re}(\mathbf{Z_{o}}),\operatorname{Im}(\mathbf{Z_{o}}))\leftarrow\mathcal{L}_{O}(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z}))$
Output FFN with nonlinearity and residual connection:
$\mathbf{Z}^{\mathrm{out}}=\mathbf{Z}_{r}+\mathrm{FFN}(\operatorname{Re}(\mathbf{Z_{o}}))$
Output: Updated sequence $\mathbf{Z}^{\mathrm{out}}\in\mathbb{R}^{d\times m}$
Algorithm 4 AFA Transformer
5 Related Work
5.1 Classical Filtering Methods
Norbert Wiener’s 1949 work on the Wiener Filter established the theoretical framework for minimum mean-square error estimation of signals from noisy observations, using second-order statistics (covariances) of the signal and noise (Wiener, 1964). Kalman introduced the discrete-time Kalman Filter, a recursive algorithm that optimally estimates the states of linear time-invariant stochastic systems from noisy measurements, achieving minimum mean-square estimation error (Kalman, 1960). Kalman & Bucy extended the approach to continuous-time systems with Gaussian noise (Kalman and Bucy, 1961). Rauch, Tung, and Striebel generalized the Kalman Filter to a batch smoothing framework, providing a maximum-likelihood estimator for the entire trajectory of a linear dynamical system by combining all measurements simultaneously, rather than recursively, which yields smoothed state estimates with minimum mean-square error over the full time horizon (Rauch et al., 1965). Extensions to nonlinear systems can be handled by the Extended Kalman Filter (EKF), which linearizes the dynamics and measurement functions about the current state estimate (Anderson and Moore, 1979). For introductions to stochastic processes and adaptive filtering, see (Jazwinski, 1970; Speyer, 2008; Haykin, 2002).
5.2 Robust and Adaptive Filtering
In many practical settings, the system and noise parameters required by the Kalman Filter are unknown or time-varying. Early work in adaptive signal processing began with Widrow & Hoff, who introduced the Least Mean Squares (LMS) algorithm, laying the foundation for recursive adaptation using the innovation signal (Widrow and Hoff, 1960). Schweppe provided early maximum likelihood foundations for estimating noise covariances in Gaussian signals, which became a cornerstone for statistically principled adaptive filtering (Schweppe, 1965). Mehra formalized adaptive Kalman Filtering by estimating unknown process and measurement noise covariances from the innovations sequence (Mehra, 1970), and compared innovation-based versus residual-based adaptation (Mehra, 1972). Iteratively reweighted least squares (IRLS) provides a framework for robust estimation by repeatedly solving weighted least-squares problems, with weights updated based on residuals; this approach was formalized in signal processing contexts by Burrus, Barreto, and Selesnick for FIR filter design (Burrus et al., 1994). Comprehensive treatments of stochastic estimation and adaptive methods include (Jazwinski, 1970) and (Maybeck and Siouris, 1980).
5.3 Neural Networks for Filtering Dynamical Systems
Kalman Filtering and Neural Networks, edited by Haykin, provides a comprehensive treatment of combining classical state estimation with neural network models (Haykin, 2001). Other work demonstrated that multilayer and recurrent neural networks can be effectively used for the identification and adaptive control of nonlinear dynamical systems, for example Narendra and Parthasarathy (Narendra and Parthasarathy, 1990).
Krishnan et al. introduced Deep Kalman Filters (DKFs), which extend classical linear Kalman Filters to high-dimensional, non-linear latent dynamics using deep neural networks. DKFs maintain a probabilistic latent state and propagate uncertainty explicitly. In this framework, the latent state is updated sequentially according to a learned, potentially nonlinear transition function, and observations are generated from the latent state through a neural network (Krishnan et al., 2015).
Liu et al. propose Kalman Filtering Attention (KFAtt) for CTR prediction, treating each user click as a noisy measurement of a hidden query-specific interest and computing a closed-form MAP estimate that combines these measurements with a global prior. Despite the “Kalman” terminology, KFAtt does not model temporal dynamics or state evolution, and effectively performs a static, precision-weighted aggregation of observations.
Goel and Bartlett investigate whether Transformers can represent classical Kalman Filters for linear time-invariant (LTI) systems. Their Transformer Filter uses softmax self-attention over nonlinear embeddings of past state estimates and observations. It effectively approximates the Kalman update indirectly, with the kernel weights mimicking the Kalman gain, and requires careful tuning of a temperature parameter to control approximation error (Goel and Bartlett, 2024). While this approach can produce state estimates arbitrarily close to the true Kalman Filter and approximate the behavior of an LQG controller, it does not naturally generalize to robust, time-varying, or non-Gaussian settings, nor does it explicitly propagate uncertainty in a structured way.
Several recent works integrate neural networks or attention into Kalman Filtering. KalmanNet uses a recurrent network to learn the Kalman gain for real-time filtering under partially known or nonlinear dynamics (Revach et al., 2022). The Neural Extended Kalman Filter parameterizes both process and observation models with neural networks for end-to-end latent state inference (Liu et al., 2023). The Attention Kalman Filter (AtKF) replaces the traditional Kalman gain with a self-attention network that learns to capture dependencies across past state and innovation sequences (Wang et al., 2024). A-KIT (2024) learns process noise covariance online to improve EKF predictions (Cohen and Klein, 2024). KalmanFormer integrates a Transformer network into the Kalman Filter framework to directly learn the Kalman gain from data. (Shen et al., 2025).
The present work is an extension of the MS Thesis, "Adaptive Filter Attention," which shows that a tensorized attention formulation exactly solves the MLE for a diagonalizable LTI system with Gaussian noise (Racioppo, 2025).
Uncertainty Propagation Networks (UPNs) extend neural ODEs to jointly model state means and covariances through coupled differential equations, allowing continuous-time propagation of uncertainty sequence modeling. UPN propagates uncertainty continuously through time by explicitly evolving the covariance of the state alongside the state mean, essentially functioning as a continuous-time Kalman Filter embedded in a learned dynamics model (Jahanshahi and Zhu, 2025).
5.4 Neural State Space Models
State space models (SSMs) are a powerful framework for modeling sequences and time series data by representing system dynamics through latent states that evolve over time. A key challenge in sequence modeling is maintaining memory over long time horizons, and avoiding the vanishing and exploding gradient problem faced by recurrent neural networks.
The HiPPO framework improves memory retention in SSMs by introducing recurrent memory with optimal polynomial projections, which allows the model to compress and retain important information from the entire history in a structured latent state space (Gu et al., 2020). The Structured State Space Sequence model (S4) designs efficient state space layers that scale linearly with sequence length, using structured representations of the system matrices (Gu et al., 2022).
Subsequent works such as Mamba further enhance efficiency by introducing selective state space components that dynamically focus computation on relevant parts of the sequence (Gu and Dao, 2024). The S4ND model applies state space models to multidimensional signals such as images and videos, interpreting them as higher-dimensional spatiotemporal signals (Nguyen et al., 2022) . Goel et al. propose a multi-scale architecture based on S4 designed specifically for raw audio waveform generation, and address instability issues in SSM sequence modeling (Goel et al., 2022). Mamba Mixer combines structured state space models with selective token and channel mixing via learned gating, enabling efficient and adaptive modeling of long sequences by focusing on relevant elements and features (Behrouz et al., 2024). Finally, simplified variants of state space layers have been proposed to reduce model complexity and improve training stability while retaining the key benefits of state space modeling in sequence tasks (Smith et al., 2023).
Both S4 and our model exploit a diagonal-plus-low-rank (DPLR) structure to reduce the cost of inverting a dense $N\times N$ matrix, using the Sherman-Morrison formula for efficiency: a diagonal part captures independent modes, and a rank-1 correction captures global interactions. In S4, the DPLR matrix models deterministic state dynamics, determining how past inputs influence future states. In our model, the DPLR structure appears in the noise covariance matrices in a stochastic differential equation (SDE) formulation, which allows us to represent how uncertainty propagates through the dynamics.
Our formulation relies on complex-valued weight matrices. Prior work on complex-valued neural networks includes studies of complex-valued activation functions, parameter initialization strategies, and the use of Wirtinger calculus, which treats complex parameters as pairs of real variables to enable effective gradient computation (Trabelsi et al., 2018), (Barrachina et al., 2023),(Hammad, 2024).
5.5 Attention as a Structured State Space Computation
The attention mechanism was first introduced in (Bahdanau et al., 2016) to improve neural machine translation by allowing models to dynamically focus on relevant input tokens, and was later generalized into the self-attention architecture of the Transformer by Vaswani et al, which became the foundation for modern large language models (Vaswani et al., 2023).
Attention mechanisms can be interpreted through the lens of structured state space computations. Several recent works have explored these connections and proprosed structured alternatives to standard Transformer attention based on state space models.
Katharopoulos et al. propose a Linear Attention framework that approximates attention weights using kernel feature maps, enabling the Transformer to be expressed in a recurrent form. While this formulation offers computational benefits, it does not explicitly connect the recurrence to an underlying system dynamics model; rather, the recurrent structure serves primarily as an efficient implementation (Katharopoulos et al., 2020).
Dao and Gu establish that a state space model (SSM) transformation is mathematically equivalent to multiplication by an $N$-sequentially semiseparable (SSS) matrix. This equivalence means that the recurrence inherent in SSMs can be expressed as a structured matrix multiplication, which directly parallels the matrix form of self-attention in Transformers. In particular, the attention matrix in a Transformer can be interpreted as a low-rank or structured matrix operating along the sequence dimension, making it amenable to the same efficient algorithms developed for SSS matrices. (Dao and Gu, 2024). While this formulation offers computational benefits and is grounded in a linear state-space model, it does not explicitly propagate uncertainty through the sequence or perform adaptive, residual-based reweighting.
Other recent works explore the integration of attention mechanisms with state space models. For example, the Dynamical Systems Framework (DSF) unifies attention, state space models, and recurrent networks under a common recurrent/convolutional form (Sieber et al., 2024). Bick et al. focus on distilling quadratic self-attention into efficient SSMs, with the goal of matching the performance of Transformers while reducing computation (Bick et al., 2025).
5.6 Dynamical Systems Perspectives on Attention
Attention mechanisms have been studied through continuous-time and dynamical systems frameworks, modeling hidden states or latent dynamics via ordinary differential equations (ODEs) or stochastic processes to capture temporal dependencies, uncertainty, and adaptivity.
Stochastic Recurrent Networks (STORNs) combine RNNs with time-dependent latent variables and variational inference to model multimodal temporal dynamics (Bayer and Osendorfer, 2015). Neural ODEs treat hidden states as continuous-time dynamics parameterized by a neural network and solved with an ODE solver (Chen et al., 2019). Neural Processes (NPs) and Neural ODE Processes (NDPs) extend this framework to learn distributions over functions or ODE trajectories, enabling uncertainty-aware predictions (Garnelo et al., 2018; Norcliffe et al., 2021). Rubanova et al. introduce ODE-RNNs and Latent ODEs, continuous-time models that handle irregularly-sampled time series without discretization or imputation (Rubanova et al., 2019). Physics-informed neural networks (PINNs) integrate differential equation constraints into neural networks to enforce physically consistent solutions (Raissi et al., 2017).
Attentive Neural Processes (ANPs) introduce attention over context sets, while Attentive Differential Networks (ADNs) handle irregularly sampled data using continuous-time attention (Kim et al., 2019; Chien and Chen, 2023). ACE-NODE and ANCDE integrate attention into ODE or controlled differential equation frameworks, learning attention dynamics continuously rather than via standard query-key comparisons (Jhin et al., 2021a, b). Neural ODE Transformers and Hopfield Networks reinterpret attention as evolving or energy-based dynamical systems (Tong et al., 2025; Ramsauer et al., 2021). Continuous-discrete state space models and Structured Inference Networks combine latent continuous dynamics with discrete observations and approximate Bayesian filtering (Ansari et al., 2023; Krishnan et al., 2016).
Geshkovski et al. model Transformers as interacting particle systems evolving on the unit sphere, and find that self-attention induces nonlinear dynamics leading to clustering of token embeddings (Zhang et al., 2021; Geshkovski et al., 2024). Molina interprets Transformers geometrically, showing that LayerNorm constrains hidden states to a hypersphere and attention moves “word particles” along this surface (Molina, 2023). Brody et al. show that LayerNorm shapes Transformer hidden states geometrically—projecting them onto a hyperplane and scaling them—which enhances the expressiveness of attention (Brody et al., 2023).
5.7 Diffusion, Score-Based Models, and Neural SDEs
Song et al. showed that diffusion models can be formulated as continuous-time stochastic differential equations (SDEs), where learning the time-dependent score allows high-quality sampling via SDE solvers (Song et al., 2021). FIM-SDE estimates drift and diffusion functions from noisy, sparse SDE observations by learning a generalizable mapping from trajectories to underlying dynamical system parameters (Seifner et al., 2025). Tzen and Raginsky show that deep latent Gaussian models converge in the infinite-layer limit to Itô SDEs with neural-network-parameterized drift and diffusion (Tzen and Raginsky, 2019). Wang and Yao combine neural processes with neural SDEs to improve uncertainty estimation on noisy or incomplete data, using attention-based NPs for multidimensional inputs (Wang and Yao, 2021).
5.8 Structuring Attention with Priors and Positional Biases
The original Transformer lacks built-in modeling of relative or absolute positions, relying instead on added absolute positional encodings (Vaswani et al., 2023).
Several methods have proposed extending self-attention to efficiently incorporate relative position information. For example, Shaw et al. introduce a mechanism to encode relative distances between tokens directly into the attention computation (Shaw et al., 2018). Transformer-XL introduces relative positional encodings, enabling the model to generalize to sequences longer than those seen during training (Dai et al., 2019). Su et al. apply a position-dependent rotation to each dimension of the query and key vectors before computing their dot products. This rotation encodes the relative distance between positions directly into the attention mechanism (Su et al., 2023).
A number of works reinterpret sequence modeling as a controlled, time-aware aggregation of past information, where learned or adaptive decay mechanisms regulate the influence of prior inputs over time. These models typically rely on heuristic, learned gating or decay functions for memory control, whereas we derive update weights from robust, data-driven precision estimates grounded in Bayesian filtering.
The Retentive Network (RetNet) introduces a retention mechanism that blends each token’s hidden state with an exponentially decaying summary of previous tokens. RetNet uses multiple attention heads, each having its own learned exponential decay rate, enabling multi-scale memory of different temporal lengths. Attention is applied with decay masks that enforce causality and decay over relative positions, while a gating mechanism modulates the retained summaries to improve expressivity and control the contribution of past information (Sun et al., 2023).
GateLoop generalizes linear RNNs with data-controlled, time-varying state transition gates and introduces a surrogate attention mechanism where each input contributes via learned, input-dependent decay factors, capturing content and relative position (Katsch, 2024). Gated Linear Attention (GLA) uses data-dependent, time-varying gates to modulate hidden state updates, combining fast, recurrent linear attention with input-adaptive, relative position-aware decay (Yang et al., 2024). The Real-Gated Linear Recurrent Unit (RG-LRU) introduces input and recurrence gates into a stable linear recurrent update, avoiding complexity from softmax attention or continuous-time dynamics. Griffin interleaves RG-LRU blocks with lightweight local attention (De et al., 2024). The Dynamically Context-Sensitive Time-Decay Attention mechanism learns a time-decay function to adjust attention weights (Su et al., 2018). This approach focuses on temporal proximity rather than integrating known system dynamics. Performer approximates softmax attention with positive-definite kernel methods to achieve scalable, linear-time self-attention for long sequences (Choromanski et al., 2022).
Several papers have explored element-wise (in the embedding dimension) attention mechanisms. McCarter investigates replacing scaled dot-product attention with negative-log Euclidean distance (McCarter, 2023). Elliptical Attention replaces the traditional dot-product attention with a Mahalanobis distance-based mechanism (Nielsen et al., 2024). Feng uses an exponentiated Euclidean distance as a similarity metric (Feng, 2025). Bacochina et al. replace matrix multiplications with learned spatial weight arrays applied directly via element-wise multiplication, focusing attention along rows and columns (Bacochina and de Souza, 2023). Wu et al. introduce the DA-Transformer, which reweights attention scores using learnable functions of token distances (Wu et al., 2021)
5.9 Modeling Uncertainty in Neural Networks
Other approaches explicitly model uncertainty in neural networks and attention mechanisms. Foundational works such as (MacKay, 1992) and (Neal, 1996) laid the groundwork for modern developments in approximate Bayesian inference in deep learning.
Chung et al. introduce the Variational Recurrent Neural Network (VRNN), which incorporates latent random variables into RNN hidden states to represent and propagate uncertainty in sequential data (Chung et al., 2016). Fraccaro et al. introduce the stochastic recurrent neural network, a sequential model combining deterministic RNN states with stochastic latent variables, using structured variational inference to propagate uncertainty over time and improve sequence modeling (Fraccaro et al., 2016).
The Bayesian Attention Mechanism (BAM) interprets attention weights as joint probabilities over token content and position, showing that positional encoding strategies correspond to specific priors and introducing a generalized Gaussian prior (GGD-BAM) that balances local and long-range dependencies (Bianchessi et al., 2025). Litman shows that scaled dot-product attention is the exact solution to a one-sided entropic optimal transport problem (Litman, 2025). The Correlated Gaussian Process Transformer models self-attention as the cross-covariance between two correlated Gaussian processes (Bui et al., 2025). Ruan and Zhang analyze self-attention as a similarity-based information propagation process converging to a drift-diffusion or heat equation on a manifold (Ruan and Zhang, 2024).
Various approaches embed structured or probabilistic biases directly into attention mechanisms. For example, Probabilistic Transformers model attention as MAP inference in a Gaussian mixture model, with each key/value pair having its own precision parameter, updated via expectation-maximization (Movellan and Gabbur, 2020). Uncertainty-Aware Attention introduces input-specific uncertainty modeling through variational inference (Heo et al., 2018). The Attention Gaussian Process (AGP) treats attention weights as probabilistic outputs from a sparse Gaussian process, with variational inference approximating the posterior (Schmidt et al., 2023). SLANG uses a diagonal plus low-rank covariance to efficiently estimate uncertainty in deep networks via natural-gradient variational inference (Mishkin et al., 2019).
Other recent probabilistic attention methods introduce uncertainty-aware mechanisms into Transformer-based architectures. These include the Uncertainty-Guided Probabilistic Transformer (UGPT) (Guo et al., 2022), Attention Gaussian Process (AGP) (Schmidt et al., 2023), Gaussian Adaptive Attention (GAAM) (Ioannides et al., 2024), the Correlated Gaussian Process Transformer (CGPT) (Bui et al., 2025), and Kernel-Eigen Pair Sparse Variational Gaussian Processes (KEP-SVGP) (Chen et al., 2024).
5.10 Structured Low-Rank Covariances for Efficient Propagation
Several methods have been developed to efficiently represent and propagate high-dimensional covariance matrices using low-rank plus diagonal approximations in Riccati-like equations. Pnevmatikakis et al. propose a low-rank perturbative approach for Kalman Filtering and smoothing, approximating the state covariance as a small-rank update to the prior (Pnevmatikakis et al., 2014). Saibaba et al. represent the covariance as a weighted sum of the process noise and a low-rank term from a generalized eigenvalue problem (Saibaba et al., 2015). Bonnabel et al. approximate large, time-dependent PSD matrices with low-rank plus diagonal structures, allowing linear-time storage, inversion, and integration in high-dimensional Riccati flows (Bonnabel et al., 2024). Schmidt et al. (2021) introduce the Rank-Reduced Kalman Filter, projecting the Lyapunov equations onto a low-rank manifold and updating only the covariance column space, reducing complexity from cubic to quadratic (or linear) while recovering the exact filter in the full-rank limit and outperforming ensemble-based approaches (Schmidt et al., 2024).
5.11 Unrolling Optimization Algorithms and Connections to Attention
Unrolling iterative optimization algorithms into trainable networks began with frameworks like LISTA, which expresses the ISTA sparse-coding algorithm as a neural network with learned parameters (Gregor and LeCun, 2010). Deep Equilibrium Models (DEQs) extend this idea by solving a single equilibrium layer via root-finding, effectively modeling infinite-depth networks with constant memory (Bai et al., 2019). Wang et al. study how attention can also be viewed as an unrolled iterative process: multi-head self-attention corresponds to iterative subspace denoising, compressing inputs into a low-rank manifold (Wang et al., 2025).
6 Discussion and Conclusion
This work explores a connection between classical control theory and attention mechanisms by framing attention through the lens of linear dynamical systems and adaptive filtering. We show that attention arises naturally as the solution to a maximum likelihood estimation problem under a linear stochastic differential equation with structured system matrices and noise covariances, and discrete observations. In this view, the attention update implements a parallelized form of adaptive Kalman Filtering: each key-value pair is interpreted as a noisy observation of a latent state, and a precision-weighted Bayesian update aggregates evidence across all observations to refine the state estimate in a parallel manner. To the best of our knowledge, this view of attention as the MLE for a structured linear SDE has not appeared in the literature on attention mechanisms.
Several promising directions remain for future work. It should be straightforward to extend the method to the LQG controller, and the framework could also be adapted to reinforcement learning settings, where the dynamics-aware attention mechanism might provide more sample-efficient policy evaluation and state estimation in partially observed or noisy environments.
The closed-form expression for the propagated precision matrix, which assumes simultaneous diagonalizability, can be computed in $\mathcal{O}(md)$ time, whereas full integration and inversion in the general case requires $\mathcal{O}(md^{3})$ time. Since our algorithm already involves $\mathcal{O}(md^{2})$ matrix multiplications, moderately less efficient integration methods introduce a limited additional overhead. In particular, if we use a diagonalizable-plus-rank-1 state matrix $\mathbf{A}$, covariance propagation can be computed efficiently by expressing the matrix exponential and integrals as a diagonal term plus a rank-1 outer-product correction, reducing complexity to $\mathcal{O}(md^{2})$, though losing a fully closed form. We could also consider a generalization using the Jordan canonical form. Alternatively, differentiable numerical solvers for the differential Lyapunov equation, or unrolled iterations of specialized low-rank methods such as Krylov subspace techniques or ADI iterations could potentially be integrated directly into a neural network layer.
To achieve the statistical accuracy of the full batch estimator while retaining the computational efficiency of a recursive filter, one could explore alternative unrolling strategies. These may include inner-loop iterations within each time step, linearizing weights around a precomputed batch solution, or hybrid schemes that propagate residual-dependent weights more globally across the trajectory.
In this work, we assume the output matrix $\mathbf{C}$ is square and invertible, so the latent state is fully observable. We can account for partially observed states with projections or pseudo-inverses to map between measurement and latent spaces.
The layer-wise updates in our model perform adaptive denoising, which parallels the iterative refinement in diffusion models. Future work could investigate these connections. Finally, it would be worthwhile to explore whether linking attention to system dynamics and measurement confidence can improve model interpretability.
Appendix A Background
A.1 Linear Systems
A continuous-time linear system is defined, in state-space form, as:
$\mathbf{\dot{x}}(t)=\mathbf{A}(t)\mathbf{x}(t)+\mathbf{B}(t)\mathbf{u}(t)$
$\mathbf{y}(t)=\mathbf{C}(t)\mathbf{x}(t)+\mathbf{D}(t)\mathbf{u}(t)$
The signals:
$\textbf{x}(t):[0,\infty)\to\mathbb{R}^{N},\quad\textbf{u}(t):[0,\infty)\to\mathbb{R}^{M},\quad\textbf{y}(t):[0,\infty)\to\mathbb{R}^{P}$
are known as the state, input, and output, respectively. In control applications, $\mathbf{u}(t)$ is usually a feedback control term. $\mathbf{A}(t),\mathbf{B}(t),\mathbf{C}(t),\mathbf{D}(t)$ are known as the state matrix, input matrix, output matrix, and direct transition matrix, respectively.
Letting $\textbf{x}(t_{0})=\textbf{x}_{0}$, the unique solution to the above system is given by:
$\textbf{x}(t)=\mathbf{\Phi}(t,t_{0})\textbf{x}_{0}+\int_{t_{0}}^{t}\mathbf{\Phi}(t,\tau)\mathbf{B}(\tau)\textbf{u}(\tau)\,d\tau,$
$\textbf{y}(t)=\mathbf{C}(t)\mathbf{\Phi}(t,t_{0})\textbf{x}_{0}+\int_{t_{0}}^{t}\mathbf{C}(t)\mathbf{\Phi}(t,\tau)\mathbf{B}(\tau)\textbf{u}(\tau)\,d\tau+\mathbf{D}(t)\textbf{u}(t),$
where $\mathbf{\Phi}(t,t_{0})$, the state transition matrix, is the unique solution to
$\frac{d}{dt}\mathbf{\Phi}(t,t_{0})=\mathbf{A}(t)\mathbf{\Phi}(t,t_{0}),$
[…]
A linear time-invariant (LTI) system is one in which the matrices $\mathbf{A},\mathbf{B},\mathbf{C},\textbf{D}$ are fixed, real-valued matrices, (i.e. not functions of time): $\mathbf{A}\in\mathbb{R}^{N\times N}$, $\mathbf{B}\in\mathbb{R}^{N\times M}$, $\mathbf{C}\in\mathbb{R}^{P\times N}$, $\mathbf{D}\in\mathbb{R}^{P\times M}$.
An autonomous LTI system $\mathbf{\dot{x}}(t)=\mathbf{A}\mathbf{x}(t)$ is asymptotically stable (i.e. $\mathbf{x}(t)$ approaches a stable equilibrium point as $t\rightarrow\infty$) if $\mathbf{A}$ is Hurwitz, that is if every eigenvalue of $\mathbf{A}$ has strictly negative real part.
In an LTI system, the state transition matrix is given by:
$\mathbf{\Phi}(t,t_{0})=e^{\mathbf{A}(t-t_{0})},$
where the matrix exponential $e^{\mathbf{M}}$ is defined as:
$e^{\mathbf{M}}=\sum_{k=0}^{\infty}\frac{1}{k!}\mathbf{M}^{k}.$
If $\mathbf{A}$ is diagonalizable, i.e. $\mathbf{A}=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1}$ for some invertible $\mathbf{S}$ and diagonal $\bm{\Lambda}$, then the matrix exponential can be computed cheaply using: $e^{\mathbf{A}}=\mathbf{S}e^{\bm{\Lambda}}\mathbf{S}^{-1}$. See also, standard references such as (Antsaklis, 1997), (Skogestad and Postlethwaite, 2005).
A.2 Stochastic Processes
Many real-world systems are subject to uncertainty, which can be modeled as stochastic disturbances or process noise. A continuous-time linear time-invariant system with additive Gaussian noise is described by the Itô stochastic differential equation (SDE):
$d\mathbf{x}(t)=\mathbf{A}\mathbf{x}(t)\,dt+\mathbf{B}\,d\mathbf{w}(t),\qquad\mathbf{z}(t_{k})=\mathbf{C}\mathbf{x}(t_{k})+\mathbf{v}(t_{k}),$
where
  • $\mathbf{x}(t)\in\mathbb{R}^{N}$ is the latent state (or just state),
  • $\mathbf{z}(t_{k})\in\mathbb{R}^{P}$ is the observation at discrete times $t_{k}$,
  • $\mathbf{A}$ is the state matrix, and $\mathbf{B}$ is the input/noise matrix,
  • $\mathbf{C}$ is the measurement matrix,
  • $d\mathbf{w}(t)\sim\mathcal{N}(\mathbf{0},\mathbf{I}\,dt)$ is a standard Wiener process,
  • $\mathbf{v}(t_{k})\sim\mathcal{N}(\mathbf{0},\mathbf{R})$ is Gaussian measurement noise.
The continuous-time process noise covariance is defined as: $\mathbf{Q}:=\mathbf{B}\mathbf{B}^{\top}.$
Discrete-time form.
Sampling the SDE at equal intervals $\Delta t$ yields a linear Gaussian state-space model:
$\mathbf{x}_{i}=\mathbf{A}_{d}\mathbf{x}_{i-1}+\bm{\eta}_{i},\qquad\mathbf{z}_{i}=\mathbf{C}\mathbf{x}_{i}+\mathbf{v}_{i},$
with $\bm{\eta}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{d})$ and $\mathbf{v}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{R})$. Here, $\mathbf{A}_{d}=e^{\mathbf{A}\Delta t}$ and $\mathbf{Q}_{d}=\int_{0}^{\Delta t}e^{\mathbf{A}\tau}\mathbf{Q}\,e^{\mathbf{A}^{\top}\tau}\,d\tau$.
This stochastic formulation provides the foundation for linear Gaussian estimation and filtering, including the classical Kalman Filter. Standard references include: (Antsaklis, 1997; Skogestad and Postlethwaite, 2005; Jazwinski, 1970).
A.3 The Kalman Filter
Filtering, or state estimation, is the task of estimating the states $\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ from noisy measurements $\{\mathbf{z}_{1},\dots,\mathbf{z}_{n}\}$.
The Kalman Filter is the optimal recursive estimator for a linear system with Gaussian noise. It maintains a mean estimate $\mathbf{\hat{x}}_{k}$ and a covariance $\mathbf{V}_{k}$ that quantifies uncertainty, and updates them sequentially as new measurements arrive.
The Kalman Filter alternates between prediction steps (propagating the state forward using the system model) and update steps (correcting the prediction using measurements).
Prediction Step:
The model dynamics $\mathbf{A}$ propagate the previous state forward, and process noise $\mathbf{Q}$ increases uncertainty:
$\mathbf{\hat{x}}_{k|k-1}=\mathbf{A}\mathbf{\hat{x}}_{k-1},\qquad\mathbf{V}_{k|k-1}=\mathbf{A}\mathbf{V}_{k-1}\mathbf{A}^{\top}+\mathbf{Q}.$
Update Step:
When a new measurement $\mathbf{z}_{k}$ is available, the filter computes a Kalman gain $\mathbf{K}_{k}$ that weights the relative uncertainty in the prediction and the measurement:
$\mathbf{K}_{k}=\mathbf{V}_{k|k-1}\mathbf{C}^{\top}(\mathbf{C}\mathbf{V}_{k|k-1}\mathbf{C}^{\top}+\mathbf{R})^{-1}.$
The state and covariance are updated to reflect the new information:
$\mathbf{\hat{x}}_{k}=\mathbf{\hat{x}}_{k|k-1}+\mathbf{K}_{k}(\mathbf{z}_{k}-\mathbf{C}\mathbf{\hat{x}}_{k|k-1}),\quad\mathbf{V}_{k}=(\mathbf{I}-\mathbf{K}_{k}\mathbf{C})\mathbf{V}_{k|k-1}.$
When the measurement is very precise (small $\mathbf{R}$), the gain $\mathbf{K}_{k}$ is large, and the estimate relies more on the observation; when the model is more certain, the estimate relies more on the prediction.
We refer to $\mathbf{\hat{x}}_{k|k-1}$ and $\mathbf{V}_{k|k-1}$ as the a priori estimates of the state and covariance (i.e. those computed before accounting for the latest measurement), and $\mathbf{\hat{x}}_{k}$ and $\mathbf{V}_{k}$ as the a posteriori estimates (which include the latest measurement).
The difference between the actual measurement and the predicted measurement, $\mathbf{y}_{k}:=\mathbf{z}_{k}-\mathbf{C}\mathbf{\hat{x}}_{k|k-1},$ is called the innovation and represents the new information that the measurement brings relative to the model prediction. The difference between the measurement and the updated estimate, $\mathbf{r}_{k}:=\mathbf{z}_{k}-\mathbf{C}\mathbf{\hat{x}}_{k},$ is called the residual and quantifies the discrepancy after incorporating the measurement. See also: (Jazwinski, 1970; Anderson and Moore, 1979; Speyer, 2008)
Continuous-time Kalman Filtering.
For continuous-time linear stochastic systems, the Kalman Filter generalizes to the Kalman–Bucy filter. Given
$d\mathbf{x}(t)=\mathbf{A}(t)\mathbf{x}(t)\,dt+\mathbf{B}(t)d\mathbf{w}(t),\quad\mathbf{z}(t)=\mathbf{C}(t)\mathbf{x}(t)+\mathbf{v}(t),$
with process noise covariance $\mathbf{Q}(t)=\mathbf{B}(t)\mathbf{B}(t)^{\top}$ and measurement noise covariance $\mathbf{R}(t)$, the state estimate $\mathbf{\hat{x}}(t)$ evolves according to:
$\frac{d}{dt}\mathbf{\hat{x}}(t)=\mathbf{A}(t)\mathbf{\hat{x}}(t)+\mathbf{K}(t)\big{(}\mathbf{z}(t)-\mathbf{C}(t)\mathbf{\hat{x}}(t)\big{)},$
where the continuous-time Kalman gain $\mathbf{K}(t)$ is given by:
$\mathbf{K}(t)=\mathbf{V}(t)\mathbf{C}(t)^{\top}\mathbf{R}^{-1}(t),$
and the covariance $\mathbf{V}(t)$ satisfies the differential Riccati equation:
$\frac{d}{dt}\mathbf{V}(t)=\mathbf{A}(t)\mathbf{V}(t)+\mathbf{V}(t)\mathbf{A}(t)^{\top}-\mathbf{V}(t)\mathbf{C}(t)^{\top}\mathbf{R}^{-1}(t)\mathbf{C}(t)\mathbf{V}(t)+\mathbf{Q}(t).$
A.4 Adaptive Filtering
Adaptive filtering is the task of estimating or tracking a signal in real time when the system dynamics or noise statistics are unknown or time-varying. Unlike the Kalman Filter, which assumes known dynamics and noise covariances, adaptive filters adjust their parameters dynamically to minimize the mean squared error (MSE) between the filter output and a desired signal. They are called “adaptive” because the filter adapts to changes in the system or environment, usually in an online fashion.
Iteratively Reweighted Least Squares (IRLS) and Robust Kalman Filtering
Iteratively Reweighted Least Squares (IRLS) is a general approach for solving weighted least squares problems, particularly when robustness to outliers is desired. Given a set of residuals $\mathbf{r}_{i}$, IRLS solves:
$\mathbf{\hat{x}}=\arg\min_{\mathbf{x}}\sum_{i}w_{i}\,\|\mathbf{r}_{i}(\mathbf{x})\|_{2}^{2},$
where the weights $w_{i}$ are updated iteratively based on the current residuals, often using an influence function $\psi(\mathbf{r}_{i})$ such that: $w_{i}=\psi(\mathbf{r}_{i})/\mathbf{r}_{i}$.
A.5 The Attention Mechanism
Self-attention (Vaswani et al., 2023) models pairwise interactions between elements in a sequence of length $m$ by computing learned similarity scores between $d$-dimensional embeddings of each token. Each token is associated with a query, key, and value, stacked as $\mathbf{X}_{Q},\mathbf{X}_{K},\mathbf{X}_{V}\in\mathbb{R}^{m\times d}$ and linearly projected:
$\mathbf{Q}=\mathbf{X}_{Q}\mathbf{W}_{Q},\quad\mathbf{K}=\mathbf{X}_{K}\mathbf{W}_{K},\quad\mathbf{V}=\mathbf{X}_{V}\mathbf{W}_{V}.$
And the output is a weighted sum:
$\mathbf{Y}=\text{softmax}(\mathbf{Q}\mathbf{K}^{\top})\mathbf{V},\quad$
A row of $\mathbf{Y}$ can be written as:
$\mathbf{y}_{i}=\sum_{j=1}^{m}p_{ij}\mathbf{v}_{j},\quad p_{ij}=\text{softmax}(\mathbf{q}_{i}^{\top}\mathbf{k}_{j}),$
Attention can be viewed as a special-case of a precision-weighted average:
$\mathbf{y}_{i}=\Big{(}\sum_{j}\mathbf{\tilde{P}}_{ij}\Big{)}^{-1}\sum_{j}\mathbf{\tilde{P}}_{ij}\mathbf{x}_{j},$
where each $\mathbf{\tilde{P}}_{ij}$ is a positive definite matrix. This tensor formulation allows the model not only to scale each input differently, but also to rotate and stretch it in feature space.
For example, if we set:
$\mathbf{\tilde{P}}_{ij}=\frac{\mathbf{P}_{ij}}{1+\mathbf{r}_{ij}^{\top}\mathbf{P}_{ij}\mathbf{r}_{ij}},$
with $\mathbf{P}_{ij}$ diagonal and $\mathbf{r}_{ij}=\mathbf{q}_{i}-\mathbf{k}_{j}$, the normalization reduces to a simple row-wise form, leading to an element-wise version of attention.
The standard dot-product attention used in Transformers corresponds to:
$\mathbf{\tilde{P}}_{ij}=\text{softmax}(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})\,\mathbf{I}.$
A.6 State Space Models
In discrete time, a linear time-invariant (LTI) system can be written as:
$\mathbf{x}_{k}=\mathbf{A}\mathbf{x}_{k-1}+\mathbf{B}\mathbf{u}_{k},\quad\mathbf{y}_{k}=\mathbf{C}\mathbf{x}_{k}+\mathbf{D}\mathbf{u}_{k},$
where $\mathbf{x}_{k}\in\mathbb{R}^{d}$ is the hidden state, $\mathbf{u}_{k}\in\mathbb{R}^{p}$ the input, and $\mathbf{y}_{k}\in\mathbb{R}^{q}$ the output.
The LTI structure allows explicit unrolling of the dynamics:
$\mathbf{y}_{k}=\mathbf{\bar{C}}\mathbf{\bar{A}}^{k}\mathbf{x}_{0}+\sum_{j=1}^{k}\mathbf{\bar{C}}\mathbf{\bar{A}}^{k-j}\mathbf{\bar{B}}\mathbf{u}_{j}=\sum_{j=0}^{k}\mathbf{\bar{h}}_{j}\mathbf{u}_{k-j},$
where $\mathbf{\bar{h}}_{j}=\mathbf{\bar{C}}\mathbf{\bar{A}}^{j}\mathbf{\bar{B}}$ is the impulse response. This shows that the system output can be viewed as a discrete convolution: $\mathbf{y}=\mathbf{\bar{h}}*\mathbf{u}$, which can be computed efficiently in the frequency domain via the Fourier or $Z$-transform.
Neural State Space Models
Recurrent models such as SSMs are prone to vanishing or exploding gradients. The HiPPO framework addresses this by constructing the state matrix $\mathbf{A}$ based on orthogonal polynomial projections, such as Legendre polynomials. The resulting HiPPO matrices optimally compress a continuous-time signal into a fixed-dimensional state while preserving information about the past. This ensures that the hidden state maintains a summary of the input sequence over long horizons, alleviating vanishing gradients (Gu et al., 2020).
Directly computing $\mathbf{y}_{k}$ involves repeated multiplications by $\mathbf{\bar{A}}$. The S4 model (Gu et al., 2022) exploits the $Z$-transform to convert the recurrence into a convolution in the frequency domain:
$\mathbf{Y}(z)=\sum_{k=0}^{\infty}\mathbf{y}_{k}z^{-k}=\sum_{i=0}^{\infty}\mathbf{h}_{i}z^{-i}\sum_{k=i}^{\infty}\mathbf{u}_{k-i}z^{-(k-i)}=\mathcal{\hat{K}}(z)\mathbf{U}(z),$
where the transfer function is:
$\mathcal{\hat{K}}(z)=\sum_{i=0}^{\infty}\mathbf{\bar{C}}\mathbf{\bar{A}}^{i}\mathbf{\bar{B}}z^{-i}.$
S4 further exploits the structure of $\mathbf{A}$ for computational efficiency. If $\mathbf{A}$ is diagonal, the transfer function simplifies:
$\mathcal{\hat{K}}_{\mathbf{\Lambda}}(z)=c(z)\sum_{i}\frac{\mathbf{C}_{i}\mathbf{B}_{i}}{g(z)-\mathbf{\Lambda}_{i}}=c(z)\cdot k_{z,\mathbf{\Lambda}}(\mathbf{C},\mathbf{B}),$
allowing parallel computation across state dimensions.
Diagonal Plus Low-Rank (DPLR) and Normal Plus Low-Rank (NPLR) Structures
To capture richer dynamics while retaining efficiency, S4 introduces a Diagonal Plus Low-Rank (DPLR) parameterization of the state matrix:
$\mathbf{A}=\bm{\Lambda}-\mathbf{P}\mathbf{Q}^{\dagger}.$
This generalizes to Normal Plus Low-Rank (NPLR) matrices:
$\mathbf{A}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\dagger}-\mathbf{P}\mathbf{Q}^{\dagger}=\mathbf{U}(\bm{\Lambda}-\mathbf{\tilde{P}}\mathbf{\tilde{Q}}^{\dagger})\mathbf{U}^{\dagger},$
showing that any NPLR matrix is unitarily similar to a DPLR matrix. This structure allows the HiPPO matrix to be embedded within S4 efficiently.
The matrix inversion lemma (Sherman–Morrison–Woodbury) enables efficient computation of $(\mathbf{\Lambda}-\mathbf{P}\mathbf{Q}^{\dagger})^{-1}$.
The Mamba models extend S4 by combining NPLR structure with careful numerical techniques to further improve stability and long-range memory (Gu and Dao, 2024). These models achieve performance competitive with Transformers on a variety of sequence modeling tasks.
Appendix B Additional Material
B.1 Separable Approximation of the Full Precision Kernel
Recall that:
$\mathbf{D}[i,j]=\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\|^{2}+\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\Big{[}\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\|^{2}$
$-2\mathrm{Re}\left(\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big{]}$
If we could factor the precision tensor into separate $i$ and $j$ terms $\mathcal{P}^{C}[:,i,j]=c+\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]$, then we could write:
$\mathbf{D}[i,j]=c\sum_{k}\big{\|}\bm{\Phi}_{\mathbf{qk}}[k,i]\big{\|}^{2}\Big{[}\big{\|}\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big{\|}^{2}+\big{\|}\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big{\|}^{2}-2\mathrm{Re}\left(\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\cdot\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big{]}$
$+\sum_{k}\big{\|}\bm{\Phi}_{\mathbf{qk}}[k,i]\big{\|}^{2}\Big{[}\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]\big{\|}\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big{\|}^{2}+\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]\big{\|}\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big{\|}^{2}$
$-2\mathrm{Re}\left(\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\cdot\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]\cdot\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big{]}$
This can now be simplified into a sum of matrix products, which can be cheaply computed without storing a $d\times m\times m$ dimensional tensor:
$\mathbf{D}=c\Big{[}\big{(}\big{|}\bm{\Phi}_{\mathbf{qk}}\big{|}^{2}\big{)}^{\top}\big{|}\mathbf{\hat{Z}}_{\mathbf{k}}\big{|}^{2}+\sum_{k}\big{(}\big{|}\bm{\Phi}_{\mathbf{qk}}\big{|}^{2}\odot\big{|}\mathbf{\hat{Z}}_{\mathbf{q}}\big{|}^{2}\big{)}-2\mathrm{Re}\Big{(}\big{(}\big{|}\bm{\Phi}_{\mathbf{qk}}\big{|}^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big{)}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\Big{)}\Big{]}$
$+\big{(}\big{|}\bm{\Phi}_{\mathbf{qk}}\big{|}^{2}\odot\mathbf{f}\big{)}^{\top}\big{(}\big{|}\mathbf{\hat{Z}}_{\mathbf{k}}\big{|}^{2}\odot\mathbf{g}\big{)}+\big{(}\big{|}\bm{\Phi}_{\mathbf{qk}}\big{|}^{2}\odot\big{|}\mathbf{\hat{Z}}_{\mathbf{q}}\big{|}^{2}\odot\mathbf{f}\big{)}^{\top}\mathbf{g}-2\mathrm{Re}\Big{(}\big{(}\big{|}\bm{\Phi}_{\mathbf{qk}}\big{|}^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\odot\mathbf{f}\big{)}^{\dagger}\big{(}\mathbf{\hat{Z}}_{\mathbf{k}}\odot\mathbf{g}\big{)}\Big{)}$
Recall that the precision is:
$\mathcal{P}^{C}[:,i,j]=\left(\bm{\lambda}_{C}^{2}\odot\bm{\lambda}_{\Omega}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda})(t_{i}-t_{j})}}{-2\mathrm{Re}(\bm{\lambda})}+\bm{\lambda}_{\Gamma}\odot e^{2\mathrm{Re}(\bm{\lambda})(t_{i}-t_{j})}\right)^{-1}$
Let $\bm{\lambda}_{eff}=\frac{\bm{\lambda}_{C}^{2}\odot\bm{\lambda}_{\Omega}}{-2\mathrm{Re}(\bm{\lambda})}>0$. Then:
$\mathcal{P}^{C}[:,i,j]=\left(\bm{\lambda}_{eff}+(-\bm{\lambda}_{eff}+\bm{\lambda}_{\Gamma})e^{2\mathrm{Re}(\bm{\lambda})(t_{i}-t_{j})}\right)^{-1}$
$\frac{1}{\bm{\lambda}_{eff}}\left(1+\bm{\lambda}_{r}e^{2\mathrm{Re}(\bm{\lambda})\Delta t}\right)^{-1},\quad\text{where: }\,\bm{\lambda}_{r}=-1+\frac{\bm{\lambda}_{\Gamma}}{\bm{\lambda}_{eff}}>-1$
We use the separable approximation:
$(1+ae^{bx})^{-1}\approx 1-ce^{-dx},$
with constants chosen to match the value and slope at $x=0$.
We obtain the following separable approximation:
$\mathcal{P}^{C}(\Delta t)\approx\frac{1}{\bm{\lambda}_{eff}}\left(1-\frac{\bm{\lambda}_{r}}{1+\bm{\lambda}_{r}}e^{\frac{2Re(\bm{\lambda})\Delta t}{1+\bm{\lambda}_{r}}}\right)=\frac{1}{\bm{\lambda}_{eff}}\left[1-\left(1-\frac{\bm{\lambda}_{eff}}{\bm{\lambda}_{\Gamma}}\right)e^{-\frac{\bm{\lambda}_{C}^{2}\bm{\lambda}_{\Omega}}{\bm{\lambda}_{\Gamma}}\Delta t}\right]$
B.2 Approximating the Batch MLE with a Recursive Filter
Define a re-weighted Kalman gain:
$\tilde{K}_{i}:=\frac{e^{\tau_{s}w_{i}}}{\tilde{p}_{i}^{+}\,\sigma_{V}^{2}}$
The precision and mean updates are:
$\tilde{p}_{i}^{+}=p_{i}^{-}+\tilde{K}_{i},\quad\mathbf{\bar{z}}_{i}^{+}=\mathbf{\bar{z}}_{i}^{-}+\tilde{K}_{i}(\mathbf{z}_{i}-\mathbf{\bar{z}}_{i}^{-})=\tilde{K}_{i}\mathbf{z}_{i}+(1-\tilde{K}_{i})\mathbf{\bar{z}}_{i}^{-}$
with the prior propagated through the dynamics:
$\mathbf{\bar{z}}_{i}^{-}=\bm{\Phi}\,\mathbf{\bar{z}}_{i-1}^{+}$
Recursive substitution gives a weighted sum of all past observations:
$\mathbf{\bar{z}}_{i}^{+}=\sum_{j=0}^{i}\left[\tilde{K}_{j}\prod_{k=j+1}^{i}(1-\tilde{K}_{k})\,\bm{\Phi}\right]\mathbf{z}_{j}$
If $\tilde{K}_{k}\ll 1$ for all $k$ and $\Delta t$ is small, this is approximately:
$\mathbf{\bar{z}}_{i}^{+}\approx\sum_{j=0}^{i}\tilde{K}_{j}\,e^{\Lambda\Delta t_{ij}}\mathbf{z}_{j}=\sum_{j=0}^{i}\tilde{K}_{j}\mathbf{\hat{z}}_{ij}$
Hence, if:
$\tilde{K}_{j}\approx\frac{w_{ij}/\sigma_{V}^{2}}{\sum_{j}w_{ij}/\sigma_{V}^{2}},$
then we recover an approximation of the batch MLE.
B.3 Normal plus Rank-1 Noise Covariance
In this section, we explore a generalization of the RT-SDE model by considering a normal-plus-rank-1 structure for $\mathbf{Q}$ and $\mathbf{R}$. Our goal is to determine whether relaxing the strict radial-tangential decomposition introduces additional expressivity in the propagated covariance or the resulting MLE.
Recall that:
$\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j}):=\int_{0}^{|\Delta t_{ij}|}e^{-\bm{\Lambda}\tau}(\mathbf{S}^{-1}\mathbf{Q}(t_{j}+\tau)\mathbf{S}^{-\dagger})e^{-\bm{\Lambda}^{\dagger}\tau}d\tau$
Let us now assume that $\mathbf{Q}$ is normal plus rank-1, i.e. $\mathbf{Q}=\mathbf{S}\left(\bm{\Lambda}_{Q}+\mathbf{q}\mathbf{q}^{\dagger}\right)\mathbf{S}^{\dagger}$. We already derived the diagonal part. The rank-1 part is:
$\int_{0}^{|\Delta t_{ij}|}e^{-\bm{\Lambda}\tau}\big{(}\mathbf{q}\mathbf{q}^{\dagger}\big{)}e^{-\bm{\Lambda}^{\dagger}\tau}d\tau$
$\int_{0}^{|\Delta t_{ij}|}\big{(}e^{-\bm{\Lambda}\tau}\mathbf{q}\big{)}\big{(}e^{-\bm{\Lambda}\tau}\mathbf{q}\big{)}^{\dagger}d\tau$
If we take: $\mathbf{q}(t_{j}+\tau)=e^{\bm{\Lambda}(\Delta t_{ij}+\tau)}\mathbf{p}$, so that $\mathbf{q}$ moves with the LTI’s flow, this becomes:
$e^{\bm{\Lambda}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}^{\dagger}\Delta t_{ij}}\int_{0}^{|\Delta t_{ij}|}d\tau=e^{\bm{\Lambda}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}^{\dagger}\Delta t_{ij}}|\Delta t_{ij}|$
If we assume that $\bm{\Lambda}=\alpha\mathbf{I}+\bm{\Omega}$, where $\bm{\Omega}^{\dagger}=-\bm{\Omega}$, then we have:
$\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}e^{-\bm{\Lambda}_{\Omega}\tau}\big{(}\mathbf{q}\mathbf{q}^{\dagger}\big{)}e^{-\bm{\Lambda}_{\Omega}^{\dagger}\tau}d\tau$
Letting $\mathbf{q}(t_{j}+\tau)=ae^{\bm{\Lambda}_{\Omega}(\Delta t_{ij}+\tau)}\mathbf{p}$ for some constant $a$, we have:
$ae^{\bm{\Lambda}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}^{\dagger}\Delta t_{ij}}\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}d\tau=a\frac{e^{-2\alpha\Delta t_{ij}}-1}{-2\alpha}e^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}$
Altogether, we have:
$\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})=\sigma^{2}\kappa_{B}(\Delta t_{ij})\Big{(}\mathbf{I}+ae^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}\Big{)}$
Assuming measurement noise with a rank-1 component aligned with the rank-1 component of the process noise, the propagated measurement covariance is:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=e^{-\bm{\Lambda}\Delta t_{ij}}\Big{[}\lambda_{C}^{2}\sigma^{2}\kappa_{B}(|\Delta t_{ij}|)\Big{(}\mathbf{I}+ae^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}\Big{)}+\eta^{2}\Big{(}\mathbf{I}+be^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}\Big{)}\Big{]}e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\lambda_{C}^{2}\sigma^{2}\kappa_{F}(|\Delta t_{ij}|)\Big{(}\mathbf{I}+a\mathbf{p}\mathbf{p}^{\dagger}\Big{)}+\eta^{2}e^{-2\alpha\Delta t_{ij}}\Big{(}\mathbf{I}+b\mathbf{p}\mathbf{p}^{\dagger}\Big{)}$
$\Big{(}\lambda_{C}^{2}\sigma^{2}\kappa_{F}(|\Delta t_{ij}|)+\eta^{2}e^{-2\alpha\Delta t_{ij}}\Big{)}\mathbf{I}+\Big{(}a\lambda_{C}^{2}\sigma^{2}\kappa_{F}(|\Delta t_{ij}|)+b\eta^{2}e^{-2\alpha\Delta t_{ij}}\Big{)}\mathbf{p}\mathbf{p}^{\dagger}$
In particular, letting $\mathbf{p}=\mathbf{u}_{\mathbf{s}i}$, and setting $\sigma=\sigma_{t}$, $\eta=\eta_{t}$, $a=\frac{\sigma_{r}^{2}-\sigma_{t}^{2}}{\sigma_{t}^{2}}$, $b=\frac{\eta_{r}^{2}-\eta_{t}^{2}}{\eta_{t}^{2}}$, this becomes:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\Big{(}\lambda_{C}^{2}\sigma_{t}^{2}\kappa_{F}(|\Delta t_{ij}|)+\eta_{t}^{2}e^{-2\alpha\Delta t_{ij}}\Big{)}\mathbf{I}+\Big{(}\lambda_{C}^{2}(\sigma_{r}^{2}-\eta_{t}^{2})\kappa_{F}(|\Delta t_{ij}|)+(\eta_{r}^{2}-\eta_{t}^{2})e^{-2\alpha\Delta t_{ij}}\Big{)}\mathbf{u}_{\mathbf{s}i}\mathbf{u}_{\mathbf{s}i}^{\dagger}$
$\sigma_{Vt}^{2}(|\Delta t_{ij}|)(\mathbf{I}-\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger})+\sigma_{Vr}^{2}(|\Delta t_{ij}|)\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$
Hence, the normal plus rank-1 process noise covariance model reduces exactly to the RT-SDE case: while $\mathbf{p}$ is not restricted to be a unit vector, any scaling of $\mathbf{p}$ can be absorbed into the radial noise parameters, so the model introduces no additional degrees of freedom beyond RT-SDE. The requirement that $\mathbf{p}$ flow with the LTI dynamics (which is necessary for the integral to be tractable), with $\mathbf{A}=\alpha\mathbf{I}+\bm{\Omega}$, constrains the rank-1 component to rotate with the system, forcing the normal-plus-rank-1 noise model to align exactly with the RT-SDE structure.
B.4 Expressivity of Radial-Tangential SDEs
Proposition 1: Expressivity of Deterministic RT-SDEs.
Let $\mathbf{A}\in\mathbb{R}^{d\times d}$ be a diagonalizable matrix, with eigendecomposition $\mathbf{A}=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1}$, where $\mathbf{S}\in\mathbb{R}^{d\times d}$, $\bm{\Lambda}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{d})\in\mathbb{C}^{d\times d}$. For any $\epsilon>0$, there exists a sequence of matrices $\mathbf{A}_{\ell}=\mathbf{S}(\alpha_{\ell}\mathbf{I}+\bm{\Lambda}_{\Omega\ell})\mathbf{S}^{-1}$, where $\alpha_{\ell}\in\mathbb{R}$, $\bm{\Lambda}_{\Omega\ell}=\mathrm{diag}(i\omega_{\ell,1},\dots,i\omega_{\ell,d})\in i\mathbb{R}^{d\times d}$, and a choice of block durations $\Delta t>0$, such that
$\left\|e^{\mathbf{A}t_{k}}-\prod_{\ell=1}^{L}e^{\mathbf{A}_{\ell}\Delta t}\right\|<\epsilon,\quad\forall k=1,\dots,m.$
Proof Sketch.
Each $\mathbf{A}_{\ell}$ generates dynamics:
$e^{\mathbf{A}_{\ell}\Delta t}=\mathbf{S}\,\mathrm{diag}\big{(}e^{(\alpha_{\ell}+i\omega_{\ell,1})\Delta t},\dots,e^{(\alpha_{\ell}+i\omega_{\ell,d})\Delta t}\big{)}\mathbf{S}^{-1}.$
Each diagonal entry corresponds to a 1D complex exponential in the eigenbasis. By choosing sufficiently many blocks $L$ and appropriately selecting $\alpha_{\ell}$ and $\omega_{\ell,k}$, one can approximate:
$e^{\lambda_{k}t_{j}}\approx\prod_{\ell=1}^{L}e^{(\alpha_{\ell}+i\omega_{\ell,k})\Delta t},\quad\forall k,j,$
to arbitrary precision. Mapping back to the original basis with $\mathbf{S}$ yields the desired approximation. Hence, for any finite set of times, a product of radial-tangential blocks can approximate the target flow arbitrarily closely.
Remark: If we have $n$ heads, each producing its own low-dimensional output, stacking them allows combinations of rotations and scalings from different subspaces to mix. The resulting flow in the original $d$-dimensional space can approximate much richer dynamics than a single high-dimensional LTI with the same number of blocks. The effective number of distinct transformations is roughly the number of heads times the number of layers.
Proposition 2: Expressivity of RT-SDE Covariance.
Let $\bm{\Sigma}^{\star}\in\mathbb{R}^{d\times d}$ be a symmetric positive definite (SPD) matrix and fix a time $t_{i}$. For any $\varepsilon>0$, there exists a number of heads $H$ and associated RT-SDE parameters:
$\{\mathbf{u}_{h}(t_{i}),\sigma_{Vr,h}^{2},\sigma_{Vt,h}^{2}\}_{h=1}^{H}$
such that the RT-SDE covariance satisfies:
$\big{\|}\bm{\Sigma}(t_{i})-\bm{\Sigma}^{\star}\big{\|}<\varepsilon,$
where
$\bm{\Sigma}(t_{i})=\sum_{h=1}^{H}\sigma_{Vt,h}^{2}\mathbf{I}+\sum_{h=1}^{H}\beta_{h}\,\mathbf{u}_{h}(t_{i})\mathbf{u}_{h}(t_{i})^{\top},\quad\beta_{h}=\sigma_{Vr,h}^{2}-\sigma_{Vt,h}^{2}.$
Proof Sketch:
Each head $h$ contributes an SPD term of the form:
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\sigma_{Vt}^{2}(|\Delta t_{ij}|)\mathbf{I}+\big{(}\sigma_{Vr}^{2}(|\Delta t_{ij}|)-\sigma_{Vt}^{2}(|\Delta t_{ij}|)\big{)}\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$
i.e. a rank-1 correction of identity aligned with the current state direction $\mathbf{u}(t_{i})$.
The total propagated covariance is the sum of the individual head covariances, assuming the uncertainty sources for each head are uncorrelated:
$\bm{\Sigma}(t_{i})\;=\;\Big{(}\sum_{h}\sigma_{Vt}^{2}(|\Delta t_{ij}|)\Big{)}\mathbf{I}+\sum_{h}\beta_{h}\,\mathbf{u}_{h}(t_{i})\mathbf{u}_{h}(t_{i})^{\top},\quad\beta_{h}=\sigma_{Vr}^{2}-\sigma_{Vt}^{2}.$
At a fixed time $t_{i}$, we can construct any target SPD matrix $\bm{\Sigma}^{\star}$ using the RT-SDE heads as follows. Let $\sigma_{\min}=\lambda_{\min}(\bm{\Sigma}^{\star}),\quad\sum_{h}\sigma_{Vt}^{2}=\sigma_{\min}$, so that the isotropic contributions sum to the minimal eigenvalue. Consider the spectral decomposition of the shifted matrix:
$\bm{\Sigma}^{\star}-\sigma_{\min}\mathbf{I}=\sum_{h}(\lambda_{h}-\sigma_{\min})\,\mathbf{v}_{h}\mathbf{v}_{h}^{\top},$
where $\{\mathbf{v}_{h}\}$ are the eigenvectors of $\bm{\Sigma}^{\star}$. Setting $\mathbf{u}_{h}(t_{i})=\mathbf{v}_{h},\,\beta_{h}=\lambda_{h}-\sigma_{\min}$, gives:
$\sum_{h}\sigma_{Vt}^{2}\mathbf{I}+\sum_{h}\beta_{h}\,\mathbf{u}_{h}(t_{i})\mathbf{u}_{h}(t_{i})^{\top}=\sigma_{\min}\mathbf{I}+(\bm{\Sigma}^{\star}-\sigma_{\min}\mathbf{I})=\bm{\Sigma}^{\star}.$
Remark: Note that this only holds at a fixed time, since $\mathbf{u}_{h}(t)$ evolves according to the underlying LTI dynamics, so we cannot freely choose the propagated covariance at all times independently. Nevertheless, this shows that the RT-SDE model is substantially more flexible than a diagonalizable LTI-SDE with isotropic noise, for which the propagated measurement covariance remains isotropic in the eigenbasis.