### v2

# Attention as an Adaptive Filter

Abstract
We introduce Adaptive Filter Attention (AFA), a novel attention mechanism that incorporates a learnable dynamics model directly into the computation of attention weights. Rather than comparing queries and keys directly, we model the input sequence as discrete observations of a linear stochastic differential equation (SDE). By imposing a linear dynamics model with simultaneously diagonalizable state matrices and noise covariances, we can make use of a closed-form solution of the differential Lyapunov equation to efficiently propagate pairwise uncertainties through the dynamics. Attention naturally arises as the maximum likelihood solution for this linear SDE, with attention weights corresponding to robust residual-based reweightings of the propagated pairwise precisions. To obtain a simplified variant with the same computational and memory complexity as standard attention, we further constrain the state matrix’s eigenvalues by tying their real parts to a shared scalar, i.e. enforcing isotropic decay. In the limit of vanishing dynamics and process noise, and using a small-angle approximation, we recover a complex-valued form of ordinary dot-product attention with rotary positional encodings.
1 Introduction
Self-attention has become the dominant paradigm for sequence modeling due to its parallelism and scalability. However, this parallelism comes at the cost of losing the recursive structure that RNNs naturally impose, which provides an implicit temporal regularization that encourages consistency over time. State space models (SSMs) offer a complementary approach: they retain temporal structure through shared linear dynamics while supporting parallel training via convolutional implementations. However, SSMs generally assume deterministic dynamics, limiting their ability to capture stochasticity and uncertainty.
Our goal is to bridge these two paradigms. We view self-attention as a form of filtering: it selectively averages across inputs, denoising by pooling information from similar states. However, unlike Kalman Filtering, standard attention does not explicitly propagate states through time. This motivates incorporating explicit temporal structure—through known or learned dynamics—into attention, restoring the inductive bias that standard attention loses by forgoing recurrence. Unlike SSMs, which are typically formulated as deterministic ODEs with inputs entering explicitly through control terms, we instead treat the entire dataset as arising from a stochastic differential equation (SDE), i.e. an autonomous system driven by noise.
This technical report is structured as follows:
Section 2: Preliminaries
We introduce the model setup and review standard results for propagating covariances through linear SDEs. We then discuss the batch maximum likelihood estimator and the closed-form solution of the differential Lyapunov equation for diagonalizable systems with simultaneously diagonalizable process and measurement noise covariances. Finally, we discuss an adaptively reweighted MLE that reweights propagated covariances using the pair-wise residuals.
Section 3: Adaptive Filter Attention
We generalize the robustified MLE to a tensor form of attention, and briefly analyze its time and memory complexity. We show that the algorithm can be simplified to a form with the same time and memory complexity as standard attention, through the use of convolutional kernels for the (diagonalized) matrix exponential and pair-wise propagated covariances, and by imposing additional structure on the linear SDE. We show that, in the case of a skew-symmetric state matrix and normalized inputs, this reduces to a complex-valued variant of standard attention with a complex-valued positional encoding equivalent to that in xPos and RetNet. We then describe an implementation of this algorithm, and an unrolled, recursive approximation for inference. We describe basic experiments on simple simulated linear dynamical systems.
Section 5: Radial-Tangential Model
We generalize the SDE model from Section 3, which assumed isotropic noise, to allow for separate noise covariances in the tangential and radial directions (which we call a radial-tangential SDE). We show that the propagated covariances in this case can be written in closed-form since the assumed diagonal plus rank-1 covariances co-rotate with the dynamics. This formulation naturally yields a maximum likelihood estimator (MLE) on the hypersphere, and a corresponding attention mechanism. From this perspective, the Transformer can be interpreted geometrically, as an approximate solution of the MLE under the constraints imposed by the radial-tangential SDE model.
Section 6: Related Work
We overview related work, including classical methods for filtering and adaptive filtering; the use of neural networks for modeling and filtering dynamical systems; neural state space models; connections between attention and structured state space models; related architectures; and various other topics.
Section 7: Discussion and Conclusion
Appendix A: Background Additional details for the material in the preliminaries section.
Appendix B: Additional Material
Extensions of AFA, a recurrent formulation, generative modeling, an extension to control, an extension to higher dimensional data, discussion of normal plus rank-1 noise covariance, and discussion of the expressivity of our structured SDE model.
Note: This technical report describes ongoing work. Evaluation on real-world datasets is in progress and will be reported in future work.
2 Preliminaries
This section introduces the setup of our filtering problem and key ideas from linear Gaussian filtering on which we base the attention mechanism developed in Section 3. Detailed derivations can be found in Appendix A.
2.1 Model Setup
Consider a linear time-invariant (LTI) Itô stochastic differential equation (SDE):
$d\mathbf{x}(t)=\mathbf{A}\mathbf{x}(t)dt+\mathbf{G}d\mathbf{w}(t),\quad\mathbf{z}(t_{k})=\mathbf{C}\mathbf{x}(t_{k})+\mathbf{v}(t_{k}),$
with the following definitions:
  • $\mathbf{x}(t)\in\mathbb{R}^{d}$: latent state at time $t$,
  • $\mathbf{z}(t_{k})\in\mathbb{R}^{d}$: $k$th measurement,
  • $\mathbf{A},\mathbf{G},\mathbf{C}\in\mathbb{R}^{d\times d}$: state, input, and output matrices,
  • $d\mathbf{w}(t)\sim\mathcal{N}(\mathbf{0},\mathbf{I}\,dt)$: standard Wiener process (Brownian motion),
  • $\mathbf{Q}:=\mathbf{G}\mathbf{G}^{\top}$: continuous-time process noise covariance,
  • $\mathbf{v}(t_{k})\sim\mathcal{N}(\mathbf{0},\mathbf{R})$: zero-mean Gaussian measurement noise with covariance $\mathbf{R}$.
Given $\mathbf{x}(t_{j})$, we can estimate $\mathbf{x}(t_{i})$ using the matrix exponential:
$\mathbf{\hat{x}}_{i}(t_{j}):=e^{-\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{j}).$
Likewise, if we measure $\mathbf{z}(t_{j})$, and assuming the measurement matrix $\mathbf{C}$ is square and invertible, we can estimate the latent state $\mathbf{x}(t_{i})$, or equivalently $\mathbf{z}^{C}(t_{i}):=\mathbf{C}\mathbf{x}(t_{i})$, by computing a "pulled-forward" measurement:
$\mathbf{\hat{z}}_{i}(t_{j}):=\mathbf{\Phi}_{\mathbf{z}}(\Delta t_{ij})\mathbf{z}(t_{j}),\quad\text{where: }\,\mathbf{\Phi}_{\mathbf{z}}(\Delta t):=\mathbf{C}e^{\mathbf{A}\Delta t}\mathbf{C}^{-1}$
In the special case that $\mathbf{A}$ and $\mathbf{C}$ commute, $\mathbf{\hat{z}}_{i}(t_{j})=\mathbf{\hat{x}}_{i}(t_{j})$.
The "pulled-forward" estimate is distributed around the true state as:
$\mathbf{\hat{x}}_{i}(t_{j})\sim\mathcal{N}\left(\mathbf{x}(t_{i}),\mathbf{V}(|\Delta t_{ij}|)\right)$
where the forward-propagated covariance is:
$\mathbf{V}(|\Delta t_{ij}|)=\int_{0}^{|\Delta t_{ij}|}e^{\mathbf{A}s}\mathbf{Q}e^{\mathbf{A}^{\top}s}\,ds.$
which is a solution of the differential Lyapunov equation:
$\frac{d}{ds}\mathbf{V}(s)=\mathbf{A}\mathbf{V}(s)+\mathbf{V}(s)\mathbf{A}^{\top}+\mathbf{Q},$
The pulled-forward measurement estimate is distributed as:
$\mathbf{\hat{z}}_{i}(t_{j})\sim\mathcal{N}\left(\mathbf{z}^{C}(t_{i}),\mathbf{V}^{C}(\Delta t_{ij})\right)$
where the forward-propagated measurement covariance is:
$\mathbf{V}^{C}(\Delta t_{ij})=\mathbf{C}\mathbf{V}(\Delta t_{ij})\mathbf{C}^{\top}+\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})\mathbf{R}\bm{\Phi}_{\mathbf{z}}^{\top}(\Delta t_{ij})$
We can also define corresponding precision matrices $\mathbf{P}(\Delta t_{ij})=\big(\mathbf{V}(\Delta t_{ij})\big)^{-1}$ and $\mathbf{P}^{C}(\Delta t_{ij})=\big(\mathbf{V}^{C}(\Delta t_{ij})\big)^{-1}$.
2.2 Batch Maximum Likelihood Estimation
We wish to estimate the trajectory of the system from $m$ measurements taken at times $\{t_{1},...,t_{m}\}$, which we denote $\{\mathbf{z}(t_{1}),...,\mathbf{z}(t_{m})\}$. We can do so by solving a separate maximum likelihood estimation (MLE) problem for each $\mathbf{x}(t_{i})$.
To simplify the notation, let $\mathbf{x}_{i}=\mathbf{x}(t_{i})$, $\mathbf{\hat{x}}_{ij}=\mathbf{\hat{x}}_{i}(t_{j})$ and $\mathbf{P}_{ij}=\mathbf{P}(\Delta t_{ij})$, etc. The (causal) MLE solution for $\mathbf{z}^{C}(t_{i})$ is a precision-weighted estimator:
$\mathbf{\bar{z}}_{i}:=\left(\sum_{j\leq i}\mathbf{P}_{ij}^{C}\right)^{-1}\sum_{j\leq i}\mathbf{P}_{ij}^{C}\mathbf{\hat{z}}_{ij},$
Repeating this computation for all $t_{i}$ produces an estimate of the full trajectory.
Computing $\mathbf{P}^{C}(\Delta t_{ij})$ is generally expensive, as it involves integration and matrix inversion. However, if we assume that $\mathbf{A}=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1},\mathbf{Q}=\mathbf{S}\bm{\Lambda}_{Q}\mathbf{S}^{\dagger}\in\mathbb{R}^{d\times d}$, where $\bm{\Lambda},\bm{\Lambda}_{Q}\in\mathbb{C}^{d\times d}$ are diagonal matrices and $\mathbf{S}\in\mathbb{C}^{d\times d}$ an invertible matrix, the backward-propagated covariance becomes:
$\mathbf{V}_{F}(|\Delta t_{ij}|)=\mathbf{S}\bm{\Lambda}_{\mathbf{V}}(|\Delta t_{ij}|)\mathbf{S}^{\dagger}$
where:
$\bm{\Lambda}_{\mathbf{V}}(|\Delta t_{ij}|)=\bm{\Lambda}_{Q}\frac{\mathbf{I}-e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}}{-2\mathrm{Re}(\bm{\Lambda})}$
(where division by matrices denotes element-wise division).
If, in addition, $\mathbf{C}=\mathbf{S}\bm{\Lambda}_{C}\mathbf{S}^{-1},\mathbf{R}=\mathbf{S}\bm{\Lambda}_{R}\mathbf{S}^{\dagger}$, then:
$\mathbf{V}^{C}(\Delta t_{ij})=\mathbf{S}\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})\mathbf{S}^{\dagger}$
where:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\big|\bm{\Lambda}_{C}\big|^{2}\bm{\Lambda}_{Q}\frac{\mathbf{I}-e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}}{-2\mathrm{Re}(\bm{\Lambda})}+e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}\bm{\Lambda}_{R}$
Finally, $\mathbf{P}^{C}(\Delta t_{ij})=\mathbf{S}^{-\dagger}\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{ij})\mathbf{S}^{-1}$, where $\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{ij})=\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})^{-1}$.
Letting $\mathbf{z}_{\mathbf{s}_{i}}=\mathbf{S}^{-1}\mathbf{z}_{i}$, and $\mathbf{\hat{z}}_{\mathbf{s}ij}:=e^{-\bm{\Lambda}\Delta t_{ij}}\mathbf{z}_{\mathbf{s}j}$, the MLE is now tractable:
$\mathbf{\bar{z}}_{i}=\mathbf{S}\left[\left(\sum_{j\leq i}\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{ij})\right)^{-1}\sum_{j\leq i}\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{ij})\mathbf{\hat{z}}_{\mathbf{s}ij}\right]\mathbf{S}^{-1}.$
2.3 Robust Maximum Likelihood Trajectory Estimation
Maximum Likelihood Estimation (MLE) of a system’s trajectory relies on accurate precision matrices, which represent the uncertainty of our estimates. One approach to mitigate model mis-specification is to adaptively adjust the precision matrix $\mathbf{P}^{C}(\Delta t_{ij})$ using observed residuals. Since $\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\left(\mathbf{z}^{C}_{i},\mathbf{V}^{C}_{ij}\right)$, it follows that the pair-wise residual has distribution $\mathbf{r}_{ij}:=\mathbf{z}_{i}^{C}-\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big(0,\mathbf{V}^{C}_{ij}\big)$. We generally can’t observe $\mathbf{z}_{i}^{C}$, but we can also define a measurement residual, $\mathbf{r_{z}}_{ij}:=\mathbf{z}_{i}-\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big(0,\mathbf{V}^{C}_{ij}+\mathbf{R})$. We can then rescale the precision matrices with an adaptive weight:
$\mathbf{P}^{C}_{ij}\leftarrow\mathbf{P}^{C}_{ij}/w_{ij},\quad\text{where: }\,w_{ij}=\frac{1}{1+\mathbf{r_{z}}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r_{z}}_{ij}/\nu},$
where $\nu\in\mathbb{R}$ is a scaling factor. This leads naturally to an iteratively-reweighted least squares (IRLS) iteration in the eigenbasis. Let $\mathbf{z}_{\mathbf{s}i}:=\mathbf{S}^{-1}\mathbf{z}_{i}$ and $\mathbf{\hat{z}}_{\mathbf{s}ij}=\mathbf{S}^{-1}\mathbf{\hat{z}}_{ij}=e^{-\bm{\Lambda}\Delta t_{ij}}$, and define a corresponding measurement residual in the eigenbasis $\mathbf{r_{zs}}_{ij}:=\mathbf{z}_{\mathbf{s}i}-\mathbf{\hat{z}}_{\mathbf{s}ij}$. Let $\bm{\lambda}_{\mathbf{P}ij}^{C}=\text{diag}(\bm{\Lambda}_{\mathbf{P}ij}^{C})$. We apply the iterative update:
$\mathbf{\hat{z}}_{\mathbf{s}ii}^{(k+1)}=\mathbf{\hat{z}}_{\mathbf{s}i}^{\text{MLE}(k)},\quad\text{where}\quad\mathbf{\hat{z}}_{\mathbf{s}i}^{\text{MLE}(k)}=\left(\sum_{j}w_{ij}^{(k)}\bm{\lambda}_{\mathbf{P}ij}^{C}\right)^{-1}\odot\sum_{j}w_{ij}^{(k)}\bm{\lambda}_{\mathbf{P}ij}^{C}\odot\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)},$
$w_{ij}^{(k)}=\frac{1}{1+(\bm{\lambda}_{\mathbf{P}ij}^{C})^{\top}\big|\mathbf{r}^{(k)}_{\mathbf{s}ij}\big|^{2}/\nu},\quad\text{where}\quad\mathbf{r}_{\mathbf{s}ij}^{(k)}:=\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ii}-\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ij}$.
and initial condition $\mathbf{\hat{z}}_{\mathbf{s}ii}^{(1)}=\mathbf{z}_{\mathbf{s}i}$. (Alternatively, if we only have a single iteration, we can replace $\bm{\lambda}_{\mathbf{P}ij}^{C}$ with $\bm{\lambda}_{\mathbf{P}ij}^{C}+\bm{\lambda}_{R}$ (the covariance of $\mathbf{z}_{\mathbf{s}i}$).) To stabilize the iteration, we can instead use an affine combination of the current and previous estimate (i.e. a residual connection):
$\mathbf{\hat{z}}^{(k+1)}_{\mathbf{s}ii}=(1-\delta)\mathbf{\hat{z}}_{\mathbf{s}ii}^{(k)}+\delta\mathbf{\hat{z}}_{\mathbf{s}i}^{\text{MLE}(k)},$
This IRLS iteration, with precision matrices derived from an LTI model, can be implemented with a multi-layer neural network with learnable parameters $\mathbf{S},\bm{\Lambda},\bm{\Lambda}_{Q},\bm{\Lambda}_{R}$ shared across layers. If $\mathbf{\hat{z}}^{(k)}_{ii}\rightarrow\mathbf{z}^{C}$ in expectation with increasing $k$, we can take $\mathbf{\bar{z}}_{\mathbf{s}}\approx\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ii}$ at the last layer, and then predict the measurement at the next time step in the eigenbasis using the (diagonalized) matrix exponential:
$\mathbf{z}^{+}_{\mathbf{s}i}=\bm{\Lambda}_{C}e^{\bm{\Lambda}(t_{i+1}-t_{i})}\bm{\Lambda}_{C}^{-1}\mathbf{\bar{z}}_{\mathbf{s}i}=e^{\bm{\Lambda}(t_{i+1}-t_{i})}\mathbf{\bar{z}}_{\mathbf{s}i}$
Finally, we compute the output prediction: $\mathbf{z}^{+}_{i}=\mathbf{S}\mathbf{z}^{+}_{\mathbf{s}i}$, which we then compare to the actual measurement at the next time step $\mathbf{z}_{i+1}$. We can then define a loss, e.g. $\mathcal{L}=\sum_{i}\big|\mathbf{z}^{+}_{i}-\mathbf{z}_{i+1}\big|_{2}^{2}$ and train our model parameters $\mathbf{S},\bm{\Lambda},\bm{\Lambda}_{Q},\bm{\Lambda}_{R}$ using backpropagation.
3 Adaptive Filter Attention
3.1 Generalizing the Adaptive Filter to a Tensor Form of Attention
We now generalize the batch MLE form of the above adaptive filter using learned, complex-valued query, key, value, and output matrices: $\mathbf{W_{q}}\in\mathbb{C}^{d_{qk}\times d_{e}}$, $\mathbf{W_{k}}\in\mathbb{C}^{d_{qk}\times d_{e}}$, $\mathbf{W_{v}}\in\mathbb{C}^{d_{v}\times d_{e}}$, $\mathbf{W_{p}}\in\mathbb{C}^{d_{e}\times d_{v}}$, where $d_{e}$ is the input embedding dimension $d_{qk}$ is the query/key embedding dimension, and $d_{v}$ is the value embedding dimension.
In this formulation, $\mathbf{W_{q}}$, $\mathbf{W_{k}}$, $\mathbf{W_{v}}$ replace $\mathbf{S}^{-1}$ and $\mathbf{W_{p}}$ replaces $\mathbf{S}$. When $\mathbf{W_{q}}=\mathbf{W_{k}}=\mathbf{W_{v}}=\mathbf{W_{p}}^{-1}$, the attention layer corresponds to an adaptive MLE. By allowing these parameter matrices to vary freely, the network can learn approximate inverses as needed, enabling attention to implement an MLE while retaining the flexibility to represent richer transformations.
We can define separate key/query and value dynamics parameters: $\bm{\lambda}_{\mathbf{qk}},\bm{\lambda}_{Q\mathbf{qk}},\bm{\lambda}_{R\mathbf{qk}},\bm{\lambda}_{C\mathbf{qk}}\in\mathbb{R}^{d_{qk}}$; $\bm{\lambda}_{\mathbf{v}},\bm{\lambda}_{Q\mathbf{v}},\bm{\lambda}_{R\mathbf{v}},\bm{\lambda}_{C\mathbf{v}}\in\mathbb{R}^{d_{v}}$. If we set $d_{qk}=d_{v}$, then we can use a shared set of parameters for the key and value dynamics: $\bm{\lambda},\bm{\lambda}_{Q},\bm{\lambda}_{R},\bm{\lambda}_{C}\in\mathbb{R}^{d_{v}}$. Note that we can absorb $\big|\bm{\lambda}_{C}\big|^{2}$ into $\bm{\lambda}_{Q}$, but we keep it here for clarity.
The inputs to the network are: $\quad\mathbf{Z}=\left[\mathbf{z}_{1},\dots,\mathbf{z}_{m}\right]\in\mathbb{C}^{d_{e}\times m}$.
We compute the following:
Query, key, and value matrices:
$\mathbf{Z_{q}}=\mathbf{W_{q}}\mathbf{Z}\in\mathbb{C}^{d_{qk}\times m},\quad\mathbf{Z_{k}}=\mathbf{W_{k}}\mathbf{Z}\in\mathbb{C}^{d_{qk}\times m},\quad\mathbf{Z_{v}}=\mathbf{W_{v}}\mathbf{Z}\in\mathbb{C}^{d_{v}\times m}$
Exponential tensors: $\mathcal{E}_{\mathbf{qk}}\in\mathbb{C}^{d_{qk}\times m\times m},\quad\mathcal{E}_{\mathbf{v}}\in\mathbb{C}^{d_{v}\times m\times m}$:
$\mathcal{E}_{\mathbf{qk}}[k,i,j]=e^{\bm{\lambda}_{\mathbf{qk},k}(t_{i}-t_{j})},\quad\mathcal{E}_{\mathbf{v}}[k,i,j]=e^{\bm{\lambda}_{\mathbf{v},k}(t_{i}-t_{j})}$
"Pulled-forward" estimates: $\mathcal{\hat{Z}}_{\mathbf{k}}\in\mathbb{C}^{d_{qk}\times m\times m},\quad\mathcal{\hat{Z}}_{\mathbf{v}}\in\mathbb{C}^{d_{v}\times m\times m}$:
$\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]=\mathcal{E}_{\mathbf{qk}}[k,i,j]\cdot\mathbf{Z_{k}}[k,j],\quad\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]=\mathcal{E}_{\mathbf{v}}[k,i,j]\cdot\mathbf{Z_{v}}[k,j],$
The residuals: $\mathcal{R}\in\mathbb{C}^{d_{qk}\times m\times m}$:
$\mathcal{R}_{\mathbf{qk}}[k,i,j]=\mathbf{Z_{q}}[k,i]-\hat{\mathcal{Z}}_{\mathbf{k}}[k,i,j]$
The norms of the residuals:
$\big|\mathcal{R}_{\mathbf{qk}}[k,i,j]\big|^{2}=\mathrm{Re}^{2}\left(\mathcal{R}_{\mathbf{qk}}[k,i,j]\right)+\mathrm{Im}^{2}\left(\mathcal{R}_{\mathbf{qk}}[k,i,j]\right)$
The key and value precision tensors: $\mathcal{P}^{C}_{\mathbf{qk}}\in\mathbb{R}^{d_{qk}\times m\times m},\quad\mathcal{P}^{C}_{\mathbf{v}}\in\mathbb{R}^{d_{v}\times m\times m}$:
$\mathcal{P}^{C}_{\mathbf{qk}}[:,i,j]=\left(\bm{\lambda}_{C\mathbf{qk}}^{2}\odot\bm{\lambda}_{Q\mathbf{qk}}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})|t_{j}-t_{i}|}}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})}+\bm{\lambda}_{R\mathbf{qk}}\odot e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})|t_{j}-t_{i}|}\right)^{-1}$
$\mathcal{P}^{C}_{\mathbf{v}}[:,i,j]=\left(\bm{\lambda}_{C\mathbf{v}}^{2}\odot\bm{\lambda}_{Q\mathbf{v}}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})|t_{j}-t_{i}|}}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})}+\bm{\lambda}_{R\mathbf{v}}\odot e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})|t_{j}-t_{i}|}\right)^{-1}$
We apply a causal attention mask to $\mathcal{P}^{C}_{\mathbf{qk}},\mathcal{P}^{C}_{\mathbf{v}}$, masking out all $j>i$.
Compute the weights: $\mathbf{W}\in\mathbb{R}^{m\times m}$:
$\mathbf{W}[i,j]=\left(1+\sum_{k}\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]\cdot\big|\mathcal{R}_{\mathbf{qk}}[k,i,j]\big|^{2}/\nu\right)^{-1}$
Here, $\nu$ is a learnable scalar. The value parameters $\mathbf{W_{v}},\bm{\lambda}_{\mathbf{v}},\bm{\lambda}_{Q\mathbf{v}},\bm{\lambda}_{R\mathbf{v}}$ are responsible for the prior, while the query and key parameters $\mathbf{W_{q}},\mathbf{W_{k}},\bm{\lambda}_{\mathbf{qk}},\bm{\lambda}_{Q\mathbf{qk}},\bm{\lambda}_{R\mathbf{qk}}$ are responsible for the adaptive reweighting.
Compute the unnormalized attention scores: $\tilde{\mathcal{A}}\in\mathbb{R}^{d_{v}\times m\times m}$:
$\tilde{\mathcal{A}}[k,i,j]=\mathbf{W}[i,j]\cdot\mathcal{P}_{\mathbf{v}}^{C}[k,i,j]$
Normalize to obtain the attention tensor: $\mathcal{A}\in\mathbb{R}^{d_{v}\times m\times m}$:
$\mathcal{A}[k,i,j]=\tilde{\mathcal{A}}[k,i,j]\big/\sum_{j^{\prime}\leq i}\tilde{\mathcal{A}}[k,i,j^{\prime}]$
Compute the estimate: $\mathbf{\bar{Z}}\in\mathbb{C}^{d_{v}\times m}$:
$\mathbf{\bar{Z}}[k,i]=\sum_{{j\leq i}}\mathcal{A}[k,i,j]\cdot\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$
Optionally, add a residual connection:
$\mathbf{\bar{Z}}[k,i]\leftarrow(\mathbf{1}-\bm{\delta})\odot\mathbf{Z}_{\mathbf{v}}[k,i]+\bm{\delta}\odot\mathbf{\bar{Z}}[k,i],\quad\bm{\delta}\in(0,1]^{d_{v}}$
After $n$ layers with shared weights, compute the prediction $\mathbf{Z}^{+}\in\mathbb{C}^{d_{e}\times m}$ using a one-step forward transition:
$\mathbf{Z}_{\mathbf{v}}^{+}[k,i]=\mathcal{E}_{\mathbf{v}}[k,i+1,i]\cdot\mathbf{\bar{Z}}[k,i]$
Map back to the original basis:
$\mathbf{Z}^{+}[k,i]=\mathbf{W_{p}}\mathbf{Z}_{\mathbf{v}}^{+}[k,i]$
This adaptive filter generalizes attention by introducing a more structured similarity measure that integrates uncertainty and time decay. By using the same (diagonal) matrix exponential to filter and to predict the next step, we enforce self-consistency in the model. One issue is that it may be preferable to do multiple steps of iterative refinement of the estimated trajectory before prediction. One way to address this is to give the model extra flexibility by outputting a gated interpolation between estimate and prediction. That is:
$\mathbf{\hat{z}}^{k+1}_{ii}=\bigl(1-\mathbf{g}^{(k)}\bigr)\,\mathbf{e}_{i}^{(k)}+\mathbf{g}^{(k)}\,\mathbf{p}_{i}^{(k)}.$
where $\mathbf{e}_{i}^{(k)}$ is the estimate, $\mathbf{p}_{i}^{(k)}$ is the prediction, and $\mathbf{g}^{(k)}\in(0,1]^{d_{v}}$ is a learnable gate. We can also make $\mathbf{g}^{(k)}$ a function of local residuals or features, for example:
$\mathbf{g}^{(k)}=\sigma\!\left(\mathbf{a}^{\top}\phi\!\left(\big|\mathbf{r}^{(k)}\big|^{2}\right)+\mathbf{G}\right),$
so the layer decides (data-dependently) whether to predict or to keep refining the estimate.
3.2 Complexity
Vanilla attention with input length $m$ and embedding dimension $d$ has time complexity $\mathcal{O}(m^{2}d)$ to compute attention scores and apply them across the sequence, plus $\mathcal{O}(md^{2})$ to transform the inputs into query, key, and value vectors and to transform the outputs through a linear layer. The storage of the query, key, and value matrices requires $\mathcal{O}(md)$ memory, and the attention matrix requires $\mathcal{O}(m^{2})$ memory.
The naive tensor implementation of attention above involves element-wise multiplication of $d\times m\times m$ tensors, resulting in time complexity $\mathcal{O}(m^{2}d+md^{2})$, which is the same as vanilla attention, but slower in practice due to less-optimized element-wise operations. The time complexity to propagate the precision matrices $\mathbf{P}_{ij}^{C}$ through the differential Lyapunov equation in the full (non-diagonalizable case) is $\mathcal{O}(d^{3})$, plus another $\mathcal{O}(d^{3})$ to invert each of these matrices, for each of the $\mathcal{O}(m^{2})$ time differences. Hence, the total time complexity is $\mathcal{O}(m^{2}d^{3})$. In contrast, for systems that admit a closed-form solution, the propagation and matrix inversions become $\mathcal{O}(d)$, so the total time complexity is $\mathcal{O}(m^{2}d)$. If the measurements are sampled at equal time intervals, the time complexity reduces to $\mathcal{O}(md)$.
Hence, the main bottleneck is not run-time, but memory, as we need to store all $m^{2}$ residuals, pulled-forward estimates, and the attention tensor, which are all tensors of size $\mathcal{O}(m^{2}d)$. In the following sections, we examine how to simplify the algorithm to avoid this bottleneck.
3.3 Convolutional Representation
If the time steps are equal, i.e. $\Delta t=t_{i+1}-t_{i}$ is fixed for all $i$, we can store all $ij$ exponentials in a kernel of size 2(m1)d2(m-1)d, or (m1)d(m-1)d in the causal case, rather than $m^{2}d$. In particular, we compute: $\mathcal{K}_{\mathbf{v}}[\tau]=e^{\bm{\lambda}_{\mathbf{v}}\Delta t\tau}$ and $\mathcal{K}_{\mathbf{qk}}[\tau]=e^{\bm{\lambda}_{\mathbf{qk}}\Delta t\tau}$, for $\tau\in[m-1,-(m-1)]$:
$\mathcal{K}_{\mathbf{v}}=\left[e^{\bm{\lambda}_{\mathbf{v}}(m-1)\Delta t},e^{\bm{\lambda}_{\mathbf{v}}(m-2)\Delta t},\dots,e^{\bm{\lambda}_{\mathbf{v}}\Delta t},e^{\mathbf{0}},e^{-\bm{\lambda}_{\mathbf{v}}\Delta t},\dots,e^{-\bm{\lambda}_{\mathbf{v}}(m-1)\Delta t},e^{-\bm{\lambda}_{\mathbf{v}}m\Delta t}\right],$
or in the causal case, $\tau\in[m-1,0]$:
$\mathcal{K}_{\mathbf{v}}=\left[e^{\bm{\lambda}_{\mathbf{v}}(m-1)\Delta t},e^{\bm{\lambda}_{\mathbf{v}}(m-2)\Delta t},\dots,e^{\bm{\lambda}_{\mathbf{v}}\Delta t},1,\dots,1\right],$
(where we pad the kernel with $m$ 1s for easier indexing.) We define $\mathcal{K}_{\mathbf{qk}}$ similarly. We can then efficiently compute the exponentials using a Toeplitz matrix to index from $\mathcal{K}_{\mathbf{v}}$ and $\mathcal{K}_{\mathbf{qk}}$. In particular, $e^{\bm{\lambda}_{\mathbf{v}k}(t_{i}-t_{j})}=\mathcal{K}_{\mathbf{v}}[m-(i-j)]$. The estimate tensors are then: $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]=\mathcal{K}_{\mathbf{v}}[m-(i-j)]\cdot\mathbf{Z}_{\mathbf{v}}[k,j]$ and $\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]=\mathcal{K}_{\mathbf{qk}}[m-(i-j)]\cdot\mathbf{Z}_{\mathbf{k}}[k,j]$.
Similarly, the precision tensor is Toeplitz, so we can again represent it with a kernel of size $md$:
$\mathcal{K}_{\mathbf{qk}}^{V}[\tau]=\bm{\lambda}_{C\mathbf{qk}}^{2}\odot\bm{\lambda}_{Q\mathbf{qk}}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})\Delta t\tau}}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})}+\bm{\lambda}_{R\mathbf{qk}}\odot e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})\Delta t\tau}$
which we can express in terms of $\mathcal{K}_{\mathbf{qk}}[\tau]$:
$\mathcal{K}_{\mathbf{qk}}^{V}[\tau]=\bm{\lambda}_{C\mathbf{qk}}^{2}\odot\bm{\lambda}_{Q\mathbf{qk}}\odot\frac{\mathbf{1}-\mathcal{K}^{\dagger}_{\mathbf{k}}[\tau]\mathcal{K}_{\mathbf{qk}}[\tau]}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})}+\bm{\lambda}_{R\mathbf{qk}}\odot\big(\mathcal{K}^{\dagger}_{\mathbf{k}}[\tau]\mathcal{K}_{\mathbf{qk}}[\tau]\big)$
The full precision tensor is then:
$\mathcal{P}_{\mathbf{k}}^{C}[k,i,j]=\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|],$
where $\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]=\mathcal{K}_{\mathbf{qk}}^{V}[k,|i-j|]^{-1}$, and similarly for $\mathcal{K}_{\mathbf{v}}^{P}[\tau]$ and $\mathcal{P}_{\mathbf{v}}^{C}$.
However, we must still store the full tensors of pulled-forward estimates $\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]$ and $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$, which are of size $d_{qk}\times m\times m$ and $d_{v}\times m\times m$, respectively.
3.4 Broadcasting the Estimates
To avoid explicitly constructing the pulled-forward estimates $\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]$ and $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$, we can define separate arrays for the $i$ and $j$ components:
$\bm{\Phi}_{\mathbf{v}}[k,i]:=e^{\bm{\lambda}_{\mathbf{v}k}t_{i}},\quad\mathbf{\hat{Z}}_{\mathbf{v}}[k,j]:=e^{-\bm{\lambda}_{\mathbf{v}k}t_{j}}\mathbf{Z_{v}}[k,j].$
We can then factorize $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$:
$\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]=\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\mathbf{\hat{Z}}_{\mathbf{v}}[k,j],$
And likewise for $\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]$. This requires $2md$ computations rather than $md$, but reduces the memory requirement from $m^{2}d$ to $2md$.
Since $\bm{\Phi}_{\mathbf{v}}[k,i]$ does not depend on $j$, we can pull it outside the sum:
$\mathbf{\bar{Z}_{v}}[k,i]:=\sum_{j\leq i}\mathcal{A}[k,i,j]\cdot\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]=\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\sum_{j\leq i}\mathcal{A}[k,i,j]\cdot\mathbf{\hat{Z}}_{\mathbf{v}}[k,j]$
So we can avoid forming the full $d_{v}\times m\times m$ tensor $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$.
3.5 Factorizing the Residuals
Recall that:
$\mathcal{R}_{\mathbf{qk}}[k,i,j]=\mathbf{Z_{q}}[k,i]-\hat{\mathcal{Z}}_{\mathbf{k}}[k,i,j]$
Letting:
$\bm{\Phi}_{\mathbf{qk}}[k,i]:=e^{\bm{\lambda}_{\mathbf{qk}}t_{i}},\quad\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]:=e^{-\bm{\lambda}_{\mathbf{qk}}t_{i}}\mathbf{Z}_{\mathbf{q}}[k,i],\quad\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]:=e^{-\bm{\lambda}_{\mathbf{qk}}t_{j}}\mathbf{Z}_{\mathbf{k}}[k,j],$
[…]
where:
[…]
We need to compute the matrix of Mahalanobis distances:
$\mathbf{D}[i,j]:=\sum_{k}\mathcal{R}^{*}_{\mathbf{qk}}[k,i,j]\cdot\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]\cdot\mathcal{R}_{\mathbf{qk}}[k,i,j]$
[…]
$\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]-\mathbf{\hat{Z}}^{*}_{\mathbf{k}}[k,j]\right)\cdot\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]\cdot\left(\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]-\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)$
[…]
$\mathbf{D}[i,j]=\sum_{k}\mathcal{K}^{C}_{\mathbf{qk}}[k,|i-j|]\cdot\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\cdot\Big[\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}+\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}-2\mathrm{Re}\big(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big)\Big]$
But computing the product $\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]$ still requires constructing a $d_{qk}\times m\times m$ tensor.
3.6 Isotropic Decay and Noise
If $\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]$ loses its $k$-dependence, each of the terms in $\mathbf{D}[i,j]$ simplifies to a matrix multiplication. In particular, let us assume that the real part of the eigenvalues is a shared scalar:
$\lambda_{k}=\alpha+i\omega_{k},\quad\alpha,\omega_{k}\in\mathbb{R}$
This corresponds to a system with an isotropic plus skew-symmetric state matrix, $\mathbf{A}=\alpha\mathbf{I}+\bm{\Omega}$, where $\bm{\Omega}\in\mathbb{R}^{d\times d}$ is skew-symmetric ($\bm{\Omega}^{\top}=-\bm{\Omega}$) and diagonalizable over $\mathbb{C}$, that is $\bm{\Omega}=\mathbf{S}\bm{\Lambda}_{\Omega}\mathbf{S}^{-1}$, where $\mathbf{S}\in\mathbb{C}^{d\times d}$, $\bm{\Lambda}_{\Omega}=\text{diag}(\omega_{1},\dots\omega_{d})$, with $\omega_{k}\in i\mathbb{R}$. We also assume that the noise is isotropic, i.e. that the noise parameters are scalar multiples of identity: $\bm{\Lambda}_{C}=\lambda_{C}\mathbf{I},\bm{\Lambda}_{Q}=\sigma^{2}\mathbf{I},\bm{\Lambda}_{R}=\eta^{2}\mathbf{I}$.
Note that we could define separate query/key and value dynamic parameters $\alpha$ and $\bm{\omega}$, but here we use shared parameters for simplicity.
Our covariance kernel becomes:
$\mathcal{K}^{V}(\tau)=\sigma_{V}^{2}(\tau)\mathbf{I}_{d},\quad\text{where:}\quad\sigma_{V}^{2}(\tau)=\big|\lambda_{C}\big|^{2}\sigma^{2}\frac{1-e^{-2\alpha\Delta t\tau}}{-2\alpha}+\eta^{2}e^{-2\alpha\Delta t\tau}$
and the $ij$th component of the precision matrix becomes:
$\mathbf{P}_{ij}^{C}=\frac{1}{\sigma_{V}^{2}(|i-j|)}.$
We can then pull $\mathbf{P}^{C}_{\mathbf{qk}}\in\mathbb{R}^{m\times m}$ outside the sum:
$\mathbf{D}[i,j]=\mathbf{P}^{C}_{\mathbf{qk}}[i,j]\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\Big[\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}+\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}-2\mathrm{Re}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
We can now express each of these three sums as matrix products:
$\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}=\sum_{k}\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\big|\mathbf{\hat{Z}}_{\mathbf{q}}\big|^{2}\big)\in\mathbb{R}^{m\times 1},$
$\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}=\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\big)^{\top}\big|\mathbf{\hat{Z}}_{\mathbf{k}}\big|^{2}\in\mathbb{R}^{m\times m},$
$\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]=\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big)^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\in\mathbb{C}^{m\times m}$
We then sum terms, where the first term is broadcast from an $m\times 1$ vector to an $m\times m$ matrix:
$\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}:=\mathbf{1}_{m}\big(\big(\big|\bm{\Phi}_{\mathbf{qk}}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big|^{2}\big)^{\top}\mathbf{1}_{d}\big)^{\top}+\big(\big|\bm{\Phi}_{\mathbf{qk}}^{\top}\odot\mathbf{\hat{Z}}_{\mathbf{k}}\big|^{2}-2\mathrm{Re}\Big(\big(\bm{\Phi}_{\mathbf{qk}}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big)^{\dagger}\big(\bm{\Phi}_{\mathbf{qk}}\odot\mathbf{\hat{Z}}_{\mathbf{k}}\big)\Big)$
We can factor out $e^{\alpha t_{i}}$. Let us define: Let $\mathbf{E_{qk}}_{ij}=e^{\alpha(t_{i}-t_{j})}$:
$\mathbf{E_{qk}}[i,j]=e^{\alpha(t_{i}-t_{j})},\quad\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}[k,i]:=e^{i\omega_{k}t_{i}},\quad\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}[k,i]:=e^{-i\omega_{k}t_{i}}$
And:
$\mathbf{\tilde{Z}_{q}}[k,i]:=e^{-i\omega_{k}t_{i}}\mathbf{Z_{q}}[k,i],\quad\mathbf{\tilde{Z}_{k}}[k,j]:=e^{-i\omega_{k}t_{j}}\mathbf{Z_{k}}[k,j].$
that is,
$e^{\alpha t_{i}}\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}[k,i]=\bm{\Phi}_{\mathbf{qk}}[k,i],\quad e^{-\alpha t_{i}}\mathbf{\tilde{Z}_{q}}[k,i]:=\mathbf{\hat{Z}_{q}}[k,i],\quad e^{-\alpha t_{j}}\mathbf{\tilde{Z}_{k}}[k,j]:=\mathbf{\hat{Z}_{k}}[k,j].$
Or in matrix form:
$\mathbf{\tilde{Z}_{q}}:=\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}\odot\mathbf{Z_{q}},\quad\mathbf{\tilde{Z}_{k}}:=\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}\odot\mathbf{Z_{k}}$
So:
$\mathcal{R}_{\mathbf{qk}}[k,i,j]=e^{i\omega_{k}t_{i}}\mathbf{\tilde{Z}_{q}}[k,i]-e^{\alpha(t_{i}-t_{j})}\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}[k,i]\cdot\mathbf{\tilde{Z}_{k}}[k,j]$
Or in matrix form:
$\mathcal{R}_{\mathbf{qk}}=\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}\odot\mathbf{\tilde{Z}_{q}}-\mathbf{E_{qk}}\odot(\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+})^{\dagger}\mathbf{\tilde{Z}_{k}}$
Hence,
$\big|\mathbf{R}_{\mathbf{qk}}[i,j]\big|^{2}=e^{2\alpha t_{i}}\sum_{k}\Big[e^{-2\alpha t_{i}}\big|\mathbf{Z_{q}}[k,i]\big|^{2}+e^{-2\alpha t_{j}}\big|\mathbf{Z_{k}}[k,j]\big|^{2}-2\mathrm{Re}\left(e^{-\alpha t_{i}}e^{-\alpha t_{j}}\mathbf{\tilde{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
$\sum_{k}\big|\mathbf{Z_{q}}[k,i]\big|^{2}+e^{2\alpha(t_{i}-t_{j})}\sum_{k}\big|\mathbf{Z_{k}}[k,j]\big|^{2}-2e^{\alpha(t_{i}-t_{j})}\sum_{k}\mathrm{Re}\left(\mathbf{\tilde{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)$
Let $\mathbf{1}_{m}\in\mathbb{R}^{m\times 1}$ and $\mathbf{1}_{d_{qk}}\in\mathbb{R}^{d_{qk}\times 1}$ be column vectors of $1$s. In matrix form:
$\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}=\mathbf{1}_{m}\big(\mathbf{1}_{d_{qk}}^{\top}|\mathbf{Z_{q}}|^{2}\big)+\mathbf{E_{qk}}^{2}\odot\big(\mathbf{1}_{d_{qk}}^{\top}|\mathbf{Z_{k}}|^{2}\big)^{\top}\mathbf{1}_{m}^{\top}-2\mathbf{E_{qk}}\odot\mathrm{Re}\big(\mathbf{\tilde{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{k}}\big)$
(where the modulus is taken element-wise).
Defining a score matrix $\mathbf{L}=\log(\mathbf{P}_{\mathbf{v}}^{C})-\log\left(\mathbf{I}+\mathbf{P}^{C}_{\mathbf{qk}}\odot\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}/\nu\right)$, and a causal mask $\mathbf{M}_{\text{causal}}$, we can then express the row-normalization using softmax:
$\mathbf{A}[i,j]=e^{\mathbf{L}[i,j]\odot\mathbf{M}_{\text{causal}}}\big/\sum_{j}e^{\mathbf{L}[i,j]\odot\mathbf{M}_{\text{causal}}}$
The estimate is:
$\mathbf{\bar{Z}_{v}}[k,i]=\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\sum_{j}\mathbf{A}[i,j]\cdot\mathbf{\hat{Z}}_{\mathbf{v}}[k,j]$
Whereas for the query-key dynamics, the isotropic decay condition $\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})=\alpha\mathbf{1}$ is necessary to reduce the memory requirement, there is no such constraint on $\bm{\lambda}_{\mathbf{v}}$. However, we face another problem: for $\mathrm{Re(}\bm{\lambda}_{\mathbf{v}})\leq\mathbf{0}$, $\mathbf{\hat{Z}_{v}}$ grows exponentially and $\bm{\Phi}_{\mathbf{v}}$ decays exponentially. However, if we again assume isotropic decay $\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})=\alpha\mathbf{1}$, we can re-factorize $\mathbf{\bar{Z}_{v}}[k,i]$ such that this problem is avoided.
Define:
$\mathbf{E_{v}}[i,j]=e^{\alpha(t_{i}-t_{j})},\quad\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}[k,i]:=e^{i\omega_{k}t_{i}},\quad\bm{\tilde{\Phi}}_{\mathbf{v}}^{-}[k,i]:=e^{-i\omega_{k}t_{i}}$
And define:
$\mathbf{\tilde{Z}_{v}}:=\bm{\tilde{\Phi}}_{\mathbf{v}}^{-}\odot\mathbf{Z_{v}}.$
That is,
$e^{\alpha t_{i}}\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}[k,i]=\bm{\Phi}_{\mathbf{v}}[k,i],\quad e^{-\alpha t_{i}}\mathbf{\tilde{Z}_{v}}[k,i]:=\mathbf{\hat{Z}_{v}}[k,i].$
Let $\mathbf{\hat{A}}=\mathbf{A}\odot\mathbf{E_{v}}$. Then,
$\mathbf{\bar{Z}_{v}}[k,i]=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}[k,i]\odot\sum_{j}\mathbf{\hat{A}}[i,j]\cdot\mathbf{\tilde{Z}}_{\mathbf{v}}[k,j]$
which, in matrix form, is:
$\mathbf{\bar{Z}_{v}}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}\odot(\mathbf{\tilde{Z}}_{\mathbf{v}}\mathbf{\hat{A}}^{\dagger})$
Or, in a form more typical for attention:
$\mathbf{\bar{Z}}_{\mathbf{v}}^{\dagger}=(\bm{\tilde{\Phi}}_{\mathbf{v}}^{+})^{\dagger}\odot(\mathbf{\hat{A}}\mathbf{\tilde{Z}}_{\mathbf{v}}^{\dagger})$
Finally, the prediction is:
$\mathbf{Z}_{\mathbf{v}}^{+}=e^{\bm{\lambda}\Delta t}\odot\mathbf{\bar{Z}_{v}}$
This factorization, which separates scalar decay/growth (magnitude) from oscillatory components (direction), reduces the memory requirement of the algorithm from $\mathcal{O}(m^{2}d)$ to $\mathcal{O}(m^{2}+md)$, as in vanilla attention, while maintaining the $\mathcal{O}(m^{2}d)$ compute time. This comes at the price of losing the second-order information provided by the full tensors $\mathcal{P}^{C}_{\mathbf{qk}}$ and $\mathcal{P}^{C}_{\mathbf{v}}$. Whereas $\mathcal{P}^{C}_{\mathbf{qk}},\mathcal{P}^{C}_{\mathbf{v}}$ represent the confidence of the $k$-th feature’s projection from time $t_{j}$ to time $t_{i}$, $\mathbf{P}^{C}_{\mathbf{qk}},\mathbf{P}^{C}_{\mathbf{v}}$ represent measurements of overall confidence between query and key.
Computing the residual matrix $\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}$ requires additional vector summations, broadcasting, and element-wise multiplications, as opposed to dot-product attention’s single matrix product. However, computing $|\mathbf{Z_{q}|}$ and $|\mathbf{Z_{k}|}$ is only $\mathcal{O}(d_{qk}m)$, so it is negligible overhead compared to the $\mathcal{O}(d_{qk}m^{2})$ dot product computation.
The assumption that the real part of the eigenvalues is a shared scalar serves two purposes: (a) it allows the cross term to be represented as a matrix multiplication, which reduces the memory requirement, and (b) it improves the exploding/vanishing gradient problem for $\alpha<0$, because rather than computing $e^{\alpha t_{i}}$ and $e^{-\alpha t_{j}}$ separately, we compute $e^{\alpha(t_{i}-t_{j})}$, which is stable, since $t_{i}>t_{j}$ for causal systems.
[…]
Practically, the differences are:
  • AFA uses complex-valued $\mathbf{W_{q}}$, $\mathbf{W_{k}}$, $\mathbf{W_{v}}$, $\mathbf{W_{p}}$, whereas RetNet uses real-valued projections.
  • AFA also applies the matrix exponential to the value projections, whereas RetNet does not, keeping its values static.
  • AFA incorporates uncertainty priors into attention using propagated covariance kernels.
3.7 Normalized Queries & Keys
If we assume normalized queries and keys, $\sum_{k}\big|\mathbf{Z_{q}}[k,i]\big|^{2}=1$, $\sum_{k}\big|\mathbf{Z_{k}}[k,i]\big|^{2}=1$, the expression for the squared residual simplifies:
$\big|\mathbf{R}_{\mathbf{qk}}[i,j]\big|^{2}=1+e^{2\alpha(t_{i}-t_{j})}-2e^{\alpha(t_{i}-t_{j})}\sum_{k}\mathrm{Re}\left(\mathbf{\tilde{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)$
Or in matrix, form:
$\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}=\mathbf{I}+\mathbf{E_{qk}}^{2}-2\mathbf{E_{qk}}\odot\mathrm{Re}\big(\mathbf{\tilde{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{k}}\big)$
If the query-key and value dynamics parameters are shared, $\mathbf{P}^{C}_{\mathbf{qk}}=\mathbf{P}^{C}_{\mathbf{v}}:=\mathbf{P}^{C}$, so:
$\mathbf{\tilde{A}}=\frac{\nu\mathbf{I}}{\nu\mathbf{V}^{C}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}}$
With both of these simplifications, the term in the denominator is:
$(\nu\mathbf{V}^{C}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2})[i,j]=\nu\big|\lambda_{C}\big|^{2}\sigma^{2}\frac{1-e^{-2\alpha\Delta t_{ij}}}{-2\alpha}+\eta^{2}e^{-2\alpha\Delta t_{ij}}+1+e^{-2\alpha\Delta t_{ij}}$
$-2e^{-\alpha\Delta t_{ij}}\sum_{k}\mathrm{Re}\left(\mathbf{\tilde{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)$
Collecting terms and letting $a=1+\frac{\nu|\lambda_{C}|^{2}\sigma^{2}}{-2\alpha}$, $b=1+\eta^{2}-\frac{\nu|\lambda_{C}|^{2}\sigma^{2}}{-2\alpha}$, this can be written as:
$(\nu\mathbf{V}^{C}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2})[i,j]=a+be^{-2\alpha\Delta t_{ij}}-2e^{-\alpha\Delta t_{ij}}\sum_{k}\mathrm{Re}\left(\mathbf{\tilde{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)$
The attention score denominator is now explicitly decomposed into a content-independent positional decay term and a content-dependent similarity term modulated by the time difference.
3.8 Unitary Dynamics
If we further constrain our state matrix $\mathbf{A}$ to have purely imaginary eigenvalues, i.e. $\alpha=0$, then $e^{\mathbf{\Lambda}}$ is orthogonal, and our dynamical system is unitary (norm preserving) in the eigenbasis.
Using L’Hôpital’s rule:
$\lim_{\alpha\to 0}\frac{1-e^{2\alpha\Delta t\tau}}{-2\alpha}=\Delta t\,\tau$
Our covariance kernel becomes:
$\sigma_{V}^{2}(\tau):=\big|\lambda_{C}\big|^{2}\sigma^{2}\Delta t\tau+\eta^{2}$
If, in addition, $\mathbf{S}$ is unitary (in which case $\mathbf{A}$ is skew-symmetric and the dynamics are unitary in the original basis), then $\big|\mathbf{Z}_{\mathbf{q}}[k,i]\big|^{2}=\big|\mathbf{Z}[k,i]\big|^{2}$, $\big|\mathbf{Z}_{\mathbf{v}}[k,i]\big|^{2}=\big|\mathbf{Z}[k,i]\big|^{2}$, so if the inputs are normalized to unit length, the queries and keys are as well: $\sum_{k}\big|\mathbf{Z}_{\mathbf{q}}[k,i]\big|^{2}=\sum_{k}\big|\mathbf{Z}_{\mathbf{k}}[k,j]\big|^{2}=1$. Hence,
$\big|\mathbf{R_{qk}}\big|^{2}=2\Big(1-\mathrm{Re}\big(\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big)\Big)$
The score matrix is:
[…]
$\log(\nu\mathbf{V}^{C}+2)-\log\left(1-\frac{2\mathrm{Re}\big(\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big)}{\nu\mathbf{V}^{C}+2}\right)$
Expanding around $\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}=\mathbf{0}$, this is approximately:
$\mathbf{L}\approx-\log(\nu\mathbf{V}^{C}+2)+\frac{2\mathrm{Re}\big(\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big)}{\nu\mathbf{V}^{C}+2}$
If we set $\sigma^{2}\rightarrow 0$ (zero process noise), i.e. assume a deterministic system, then $\mathbf{V}^{C}=\eta^{2}\mathbf{1}$ becomes a simple scalar bias. Letting $\tau=2/(\nu\eta^{2}+2)$, the unnormalized softmax becomes:
$\frac{\tau}{2}\cdot\exp(\tau\mathrm{Re}\big(\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}))$
The constant term $\frac{\tau}{2}$ outside the exponential cancels in the normalization. The imaginary part of the matrix exponential becomes a multiplicative relative positional encoding, of the form of xPos (Sun et al., 2022) and RetNet (Sun et al., 2023). Under the assumption $\alpha=0$, the propagated covariance $\mathbf{V}^{C}$ is a bias of the form a|ij|+ba|i-j|+b, as in ALiBi (Press et al., 2022).
In the above formulation, if the inputs are normalized beforehand, then the projection matrices $\mathbf{W_{q}}$ and $\mathbf{W_{k}}$ must be unitary to ensure that $|\mathbf{\hat{Z}}_{q}|$ and $|\mathbf{\hat{Z}}_{k}|$ remain 1, simplifying the computation. By contrast, normalizing the queries and keys after the projections removes this constraint, since $\mathbf{W_{q}}$ and $\mathbf{W_{k}}$ no longer need to preserve norms. This motivates the choice of post-projection normalization, a strategy explored in prior work such as Query-Key Normalization for Transformers (Henry et al., 2020).
3.9 The Inhomogeneous Case
Consider the inhomogeneous SDE:
$d\mathbf{x}(t)=\big(\mathbf{A}\mathbf{x}(t)+\mathbf{u}\big)\,dt+\mathbf{G}\,d\mathbf{w}(t),\qquad\mathbf{x},\mathbf{u}\in\mathbb{R}^{d};\,\mathbf{A},\mathbf{G}\in\mathbb{R}^{d\times d}$
where $\mathbf{u}$ is a constant drift term.
We can represent this as a homogeneous SDE with one additional dimension:
$d\mathbf{x}_{\mathrm{aug}}(t)=\mathbf{A}_{\mathrm{aug}}\,\mathbf{x}_{\mathrm{aug}}(t)\,dt+\mathbf{G}_{\mathrm{aug}}\,d\mathbf{w}(t),$
$\mathbf{x}_{\mathrm{aug}}(t)=\begin{bmatrix}\mathbf{x}(t)\\[2.0pt] 1\end{bmatrix}\in\mathbb{R}^{d+1},\quad\mathbf{A}_{\mathrm{aug}}=\begin{pmatrix}\mathbf{A}&\mathbf{u}\\[4.0pt] \mathbf{0}^{\top}&0\end{pmatrix},\quad\mathbf{G}_{\mathrm{aug}}=\begin{pmatrix}\mathbf{G}\\[4.0pt] \mathbf{0}^{\top}\end{pmatrix}\in\mathbb{R}^{(d+1)\times(d+1)}.$
However, $\mathbf{A}_{\mathrm{aug}}$ and $\mathbf{Q}_{\mathrm{aug}}=\mathbf{G}_{\mathrm{aug}}\mathbf{G}_{\mathrm{aug}}^{\top}$ are now singular. Hence, we must deal with the inhomogeneous case directly.
The solution to the above inhomogeneous SDE is:
$\mathbf{x}(t_{j})=e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i})+\left(\int_{0}^{\Delta t_{ij}}e^{\mathbf{A}(\Delta t_{ij}-\tau)}d\tau\right)\mathbf{u}+\int_{0}^{\Delta t_{ij}}e^{\mathbf{A}(\Delta t_{ij}-\tau)}\mathbf{G}d\mathbf{w}(\tau)$
where $\Delta t=t_{j}-t_{i}$. Letting $\mathbf{G}_{u}(\Delta t_{ij}):=\int_{0}^{\Delta t_{ij}}e^{\mathbf{A}(\Delta t_{ij}-\tau)}d\tau$, our pulled-back/forward estimate of $\mathbf{x}(t_{i})$ from $\mathbf{x}(t_{j})$ is then:
$\mathbf{\hat{x}}_{i}(t_{j})=e^{-\mathbf{A}\Delta t_{ij}}\big(\mathbf{x}(t_{j})-\mathbf{G}_{u}(\Delta t_{ij})\mathbf{u}\big)$
Letting $\mathbf{u_{s}}:=\mathbf{S}\mathbf{u}\mathbf{S}^{-1}$, the drift term is:
$\mathbf{G}_{u}(\Delta t_{ij})\mathbf{u}=\mathbf{S}\Big(e^{\bm{\Lambda}\Delta t_{ij}}\int_{0}^{\Delta t_{ij}}e^{-\bm{\Lambda}\tau}d\tau\Big)\mathbf{u}_{\mathbf{s}}\mathbf{S}^{-1}=\mathbf{S}\Big(\frac{e^{\bm{\Lambda}\Delta t_{ij}}-\mathbf{I}}{\bm{\Lambda}}\Big)\mathbf{u}_{\mathbf{s}}\mathbf{S}^{-1}$
In the eigenbasis,
$\mathbf{\hat{x}}_{\mathbf{s}i}(t_{j})=e^{-\bm{\Lambda}\Delta t_{ij}}\Big(\mathbf{x}_{\mathbf{s}}(t_{j})-\frac{e^{\bm{\Lambda}\Delta t_{ij}}-\mathbf{I}}{\bm{\Lambda}}\Big)\mathbf{u}_{\mathbf{s}}$
$e^{-\bm{\Lambda}\Delta t_{ij}}\mathbf{x}_{\mathbf{s}}(t_{j})-\Big(\frac{\mathbf{I}-e^{-\bm{\Lambda}\Delta t_{ij}}}{\bm{\Lambda}}\Big)\mathbf{u}_{\mathbf{s}}$
$e^{-\bm{\Lambda}\Delta t_{ij}}\Big(\mathbf{x}_{\mathbf{s}}(t_{j})+\frac{\mathbf{u}_{\mathbf{s}}}{\bm{\Lambda}}\Big)-\frac{\mathbf{u}_{\mathbf{s}}}{\bm{\Lambda}}$
(where, as always, division is element-wise).
Including the measurement model:
$\mathbf{\hat{z}}_{\mathbf{s}i}(t_{j})=e^{-\bm{\Lambda}\Delta t_{ij}}\Big(\mathbf{z}_{\mathbf{s}}(t_{j})+\bm{\Lambda}_{C}\frac{\mathbf{u}_{\mathbf{s}}}{\bm{\Lambda}}\Big)-\bm{\Lambda}_{C}\frac{\mathbf{u}_{\mathbf{s}}}{\bm{\Lambda}}$
Since $\mathbf{u}$ is deterministic, the propagation of uncertainty is unchanged.
We will now examine how this modifies AFA. Define a bias:
$\mathbf{b}_{\mathbf{qk}}[k]:=\bm{\lambda}_{C,k}\frac{\mathbf{u}_{\mathbf{qk}}[k]}{\bm{\lambda}_{\mathbf{qk},k}}$
Our key estimate tensor is:
$\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]=\mathcal{E}_{\mathbf{qk}}[k,i,j]\cdot\Big(\mathbf{Z_{k}}[k,j]+\mathbf{b}_{\mathbf{qk}}[k]\Big)-\mathbf{b}_{\mathbf{qk}}[k]$
So the residual tensor is:
$\mathcal{R}_{\mathbf{qk}}[k,i,j]=\mathbf{Z_{q}}[k,i]-\hat{\mathcal{Z}}_{\mathbf{k}}[k,i,j]$
$\Big(\mathbf{Z_{q}}[k,i]+\mathbf{b}_{\mathbf{qk}}[k]\Big)-\mathcal{E}_{\mathbf{qk}}[k,i,j]\cdot\Big(\mathbf{Z_{k}}[k,j]+\mathbf{b}_{\mathbf{qk}}[k]\Big)$
Define biased queries and keys:
$\mathbf{Z}^{u}_{\mathbf{q}}[k,j]:=\mathbf{Z_{q}}[k,i]+\mathbf{b}_{\mathbf{qk}}[k],\quad\mathbf{Z}^{u}_{\mathbf{k}}[k,j]:=\mathbf{Z_{k}}[k,j]+\mathbf{b}_{\mathbf{qk}}[k].$
The residual becomes:
$\mathcal{R}_{\mathbf{qk}}[k,i,j]=\mathbf{Z}^{u}_{\mathbf{q}}[k,j]-\mathcal{E}_{\mathbf{qk}}[k,i,j]\cdot\mathbf{Z}^{u}_{\mathbf{k}}[k,j]$
$\bm{\Phi}_{\mathbf{qk}}[k,i]\cdot\left(\mathbf{\hat{Z}}^{u}_{\mathbf{q}}[k,i]-\mathbf{\hat{Z}}_{\mathbf{k}}^{u}[k,j]\right)$
We can then proceed exactly as before, using $\mathbf{\hat{Z}}^{u}_{\mathbf{q}}[k,i]$ and $\mathbf{\hat{Z}}_{\mathbf{k}}^{u}[k,j]$ in place of $\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]$ and $\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]$.
The value estimate tensor is:
$\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]=\mathcal{E}_{\mathbf{v}}[k,i,j]\cdot\Big(\mathbf{Z_{v}}[k,j]+\mathbf{b}_{\mathbf{v}}[k]\Big)-\mathbf{b}_{\mathbf{v}}[k]$
So the MLE is:
$\mathbf{\bar{Z}_{v}}[k,i]:=\sum_{j\leq i}\mathcal{A}[k,i,j]\cdot\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$
$\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\sum_{j\leq i}\mathcal{A}[k,i,j]\cdot\mathbf{\hat{Z}}^{u}_{\mathbf{v}}[k,j]-\mathbf{b}_{\mathbf{v}}[k]\cdot\sum_{j\leq i}\mathcal{A}[k,i,j]$
$\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\sum_{j\leq i}\mathbf{A}[i,j]\cdot\mathbf{\hat{Z}}^{u}_{\mathbf{v}}[k,j]-\mathbf{b}_{\mathbf{v}}[k]$
Finally, the prediction step becomes:
$\mathbf{Z}_{\mathbf{v}}^{+}[k,i]=e^{\bm{\lambda}_{\mathbf{v},k}\Delta t}\cdot\mathbf{\bar{Z}}[k,i]+\mathbf{b}_{\mathbf{v}}[k]\Big(e^{\bm{\lambda}_{\mathbf{v},k}\Delta t}-1\Big)$
$e^{\bm{\lambda}_{\mathbf{v},k}\Delta t}\cdot\Big(\mathbf{\bar{Z}}[k,i]+\mathbf{b}_{\mathbf{v}}[k]\Big)-\mathbf{b}_{\mathbf{v}}[k]$
Hence,
$\mathbf{Z}_{\mathbf{v}}^{+}[k,i]=e^{\bm{\lambda}_{\mathbf{v},k}\Delta t}\cdot\Big(\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\sum_{j\leq i}\mathbf{A}[i,j]\cdot\mathbf{\hat{Z}}^{u}_{\mathbf{v}}[k,j]\Big)-\mathbf{b}_{\mathbf{v}}[k]$
Hence, we can handle the inhomogeneous case simply by including bias terms in the definitions of the queries, keys, values, and predictions:
$\mathbf{Z_{q}}=\mathbf{W_{q}}\mathbf{Z}+\mathbf{b_{q}},\quad\mathbf{Z_{k}}=\mathbf{W_{k}}\mathbf{Z}+\mathbf{b_{k}},\quad\mathbf{Z_{v}}=\mathbf{W_{v}}\mathbf{Z}+\mathbf{b_{v}},\quad\mathbf{Z}^{+}=\mathbf{W_{p}}\mathbf{Z}_{\mathbf{v}}^{+}+\mathbf{b_{p}}.$
3.10 Measurement Time Delay and Bias
Let us now consider the effect of a time delay $\phi_{t}$ and a bias term $\mathbf{b_{z}}$ in our measurement model:
$d\mathbf{x}(t)=\big(\mathbf{A}\mathbf{x}(t)+\mathbf{u}\big)dt+\mathbf{G}d\mathbf{w}(t),\quad\mathbf{z}(t_{k})=\mathbf{C}\mathbf{x}(t_{k}+\phi_{t})+\mathbf{b}_{\mathbf{z}}+\mathbf{v}(t_{k}),$
Our estimate of the state at time $t_{j}$ from the measurement at time $t_{j}$ is now:
$\mathbf{x}^{\text{est}}(t_{j}+\phi_{t})=\mathbf{C}^{-1}\big(\mathbf{z}(t_{j})-\mathbf{b}_{\mathbf{z}}\big)$
We can then estimate the state at time $t_{j}$ from the estimate of the state at time $t_{j}+\phi_{t}$:
$\mathbf{x}^{\text{est}}(t_{j})=e^{-\mathbf{A}\phi_{t}}\big(\mathbf{x}^{\text{est}}(t_{j}+\phi_{t})-\mathbf{G}_{u}(\phi_{t})\mathbf{u}\big)$
We can then estimate the state at time $t_{i}$ from the state at time $t_{j}$:
$\mathbf{x}^{\text{est}}(t_{i})=e^{\mathbf{-A}\Delta t_{ij}}\big(\mathbf{x}^{\text{est}}(t_{j})-\mathbf{G}_{u}(\Delta t_{ij})\mathbf{u}\big)$
We then estimate the state at time $t_{i}+\phi_{t}$ from the estimate of the state at time $t_{i}$:
$\mathbf{x}^{\text{est}}(t_{i})=e^{-\mathbf{A}\phi_{t}}\big(\mathbf{x}^{\text{est}}(t_{i}+\phi_{t})-\mathbf{G}_{u}(\phi_{t})\mathbf{u}\big)$
Hence,
$\mathbf{x}^{\text{est}}(t_{i}+\phi_{t})=e^{\mathbf{A}\phi_{t}}\mathbf{x}^{\text{est}}(t_{i})+\mathbf{G}_{u}(\phi_{t})\mathbf{u}$
Substituting the expression for $\mathbf{x}^{\text{est}}(t_{i})$:
$\mathbf{x}^{\text{est}}(t_{i}+\phi_{t})=e^{\mathbf{A}\phi_{t}}\Big(e^{\mathbf{-A}\Delta t_{ij}}\big(\mathbf{x}^{\text{est}}(t_{j})-\mathbf{G}_{u}(\Delta t_{ij})\mathbf{u}\big)\Big)+\mathbf{G}_{u}(\phi_{t})\mathbf{u}$
$e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{x}^{\text{est}}(t_{j})+\Big(\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}$
Substituting the expression for $\mathbf{x}^{\text{est}}(t_{j})$:
$\mathbf{x}^{\text{est}}(t_{i}+\phi_{t})=e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\Big(e^{-\mathbf{A}\phi_{t}}\big(\mathbf{x}^{\text{est}}(t_{j}+\phi_{t})-\mathbf{G}_{u}(\phi_{t})\mathbf{u}\big)\Big)+\Big(\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}$
$e^{\mathbf{-A}\Delta t_{ij}}\mathbf{x}^{\text{est}}(t_{j}+\phi_{t})+\Big(\big(\mathbf{I}-e^{\mathbf{-A}\Delta t_{ij}}\big)\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}$
Substituting the expression for $\mathbf{x}^{\text{est}}(t_{j}+\phi_{t})$:
$\mathbf{x}^{\text{est}}(t_{i}+\phi_{t})=e^{\mathbf{-A}\Delta t_{ij}}\Big(\mathbf{C}^{-1}\big(\mathbf{z}(t_{j})-\mathbf{b}_{\mathbf{z}}\big)\Big)+\Big(\big(\mathbf{I}-e^{\mathbf{-A}\Delta t_{ij}}\big)\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}$
Finally, we estimate the measurement at time $t_{i}$ from the estimated state at time $t_{i}+\phi_{t}$:
$\mathbf{\hat{z}}(t_{i})=\mathbf{C}\mathbf{x}^{\text{est}}(t_{i}+\phi_{t})+\mathbf{b}_{\mathbf{z}}$
$\mathbf{C}\Big[e^{\mathbf{-A}\Delta t_{ij}}\mathbf{C}^{-1}\big(\mathbf{z}(t_{j})-\mathbf{b}_{\mathbf{z}}\big)+\Big(\big(\mathbf{I}-e^{\mathbf{-A}\Delta t_{ij}}\big)\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}\Big]+\mathbf{b}_{\mathbf{z}}$
$\underbrace{\mathbf{C}e^{\mathbf{-A}\Delta t_{ij}}\mathbf{C}^{-1}\mathbf{z}(t_{j})}_{\text{{Propagated measurement}}}+\underbrace{\mathbf{C}\Big(\big(\mathbf{I}-e^{\mathbf{-A}\Delta t_{ij}}\big)\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}}_{\text{Propagated drift}}+\underbrace{\Big(\mathbf{I}-\mathbf{C}e^{\mathbf{-A}\Delta t_{ij}}\mathbf{C}^{-1}\Big)\mathbf{b}_{\mathbf{z}}}_{\text{Propagated bias}}$
If $\mathbf{A}$ and $\mathbf{C}$ commute, this becomes:
$\mathbf{\hat{z}}(t_{i})=e^{\mathbf{-A}\Delta t_{ij}}\mathbf{z}(t_{j})+\mathbf{C}\Big(\big(\mathbf{I}-e^{\mathbf{-A}\Delta t_{ij}}\big)\mathbf{G}_{u}(\phi_{t})-e^{\mathbf{-A}(\Delta t_{ij}-\phi_{t})}\mathbf{G}_{u}(\Delta t_{ij})\Big)\mathbf{u}+\Big(\mathbf{I}-e^{\mathbf{-A}\Delta t_{ij}}\Big)\mathbf{b}_{\mathbf{z}}$
Let’s examine the drift term (in the eigenbasis):
$\big(\mathbf{I}-e^{\bm{-\Lambda}\Delta t_{ij}}\big)\Big(\frac{e^{\bm{\Lambda}\phi_{t}}-\mathbf{I}}{\bm{\Lambda}}\Big)-e^{\bm{-\Lambda}(\Delta t_{ij}-\phi_{t})}\Big(\frac{e^{\bm{\Lambda}\Delta t_{ij}}-\mathbf{I}}{\bm{\Lambda}}\Big)$
$\Big(\frac{e^{\bm{\Lambda}\phi_{t}}-\mathbf{I}}{\bm{\Lambda}}\Big)-e^{\bm{-\Lambda}\Delta t_{ij}}\Big(\frac{e^{\bm{\Lambda}\phi_{t}}-\mathbf{I}}{\bm{\Lambda}}\Big)-e^{\bm{\Lambda}(\phi_{t})}\Big(\frac{\mathbf{I}-e^{\bm{-\Lambda}\Delta t_{ij}}}{\bm{\Lambda}}\Big)$
$\frac{\mathbf{I}-e^{\bm{-\Lambda}\Delta t_{ij}}}{\bm{\Lambda}}$
Hence,
$\mathbf{\hat{z}}_{\mathbf{s}}(t_{i})=e^{-\bm{\Lambda}\Delta t_{ij}}\mathbf{z}_{\mathbf{s}}(t_{j})+\Big(\mathbf{I}-e^{-\bm{\Lambda}\Delta t_{ij}}\Big)\Big(\mathbf{b}_{\mathbf{z,s}}-\bm{\Lambda}_{C}\frac{\mathbf{u}_{\mathbf{s}}}{\bm{\Lambda}}\Big)$
So the time delay term disappears and we just get an extra constant bias term, which can be absorbed into the query, key, value, and prediction biases, as before.
4 Implementation
4.1 Algorithm
We now present an implementation of Adaptive Filter Attention, for the case of isotropic noise and decay.
We use scalar noise parameters $\sigma^{2},\eta^{2}$ and dynamics parameters $\alpha,\bm{\omega}$, where the real part of the eigenvalues is set to $\alpha^{2}$ to ensure stability of our LTI model. Note that the eigenvalues of $\bm{\Omega}$ appear in complex conjugate pairs, as a necessary condition for $\mathbf{A}$ to be real, so we can write the vector of eigenvalues as:
$\bm{\lambda}_{\Omega}=\big\{i\omega_{1},-i\omega_{1},\dots,i\omega_{d/2},-i\omega_{d/2}\big\},\,\omega_{k}\in\mathbb{R}$
We switch from the rational weight $\tfrac{1}{1+r_{ij}^{2}/\nu}$ to the equivalent exponential–log form $\exp\!\big(-\tau_{s}\log(1+r_{ij}^{2}/\nu)\big)$, where $\tau_{s}$ is the usual temperature parameter that controls the sharpness of the weighting, with the aim of improving numerical stability; this reduces exactly to the rational form when $\tau_{s}=1$. We drop $\big|\lambda_{C}\big|^{2}$ since we can absorb it into $\sigma^{2}$, and drop the Mahalanobis scale $\nu$, since we can absorb it into $\sigma,\eta,\tau_{s},\mathbf{W}_{Q},\mathbf{W}_{K}$.
Note that $\bm{\Omega}$ is similar to a block diagonal matrix of $2\times 2$ rotation matrices, i.e. there exists some $\mathbf{P}\in\mathbb{R}^{d\times d}$ such that:
[…]
However, to retain our fully diagonalized representation, we must work with complex numbers. We define the following linear layer variant to reproduce the effect of complex multiplication:
Complex Linear Layer. We denote a complex linear layer by $\mathcal{L}(\cdot)$. Instead of operating directly on complex vectors, we represent inputs and outputs by stacking their real and imaginary parts. For an input $\mathbf{x}=\left[\mathbf{x}_{r},\mathbf{x}_{i}\right]^{\top}$ with $\mathbf{x}_{r},\mathbf{x}_{i}\in\mathbb{R}^{d}$, the layer computes:
$\mathcal{L}(\mathbf{x})=\begin{bmatrix}\mathbf{W}_{r}&-\mathbf{W}_{i}\\[4.0pt] \mathbf{W}_{i}&\mathbf{W}_{r}\end{bmatrix}\mathbf{x}+\begin{bmatrix}\mathbf{b}_{r}\\[4.0pt] \mathbf{b}_{i}\end{bmatrix}$,
where $\mathbf{W}_{r},\mathbf{W}_{i}\in\mathbb{R}^{d_{e}\times d}$ are the real and imaginary weight matrices, and $\mathbf{b}_{r},\mathbf{b}_{i}\in\mathbb{R}^{d_{e}}$ are real-valued biases (where $d_{e}$ is the embedding dimension). This construction is equivalent to multiplication by a complex weight matrix $\mathbf{W}_{r}+i\mathbf{W}_{i}$ with bias $\mathbf{b}_{r}+i\mathbf{b}_{i}$, while remaining entirely in the real domain.
We initialize the complex linear layer weights isotropically in the complex plane as:
$\mathbf{W}_{ij}:=\mathbf{W}_{r,ij}+i\mathbf{W}_{i,ij}=\alpha_{ij}\begin{bmatrix}\cos(\phi_{ij})\\ \sin(\phi_{ij})\end{bmatrix},\quad\alpha_{ij}\sim\sqrt{\tfrac{2}{d_{\text{in}}+d_{\text{{out}}}}}\,\mathcal{N}(0,1),\;\;\phi_{ij}\sim\mathcal{U}(0,2\pi).$
Since the inputs to the network layer are real-valued, we can concatenate a vector of zeros to the inputs to represent the imaginary part. In general, the outputs of the network layer may be complex-valued. We pass the real part of the output to the next layer and discard the imaginary part. Alternatively, we can design the input and output projections to map between real and complex inputs/outputs. We pursue the former option here, for simplicity.
Another option to enforce real inputs is to impose a penalty on $\big|\operatorname{Im}(\mathbf{Z_{p}})\big|$ or on $\big|\mathbf{W_{v}}\mathbf{W_{p}}-\mathbf{I}^{d_{v}\times d_{v}}\big|$. For simplicity, we do not impose penalties and allow these matrices to freely vary.
If our data has a known time scale, we can set $\Delta t$ as a known constant, otherwise we can treat it as a learned parameter.
In our implementation, queries and keys are computed from the current values rather than directly from the raw inputs. This design enables a natural recursive approximation of full batch attention: as the value estimates are iteratively updated at each step, the queries and keys reflect the latest state, allowing each iteration to propagate residual information through the trajectory. This is optional.
For clarity, we present the single-head version. A multi-head extension may be obtained in the usual way by applying the procedure in parallel with separate projections and concatenating the outputs, with each head having its own set of complex linear layers $\mathcal{L}_{Q},\mathcal{L}_{K},\mathcal{L}_{V},\mathcal{L}_{P}$ and dynamics parameters $\alpha,\bm{\omega},\sigma,\eta$.
For simplicity, we use shared dynamic parameters $\alpha,\bm{\omega}$ for the queries/keys and values, so $\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}$ and $\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{-}$. We define these as separate variables however, since in general they can be different.
Input: Input sequence $\mathbf{Z}_{r}\in\mathbb{R}^{d\times m}$
Complex linear layers: $\mathcal{L}_{Q},\mathcal{L}_{K},\mathcal{L}_{V},\mathcal{L}_{P}$; dynamics/noise parameters: $\alpha,\bm{\omega},\sigma,\eta$; time step $\Delta t$; softmax temperature $\tau_{s}$; residual weight $\delta$.
Constants: Causal mask $\mathbf{M}_{\text{causal}}\in\{0,1\}^{m\times m}$; stability constant $\epsilon$.
(Define $t_{i}:=\Delta t\cdot i,\quad t_{j}:=\Delta t\cdot j\$.)
1exComplex-valued Projections:
Lift real inputs to complex by concatenating zeros: $\mathbf{Z}=\left[\mathbf{Z}_{r},\mathbf{0}\right]^{\top}\in\mathbb{C}^{2\times d\times m}$
$\big(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big)\leftarrow\mathcal{L}_{V}\big(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z})\big).$
$\big(\operatorname{Re}(\mathbf{Z_{q}}),\operatorname{Im}(\mathbf{Z_{q}})\big)\leftarrow\mathcal{L}_{Q}\big(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big)$,
$\big(\operatorname{Re}(\mathbf{Z_{k}}),\operatorname{Im}(\mathbf{Z_{k}})\big)\leftarrow\mathcal{L}_{K}\big(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big)$.
Exponential kernels:
$\mathcal{K}_{\alpha}[\tau]=\mathbf{1}_{\{\tau>0\}}\odot e^{-\alpha^{2}\Delta t\tau}+\mathbf{1}_{\{\tau\leq 0\}},\quad\mathcal{K}_{\bm{\omega}}[\tau]=\mathbf{1}_{\{\tau>0\}}\odot e^{\bm{\omega}\Delta t\tau}+\mathbf{1}_{\{\tau\leq 0\}}$
Decay kernel: $\kappa[\tau]=\frac{1-\mathcal{K}^{2}_{\alpha}[\tau]+\epsilon}{2\alpha^{2}+\epsilon}\mathbf{1}_{\{\alpha^{2}>\epsilon\}}+\tau\mathbf{1}_{\{\alpha^{2}\leq\epsilon\}}$
Covariance kernel: $\sigma^{2}_{V}(\tau)=\sigma^{2}\kappa[|\tau|]+\eta^{2}\mathcal{K}^{2}_{\alpha}[|\tau|],\quad\mathbf{V}^{C}[i,j]=\sigma^{2}_{V}(i-j)$
Query/Key dynamics:
$\mathbf{E_{qk}}[i,j]=\mathcal{K}_{\alpha}[i-j]=\mathcal{K}_{\alpha}[i]\mathcal{K}_{\alpha}[j]^{-\top},\quad\bm{\tilde{\Phi}}^{+}_{\mathbf{qk}}[k,i]=e^{i\bm{\omega}_{k}t_{i}},\quad\bm{\tilde{\Phi}}^{-}_{\mathbf{qk}}[k,i]=e^{-i\bm{\omega}_{k}t_{i}}$
$\mathbf{\tilde{Z}_{q}}[k,i]=\bm{\tilde{\Phi}}^{-}_{\mathbf{qk}}\odot\mathbf{Z_{q}}[k,i],\quad\mathbf{\tilde{Z}_{k}}[k,j]=\bm{\tilde{\Phi}}^{-}_{\mathbf{qk}}\odot\mathbf{Z_{k}}[k,j]$
Value dynamics:
$\mathbf{E_{v}}[i,j]=\mathcal{K}_{\alpha}[i-j]=\mathcal{K}_{\alpha}[i]\mathcal{K}_{\alpha}[j]^{-\top},\quad\bm{\tilde{\Phi}}^{+}_{\mathbf{v}}[k,i]=e^{i\bm{\omega}_{k}t_{i}},\quad\bm{\tilde{\Phi}}^{-}_{\mathbf{v}}[k,i]=e^{-i\bm{\omega}_{k}t_{i}}$
$\mathbf{\tilde{Z}_{v}}[k,i]=\bm{\tilde{\Phi}}_{\mathbf{v}}^{-}\odot\mathbf{Z_{v}}[k,i]$
Squared Residuals:
$\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}=\mathbf{1}_{m}(\mathbf{1}_{d}^{\top}\big|\mathbf{Z_{q}}\big|^{2})+\mathbf{E_{qk}}^{2}\odot(\mathbf{1}_{d}^{\top}\big|\mathbf{Z_{k}}\big|^{2})^{\top}\mathbf{1}_{m}^{\top}-2\mathbf{E_{qk}}\odot\mathrm{Re}(\mathbf{\tilde{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{k}})$
1exAttention matrix:
$\mathbf{L}=-\log\left(\mathbf{V}^{C}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}+\epsilon\right)\odot\mathbf{M}_{\text{causal}},$
$\mathbf{A}[i,j]=\text{softmax}_{j}\big(\tau_{s}\mathbf{L}[i,j]\big),\quad\mathbf{\hat{A}}=\mathbf{E_{v}}\odot\mathbf{A}$
1exOutput with value dynamics: $\mathbf{\bar{Z}_{v}}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}\odot(\mathbf{\tilde{Z}}_{\mathbf{v}}\mathbf{\hat{A}}^{\dagger})$
Residual connection: $\mathbf{\bar{Z}}=(1-\text{sigmoid}(\delta))\mathbf{Z_{v}}+\text{sigmoid}(\delta)\mathbf{\bar{Z}_{v}}$
Prediction: $\mathbf{Z}_{\mathbf{v}}^{+}=(e^{\alpha\Delta t}e^{\bm{\lambda}_{\Omega}\Delta t})\odot\mathbf{\bar{Z}_{v}}$
Output Projection: $\big(\operatorname{Re}(\mathbf{Z_{p}}),\operatorname{Im}(\mathbf{Z_{p}})\big)\leftarrow\mathcal{L}_{P}\big(\operatorname{Re}(\mathbf{Z}_{\mathbf{v}}^{+}),\operatorname{Im}(\mathbf{Z}_{\mathbf{v}}^{+})\big)$
Return: $\operatorname{Re}(\mathbf{Z_{p}})$ ;
// We discard $\operatorname{Im}(\mathbf{Z_{p}})$.
Algorithm 1 Adaptive Filter Attention (Isotropic; Single Head)
This defines a single neural network layer, which we can train with backpropagation, as usual.
In summary, this formulation modifies standard attention by including:
  1. 1.
    Complex-valued linear layers;
  2. 2.
    Complex-valued multiplicative positional encodings with learned parameters;
  3. 3.
    Scalar decay constants (one per head) with pair-wise uncertainty propagation via the closed-form solution of the differential Lyapunov equation;
  4. 4.
    A similarity function based on the Mahalanobis-distance;
  5. 5.
    A shared linear dynamics model for state estimation and prediction, to enforce model self-consistency.
A multi-head implementation of Algorithm 1 is available at the following: Github Repository. (This repository also contains an implementation of the tensor form discussed in Section 3.1.)
4.2 Basic Experiments
We tested our model (Algorithm 1) on a diagonalizable 2D linear system with zero-mean Gaussian process and measurement noise. For reproducibility, we used a fixed state matrix $\mathbf{A}=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1}$, where:
$\bm{\Lambda}=\begin{bmatrix}-0.1-i&0\\ 0&-0.1+i\end{bmatrix},\quad\mathbf{S}=\begin{bmatrix}1-i&1+i\\ 1&1\end{bmatrix},\quad\mathbf{S}^{-1}=\frac{1}{2}\begin{bmatrix}i&1-i\\ -i&1+i\end{bmatrix}$
We simulated this linear system starting from a random initial condition, over 1,000 time steps. We use simple Euler updates, with additive zero-mean Gaussian process noise scaled by the square root of the time step injected between time steps:
$\mathbf{x}_{t+\Delta t}=\mathbf{A}\mathbf{x}_{t}\,\Delta t+\mathbf{x}_{t}+\sqrt{\Delta t}\,\bm{\epsilon}_{t},\quad\bm{\epsilon}_{t}\sim\mathcal{N}(0,\sigma^{2})$
We sampled measurements at regular intervals (we chose every 10 time steps), added zero-mean Gaussian measurement noise, and then projected the data into an embedding dimension of $d_{e}=128$ (an effective embedding dimension of 256 since the real/imag stacking doubles the input dimension). We trained the model on one-step prediction, using only the noisy measurements, with the model having no access to the ground-truth trajectory. After learning, we projected the model’s output back into 2D to assess its performance.
The ground-truth and measured trajectories and the model predictions with a single layer of AFA after training are shown in Fig. 1.
Figure 1: Filter performance on different 2D systems: ground-truth trajectory (black), measured (blue), and predicted (red). (a) system with only measurement noise ($\sigma^{2}=0.0,\,\eta^{2}=1.0$). (b) system with both process noise and measurement noise ($\sigma^{2}=0.3,\,\eta^{2}=0.5$). (c) higher noise ($\sigma^{2}=0.5,\,\eta^{2}=2.0$).
To check that the model has learned the right dynamics, we compute "pulled-forward" estimates at four time points. As training progresses, the "pulled-forward" estimates converge to point clouds near the ground truth, as shown in Fig. 2.
Figure 2: Comparison of an AFA layer’s "pulled-forward" state estimates at different stages of training. The true trajectory is shown as a solid black line, and the pulled-forward estimates as colored point clouds. (a) State estimates early in training. (b) State estimates midway through training. (c) State estimates after training is complete.
While standard attention is also able to filter this noisy trajectory, it is not able to do so effectively in a single layer, and the resulting attention matrices lack the consistent and interpretable attention matrices of AFA. We compare the attention matrices from a trained model comprised of two layers of standard attention, with an additive positional embedding, to a trained model comprised of a single layer of AFA in Fig. 3. The AFA attention matrix displays a clear periodic structure, corresponding to pair-wise precisions propagated through the underlying oscillatory dynamics.
Figure 3: Attention matrices produced by training standard attention and AFA on a 2D LTI with process and measurement noise. (a) First layer of standard attention (b) Second layer of standard attention. (c) Single layer of AFA.
These experiments, along with animations of the predicted trajectories, pulled-forward estimates, and attention matrices during the course of training are available at the same Github link provided above.
Testing the model on real-world language and time series data will be the subject of future work.
5 Radial-Tangential Model
As we saw in Section 3, we were able to significantly simplify the adaptive filter by assuming that the dynamics decouple into separate magnitude and angular components (i.e. an isotropic plus skew-symmetric state matrix), and that the process and measurement noise are isotropic. Here, we will slightly relax these assumptions by assuming that the process and measurement noise also decompose into separate radial and tangential components, and derive a modification of the above algorithm.
5.1 Itô SDE with Radial–Tangential Dynamics and Noise
Consider the linear stochastic differential equation:
$d\mathbf{x}(t)=\mathbf{A}\mathbf{x}(t)dt+\mathbf{G}(t)d\mathbf{w}(t),\quad\mathbf{z}(t_{k})=\mathbf{C}\mathbf{x}(t_{k})+\mathbf{v}(t_{k}),$
Let $\mathbf{A}(t)=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1}$, with $\bm{\Lambda}=\alpha\mathbf{I}+\bm{\Lambda}_{\Omega}$, where $\bm{\Lambda}_{\Omega}\in i\mathbb{R}^{d\times d}$ is diagonal, and $\mathbf{G}(t)=\mathbf{S}\bm{\Lambda}_{B}(t)\mathbf{S}^{\dagger}$, $\mathbf{C}(t)=\lambda_{C}\mathbf{I}$. We can also write the eigenvalues as $\bm{\lambda}=\alpha\mathbf{1}+\bm{\omega}$, where $\alpha\in\mathbb{R}$, $\bm{\omega}=\text{diag}(\bm{\Lambda}_{\Omega})\in i\mathbb{R}^{d}$.
We transform the state and measurements into the eigenbasis:
$\mathbf{x}_{\mathbf{s}}(t)=\mathbf{S}^{-1}\mathbf{x}(t)\in\mathbb{C}^{d},\quad\mathbf{z}_{\mathbf{s}}(t_{k})=\mathbf{S}^{-1}\mathbf{z}(t_{k})\in\mathbb{C}^{d}.$
And we decompose the state and measurement into radial and tangential components:
$\mathbf{x}_{\mathbf{s}}(t)=m_{\mathbf{s}}(t)\mathbf{u}_{\mathbf{s}}(t),\quad\mathbf{z}_{\mathbf{s}}(t)=m_{\mathbf{zs}}(t)\mathbf{u}_{\mathbf{zs}}(t)\quad\text{where:}$
$m_{\mathbf{s}}(t)=\big|\mathbf{x}_{\mathbf{s}}(t)\big|_{2}=\sqrt{\mathbf{x}_{\mathbf{s}}^{\dagger}(t)\mathbf{x}_{\mathbf{s}}(t)}\in\mathbb{R},\quad\mathbf{u}_{\mathbf{s}}(t)=\mathbf{x}_{\mathbf{s}}(t)/\big|\mathbf{x}_{\mathbf{s}}(t)\big|_{2}\in\mathbb{S}^{d-1}.$
$m_{\mathbf{zs}}(t)=\big|\mathbf{z}_{\mathbf{s}}(t)\big|_{2}=\sqrt{\mathbf{z}_{\mathbf{s}}^{\dagger}(t)\mathbf{z}_{\mathbf{s}}(t)}\in\mathbb{R},\quad\mathbf{u}_{\mathbf{zs}}(t)=\mathbf{z}_{\mathbf{s}}(t)/\big|\mathbf{z}_{\mathbf{s}}(t)\big|_{2}\in\mathbb{S}^{d-1}.$
The dynamics in the eigenbasis are:
$d\mathbf{x_{s}}(t)=\bm{\Lambda}\mathbf{x_{s}}(t)dt+\bm{\Lambda}_{B}d\mathbf{w_{s}}(t)$
We assume independent process noise acting in the radial and tangential directions in the eigenbasis:
$d\mathbf{x}_{\mathbf{s}}=(\alpha\mathbf{I}+\bm{\Lambda}_{\Omega})\mathbf{x}_{\mathbf{s}}\,dt+\sigma_{r}\mathbf{u}_{\mathbf{s}}\,dw_{r}+\sigma_{t}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,d\mathbf{w}_{t},$
where $\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}):=\mathbf{I}-\mathbf{u}_{\mathbf{s}}\mathbf{u}_{\mathbf{s}}^{\dagger}$ projects onto the tangent space, $\mathbf{P}_{R}(\mathbf{u}_{\mathbf{s}}):=\mathbf{u}_{\mathbf{s}}\mathbf{u}_{\mathbf{s}}^{\dagger}$ projects into the radial direction, and $w_{r},\mathbf{w}_{t}$ are independent Wiener processes.
Hence, the process noise covariance is:
$\bm{\Lambda}_{Q}(t)=\bm{\Lambda}_{B}(t)\bm{\Lambda}_{B}(t)^{\dagger}=\sigma_{r}^{2}\mathbf{u_{s}}(t)\mathbf{u_{s}}(t)^{\dagger}+\sigma_{t}^{2}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}).$
The process noise rotates with the dynamics when viewed from the global frame, which corresponds to a fixed anisotropic noise applied in the token’s frame.
Applying Itô’s lemma for $m_{\mathbf{s}}=\sqrt{s}$,
$dm_{\mathbf{s}}=f^{\prime}(s)\,ds+\tfrac{1}{2}f^{\prime\prime}(s)\,d\langle s\rangle=\frac{1}{2m_{\mathbf{s}}}\,ds-\frac{1}{8m_{\mathbf{s}}^{3}}\,d\langle s\rangle.$
Let $s=\mathbf{x}_{\mathbf{s}}^{\dagger}\mathbf{x}_{\mathbf{s}}$. Then:
$dm_{\mathbf{s}}=\frac{1}{2m_{\mathbf{s}}}d(\mathbf{x}_{\mathbf{s}}^{\dagger}\mathbf{x}_{\mathbf{s}})=\frac{1}{2m_{\mathbf{s}}}\big(d\mathbf{x}_{\mathbf{s}}^{\dagger}\mathbf{x}_{\mathbf{s}}+\mathbf{x}_{\mathbf{s}}^{\dagger}d\mathbf{x}_{\mathbf{s}}+d\mathbf{x}_{\mathbf{s}}^{\dagger}d\mathbf{x}_{\mathbf{s}}\big).$
Substituting the SDE,
$\mathbf{x}_{\mathbf{s}}^{\dagger}d\mathbf{x}_{\mathbf{s}}=\mathbf{x}_{\mathbf{s}}^{\dagger}\Big((\alpha\mathbf{I}+\bm{\Lambda}_{\Omega})\mathbf{x}_{\mathbf{s}}\,dt+\sigma_{r}\mathbf{u}_{\mathbf{s}}\,dw_{r}+\sigma_{t}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,d\mathbf{w}_{t}\Big)$
$\mathbf{x}_{\mathbf{s}}^{\dagger}(\alpha\mathbf{I}+\bm{\Lambda}_{\Omega})\mathbf{x}_{\mathbf{s}}\,dt+\sigma_{r}m_{\mathbf{s}}\,dw_{r}$
so:
$\mathbf{x}_{\mathbf{s}}^{\dagger}d\mathbf{x}_{\mathbf{s}}+d\mathbf{x}_{\mathbf{s}}d\mathbf{x}_{\mathbf{s}}^{\dagger}=2\alpha m_{\mathbf{s}}^{2}dt+2\sigma_{r}m_{\mathbf{s}}\,dw_{r}$
The quadratic term is: $d\mathbf{x}_{\mathbf{s}}^{\dagger}d\mathbf{x}_{\mathbf{s}}=\sigma_{r}^{2}dt+\sigma_{t}^{2}\mathbf{u}_{\mathbf{s}}^{\dagger}\mathbf{P}_{T}\mathbf{u}_{\mathbf{s}}dt=\sigma_{r}^{2}dt,$ since $\mathbf{u}_{\mathbf{s}}^{\dagger}\mathbf{P}_{T}\mathbf{u}_{\mathbf{s}}=0$.
Altogether, we have:
$dm_{\mathbf{s}}=\frac{1}{2m_{\mathbf{s}}}\big(\alpha m_{\mathbf{s}}^{2}dt+\sigma_{r}m_{\mathbf{s}}dw_{r}+\sigma_{r}^{2}dt\big)-\frac{1}{8m_{\mathbf{s}}^{3}}(2\sigma_{r}m_{\mathbf{s}})^{2}dt$
$\alpha m_{\mathbf{s}}\,dt+\sigma_{r}\,dw_{r}+\frac{1}{2m_{\mathbf{s}}}\sigma_{r}^{2}dt-\frac{1}{2m_{\mathbf{s}}}\sigma_{r}^{2}dt$
Hence the Itô correction cancels, and the radial SDE simplifies to:
$dm_{\mathbf{s}}=\alpha m_{\mathbf{s}}\,dt+\sigma_{r}\,dw_{r}.$
For the unit direction $\mathbf{u}_{\mathbf{s}}=\mathbf{x}_{\mathbf{s}}/m_{\mathbf{s}}$, the projected Itô SDE is:
$d\mathbf{u}_{\mathbf{s}}=\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,d\mathbf{x}_{\mathbf{s}}/m_{\mathbf{s}}.$
Substituting $d\mathbf{x}_{\mathbf{s}}$ and using the fact that $\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\mathbf{u}_{\mathbf{s}}=\mathbf{0}$ gives:
$d\mathbf{u}_{\mathbf{s}}=\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,\bm{\Lambda}_{\Omega}\mathbf{u}_{\mathbf{s}}\,dt+\frac{\sigma_{t}}{m_{\mathbf{s}}}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,d\mathbf{w}_{t}.$
Finally, we assume that measurement noise in the eigenbasis also has a radial–tangential decomposition:
$\mathbf{z}_{\mathbf{s}}(t_{k})=\lambda_{C}\mathbf{x}_{\mathbf{s}}(t_{k})+v_{r}^{(k)}\mathbf{u}_{\mathbf{s}}(t_{k})+\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))\mathbf{v}_{t}^{(k)}$
$v_{r}^{(k)}\sim\mathcal{N}(0,\eta_{r}^{2}),\quad\mathbf{v}_{t}^{(k)}\sim\mathcal{N}(\mathbf{0},\eta_{t}^{2}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))).$
So, the measurement noise covariance is:
$\bm{\Lambda}_{R}(t)=\eta_{r}^{2}\mathbf{u_{s}}(t)\mathbf{u_{s}}(t)^{\dagger}+\eta_{t}^{2}\big(\mathbf{I}-\mathbf{u_{s}}(t)\mathbf{u_{s}}(t)^{\dagger}\big).$
The radial measurement is obtained by projecting $\mathbf{z}_{\mathbf{s}}(t_{k})$ onto the state direction $\mathbf{u}_{\mathbf{s}}(t_{k})$:
$m_{\mathbf{zs}}(t_{k})$ $=\mathbf{u}_{\mathbf{s}}^{\dagger}(t_{k})\Big(\lambda_{C}\mathbf{x}_{\mathbf{s}}(t_{k})+v_{r}^{(k)}\mathbf{u}_{\mathbf{s}}(t_{k})+\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))\mathbf{v}_{t}^{(k)}\Big)$
$=\lambda_{C}\mathbf{u}_{\mathbf{s}}^{\dagger}\mathbf{x}_{\mathbf{s}}+v_{r}^{(k)}\mathbf{u}_{\mathbf{s}}^{\dagger}\mathbf{u}_{\mathbf{s}}+\mathbf{u}_{\mathbf{s}}^{\dagger}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\mathbf{v}_{t}^{(k)}$
$=\lambda_{C}m_{\mathbf{s}}(t_{k})+v_{r}^{(k)}.$
Hence, the radial measurement is 1D Gaussian:
$m_{\mathbf{zs}}(t_{k})\sim\mathcal{N}\big(\lambda_{C}m_{\mathbf{s}}(t_{k}),\eta_{r}^{2}\big).$
The tangential (unit-direction) measurement is:
$\mathbf{u}_{\mathbf{zs}}=\exp_{\mathbf{u}_{\mathbf{s}}}\!\Big(\mathbf{v}_{t}/m_{\mathbf{s}}\Big)$
where $\exp_{\mathbf{u}}$ is the spherical exponential map, which maps a vector in the tangent space of the sphere at $\mathbf{u}\in\mathbb{S}^{d-1}$ to a point on the sphere along the geodesic defined by that tangent vector $\mathbf{v}\in T_{\mathbf{u}}\mathbb{S}^{d-1}$, i.e. $\mathbf{u}^{\top}\mathbf{v}=0$. The spherical exponential map is defined as:
$\exp_{\mathbf{u}}:T_{\mathbf{u}}\mathbb{S}^{d-1}\to\mathbb{S}^{d-1},\quad\exp_{\mathbf{u}}(\mathbf{v})=\cos\big(\big|\mathbf{v}\big|\big.)\,\mathbf{u}+\sin\big(\big|\mathbf{v}\big|\big.)\,\frac{\mathbf{v}}{\big|\mathbf{v}\big|},$
For small tangential noise ($\eta_{t}\ll m_{\mathbf{s}}$), we can linearize on the tangent plane:
$\mathbf{u}_{\mathbf{zs}}(t_{k})\approx\mathbf{u}_{\mathbf{s}}(t_{k})+\frac{1}{m_{\mathbf{s}}(t_{k})}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))\mathbf{v}_{t}^{(k)}.$
(The $1/m_{\mathbf{s}}$ factor comes from projecting the noise onto the unit sphere (radial normalization), and $\mathbf{P}_{T}\mathbf{v}_{t}^{(k)}$ ensures the noise is tangential.) Hence, the tangential distribution is approximately Gaussian:
$\mathbf{u}_{\mathbf{zs}}(t_{k})\sim\mathcal{N}\Big(\mathbf{u}_{\mathbf{s}}(t_{k}),\frac{\eta_{t}^{2}}{m_{\mathbf{s}}^{2}(t_{k})}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))\Big).$
In summary, our radial-tangential SDE (RT-SDE) model is:
$d\mathbf{x}_{\mathbf{s}}$ $=(\alpha\mathbf{I}+\bm{\Lambda}_{\Omega})\mathbf{x}_{\mathbf{s}}\,dt+\sigma_{r}\mathbf{u}_{\mathbf{s}}\,dw_{r}+\sigma_{t}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,d\mathbf{w}_{t},$
$\mathbf{z}_{\mathbf{s}}(t_{k})$ $\approx\lambda_{C}\mathbf{x}_{\mathbf{s}}(t_{k})+v_{r}^{(k)}\mathbf{u}_{\mathbf{s}}(t_{k})+\frac{1}{m_{\mathbf{s}}(t_{k})}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))\mathbf{v}_{t}^{(k)},$
$v_{r}^{(k)}$ $\sim\mathcal{N}(0,\eta_{r}^{2}),\quad\mathbf{v}_{t}^{(k)}\sim\mathcal{N}(\mathbf{0},\eta_{t}^{2}\mathbf{I}).$
Or, in the decoupled radial-tangential coordinates:
$dm_{\mathbf{s}}$ $=\alpha m_{\mathbf{s}}\,dt+\sigma_{r}\,dw_{r},$
$d\mathbf{u}_{\mathbf{s}}$ $=\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,\bm{\Lambda}_{\Omega}\mathbf{u}_{\mathbf{s}}\,dt+\frac{\sigma_{t}}{m_{\mathbf{s}}}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}})\,d\mathbf{w}_{t},$
$m_{\mathbf{zs}}(t_{k})$ $=\lambda_{C}\,m_{\mathbf{s}}(t_{k})+v_{r}^{(k)},$ $v_{r}^{(k)}$ $\sim\mathcal{N}(0,\eta_{r}^{2}),$
$\mathbf{u}_{\mathbf{zs}}(t_{k})$ $\approx\mathbf{u}_{\mathbf{s}}(t_{k})+\frac{1}{m_{\mathbf{s}}(t_{k})}\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{k}))\mathbf{v}_{t}^{(k)},$ $\mathbf{v}_{t}^{(k)}$ $\sim\mathcal{N}(\mathbf{0},\eta_{t}^{2}\mathbf{I}).$
Remark: Polar vs Cartesian parameterizations.
Although the radial and angular dynamics decouple, operating in polar coordinates $(r,\theta)$ makes the MLE nonlinear. Even if $\mathbf{x}(t)$ evolves linearly, the radial and angular dynamics become state-dependent, and the noise covariance depends on the current state. This makes the log-likelihood non-quadratic, so the MLE cannot be expressed in closed form and requires iterative optimization. Cartesian coordinates are therefore preferred, as they preserve linear dynamics with Gaussian noise and allow exact, closed-form propagation of covariances and MLE computation.
5.2 Propagation of Uncertainty through the RT-SDE
If $\mathbf{Q}$ is a function of $t$, the state evolution (for the causal case) is governed by:
$\mathbf{x}(t_{i})=e^{\mathbf{A}(t_{i}-t_{j})}\mathbf{x}(t_{j})+\int_{t_{j}}^{t_{i}}e^{\mathbf{A}(t_{i}-s)}\mathbf{G}(s)d\mathbf{w}(s)$
Letting $\tau=s-t_{j}$, this is:
$\mathbf{x}(t_{i})=e^{-\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{j})+\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}(t_{j}+\tau)d\mathbf{w}(\tau)\right).$
So the backward-propagated covariance becomes:
$\mathbf{V}_{B}(t_{i},t_{j})=\int_{0}^{|\Delta t_{ij}|}e^{-\mathbf{A}\tau}\mathbf{Q}(t_{j}+\tau)e^{-\mathbf{A}^{\top}\tau}\,d\tau.$
Plugging in $\mathbf{A}=\mathbf{S}(\alpha\mathbf{I}+\bm{\Lambda}_{\Omega})\mathbf{S}^{-1}$,
$\mathbf{V}_{B}(t_{i},t_{j})=\mathbf{S}\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})\mathbf{S}^{-1},\quad\text{where:}\quad\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})=\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}e^{-\bm{\Lambda}_{\Omega}\tau}\bm{\Lambda}_{Q}(t_{j}+\tau)e^{-\bm{\Lambda}_{\Omega}^{\dagger}\tau}\,d\tau$
where the process noise is:
$\bm{\Lambda}_{Q}(t)=\sigma_{r}^{2}\mathbf{u_{s}}(t)\mathbf{u_{s}}(t)^{\dagger}+\sigma_{t}^{2}\big(\mathbf{I}-\mathbf{u_{s}}(t)\mathbf{u_{s}}(t)^{\dagger}\big).$
Note that $\mathbf{u_{s}}(t_{j}+\tau)=e^{\bm{\Lambda}_{\Omega}\tau}\mathbf{u_{s}}(t_{j})$. Hence, the rotation cancels inside the integral:
$\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})=\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}e^{-\bm{\Lambda}_{\Omega}\tau}\Big(\sigma_{r}^{2}\mathbf{u_{s}}(t_{j}+\tau)\mathbf{u_{s}}(t_{j}+\tau)^{\dagger}+\sigma_{t}^{2}\big(\mathbf{I}-\mathbf{u_{s}}(t_{j}+\tau)\mathbf{u_{s}}(t_{j}+\tau)^{\dagger}\big)\Big)e^{-\bm{\Lambda}_{\Omega}^{\dagger}\tau}\,d\tau$
$\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}\Big(\sigma_{r}^{2}\mathbf{u_{s}}(t_{j})\mathbf{u_{s}}(t_{j})^{\dagger}+\sigma_{t}^{2}\big(\mathbf{I}-\mathbf{u_{s}}(t_{j})\mathbf{u_{s}}(t_{j})^{\dagger}\big)\Big)\,d\tau$
$\bm{\Lambda}_{Q}(t_{j})\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha s}\,ds=\kappa_{B}(|\Delta t_{ij}|)\,\bm{\Lambda}_{Q}(t_{j}),$
where:
$\kappa_{B}(\Delta t)=\begin{cases}\dfrac{e^{-2\alpha\Delta t}-1}{-2\alpha},&\alpha\neq 0,\\ \Delta t,&\alpha=0.\end{cases}$
The forward-propagated covariance is then:
$\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})=e^{\bm{\Lambda}|\Delta t_{ij}|}\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})e^{\bm{\Lambda}^{\dagger}|\Delta t_{ij}|}$
$e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\kappa_{F}(|\Delta t_{ij}|)e^{-\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{\Lambda}_{Q}(t_{j})e^{-\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}$
where:
$\kappa_{F}(\Delta t)=e^{2\alpha\Delta t}\kappa_{B}(\Delta t)=\begin{cases}\dfrac{1-e^{2\alpha\Delta t}}{-2\alpha},&\alpha\neq 0,\\ \Delta t,&\alpha=0.\end{cases}$
(Note: we prefer to keep things in terms of $\kappa_{F}(\Delta t)$ rather than $\kappa_{B}(\Delta t)$, because the former is bounded for causal models, while the latter is unbounded.)
Plugging in $\bm{\Lambda}_{Q}(t_{j})$:
$\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})=\kappa_{F}(|\Delta t_{ij}|)e^{-\bm{\Lambda}_{\Omega}\Delta t_{ij}}\Big(\sigma_{r}^{2}\mathbf{u_{s}}(t_{j})\mathbf{u_{s}}(t_{j})^{\dagger}+\sigma_{t}^{2}\big(\mathbf{I}-\mathbf{u_{s}}(t_{j})\mathbf{u_{s}}(t_{j})^{\dagger}\big)\Big)e^{-\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}$
$\kappa_{F}(|\Delta t_{ij}|)\Big(\sigma_{r}^{2}\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}+\sigma_{t}^{2}\big(\mathbf{I}-\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}\big)\Big)$
$\kappa_{F}(|\Delta t_{ij}|)\Big(\sigma_{t}^{2}\mathbf{I}+(\sigma_{r}^{2}-\sigma_{t}^{2})\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}\Big)$
Finally,
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\big|\lambda_{C}\big|^{2}\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})+e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{R}(t_{j})e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\big|\lambda_{C}\big|^{2}\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})+e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{R}(t_{j})e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\big|\lambda_{C}\big|^{2}\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})+e^{-2\alpha\Delta t_{ij}}\bm{\Lambda}_{R}(t_{i})$
Altogether, the propagated measurement covariance is:
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})$ $=\Big(\big|\lambda_{C}\big|^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma_{t}^{2}+e^{-2\alpha\Delta t_{ij}}\eta_{t}^{2}\Big)\mathbf{I}$
$\quad+\Big(\big|\lambda_{C}\big|^{2}\kappa_{F}(|\Delta t_{ij}|)(\sigma_{r}^{2}-\sigma_{t}^{2})+e^{-2\alpha\Delta t_{ij}}(\eta_{r}^{2}-\eta_{t}^{2})\Big)\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$
$\sigma_{Vr}^{2}(|\Delta t_{ij}|)\mathbf{I}+\Big(\sigma_{Vr}^{2}(|\Delta t_{ij}|)-\sigma_{Vt}^{2}(|\Delta t_{ij}|)\Big)\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$
where:
$\sigma_{Vr}^{2}(|\Delta t_{ij}|)$ $=\big|\lambda_{C}\big|^{2}\,\kappa_{F}(|\Delta t_{ij}|)\,\sigma_{r}^{2}+e^{-2\alpha\Delta t_{ij}}\,\eta_{r}^{2},$
$\sigma_{Vt}^{2}(|\Delta t_{ij}|)$ $=\big|\lambda_{C}\big|^{2}\,\kappa_{F}(|\Delta t_{ij}|)\,\sigma_{t}^{2}+e^{-2\alpha\Delta t_{ij}}\,\eta_{t}^{2}.$
Rearranging, this is:
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\sigma_{Vt}^{2}(|\Delta t_{ij}|)(\mathbf{I}-\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger})+\sigma_{Vr}^{2}(|\Delta t_{ij}|)\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$
If $\sigma_{r}=\sigma_{t}:=\sigma$ and $\eta_{r}=\eta_{t}:=\eta$, then this simplifies to:
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\Big(\big|\lambda_{C}\big|^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma^{2}+e^{-2\alpha\Delta t_{ij}}\eta^{2}\Big)\mathbf{I}$
which was exactly the result for the case of isotropic noise.
Since the propagated covariance is of the form $\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=a\mathbf{I}+b\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$, where $a=\sigma_{Vt}^{2},\,b=\sigma_{Vr}^{2}-\sigma_{Vt}^{2}$, i.e. a rank-1 correction of identity, we can invert it with the Sherman-Morrison formula:
$\bm{\Lambda}_{\mathbf{P}}^{C}=\big(\bm{\Lambda}_{\mathbf{V}}^{C}\big)^{-1}=\frac{1}{a}\Big(\mathbf{I}+\frac{b}{a}\mathbf{u_{s}}\mathbf{u_{s}}^{\dagger}\Big)^{-1}=\frac{1}{a}\Big(\mathbf{I}-\frac{b}{a}\frac{\mathbf{u_{s}}\mathbf{u_{s}}^{\dagger}}{1+\frac{b}{a}\mathbf{u_{s}}^{\dagger}\mathbf{u_{s}}}\Big)$
$\frac{1}{a}\Big(\mathbf{I}-\mathbf{u_{s}}\mathbf{u_{s}}^{\dagger}\Big)+\Big(-\frac{b}{a(a+b)}+\frac{1}{a}\Big)\mathbf{u_{s}}\mathbf{u_{s}}^{\dagger},$
$\frac{1}{a}\Big(\mathbf{I}-\mathbf{u_{s}}\mathbf{u_{s}}^{\dagger}\Big)+\Big(\frac{1}{a+b}\Big)\mathbf{u_{s}}\mathbf{u_{s}}^{\dagger},$
Hence:
$\bm{\Lambda}_{\mathbf{P}}^{C}(t_{i},t_{j})=\frac{1}{\sigma_{Vt}^{2}(|\Delta t_{ij}|)}\Big(\mathbf{I}-\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}^{\dagger}(t_{i})\Big)+\frac{1}{\sigma_{Vr}^{2}(|\Delta t_{ij}|)}\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}^{\dagger}(t_{i}).$
(which we could also have obtained by just inverting each term, since they are along orthogonal subspaces).
5.3 Radial–Tangential Kalman IRLS
We now seek to estimate the latent states $\{\mathbf{z}_{i}^{C}\}_{i=1}^{m}\in\mathbb{R}^{d}$ from the observations $\{\mathbf{z}_{i}\}_{i=1}^{m}\in\mathbb{R}^{d}$, assuming the latent states were generated by an RT-SDE.
The joint log-likelihood of the measurements factors into radial and tangential parts:
$\log p\big(\{\mathbf{z}_{s}(t_{j})\}\mid m_{s}(\cdot),\mathbf{u}_{s}(\cdot)\big)=\sum_{j}\log p\!\left(m_{\mathbf{z}s}(t_{j})\mid m_{s}(t_{j})\right)\,+\,\sum_{j}\log p\!\left(\mathbf{u}_{\mathbf{z}s}(t_{j})\mid\mathbf{u}_{s}(t_{j}),m_{s}(t_{j})\right).$
$\text{where:}\quad m_{\mathbf{zs}}(t_{j})\sim\mathcal{N}\big(\lambda_{C}m_{\mathbf{s}}(t_{j}),\eta_{r}^{2}\big),\quad\mathbf{u}_{\mathbf{zs}}(t_{j})\sim\mathcal{N}\Big(\mathbf{u}_{\mathbf{s}}(t_{j}),\frac{\eta_{t}^{2}}{m_{\mathbf{s}}^{2}(t_{j})}\,\mathbf{P}_{T}(\mathbf{u}_{\mathbf{s}}(t_{j}))\Big).$
Hence, we can treat the estimation problem as two decoupled weighted least squares problems: one for magnitude and one for direction. This motivates the following iteration:
Let $\mathbf{z}_{\mathbf{s}i}:=\mathbf{z}_{\mathbf{s}}(t_{i})$. The initial condition is: $\mathbf{z}_{\mathbf{s}i}^{(0)}=\mathbf{z}_{\mathbf{s}i}$.
At iteration $k$, we compute the magnitudes and directions:
$m_{\mathbf{zs}i}^{(k)}=\big|\mathbf{z}_{\mathbf{s}i}^{(k)}\big|_{2},\quad\mathbf{u}_{\mathbf{zs}i}^{(k)}=\mathbf{z}_{\mathbf{s}i}^{(k)}/\big|\mathbf{z}_{\mathbf{s}i}^{(k)}\big|_{2}.$
And the estimates:
$\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}=e^{-\bm{\lambda}\Delta t_{ij}}\mathbf{z_{s}}_{j}^{(k)}=\Big(e^{-\alpha\Delta t_{ij}}m_{\mathbf{zs}j}^{(k)}\Big)\Big(e^{-\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{u}_{\mathbf{zs}j}^{(k)}\Big)$
We showed above that:
$\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}\sim\mathcal{N}\Big(\mathbf{z}_{\mathbf{s}i}^{C},\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})\Big)$
where:
$\begin{gathered}\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\sigma_{Vt}^{2}(|\Delta t_{ij}|)\big(\mathbf{I}-\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}\big)+\sigma_{Vr}^{2}(|\Delta t_{ij}|)\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger},\\[5.69054pt] \sigma_{Vr}^{2}(|\Delta t_{ij}|)=\big|\lambda_{C}\big|^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma_{r}^{2}+e^{-2\alpha\Delta t_{ij}}\eta_{r}^{2},\\[2.84526pt] \sigma_{Vt}^{2}(|\Delta t_{ij}|)=\big|\lambda_{C}\big|^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma_{t}^{2}+e^{-2\alpha\Delta t_{ij}}\eta_{t}^{2}.\end{gathered}$
It follows that:
$(\mathbf{u}_{\mathbf{s}i}^{(k)})^{\dagger}\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}\sim\mathcal{N}\left((\mathbf{u}_{\mathbf{s}i}^{(k)})^{\dagger}\mathbf{z}_{\mathbf{s}i}^{C},\sigma_{Vr}^{2}(|\Delta t_{ij}|)\right),\quad\mathbf{P}_{T\mathbf{zs}i}\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}\sim\mathcal{N}\left(\mathbf{P}_{T\mathbf{zs}i}\mathbf{z}_{\mathbf{s}i}^{C},\sigma_{Vt}^{2}(|\Delta t_{ij}|)\mathbf{P}_{T\mathbf{zs}i}^{(k)}\right)$
We do not have access to the true tangent projection matrix at $\mathbf{x}_{\mathbf{s}i}$, $\mathbf{P}_{T\mathbf{s}i}:=\mathbf{I}-\mathbf{u}_{\mathbf{s}i}\mathbf{u}_{\mathbf{s}i}^{\dagger}$, but we can approximate it using the tangent projection matrix at $\mathbf{z}_{\mathbf{s}i}^{(k)}/\lambda_{C}$:
$\mathbf{P}_{T\mathbf{zs}i}^{(k)}\Big(\mathbf{z}_{\mathbf{s}i}^{(k)}/\lambda_{C}\Big):=\mathbf{I}-\frac{\mathbf{z}_{\mathbf{s}i}^{(k)}\Big(\mathbf{z}_{\mathbf{s}i}^{(k)}\Big)^{\dagger}/\big|\lambda_{C}\big|^{2}}{\Big(\mathbf{z}_{\mathbf{s}i}^{(k)}\Big)^{\dagger}\Big(\mathbf{z}_{\mathbf{s}i}^{(k)}\Big)/\big|\lambda_{C}\big|^{2}}=\mathbf{I}-\mathbf{u}_{\mathbf{zs}i}^{(k)}(\mathbf{u}_{\mathbf{zs}i}^{(k)})^{\dagger}.$
Define separate radial and tangential estimates:
$\hat{m}_{\mathbf{zs}ij}^{(k)}:=(\mathbf{u}_{\mathbf{zs}i}^{(k)})^{\dagger}\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)},\qquad\mathbf{\hat{v}}_{\mathbf{zs}ij}^{(k)}:=\mathbf{P}_{T\mathbf{zs}i}^{(k)}\mathbf{\hat{z}}_{\mathbf{zs}ij}^{(k)}.$
We assume:
$\hat{m}_{\mathbf{zs}ij}^{(k)}\overset{\text{approx}}{\sim}\mathcal{N}\left((\mathbf{u}_{\mathbf{zs}i}^{(k)})^{\dagger}\mathbf{z}_{\mathbf{s}i}^{C},\sigma_{Vr}^{2}(|\Delta t_{ij}|)\right),\quad\mathbf{\hat{v}}_{\mathbf{zs}ij}^{(k)}\overset{\text{approx}}{\sim}\mathcal{N}\left(\mathbf{P}_{T\mathbf{zs}i}^{(k)}\mathbf{z}_{\mathbf{s}i}^{C},\sigma_{Vt}^{2}(|\Delta t_{ij}|)\mathbf{P}_{T\mathbf{zs}i}^{(k)}\right)$
And define corresponding radial and tangential residuals:
$\mathbf{r}_{\mathbf{zs}ij}=\mathbf{z}_{\mathbf{s}i}^{(k)}-\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)},\quad\ell_{\mathbf{zs}ij}^{(k)}:=(\mathbf{u}_{\mathbf{zs}i}^{(k)})^{\dagger}\mathbf{r}_{\mathbf{zs}ij},\quad\mathbf{d}_{\mathbf{zs}ij}^{(k)}:=\mathbf{P}_{T\mathbf{zs}i}^{(k)}\mathbf{r}_{\mathbf{zs}ij}$
where:
$\ell_{\mathbf{zs}ij}^{(k)}\overset{\text{approx}}{\sim}\mathcal{N}\left(0,\sigma_{Vr}^{2}(|\Delta t_{ij}|)\right),\quad\mathbf{d}_{\mathbf{zs}ij}^{(k)}\overset{\text{approx}}{\sim}\mathcal{N}\left(\mathbf{0},\sigma_{Vt}^{2}(|\Delta t_{ij}|)\mathbf{P}_{T\mathbf{zs}i}^{(k)}\right)$
Defining robust weights:
$w_{r,ij}^{(k)}=\frac{1}{1+(\ell_{ij}^{(k)})^{2}/(\nu_{r}\sigma_{Vr,ij}^{2})},\quad w_{t,ij}^{(k)}=\frac{1}{1+(\mathbf{d}_{\mathbf{zs}ij}^{(k)})^{\dagger}\mathbf{d}_{\mathbf{zs}ij}^{(k)}/(\nu_{t}\sigma_{Vt,ij}^{2})}.$
the IRLS update for the magnitude is:
$\bar{m}_{\mathbf{zs}i}^{(k)}:=\frac{\sum_{j\leq i}\big(w_{r,ij}^{(k)}/\sigma_{Vr,ij}^{2}\big)\hat{m}_{\mathbf{zs}ij}^{(k)}}{\sum_{j\leq i}\big(w_{r,ij}^{(k)}/\sigma_{Vr,ij}^{2}\big)}$
And we can update the magnitude with the MLE:
$m_{\mathbf{zs}i}^{(k+1)}=\bar{m}_{\mathbf{zs}i}^{(k)}$
Letting $\mathbf{\Sigma}_{Tij}^{+(k)}=\mathbf{P}_{T\mathbf{zs}i}^{(k)}/\sigma_{Vt,ij}^{2}$, we compute a tangent-space precision-weighted average:
$\mathbf{\bar{v}}_{i}^{(k)}:=\left(\sum_{j\leq i}w_{t,ij}^{(k)}\mathbf{\Sigma}_{Tij}^{+(k)}\right)^{-1}\sum_{j\leq i}w_{t,ij}^{(k)}\mathbf{\Sigma}_{Tij}^{+(k)}\mathbf{\hat{v}}_{\mathbf{zs}ij}^{(k)}$
$\left(\sum_{j\leq i}w_{t,ij}^{(k)}/\sigma_{Vt,ij}^{2}\right)^{-1}\sum_{j\leq i}\big(w_{t,ij}^{(k)}/\sigma_{Vt,ij}^{2}\big)\mathbf{\hat{v}}_{\mathbf{zs}ij}^{(k)}$
We can then update the full state estimate with a simple first-order update:
$\mathbf{z}_{\mathbf{s}i}^{(k+1)}=m_{\mathbf{zs}i}^{(k+1)}\mathbf{u}_{\mathbf{zs}i}^{(k)}+\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}$
We can do better using the (spherical) exponential map. Let:
$\theta_{i}^{(k)}:=\big|\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}\big|,\qquad\mathbf{a}_{i}^{(k)}:=\frac{\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}}{\big|\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}\big|},$
The exact geodesic step is:
$\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\begin{cases}\cos\theta_{i}^{(k)}\,\mathbf{u}_{\mathbf{zs}i}^{(k)}+\sin\theta_{i}^{(k)}\,\mathbf{a}_{i}^{(k)},&\text{if }\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}\neq 0,\\ \mathbf{u}_{\mathbf{zs}i}^{(k)},&\text{if }\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}=0.\end{cases}$
A first-order approximation is:
$\mathbf{y}_{\mathbf{zs}i}^{(k+1)}=\mathbf{u}_{\mathbf{zs}i}^{(k)}+\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}/m_{\mathbf{zs}i}^{(k)},\qquad\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\frac{\mathbf{y}_{\mathbf{zs}i}^{(k+1)}}{\big|\mathbf{y}_{\mathbf{zs}i}^{(k+1)}\big|},$
Or, equivalently:
$\mathbf{y}_{\mathbf{zs}i}^{(k+1)}=\mathbf{z}_{\mathbf{s}i}^{(k)}+\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)},\qquad\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\frac{\mathbf{y}_{\mathbf{zs}i}^{(k+1)}}{\big|\mathbf{y}_{\mathbf{zs}i}^{(k+1)}\big|},$
Having updated the magnitude and direction, set:
$\mathbf{z}_{\mathbf{s}i}^{(k+1)}=m_{\mathbf{zs}i}^{(k+1)}\,\mathbf{u}_{\mathbf{zs}i}^{(k+1)}.$
Or we can use a conservative update:
$\mathbf{z}_{\mathbf{s}i}^{(k+1)}=\gamma\,\mathbf{z}_{\mathbf{s}i}^{(k)}+(1-\gamma)\,m_{\mathbf{zs}i}^{(k+1)}\mathbf{u}_{\mathbf{zs}i}^{(k+1)}.$
5.4 Algorithm
In this section, we present a vectorized implementation of the RT–Kalman IRLS update. The algorithm combines scalar precision-weighted updates for radial components with a separate precision-weighted update followed by a geodesic step for tangential components.
Input sequence: $\mathbf{Z}_{r}\in\mathbb{R}^{d\times m}$
Complex linear layers: $\mathcal{L}_{Q},\mathcal{L}_{K},\mathcal{L}_{V},\mathcal{L}_{P}$
Parameters: Dynamics and noise $\{\alpha,\bm{\omega},\sigma_{r},\sigma_{t},\eta_{r},\eta_{t}\}$; temperatures $\{\tau_{r},\tau_{t}\}$; timestep $\Delta t$; residual weight $\delta$. Define $t_{i}=\Delta t\cdot i;$$t_{j}=\Delta t\cdot j.$
Constants: stability $\epsilon$, causal mask $\mathbf{M}_{\text{causal}}$
1exComplex-valued projections:
Lift inputs to complex form: $\mathbf{Z}=[\mathbf{Z}_{r},0]^{\top}$
$\big(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big)\leftarrow\mathcal{L}_{V}(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z}))$
$\big(\operatorname{Re}(\mathbf{Z_{q}}),\operatorname{Im}(\mathbf{Z_{q}})\big)\leftarrow\mathcal{L}_{Q}(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}}))$
$\big(\operatorname{Re}(\mathbf{Z_{k}}),\operatorname{Im}(\mathbf{Z_{k}})\big)\leftarrow\mathcal{L}_{K}(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}}))$
Exponential kernels:
$\mathcal{K}_{\alpha}[\tau]=\mathbf{1}_{\{\tau>0\}}\odot e^{-\alpha^{2}\Delta t\tau}+\mathbf{1}_{\{\tau\leq 0\}},\quad\mathcal{K}_{\bm{\omega}}[\tau]=\mathbf{1}_{\{\tau>0\}}\odot e^{\bm{\omega}\Delta t\tau}+\mathbf{1}_{\{\tau\leq 0\}}$
Decay kernel: $\kappa[\tau]=\frac{1-\mathcal{K}_{\alpha}^{2}[\tau]+\epsilon}{2\alpha^{2}+\epsilon}\mathbf{1}_{\{\alpha^{2}>\epsilon\}}+\tau\mathbf{1}_{\{\alpha^{2}\leq\epsilon\}}$
Covariance kernels:
$\mathbf{V}^{C}_{r}[i,j]=\kappa(|i-j|)\sigma_{r}^{2}+e^{-2\alpha^{2}\Delta t|i-j|}\eta_{r}^{2}$
$\mathbf{V}^{C}_{t}[i,j]=\kappa(|i-j|)\sigma_{t}^{2}+e^{-2\alpha^{2}\Delta t|i-j|}\eta_{t}^{2}$
Precompute kernels:
$\mathbf{E}_{\mathbf{qk}}[i,j]=e^{-\alpha^{2}(t_{i}-t_{j})},\quad\mathbf{E}_{\mathbf{v}}[i,j]=e^{-\alpha^{2}(t_{i}-t_{j})},$
$\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}[k,j]=e^{-i\omega_{k}t_{j}},\quad\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}[k,j]=e^{i\omega_{k}t_{j}},$
$\bm{\tilde{\Phi}}_{\mathbf{v}}^{-}[k,j]=e^{-i\omega_{k}t_{j}},\quad\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}[k,j]=e^{i\omega_{k}t_{j}},$
$\mathbf{\tilde{Z}_{q}}=\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}\odot\mathbf{Z_{q}},\quad\mathbf{\tilde{Z}_{k}}=\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}\odot\mathbf{Z_{k}}$
Compute normalized value directions: $\mathbf{U}_{\mathbf{v}}[:,i]=\mathbf{Z_{v}}[:,i]\big/\left(\big|\mathbf{Z_{v}}[:,i]\big|+\epsilon\right)$
Radial and tangential decompositions:
$\mathbf{\tilde{m}}_{\mathbf{q}}=\mathbf{U}_{\mathbf{v}}^{\dagger}\mathbf{\tilde{Z}_{q}},\quad\mathbf{\tilde{m}}_{\mathbf{k}}=\mathbf{U}_{\mathbf{v}}^{\dagger}\mathbf{\tilde{Z}_{k}},\quad\mathbf{\tilde{m}}_{\mathbf{v}}=\mathbf{U}_{\mathbf{v}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{v}})$
$\mathbf{\tilde{V}}_{\mathbf{q}}=\mathbf{\tilde{Z}_{q}}-\mathbf{U}_{\mathbf{v}}\mathbf{\tilde{m}}_{\mathbf{q}},\quad\mathbf{\tilde{V}}_{\mathbf{k}}=\mathbf{\tilde{Z}_{k}}-\mathbf{U}_{\mathbf{v}}\mathbf{\tilde{m}}_{\mathbf{k}},\quad\mathbf{\tilde{V}}_{\mathbf{v}}=\mathbf{\tilde{Z}_{v}}-\mathbf{U}_{\mathbf{v}}\mathbf{\tilde{m}}_{\mathbf{v}}$
Compute residuals:
Tangential: $\mathbf{R}_{\text{tan}}^{2}=\big|\mathbf{\tilde{V}}_{\mathbf{q}}\big|_{2}^{2}\mathbf{1}_{m}^{\top}+\mathbf{1}_{m}\big|\mathbf{\tilde{V}}_{\mathbf{k}}\big|_{2}^{2}{}^{\top}-2\mathrm{Re}(\mathbf{E}_{\mathbf{qk}}\odot\mathbf{\tilde{V}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{V}}_{\mathbf{k}})$
Radial: $\mathbf{R}_{\text{rad}}^{2}=(\mathbf{\tilde{m}}_{\mathbf{q}}\mathbf{\tilde{m}}_{\mathbf{q}}^{\dagger})\mathbf{1}_{m}^{\top}+\mathbf{1}_{m}(\mathbf{\tilde{m}}_{\mathbf{k}}\mathbf{\tilde{m}}_{\mathbf{k}}^{\dagger})^{\top}-2\mathrm{Re}(\mathbf{E}_{\mathbf{qk}}\odot\mathbf{\tilde{m}}_{\mathbf{q}}\mathbf{\tilde{m}}_{\mathbf{k}}^{\dagger})$
Dynamic precision kernels:
$\mathbf{P}^{C}_{\text{rad}}[i,j]=1/\sigma_{V_{r}}^{2}(|t_{i}-t_{j}|),\quad\mathbf{P}^{C}_{\text{tan}}[i,j]=1/\sigma_{V_{t}}^{2}(|t_{i}-t_{j}|)$
Robust weights and attention:
$\mathbf{\tilde{A}}_{\text{rad}}[i,j]=-\log(\mathbf{V}^{C}_{\text{rad}}[i,j]+\mathbf{R}_{\text{rad}}^{2}[i,j]+\epsilon),$
$\mathbf{\tilde{A}}_{\text{tan}}[i,j]=-\log(\mathbf{V}^{C}_{\text{tan}}[i,j]+\mathbf{R}_{\text{tan}}^{2}[i,j]+\epsilon)$
$\mathbf{A}_{\text{rad}}=\operatorname{softmax}\Big(\tau_{r}\,\mathbf{\tilde{A}}_{\text{rad}}\odot\mathbf{M}_{\text{causal}},\;\text{dim}=1\Big),$
$\mathbf{A}_{\text{tan}}=\operatorname{softmax}\Big(\tau_{t}\,\mathbf{\tilde{A}}_{\text{tan}}\odot\mathbf{M}_{\text{causal}},\;\text{dim}=1\Big)$
Apply value decay and compute outputs:
$\mathbf{\hat{A}}_{\text{rad}}=\mathbf{E}_{\mathbf{v}}\odot\mathbf{A}_{\text{rad}},\quad\mathbf{\hat{A}}_{\text{tan}}=\mathbf{E}_{\mathbf{v}}\odot\mathbf{A}_{\text{tan}}$
$\mathbf{\bar{m}}_{\mathbf{v}}=\mathbf{\hat{A}}_{\text{rad}}\,\mathbf{\tilde{m}}_{\mathbf{v}},\quad\mathbf{\bar{V}}_{\mathbf{v}}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}\odot(\mathbf{\tilde{V}}_{\mathbf{v}}\,(\mathbf{\hat{A}}_{\text{tan}})^{\dagger})$
Geodesic step: $\mathbf{\bar{U}}_{\mathbf{v}}[:,i]=\left(\mathbf{Z}_{\mathbf{v}}[:,i]+\mathbf{\bar{V}}_{\mathbf{v}}[:,i]\right)\Big/\left(\big|\mathbf{Z}_{\mathbf{v}}[:,i]+\mathbf{\bar{V}}_{\mathbf{v}}[:,i]\big|+\epsilon\right)$
Reconstruct output: $\mathbf{\bar{Z}}_{\mathbf{v}}[:,i]=\mathbf{\bar{m}}_{\mathbf{v}}[i]\cdot\mathbf{\bar{U}}_{\mathbf{v}}[:,i]$
Residual connection and prediction:
$\mathbf{\bar{Z}_{v}}\leftarrow(1-\text{sigmoid}(\delta))\mathbf{Z_{v}}+\text{sigmoid}(\delta)\mathbf{\bar{Z}_{v}},\quad\mathbf{Z}_{\mathbf{v}}^{+}=\mathbf{E_{v}}[:,1]\odot\mathbf{\bar{Z}}_{\mathbf{v}}$
Output projection:
$\big(\operatorname{Re}(\mathbf{Z_{p}}),\operatorname{Im}(\mathbf{Z_{p}})\big)\leftarrow\mathcal{L}_{P}(\operatorname{Re}(\mathbf{Z}_{\mathbf{v}}^{+}),\operatorname{Im}(\mathbf{Z}_{\mathbf{v}}^{+}))$
Return: $\operatorname{Re}(\mathbf{Z_{p}})$
Algorithm 2 RT Adaptive Filter Attention
5.5 Interpreting the Transformer as an MLE on a Hypersphere
The algorithm above faithfully implements RT-MLE with geometric constraints, but is somewhat intricate. We propose the following simplifications for clarity and practical use.
Recall that, in RT-Kalman IRLS our estimated state decomposes as:
$\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}=e^{-\bm{\lambda}\Delta t_{ij}}\mathbf{z_{s}}_{j}^{(k)}=\Big(e^{-\alpha\Delta t_{ij}}m_{\mathbf{zs}j}^{(k)}\Big)\Big(e^{-\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{u}_{\mathbf{zs}j}^{(k)}\Big)$
where $\mathbf{\hat{u}}_{\mathbf{zs}ij}^{(k)}=\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}/\big|\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}\big|=e^{-\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{u}_{\mathbf{zs}j}^{(k)}$. Assuming that the transformed magnitude is approximately constant across time steps $e^{-\alpha\Delta t_{ij}}m_{\mathbf{zs}j}^{(k)}\approx m_{\mathbf{zs}i}^{(k)}$, gives a simple precision-weighted average of $\mathbf{\hat{u}}$:
$\mathbf{\bar{v}}_{\mathbf{s}i}^{(k)}\approx m_{\mathbf{zs}i}^{(k)}\mathbf{\bar{u}}_{\mathbf{zs}i}^{(k)},\quad\text{where:}\quad\mathbf{\bar{u}}_{\mathbf{zs}i}^{(k)}\approx\left(\sum_{j\leq i}w_{t,ij}^{(k)}/\sigma_{Vt,ij}^{2}\right)^{-1}\sum_{j\leq i}\big(w_{t,ij}^{(k)}/\sigma_{Vt,ij}^{2}\big)\mathbf{\hat{u}}_{\mathbf{zs}ij}^{(k)}$
For small angles, the attention update is largely aligned with the tangent space and any component along $u$ is partially corrected by the subsequent geodesic normalization, so we omit the explicit tangent projection for simplicity.
The geodesic update step becomes:
$\mathbf{y}_{\mathbf{zs}i}^{(k)}\approx\mathbf{u}_{\mathbf{zs}i}^{(k)}+\mathbf{\bar{u}}_{\mathbf{zs}i}^{(k)},\qquad\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\frac{\mathbf{y}_{\mathbf{zs}i}^{(k)}}{\big|\mathbf{y}_{\mathbf{zs}i}^{(k)}\big|},$
And the output is again:
$\mathbf{z}_{\mathbf{s}i}^{(k+1)}=m_{\mathbf{zs}i}^{(k+1)}\,\mathbf{u}_{\mathbf{zs}i}^{(k+1)}.$
This structure — Norm, Attention, Add & Norm — closely resembles a Transformer, except that the normalization is applied in the eigenbasis (after the query/key/value projections), to complex-valued vectors.
This suggests that the Transformer’s effectiveness may stem, at least in part, from its ability to approximate a principled filtering mechanism, with attention functioning as a generalized MLE for an SDE with unitary (or norm-preserving) linear dynamics, and normalization layers realizing approximate geodesic steps along the hypersphere.
We propose the following Transformer-inspired simplification of RT-AFA:
Input: Input sequence $\mathbf{Z}_{r}\in\mathbb{R}^{d\times m}$
Parameters: AFA layer $\mathrm{AFA}_{\mathrm{dir}}$, scalar AFA layer $\mathrm{AFA}_{\mathrm{mag}}$, output complex linear layer $\mathcal{L}_{O}$, real-valued FFN layers: $\mathrm{FFN}=\mathcal{L}_{2}\circ\mathrm{ReLU}\circ\mathcal{L}_{1}$, stability const. $\epsilon$.
Lift real inputs to complex by concatenating zeros: $\mathbf{Z}=\left[\mathbf{Z}_{r},\mathbf{0}\right]^{\top}\in\mathbb{C}^{2\times d\times m}$
Complex projections:
$\big(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big)\leftarrow\mathcal{L}_{V}\big(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z})\big).$
$\big(\operatorname{Re}(\mathbf{Z_{q}}),\operatorname{Im}(\mathbf{Z_{q}})\big)\leftarrow\mathcal{L}_{Q}\big(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z})\big)$,
$\big(\operatorname{Re}(\mathbf{Z_{k}}),\operatorname{Im}(\mathbf{Z_{k}})\big)\leftarrow\mathcal{L}_{K}\big(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z})\big)$.
Magnitudes and unit directions (complex-valued):
$\mathbf{m}_{\mathbf{q}}=\big|\mathbf{Z}_{\mathbf{q}}\big|_{\text{rowwise}};\,\mathbf{m}_{\mathbf{k}}=\big|\mathbf{Z}_{\mathbf{k}}\big|_{\text{rowwise}};\,\mathbf{m}_{\mathbf{v}}=\big|\mathbf{Z}_{\mathbf{v}}\big|_{\text{rowwise}}\in\mathbb{R}^{1\times m}$
$\mathbf{U}_{\mathbf{q}}=\gamma_{1}\mathbf{Z}_{\mathbf{q}}\big/(\mathbf{m}_{\mathbf{q}}+\epsilon);\,\mathbf{U}_{\mathbf{k}}=\gamma_{1}\mathbf{Z}_{\mathbf{k}}\big/(\mathbf{m}_{\mathbf{k}}+\epsilon);\,\mathbf{U}_{\mathbf{v}}=\gamma_{1}\mathbf{Z}_{\mathbf{v}}\big/(\mathbf{m}_{\mathbf{v}}+\epsilon)$
Apply AFA attention (excluding projections):
$\mathbf{\hat{m}}=\mathrm{AFA}_{\mathrm{mag}}(\mathbf{m_{q}},\mathbf{m_{k}},\mathbf{m_{v}})$
$\mathbf{\hat{U}}=\mathrm{AFA}_{\mathrm{dir}}(\mathbf{U_{q}},\mathbf{U_{k}},\mathbf{U_{v}})$
Geodesic step:
$\mathbf{Y}=\mathbf{U}+\mathbf{\hat{U}}$
$\mathbf{m}_{y}=\big|\mathbf{Y}\big|_{\text{row-wise}}$
$\mathbf{\tilde{U}}=\gamma_{2}\mathbf{Y}\big/(\mathbf{m}_{y}+\epsilon)\quad(\text{element-wise division})$
Combine magnitude and direction estimates:
$\mathbf{Z}=\mathbf{\hat{m}}\odot\mathbf{\tilde{U}}$
Output complex linear projection:
$(\operatorname{Re}(\mathbf{Z_{o}}),\operatorname{Im}(\mathbf{Z_{o}}))\leftarrow\mathcal{L}_{O}(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z}))$
Output FFN with nonlinearity and residual connection:
$\mathbf{Z}^{\mathrm{out}}=\mathbf{Z}_{r}+\mathrm{FFN}(\operatorname{Re}(\mathbf{Z_{o}}))$
Output: Updated sequence $\mathbf{Z}^{\mathrm{out}}\in\mathbb{R}^{d\times m}$
Algorithm 3 AFA Transformer
6 Related Work
6.1 Classical Filtering Methods
Norbert Wiener’s 1949 work on the Wiener Filter established the theoretical framework for minimum mean-square error estimation of signals from noisy observations, using second-order statistics (covariances) of the signal and noise (Wiener, 1964). Kalman introduced the discrete-time Kalman Filter, a recursive algorithm that optimally estimates the states of linear time-invariant stochastic systems from noisy measurements, achieving minimum mean-square estimation error (Kalman, 1960). Kalman & Bucy extended the approach to continuous-time systems with Gaussian noise (Kalman and Bucy, 1961). Rauch, Tung, and Striebel generalized the Kalman Filter to compute smoothed state estimates over the full trajectory of a linear dynamical system. Their method recursively combines forward-filtered estimates with backward-time updates to produce maximum-likelihood estimates of the states with minimum mean-square error over the entire time horizon (Rauch et al., 1965). Extensions to nonlinear systems can be handled by the Extended Kalman Filter (EKF), which linearizes the dynamics and measurement functions about the current state estimate (Anderson and Moore, 1979).
6.2 Robust and Adaptive Filtering
In many practical settings, the system and noise parameters required by the Kalman Filter are unknown or time-varying. Early work in adaptive signal processing began with Widrow & Hoff, who introduced the Least Mean Squares (LMS) algorithm, laying the foundation for recursive adaptation using the innovation signal (Widrow and Hoff, 1960). Schweppe provided early maximum likelihood foundations for estimating noise covariances in Gaussian signals (Schweppe, 1965). Mehra formalized adaptive Kalman Filtering by estimating unknown process and measurement noise covariances from the innovations sequence (Mehra, 1970). Iteratively reweighted least squares (IRLS) provides a framework for robust estimation by repeatedly solving weighted least-squares problems, with weights updated based on residuals; this approach was formalized in signal processing contexts by Burrus, Barreto, and Selesnick for FIR filter design (Burrus et al., 1994).
6.3 Neural Networks for Filtering Dynamical Systems
Kalman Filtering and Neural Networks, edited by Haykin, provides a comprehensive treatment of combining classical state estimation with neural network models (Haykin, 2001). Other work demonstrated that multilayer and recurrent neural networks can be effectively used for the identification and adaptive control of nonlinear dynamical systems, for example Narendra and Parthasarathy (Narendra and Parthasarathy, 1990).
Krishnan et al. introduced Deep Kalman Filters (DKFs), which extend classical linear Kalman Filters to high-dimensional, non-linear latent dynamics using deep neural networks. DKFs maintain a probabilistic latent state and propagate uncertainty explicitly. In this framework, the latent state is updated sequentially according to a learned, potentially nonlinear transition function, and observations are generated from the latent state through a neural network (Krishnan et al., 2015).
Liu et al. propose Kalman Filtering Attention (KFAtt) for CTR prediction, treating each user click as a noisy measurement of a hidden query-specific interest and computing a closed-form MAP estimate that combines these measurements with a global prior. Despite the “Kalman” terminology, KFAtt does not model temporal dynamics or state evolution, and effectively performs a static, precision-weighted aggregation of observations.
Goel and Bartlett investigate whether Transformers can represent classical Kalman Filters for linear time-invariant (LTI) systems. Their Transformer Filter uses softmax self-attention over nonlinear embeddings of past state estimates and observations. It effectively approximates the Kalman update indirectly, with the kernel weights mimicking the Kalman gain, and requires careful tuning of a temperature parameter to control approximation error (Goel and Bartlett, 2024).
Several recent works integrate neural networks or attention into Kalman Filtering (Shlezinger et al., 2025). KalmanNet uses a recurrent network to learn the Kalman gain for real-time filtering under partially known or nonlinear dynamics (Revach et al., 2022). The Neural Extended Kalman Filter parameterizes both process and observation models with neural networks for end-to-end latent state inference (Liu et al., 2023). The Attention Kalman Filter (AtKF) replaces the traditional Kalman gain with a self-attention network that learns to capture dependencies across past state and innovation sequences (Wang et al., 2024). A-KIT (2024) learns process noise covariance online to improve EKF predictions (Cohen and Klein, 2024). KalmanFormer integrates a Transformer network into the Kalman Filter framework to directly learn the Kalman gain from data. (Shen et al., 2025).
The present work extends the MS Thesis, "Adaptive Filter Attention," which demonstrates that a tensorized, residual- and precision-weighted attention formulation implements a maximum-likelihood estimator for a diagonalizable linear time-invariant SDE (Racioppo, 2025).
Uncertainty Propagation Networks (UPNs) extend Neural ODEs to jointly model state means and covariances through coupled differential equations, allowing continuous-time propagation of uncertainty sequence modeling. UPN propagates uncertainty continuously through time by explicitly evolving the covariance of the state alongside the state mean, essentially functioning as a continuous-time Kalman Filter embedded in a learned dynamics model (Jahanshahi and Zhu, 2025).
6.4 Neural State Space Models
State space models (SSMs) are a powerful framework for modeling sequences and time series data by representing system dynamics through latent states that evolve over time. A key challenge in sequence modeling is maintaining memory over long time horizons, and avoiding the vanishing and exploding gradient problem faced by recurrent neural networks (Somvanshi et al., 2025).
The HiPPO framework improves memory retention in SSMs by introducing recurrent memory with optimal polynomial projections, which allows the model to compress and retain important information from the entire history in a structured latent state space (Gu et al., 2020). The Structured State Space Sequence model (S4) designs efficient state space layers that scale linearly with sequence length, using structured representations of the system matrices (Gu et al., 2022). The Toeplitz Neural Network (TNN) efficiently models sequences using relative position–encoded Toeplitz matrices and a lightweight Relative Position Encoder, enabling fast token mixing and extrapolation to much longer sequences with log-linear complexity (Qin et al., 2023).
Subsequent works such as Mamba further improve efficiency by introducing selective state space components that dynamically focus computation on relevant parts of the sequence (Gu and Dao, 2024). Mamba Mixer combines structured state space models with selective token and channel mixing via learned gating, enabling efficient and adaptive modeling of long sequences by focusing on relevant elements and features (Behrouz et al., 2024).
Both S4 and our model exploit a diagonal-plus-low-rank (DPLR) structure to reduce the cost of inverting a dense $d\times d$ matrix, using the Sherman-Morrison formula for efficiency: a diagonal part captures independent modes, and a rank-1 correction captures global interactions. In S4, the DPLR matrix models deterministic state dynamics, determining how past inputs influence future states. In our model, the DPLR structure appears in the noise covariance matrices in a stochastic differential equation (SDE) formulation, which allows us to represent how uncertainty propagates through the dynamics.
6.5 Complex-Valued Neural Networks
Complex-valued neural networks (CVNNs) extend traditional real-valued ANNs by using complex numbers in weights, activations, and computations, naturally capturing rotations, phase relationships, and oscillatory dynamics that arise in linear algebra. Prior work includes studies of complex-valued activation functions, parameter initialization strategies, and the use of Wirtinger calculus, which treats complex parameters as pairs of real variables to enable effective gradient computation (Hirose, 2012; Bassey et al., 2021; Barrachina et al., 2023; Hammad, 2024).
Trabelsi et al. (2018) introduced practical components for training deep complex-valued neural networks—complex convolutions, batch normalization, weight initialization, and activation functions—and demonstrated their effectiveness on vision and audio tasks, showing competitive performance with real-valued networks (Trabelsi et al., 2018). Eilers and Jiang (2023) developed core components for a complex-valued transformer, including complex Scaled Dot-Product Attention and layer normalization, enabling fully complex-valued models that match real-valued performance while improving robustness to overfitting on complex-valued tasks (Eilers and Jiang, 2023). Ran-Milo et al. (2024) show that complex-valued SSMs can express oscillatory and long-range mappings with moderate dimensions and parameters, while real SSMs require much larger or exponentially scaled parameters for the same expressivity (Ran-Milo et al., 2024). Leng et al. develop a framework for complex-valued neural networks and transformers in wireless communications, showing that CVNNs require fewer layers than real-valued networks and introducing complex-valued embedding, encoding, decoding, and output modules (Leng et al., 2025).
6.6 Attention as a Structured State Space Computation
The attention mechanism was first introduced in (Bahdanau et al., 2016) to improve neural machine translation by allowing models to dynamically focus on relevant input tokens, and was later generalized into the self-attention architecture of the Transformer by Vaswani et al, which became the foundation for modern large language models (Vaswani et al., 2023).
Attention mechanisms can be interpreted through the lens of structured state space computations. Several recent works have explored these connections and proprosed structured alternatives to standard Transformer attention based on state space models.
Katharopoulos et al. propose a Linear Attention framework that approximates attention weights using kernel feature maps, enabling the Transformer to be expressed in a recurrent form. While this formulation offers computational benefits, it does not explicitly connect the recurrence to an underlying system dynamics model; rather, the recurrent structure serves primarily as an efficient implementation (Katharopoulos et al., 2020).
Dao and Gu establish a formal equivalence between the convolution kernel of a recurrent linear time-invariant (LTI) system and the attention weights in causal linear attention. They show that the state-space recurrence can be expressed as multiplication by an $N$-sequentially semiseparable (SSS) matrix, revealing that the attention matrix in a Transformer is effectively a low-rank, structured operator along the sequence dimension. This connection enables the use of fast algorithms developed for SSS matrices to accelerate attention. (Dao and Gu, 2024).
Other recent works explore the integration of attention mechanisms with state space models. For example, the Dynamical Systems Framework (DSF) unifies attention, state space models, and recurrent networks under a common recurrent/convolutional form (Sieber et al., 2024). Bick et al. focus on distilling quadratic self-attention into efficient SSMs, with the goal of matching the performance of Transformers while reducing computation (Bick et al., 2025).
6.7 Structuring Attention with Priors and Positional Biases
The original Transformer lacks built-in modeling of relative or absolute positions, relying instead on fixed additive sinusoidal positional encodings (Vaswani et al., 2023).
Many methods have proposed extending self-attention to efficiently incorporate relative position information. For example, Shaw et al. introduce a mechanism to encode relative distances between tokens directly into the attention computation (Shaw et al., 2018). Transformer-XL introduces relative positional encodings, enabling the model to generalize to sequences longer than those seen during training (Dai et al., 2019). Liu et al. treat the sequence of positional encodings as samples from a continuous trajectory governed by a Neural ODE. This formulation enforces smoothness, allows irregularly spaced positions via arbitrary sampling of time, and recovers sinusoidal encodings as a special case (Liu et al., 2020).
Su et al. introduce rotary positional encodings (RoPE), which encode relative position by applying a complex rotation to each query–key pair before computing their dot product. This makes attention scores depend only on relative displacement, yielding translation-equivariant attention (Su et al., 2023).
LeX Transformer extends RoPE with xPos, an exponential-decay stabilized rotary embedding, and blockwise masking, enabling Transformers to generalize smoothly to longer sequences without retraining (Sun et al., 2022). The Retentive Network (RetNet) combines each token’s hidden state with an exponentially decaying summary of previous tokens, using decay-masked attention to enforce causality and a gating mechanism to control past contributions and increase expressivity (Sun et al., 2023). Other models, including Dynamically Context-Sensitive Time-Decay Attention (Su et al., 2018) and RWKV also learn decay functions based on temporal proximity (Peng et al., 2023).
Some recent works focus on understanding the mathematical structure of rotary embeddings and extending them to higher dimensions or more flexible parameterizations. Yu et al. introduce ComRoPE, which replaces fixed 2D rotation matrices with trainable commuting angle matrices (Yu et al., 2025). Liu et al. use commuting skew-symmetric generators to define an $N$-dimensional RoPE, generalizing 1D and 2D forms and allowing learnable inter-dimensional interactions (Liu et al., 2025). CoPE (Amballa, 2025), and ComplexFormer (Shao et al., 2025) extend RoPE into the complex domain. Wu et al. observe that the same rotation trick used in Transformers (on queries and keys) can be applied to the matrices in an SSM formulation (Wu et al., 2025).
Press et al. introduce Attention with Linear Biases (ALiBi), which adds a distance-dependent bias to attention logits to enable length extrapolation (Press et al., 2022).
Several papers have explored element-wise (in the embedding dimension) attention mechanisms. Elliptical Attention replaces the traditional dot-product attention with a Mahalanobis distance-based mechanism (Nielsen et al., 2024). McCarter investigates replacing scaled dot-product attention with negative-log Euclidean distance (McCarter, 2023). Feng uses an exponentiated Euclidean distance as a similarity metric (Feng, 2025). Bacochina et al. replace matrix multiplications with learned spatial weight arrays applied directly via element-wise multiplication, focusing attention along rows and columns (Bacochina and de Souza, 2023). Wu et al. introduce the DA-Transformer, which reweights attention scores using learnable functions of token distances (Wu et al., 2021)
6.8 Dynamical Systems Perspectives on Attention
Attention mechanisms have been studied through continuous-time and dynamical systems frameworks, modeling hidden states or latent dynamics via ordinary differential equations (ODEs) or stochastic processes to capture temporal dependencies, uncertainty, and adaptivity.
Stochastic Recurrent Networks (STORNs) combine RNNs with time-dependent latent variables and variational inference to model multimodal temporal dynamics (Bayer and Osendorfer, 2015). Neural ODEs treat hidden states as continuous-time dynamics parameterized by a neural network and solved with an ODE solver (Chen et al., 2019). Neural Processes (NPs) (Garnelo et al., 2018) and Neural ODE Processes (NDPs) (Norcliffe et al., 2021) extend this framework to learn distributions over functions or ODE trajectories, enabling uncertainty-aware predictions. Rubanova et al. introduce ODE-RNNs and Latent ODEs, continuous-time models that handle irregularly-sampled time series without discretization or imputation (Rubanova et al., 2019).
Chien and Chen (2021) introduce continuous-time attention, integrating attention mechanisms with neural differential equations to operate over latent trajectories rather than discrete hidden states, enabling effective modeling of irregularly-sampled sequences (Chien and Chen, 2021). Attentive Neural Processes (ANPs) extend Neural Processes by incorporating attention to allow each input to selectively attend to relevant context points (Kim et al., 2019). Chen at al. introduce self-modulating attention , which extends classical attention to continuous-time sequences by adaptively reweighting based on both sequential order and temporal distance (Chen et al., 2022). ACE-NODE and ANCDE integrate attention into ODE or controlled differential equation frameworks, learning attention dynamics continuously rather than via standard query-key comparisons (Jhin et al., 2021a, b). Neural ODE Transformers and Hopfield Networks reinterpret attention as evolving or energy-based dynamical systems (Tong et al., 2025; Ramsauer et al., 2021). Continuous-discrete state space models and Structured Inference Networks combine latent continuous dynamics with discrete observations and approximate Bayesian filtering (Ansari et al., 2023; Krishnan et al., 2016).
Wang and Yao combine neural processes with neural SDEs to improve uncertainty estimation on noisy or incomplete data, using attention-based NPs for multidimensional inputs (Wang and Yao, 2021). FIM-SDE estimates drift and diffusion functions from noisy, sparse SDE observations by learning a generalizable mapping from trajectories to underlying dynamical system parameters (Seifner et al., 2025). Shen and Cheng propose a Neural SDE framework that models continuous-time sequences with neural network–parameterized drift and diffusion, using a decoupled maximum likelihood scheme for efficient, robust training (Shen and Cheng, 2025).
6.9 Modeling Uncertainty in Neural Networks
Other approaches explicitly model uncertainty in neural networks and attention mechanisms. Chung et al. introduce the Variational Recurrent Neural Network (VRNN), which incorporates latent random variables into RNN hidden states to represent and propagate uncertainty in sequential data (Chung et al., 2016). Fraccaro et al. introduce the stochastic recurrent neural network, a sequential model combining deterministic RNN states with stochastic latent variables, using structured variational inference to propagate uncertainty over time and improve sequence modeling (Fraccaro et al., 2016).
The Bayesian Attention Mechanism (BAM) interprets attention weights as joint probabilities over token content and position, showing that positional encoding strategies correspond to specific priors and introducing a generalized Gaussian prior (GGD-BAM) that balances local and long-range dependencies (Bianchessi et al., 2025). Litman shows that scaled dot-product attention is the exact solution to a one-sided entropic optimal transport problem (Litman, 2025). The Correlated Gaussian Process Transformer models self-attention as the cross-covariance between two correlated Gaussian processes (Bui et al., 2025). Ruan and Zhang analyze self-attention as a similarity-based information propagation process converging to a drift-diffusion or heat equation on a manifold (Ruan and Zhang, 2024).
Various approaches embed structured or probabilistic biases directly into attention mechanisms. For example, Probabilistic Transformers model attention as MAP inference in a Gaussian mixture model, with each key/value pair having its own precision parameter, updated via expectation-maximization (Movellan and Gabbur, 2020). Uncertainty-Aware Attention introduces input-specific uncertainty modeling through variational inference (Heo et al., 2018). The Attention Gaussian Process (AGP) treats attention weights as probabilistic outputs from a sparse Gaussian process, with variational inference approximating the posterior (Schmidt et al., 2023). SLANG uses a diagonal plus low-rank covariance to efficiently estimate uncertainty in deep networks via natural-gradient variational inference (Mishkin et al., 2019).
Other recent probabilistic attention methods introduce uncertainty-aware mechanisms into Transformer-based architectures. These include the Uncertainty-Guided Probabilistic Transformer (UGPT) (Guo et al., 2022), Attention Gaussian Process (AGP) (Schmidt et al., 2023), Gaussian Adaptive Attention (GAAM) (Ioannides et al., 2024), the Correlated Gaussian Process Transformer (CGPT) (Bui et al., 2025), and Kernel-Eigen Pair Sparse Variational Gaussian Processes (KEP-SVGP) (Chen et al., 2024).
6.10 Structured Low-Rank Covariances for Efficient Propagation
Several methods have been developed to efficiently represent and propagate high-dimensional covariance matrices using low-rank plus diagonal approximations in Riccati-like equations. Pnevmatikakis et al. propose a low-rank perturbative approach for Kalman Filtering and smoothing, approximating the state covariance as a small-rank update to the prior (Pnevmatikakis et al., 2014). Saibaba et al. represent the covariance as a weighted sum of the process noise and a low-rank term from a generalized eigenvalue problem (Saibaba et al., 2015). Bonnabel et al. approximate large, time-dependent PSD matrices with low-rank plus diagonal structures, allowing linear-time storage, inversion, and integration in high-dimensional Riccati flows (Bonnabel et al., 2024). Schmidt et al. (2021) introduce the Rank-Reduced Kalman Filter, projecting the Lyapunov equations onto a low-rank manifold and updating only the covariance column space, reducing complexity from cubic to quadratic (or linear) while recovering the exact filter in the full-rank limit and outperforming ensemble-based approaches (Schmidt et al., 2024).
6.11 Neural Network Layers as Unrolled Iterative Algorithms
Unrolling iterative optimization algorithms into trainable networks began with frameworks like LISTA, which expresses the ISTA sparse-coding algorithm as a neural network with learned parameters (Gregor and LeCun, 2010). Deep Equilibrium Models (DEQs) extend this idea by solving a single equilibrium layer via root-finding, effectively modeling infinite-depth networks with constant memory (Bai et al., 2019). Wang et al. study how attention can also be viewed as an unrolled iterative process: multi-head self-attention corresponds to iterative subspace denoising, compressing inputs into a low-rank manifold (Wang et al., 2025). This is analogous to our framing of attention as a step of iteratively-reweighted least squares.
6.12 Geometric Views of Transformers
Geshkovski et al. model Transformers as interacting particle systems evolving on the unit sphere, and find that self-attention induces nonlinear dynamics leading to clustering of token embeddings (Zhang et al., 2021; Geshkovski et al., 2024). Molina interprets Transformers geometrically, showing that LayerNorm constrains hidden states to a hypersphere and attention moves “word particles” along this surface (Molina, 2023). Brody et al. show that LayerNorm shapes Transformer hidden states geometrically — projecting them onto a hyperplane and scaling them — which enhances the expressiveness of attention (Brody et al., 2023). Loshichilov et al. introduce nGPT, a Transformer variant that normalizes all embeddings and network weights on the hypersphere, enabling cosine-similarity-based updates with learnable per-dimension step sizes, resulting in faster convergence and more stable training (Loshchilov et al., 2025). Gupta et al. observe that LayerNorm can be interpreted geometrically as projecting a vector orthogonally to the uniform vector, normalizing it, and scaling by the dimensionality, making it irreversible and largely redundant compared to RMSNorm, which achieves similar hidden representations more efficiently (Gupta et al., 2025).
7 Discussion and Conclusion
This work explores a connection between classical control theory and attention mechanisms by framing attention through the lens of linear dynamical systems and adaptive filtering. We show that attention arises naturally as the solution to a maximum likelihood estimation problem under a linear stochastic differential equation with structured system matrices and noise covariances, and discrete observations. Unlike approaches that treat attention as a static kernel or a matrix factorization, we view attention as a parallelized form of adaptive Kalman Filtering: each key-value pair is interpreted as a noisy observation of a latent state, and a precision-weighted Bayesian update aggregates evidence across all observations to refine the state estimate in a parallel manner. To the best of our knowledge, this view of attention as the MLE for a structured linear SDE has not appeared in the literature on attention mechanisms.
Imposing an isotropic-plus-skew-symmetric state matrix and isotropic noise, our model achieves the same asymptotic time and memory complexity as standard attention. This result in a form of complex-valued rotary positional encoding equivalent to that found in xPos and RetNet, along with a query-key precision kernel that quantifies an uncertainty prior under the learned dynamics model.
We also introduce a Radial–Tangential SDE model, whose isotropic-plus-skew-symmetric state matrix and anisotropic noise in the token frame define the most expressive SDE structure that remains analytically tractable. This formulation leads directly to a maximum likelihood estimator on the hypersphere, which provides a novel, principled foundation for understanding norm-preserving attention mechanisms, shedding new light on the architectural properties of the Transformer.
Several promising directions remain for future work. A natural extension is to allow the state matrix and noise to depend on the state. Although a fully time-varying dynamics model would break the linear time-invariant (LTI) structure and would no longer allow a closed-form solution to the Differential Lyapunov Equation (DLE), this issue can be avoided by fixing the dynamics within each context window. In this way, the parameters may adapt across windows while remaining constant within a window, enabling content adaptivity without sacrificing computational efficiency.
Future work should also investigate the relationship between the precision-weighted update and concepts from robust and adaptive control, such as $H_{\infty}$ filtering. This could lead to attention mechanisms with inherent guarantees of stability and robustness against dynamic model mismatch, which is critical for trustworthy AI applications.
The closed-form expression for the propagated precision matrix, which assumes simultaneous diagonalizability, can be computed in $\mathcal{O}(md)$ time, whereas full integration and inversion in the general case requires $\mathcal{O}(md^{3})$ time. Since our algorithm already involves $\mathcal{O}(md^{2})$ matrix multiplications, moderately less efficient integration methods introduce a limited additional overhead. In particular, if we use a diagonalizable-plus-rank-1 state matrix $\mathbf{A}$, covariance propagation can be computed efficiently by expressing the matrix exponential and integrals as a diagonal term plus a rank-1 outer-product correction, reducing complexity to $\mathcal{O}(md^{2})$, though losing a fully closed form. We could also consider a generalization using the Jordan canonical form. Alternatively, differentiable numerical solvers for the differential Lyapunov equation, or unrolled iterations of specialized low-rank methods such as Krylov subspace techniques or ADI iterations could potentially be integrated directly into a neural network layer.
To achieve the statistical accuracy of the full batch estimator while retaining the computational efficiency of a recursive filter, one could explore alternative unrolling strategies. These may include inner-loop iterations within each time step, linearizing weights around a precomputed batch solution, or hybrid schemes that propagate residual-dependent weights more globally across the trajectory.
Finally, the duality between attention weights and the Kalman gain suggests a path toward dynamics-aware neural control policies, where the separation principle can be leveraged to achieve more sample-efficient policy evaluation and state estimation in noisy or partially observed reinforcement learning settings.
Appendix A Background
A.1 Propagation of Uncertainty
Recall that the state at time $t_{j}$ can be written as:
$\mathbf{x}(t_{j})=e^{\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}\,d\mathbf{w}(\tau)\right).$
The covariance of the integral term is:
$\mathbf{V}_{B}(\Delta t_{ij}):=\mathbb{E}\Bigg[\left(\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau_{1}}\mathbf{G}\,d\mathbf{w}(\tau_{1})\right)\left(\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau_{2}}\mathbf{G}\,d\mathbf{w}(\tau_{2})\right)^{\top}\Bigg]$
$\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{Q}e^{-\mathbf{A}^{\top}\tau}\,d\tau,\quad\mathbf{Q}=\mathbf{G}\mathbf{G}^{\top}$
We will refer to this as the backward-propagated state covariance. As the system evolves, the state distribution evolves as:
$\mathbf{x}(t_{j})\sim\mathcal{N}\left(e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i}),\mathbf{V}_{F}(\Delta t_{ij})\right)$
where the forward-propagated state covariance is:
$\mathbf{V}_{F}(\Delta t_{ij})=e^{\mathbf{A}\Delta t_{ij}}\mathbf{V}_{B}(\Delta t_{ij})e^{\mathbf{A}^{\top}\Delta t_{ij}}$
$e^{\mathbf{A}\Delta t_{ij}}\left[\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{Q}e^{-\mathbf{A}^{\top}\tau}\,d\tau\right]e^{\mathbf{A}^{\top}\Delta t_{ij}}.$
$\int_{0}^{\Delta t_{ij}}e^{\mathbf{A}(\Delta t_{ij}-\tau)}\mathbf{Q}e^{\mathbf{A}^{\top}(\Delta t_{ij}-\tau)}\,d\tau.$
Using the change of variables $s=\Delta t_{ij}-\tau\implies ds=-d\tau$ and flipping the limits, we get:
$\mathbf{V}_{F}(\Delta t_{ij})=\int_{0}^{\Delta t_{ij}}e^{\mathbf{A}s}\mathbf{Q}e^{\mathbf{A}^{\top}s}\,ds.$
which is a solution of the differential Lyapunov equation:
$\frac{d}{ds}\mathbf{V}_{F}(s)=\mathbf{A}\mathbf{V}_{F}(s)+\mathbf{V}_{F}(s)\mathbf{A}^{\top}+\mathbf{Q},$
with initial condition $\mathbf{V}_{F}(0)=\mathbf{0}$.
Recall that the estimate of the state is:
$\mathbf{\hat{x}}_{i}(t_{j}):=e^{\mathbf{-A}\Delta t_{ij}}\mathbf{x}(t_{j})=\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}d\mathbf{w}(\tau).$
Hence, $\mathbf{\hat{x}}_{i}(t_{j})\sim\mathcal{N}\left(\mathbf{x}(t_{i}),\mathbf{V}_{B}(\Delta t_{ij})\right)$, where: $\mathbf{V}_{B}(\Delta t_{ij})$ quantifies the uncertainty of the estimate $\mathbf{\hat{x}}_{i}(t_{j})$ of the state $\mathbf{x}(t_{i})$ at an earlier time $t_{i}$, given the state $\mathbf{x}(t_{j})$ at a later time $t_{j}$.
Conversely, if we observe the state at an earlier time $t_{j}$ and wish to quantify our uncertainty about an estimate of its state at a later time $t_{i}$ (i.e. the causal case), we can switch indices $i$ and $j$ and solve for $\mathbf{x}(t_{j})$:
$\mathbf{x}(t_{i})=e^{\mathbf{A}(t_{i}-t_{j})}\mathbf{x}(t_{j})+\int_{t_{j}}^{t_{i}}e^{\mathbf{A}(t_{i}-\tau)}\mathbf{G}d\mathbf{w}(\tau)$
$e^{-\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{j})+\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}d\mathbf{w}(\tau)\right).$
$\mathbf{x}(t_{j})=e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i})-\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}d\mathbf{w}(\tau)$
So:
$\mathbf{x}(t_{j})\sim\mathcal{N}\left(e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i}),\mathbf{V}_{B}(-\Delta t_{ij})\right)$
(Note that, according to our definition, $\Delta t_{ij}<0$ in the causal case, so $\Delta t_{ij}=|\Delta t_{ij}|$.)
Hence, in an LTI with Gaussian noise, the covariance of the state propagated forward in time under $\mathbf{A}$ is equal to the covariance of the state propagated backward in time under $\mathbf{A}$.
We again form an estimate:
$\mathbf{\hat{x}}_{i}(t_{j})=e^{-\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{j})=\mathbf{x}(t_{i})-e^{-\mathbf{A}\Delta t_{ij}}\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}d\mathbf{w}(\tau)$
which has distribution: $\mathbf{\hat{x}}_{i}(t_{j})\sim\mathcal{N}\left(\mathbf{x}(t_{i}),\mathbf{V}_{F}(|\Delta t_{ij}|)\right)$,
where $\mathbf{V}_{F}(|\Delta t_{ij}|)=e^{-\mathbf{A}\Delta t_{ij}}\mathbf{V}_{B}(|\Delta t_{ij}|)e^{-\mathbf{A}^{\top}\Delta t_{ij}}=e^{\mathbf{A}|\Delta t_{ij}|}\mathbf{V}_{B}(|\Delta t_{ij}|)e^{\mathbf{A}^{\top}|\Delta t_{ij}|}$.
So, in general, $\mathbf{\hat{x}}_{i}(t_{j})\sim\mathcal{N}\left(\mathbf{x}(t_{i}),\mathbf{V}(\Delta t_{ij})\right)$, where the propagated state covariance is:
[…]
Recall that the measurement at time $t_{j}$ is:
$\mathbf{z}(t_{j})=\mathbf{C}e^{\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{i})+\int_{0}^{\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}d\mathbf{w}(\tau)\right)+\mathbf{v}(t_{j}).$
The measurement process is distributed as:
$\mathbf{z}(t_{j})\sim\mathcal{N}\left(\mathbf{C}e^{\mathbf{A}\Delta t_{ij}}\mathbf{x}(t_{i}),\mathbf{C}\mathbf{V}_{F}(\Delta t_{ij})\mathbf{C}^{\top}+\mathbf{R}\right)$
Recall that our "pulled-forward" estimated measurement is defined as:
$\mathbf{\hat{z}}_{i}(t_{j}):=\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})\mathbf{z}(t_{j})\quad\text{where}\quad\bm{\Phi}_{\mathbf{z}}(\Delta t):=\mathbf{C}e^{\mathbf{-A}\Delta t}\mathbf{C}^{-1}.$
In terms of the "pulled-forward" state estimate, this is:
$\mathbf{\hat{z}}_{i}(t_{j})=\mathbf{C}e^{\mathbf{-A}\Delta t_{ij}}\mathbf{x}(t_{j})+\mathbf{C}e^{\mathbf{-A}\Delta t_{ij}}\mathbf{C}^{-1}\mathbf{v}(t_{j})$
$\mathbf{C}\mathbf{\hat{x}}_{i}(t_{j})+\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})\mathbf{v}(t_{j})$
Hence,
$\mathbf{\hat{z}}_{i}(t_{j})\sim\mathcal{N}\left(\mathbf{z}^{C}(t_{i}),\mathbf{V}^{C}(\Delta t_{ij})\right)$
where the propagated measurement covariance is:
$\mathbf{V}^{C}(\Delta t_{ij})=\mathbf{C}\mathbf{V}(\Delta t_{ij})\mathbf{C}^{\top}+\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})\mathbf{R}\bm{\Phi}_{\mathbf{z}}^{\top}(\Delta t_{ij})$
We can also define a propagated state precision matrix $\mathbf{P}(\Delta t_{ij})=\big(\mathbf{V}(\Delta t_{ij})\big)^{-1}$ and a propagated measurement precision matrix $\mathbf{P}^{C}(\Delta t_{ij})=\big(\mathbf{V}^{C}(\Delta t_{ij})\big)^{-1}$. Note that $\mathbf{V}(\Delta t)$ and $\mathbf{V}^{C}(\Delta t)$ and their inverses are positive definite (PD) and symmetric for all $\Delta t$, provided that $\mathbf{Q},\mathbf{R}$ are PD and symmetric.
A.2 Batch Maximum Likelihood Trajectory Estimation
We wish to estimate the trajectory of the system from $m$ measurements taken at times $\{t_{1},...,t_{m}\}$, which we denote $\{\mathbf{z}(t_{1}),...,\mathbf{z}(t_{m})\}$. We can do so by solving a separate maximum likelihood estimation (MLE) problem for each $\mathbf{x}_{i}$.
Let $\mathbf{x}_{i}=\mathbf{x}(t_{i})$, $\mathbf{\hat{x}}_{ij}=\mathbf{\hat{x}}_{i}(t_{j})$ and $\mathbf{V}_{ij}=\mathbf{V}(\Delta t_{ij})$. The log-likelihood is:
$\ell=-\frac{d_{x}m}{2}\log(2\pi)-\frac{1}{2}\sum_{j=1}^{m}\log|\det(\mathbf{V}_{ij})|-\frac{1}{2}\sum_{j=1}^{m}\left(\mathbf{\hat{x}}_{ij}-\mathbf{x}_{i}\right)^{\top}\mathbf{V}_{ij}^{-1}\left(\mathbf{\hat{x}}_{ij}-\mathbf{x}_{i}\right)$
Taking the derivative with respect to $\mathbf{x}_{i}$:
$\frac{\partial\ell}{\partial\mathbf{x}_{i}}=\sum_{j}\left(\mathbf{x}_{i}^{\top}\left(\mathbf{V}_{ij}^{-1}+\mathbf{V}_{ij}^{-\top}\right)-2\left(\mathbf{V}_{ij}^{-1}\mathbf{\hat{x}}_{ij}\right)^{\top}\right)=\mathbf{0}.$
Defining a precision matrix $\mathbf{P}_{ij}=\mathbf{V}_{ij}^{-1}$, the MLE solution for $\mathbf{x}(t_{i})$ is a precision-weighted estimator:
$\mathbf{\bar{x}}_{i}:=\left(\sum_{j}\mathbf{P}_{ij}\right)^{-1}\sum_{j}\mathbf{P}_{ij}\mathbf{\hat{x}}_{ij}$
Letting $\mathbf{z}_{i}=\mathbf{z}(t_{i})$, $\mathbf{\hat{z}}_{ij}=\mathbf{\hat{z}}_{i}(t_{j})$ and $\mathbf{V}^{C}_{ij}=\mathbf{V}^{C}(\Delta t_{ij})$, and repeating the above with the PDF for $\mathbf{\hat{z}}_{ij}$, the MLE solution for $\mathbf{z}^{C}(t_{i})$ is:
$\mathbf{\bar{z}}_{i}:=\left(\sum_{j}\mathbf{P}_{ij}^{C}\right)^{-1}\sum_{j}\mathbf{P}_{ij}^{C}\mathbf{\hat{z}}_{ij},\quad\text{where}\quad\mathbf{P}_{ij}^{C}=\mathbf{V}_{ij}^{C-1}.$
For the causal case:
$\mathbf{\bar{z}}_{i}:=\left(\sum_{j\leq i}\mathbf{P}_{ij}^{C}\right)^{-1}\sum_{j\leq i}\mathbf{P}_{ij}^{C}\mathbf{\hat{z}}_{ij}$
Repeating this computation for all $t_{i}$ produces an estimate of the full trajectory.
After estimating $\mathbf{\bar{z}}_{i}$, we can predict the next measurement $\mathbf{z}(t_{i+1})$ using the system dynamics:
$\mathbf{z}(t_{i+1})=\mathbf{C}e^{\mathbf{A}\Delta t}\mathbf{C}^{-1}\mathbf{z}^{C}(t_{i})+\mathbf{C}\int_{0}^{\Delta t}e^{-\mathbf{A}\tau}\mathbf{G}d\mathbf{w}(\tau)+\mathbf{v}(t_{i+1}),$
where $\Delta t=t_{i+1}-t_{i}$. Since the noise terms are zero mean and independent, the minimum mean squared error (MMSE) prediction given $\mathbf{\bar{z}}_{i}$ is:
$\mathbf{z}^{+}(t_{i+1}):=\mathbb{E}[\mathbf{z}(t_{i+1})\mid\mathbf{\bar{z}}_{i}]=\mathbf{C}e^{\mathbf{A}\Delta t}\mathbf{C}^{-1}\mathbf{\bar{z}}_{i}.$
which matches the standard approach in Kalman Filtering of propagating the current estimate forward through the deterministic part of the dynamics.
A.3 Closed-form Precision Matrix Calculation for Diagonalizable Matrices
The estimation procedure above requires computing $\mathbf{P}^{C}(\Delta t_{ij})$, which is generally expensive, as it involves integration and matrix inversion. However, if we assume that $\mathbf{A}$ is diagonalizable, that is we can write it in the form $\mathbf{A}=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1}\in\mathbb{R}^{d\times d}$, where $\bm{\Lambda}\in\mathbb{C}^{d\times d}$ is a diagonal matrix and $\mathbf{S}\in\mathbb{C}^{d\times d}$ an invertible matrix, the backward-propagated covariance becomes:
$\mathbf{V}_{B}(\Delta t)=\int_{0}^{|\Delta t|}e^{\mathbf{-A}\tau}\mathbf{Q}e^{\mathbf{-A}^{\dagger}\tau}d\tau$
$\mathbf{S}\left(\int_{0}^{|\Delta t|}e^{-\bm{\Lambda}\tau}(\mathbf{S}^{-1}\mathbf{Q}\mathbf{S}^{-\dagger})e^{-\bm{\Lambda}^{\dagger}\tau}d\tau\right)\mathbf{S}^{\dagger}$
Let $\mathbf{G}:=\mathbf{S}^{-1}\mathbf{Q}\mathbf{S}^{-\dagger}$ and $\bm{\Lambda}_{\mathbf{V}B}(\Delta t):=\int_{0}^{|\Delta t|}e^{-\bm{\Lambda}\tau}(\mathbf{S}^{-1}\mathbf{Q}\mathbf{S}^{-\dagger})e^{-\bm{\Lambda}^{\dagger}\tau}d\tau$.
Note that $(e^{-\bm{\Lambda}t}\mathbf{G}e^{-\bm{\Lambda}^{\dagger}t})_{kl}=\mathbf{G}_{kl}e^{-\lambda_{k}t}e^{-\lambda^{*}_{l}t}=\mathbf{G}_{kl}e^{-(\lambda_{k}+\lambda^{*}_{l})t}$. Hence, the $kl$th component of $\bm{\Lambda}_{\mathbf{V}B}(\Delta t)$ is:
$\bm{\Lambda}_{\mathbf{V}B}(\Delta t)_{kl}=\mathbf{G}_{kl}\int_{0}^{|\Delta t|}e^{-(\lambda_{k}+\lambda^{*}_{l})\tau}d\tau$
$\begin{cases}\mathbf{G}_{kl}\,\dfrac{1-e^{-(\lambda_{k}+\lambda^{*}_{l})|\Delta t|}}{\lambda_{k}+\lambda^{*}_{l}},&\lambda_{k}+\lambda^{*}_{l}\neq 0\\[5.69054pt] \mathbf{G}_{kl}\,|\Delta t|,&\lambda_{k}+\lambda^{*}_{l}=0\end{cases}$
If we assume that $\mathbf{Q}=\mathbf{S}\bm{\Lambda}_{Q}\mathbf{S}^{\dagger}$, i.e. that the process noise acts independently along the decoupled modes of the system, then $\mathbf{G}=\bm{\Lambda}_{Q}$ and $\bm{\Lambda}_{\mathbf{V}B}$ becomes diagonal:
$\bm{\Lambda}_{\mathbf{V}B}(|\Delta t_{ij}|)=\bm{\Lambda}_{Q}\frac{e^{-2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}-\mathbf{I}}{-2\mathrm{Re}(\bm{\Lambda})}$
where division is taken element-wise (ignoring, for now, the case $\mathrm{Re}(\lambda_{k})=0$).
For the causal case, the forward-propagated covariance is:
$\mathbf{V}_{F}(|\Delta t_{ij}|)=e^{\mathbf{A}|\Delta t_{ij}|}\mathbf{V}_{B}(|\Delta t_{ij}|)e^{\mathbf{A}^{\top}|\Delta t_{ij}|}$
$\mathbf{S}e^{\bm{\Lambda}|\Delta t_{ij}|}\bm{\Lambda}_{\mathbf{V}B}(|\Delta t_{ij}|)e^{\bm{\Lambda}^{\dagger}|\Delta t_{ij}|}\mathbf{S}^{\dagger}$
$\mathbf{S}e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}\left(\bm{\Lambda}_{Q}\frac{e^{-2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}-\mathbf{I}}{-2\mathrm{Re}(\bm{\Lambda})}\right)\mathbf{S}^{\dagger}$
$\mathbf{S}\bm{\Lambda}_{\mathbf{V}}(|\Delta t_{ij}|)\mathbf{S}^{\dagger}$
where
$\bm{\Lambda}_{\mathbf{V}}(|\Delta t_{ij}|)=\bm{\Lambda}_{Q}\frac{\mathbf{I}-e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}}{-2\mathrm{Re}(\bm{\Lambda})}$
$\bm{\Lambda}_{\mathbf{V}}(|\Delta t_{ij}|)$ is positive semi-definite provided $\bm{\Lambda}_{Q}\geq\mathbf{0}$ (which is true by definition). (Here, the inequalities are applied element-wise.)
The propagated precision matrix $\mathbf{P}(t)$ can be easily obtained by inverting $\bm{\Lambda}_{\mathbf{V}}(\Delta t)$ element-wise: $\mathbf{P}(\Delta t)=\mathbf{S}^{-\dagger}\bm{\Lambda}_{\mathbf{P}}(\Delta t)\mathbf{S}^{-1}$, where $\bm{\Lambda}_{\mathbf{P}}(\Delta t)=\bm{\Lambda}_{\mathbf{V}}(\Delta t)^{-1}$.
To include the measurement model, recall that the propagated measurement covariance is:
$\mathbf{V}_{C}(\Delta t_{ij})=\mathbf{C}\mathbf{V}(\Delta t_{ij})\mathbf{C}^{\dagger}+\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})\mathbf{R}\mathbf{M}^{\dagger}(\Delta t_{ij}),\quad\text{where}\quad\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})=\mathbf{C}e^{\mathbf{-A}\Delta t_{ij}}\mathbf{C}^{-1}.$
If $\mathbf{C}=\mathbf{S}\bm{\Lambda}_{C}\mathbf{S}^{-1}$, this becomes: $\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})=\mathbf{S}\bm{\Lambda}_{C}e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{C}^{-1}\mathbf{S}=\mathbf{S}e^{-\bm{\Lambda}\Delta t_{ij}}\mathbf{S}^{-1}$. If, in addition, $\mathbf{R}=\mathbf{S}\bm{\Lambda}_{R}\mathbf{S}^{\dagger}$, then:
$\mathbf{V}^{C}(\Delta t_{ij})=\mathbf{S}\bm{\Lambda}_{C}\bm{\Lambda}_{\mathbf{V}}(\Delta t_{ij})\bm{\Lambda}_{C}^{\dagger}\mathbf{S}^{\dagger}+\mathbf{S}e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{R}e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}\mathbf{S}^{\dagger}$
$\mathbf{S}\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})\mathbf{S}^{\dagger}$
where:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\big|\bm{\Lambda}_{C}\big|^{2}\bm{\Lambda}_{\mathbf{V}}(\Delta t_{ij})+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R}$
$\big|\bm{\Lambda}_{C}\big|^{2}\bm{\Lambda}_{Q}\frac{\mathbf{I}-e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}}{-2\mathrm{Re}(\bm{\Lambda})}+e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}\bm{\Lambda}_{R}$
(Stable dynamics attenuate the effect of measurement noise.)
Finally, $\mathbf{P}^{C}(\Delta t_{ij})=\mathbf{S}^{-\dagger}\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{ij})\mathbf{S}^{-1}$, where $\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{ij})=\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})^{-1}$.
Letting $\mathbf{z}_{\mathbf{s}_{i}}=\mathbf{S}^{-1}\mathbf{z}_{i}$, and $\mathbf{\hat{z}}_{\mathbf{s}ij}:=e^{-\bm{\Lambda}\Delta t_{ij}}\mathbf{z}_{\mathbf{s}j}$, the MLE is now tractable:
$\mathbf{\bar{z}}_{i}=\mathbf{S}\left[\left(\sum_{j\leq i}\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{ij})\right)^{-1}\sum_{j\leq i}\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{ij})\mathbf{\hat{z}}_{\mathbf{s}ij}\right]\mathbf{S}^{-1}.$
For the non-causal case,
$\bm{\Lambda}_{\mathbf{V}}(\Delta t)=\bm{\Lambda}_{\mathbf{V}B}(\Delta t)=\bm{\Lambda}_{Q}\frac{e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t}-\mathbf{I}}{-2\mathrm{Re}(\bm{\Lambda})}$
So,
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\big|\bm{\Lambda}_{C}\big|^{2}\bm{\Lambda}_{\mathbf{V}}(\Delta t_{ij})+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R}$
$\big|\bm{\Lambda}_{C}\big|^{2}\bm{\Lambda}_{Q}\frac{e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}-\mathbf{I}}{-2\mathrm{Re}(\bm{\Lambda})}+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R}$
In summary, the general expression for the propagated measurement covariance, including both the causal and non-causal cases, and all values of $\operatorname{Re}(\lambda)$ is:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\text{sign}(\Delta t_{ij})\big|\bm{\Lambda}_{C}\big|^{2}\bm{\Lambda}_{Q}\mathcal{K}(\bm{\Lambda},\Delta t_{ij})+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R},$
$\mathcal{K}(\bm{\Lambda},\Delta t):=\operatorname{diag}\!\big(\varphi(\lambda_{k},\Delta t)\big)_{k=1}^{d},\qquad\varphi(\lambda,\Delta t):=\begin{cases}\dfrac{e^{-2\operatorname{Re}(\lambda)\,\Delta t}-1}{-2\operatorname{Re}(\lambda)},&\operatorname{Re}(\lambda)\neq 0,\\[6.0pt] \Delta t,&\operatorname{Re}(\lambda)=0,\end{cases}$
This is bounded for all $\Delta t_{ij}$ in the causal case if and only if every component of $\mathrm{Re}(\bm{\Lambda})$ is negative.
In particular, consider $\mathbf{A}=\alpha\mathbf{I}+\bm{\Omega}\in\mathbb{R}^{d\times d}$ where $\alpha\in\mathbb{R}$ and $\bm{\Omega}\in\mathbb{R}^{d\times d}$ is diagonalizable over $\mathbb{C}$, and has strictly imaginary eigenvalues, i.e. $\bm{\Omega}=\mathbf{S}\bm{\Lambda}_{\Omega}\mathbf{S}^{-1}$, where $\mathbf{S}\in\mathbb{C}^{d\times d}$, $\bm{\Lambda}_{\Omega}=\text{diag}(\lambda_{\Omega,1},\dots\lambda_{\Omega,d})$, with $\lambda_{\Omega,k}\in i\mathbb{R}$. If we also assume that the process and measurement noise are isotropic $\bm{\Lambda}_{Q}=\sigma^{2}\mathbf{I}$, $\bm{\Lambda}_{R}=\eta^{2}\mathbf{I}$, and $\bm{\Lambda}_{C}=\lambda_{C}\mathbf{I}$, then we have:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\sigma_{V}^{2}(\Delta t_{ij})\mathbf{I},\quad\text{where}\quad\sigma_{V}^{2}(\Delta t_{ij}):=\text{sign}(\Delta t_{ij})\big|\lambda_{C}\big|^{2}\sigma^{2}\frac{e^{-2\alpha\Delta t_{ij}}-1}{-2\alpha}+e^{-2\alpha\Delta t_{ij}}\eta^{2}$
We can also generalize the above to the case where continuous-time measurements are incorporated. In this case, the covariance evolution is governed by the differential Riccati equation, which modifies the differential Lyapunov equation to include the measurement update:
$\mathbf{\dot{V}}_{F}(t)=\mathbf{A}\mathbf{V}_{F}(t)+\mathbf{V}_{F}(t)\mathbf{A}^{\top}+\mathbf{Q}-\mathbf{V}_{F}(t)\mathbf{C}^{\top}\mathbf{R}^{-1}\mathbf{C}\mathbf{V}_{F}(t)$
Again assuming simultaneously diagonalizble $\mathbf{A},\mathbf{C},\mathbf{Q},\mathbf{R}$, the matrix equations decouple into separate scalar Riccati differential equations for each diagonal element, with $k$th diagonal term given by:
$\dot{\lambda}_{vk}(t)=2\mathrm{Re}(\lambda_{i})\lambda_{vk}(t)+\lambda_{Qk}-\frac{\big|\lambda_{Ck}\big|^{2}}{\lambda_{Rk}}\lambda_{vk}(t)^{2}$
Letting $\alpha_{k}=\mathrm{Re}(\lambda_{k})$, $\beta_{k}=\sqrt{\alpha_{k}^{2}+\frac{\lambda_{Qk}}{\lambda_{Rk}}\big|\lambda_{Ck}\big|^{2}}$, $\gamma_{k}=\frac{1}{\beta_{k}}\log\left(\frac{\beta_{k}-\alpha_{k}}{\beta_{k}+\alpha_{k}}\right)$, the $k$th element of $\bm{\Lambda}_{\mathbf{V}}$ becomes:
$\lambda_{vk}(\Delta t_{ij})=\frac{\lambda_{Rk}}{\big|\lambda_{Ck}\big|^{2}}\left(\beta_{k}\tanh\left(\frac{\beta_{k}}{2}\left(\Delta t_{ij}-\gamma_{k}\right)\right)+\alpha_{k}\right)$
A.4 Variational Formulation of Robust Estimation
Suppose we have vector-valued measurements:
$\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big(\bm{\mu}_{i},\mathbf{\Sigma}_{ij}\big),$
where $\bm{\mu}_{i}\in\mathbb{R}^{d}$ is the output of the true latent state at time $i$ (which we distinguish from $\mathbf{z}^{C}_{i}$, which denotes the output under the latent state generated by our LTI model), and $\mathbf{\Sigma}_{ij}$ is the unknown measurement covariance. Our objective is to maximize the marginal log-likelihood:
$\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i})\;=\;\log\int p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{ij})\,p(\mathbf{\Sigma}_{ij})\,d\mathbf{\Sigma}_{ij}.$
i.e. a Gaussian scale mixture. Introducing a variational distribution $q(\mathbf{\Sigma}_{ij})$ over the unknown covariance and applying Jensen’s inequality yields the Evidence Lower Bound (ELBO):
$\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i})\geq\mathcal{L}(q,\bm{\mu}_{i}):=\mathbb{E}_{q(\mathbf{\Sigma}_{ij})}\big[\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{ij})\big]-\mathrm{KL}\big(q(\mathbf{\Sigma}_{ij})\,\big|\,p(\mathbf{\Sigma}_{ij})\big).$
Assuming $\mathbf{\Sigma}_{ij}$ is diagonal, the expected log-likelihood term becomes:
$\mathbb{E}_{q}\big[\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{ij})\big]=-\frac{1}{2}\mathbb{E}_{q}[\log\det\mathbf{\Sigma}_{ij}]-\frac{1}{2}(\mathbf{\hat{z}}_{ij}-\bm{\mu}_{i})^{\top}\mathbb{E}_{q}[\mathbf{\Sigma}_{ij}^{-1}](\mathbf{\hat{z}}_{ij}-\bm{\mu}_{i})+\text{const}.$
Define the expected precision: $\mathbf{W}_{ij}:=\mathbb{E}_{q}\big[\mathbf{\Sigma}_{ij}^{-1}\big].$
Then, the expected log-likelihood is proportional to a weighted quadratic form:
$\mathbb{E}_{q}\big[\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{ij})\big]\;\propto\;-\frac{1}{2}(\mathbf{\hat{z}}_{ij}-\bm{\mu}_{i})^{\top}\mathbf{W}_{ij}(\mathbf{\hat{z}}_{ij}-\bm{\mu}_{i}).$
We model the unknown measurement precision as a scaled version of the dynamics-based precision:
$\mathbf{\Sigma}_{ij}^{-1}=\tau_{ij}\,\mathbf{P}^{C}_{ij},$
where $\mathbf{P}^{C}_{ij}$ is the precision predicted by the LTI model and $\tau_{ij}>0$ is a latent scalar that modulates the overall confidence in this observation.
Conditional on $\tau_{ij}$, the Gaussian likelihood for the measurement $\mathbf{\hat{z}}_{ij}$ is:
$p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\tau_{ij})=\mathcal{N}\big(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},(\tau_{ij}\mathbf{P}^{C}_{ij})^{-1}\big).$
The optimal variational distribution over $\tau_{ij}$ is proportional to the prior times the likelihood:
$q^{*}(\tau_{ij})\propto p(\tau_{ij})\,p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\tau_{ij}).$
To enable Bayesian inference, we place a conjugate prior on $\tau_{ij}$. Using a Gamma prior:
$\tau_{ij}\sim\mathrm{Gamma}(\alpha,\beta).$
Since the prior is Gamma and the likelihood is Gaussian with a precision multiplier, the posterior is again a Gamma distribution:
$q^{*}(\tau_{ij})=\mathrm{Gamma}\!\Big(\alpha+\frac{d}{2},\;\beta+\frac{1}{2}\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}\Big),\quad\mathbf{r}_{ij}:=\mathbf{\hat{z}}_{ij}-\bm{\mu}_{i}.$
The expected precision under the posterior is:
$\mathbf{W}_{ij}:=\mathbb{E}_{q^{*}}[\mathbf{\Sigma}_{ij}^{-1}]=\mathbb{E}_{q^{*}}[\tau_{ij}]\,\mathbf{P}^{C}_{ij}=w_{ij}\,\mathbf{P}^{C}_{ij},$
where:
$w_{ij}=\frac{\alpha+d/2}{\beta+\frac{1}{2}\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}}.$
The corresponding (marginal) likelihood $p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i})$ is a Student’s $t$. This is equivalent to an M-estimator with loss:
$\rho(\mathbf{r}_{ij})\;=\;\Big(\alpha+\tfrac{d}{2}\Big)\,\log\!\Big(\beta+\tfrac{1}{2}\,\mathbf{r}_{ij}^{\top}\mathbf{P}_{ij}\,\mathbf{r}_{ij}\Big),\qquad\mathbf{r}_{ij}:=\bm{\mu}_{i}-\mathbf{\hat{z}}_{ij}.$
Dividing numerator and denominator by $\beta$ and setting the numerator to $1$ (since it is canceled by the normalization), there is only one independent constant, $\nu=2\beta$, which controls the residual scaling:
$w_{ij}=\frac{1}{1+\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}/\nu}.$
The term $\mathbf{r}^{\top}\mathbf{V}^{-1}\mathbf{r}$ is known as the Mahalanobis distance, which quantifies the difference between a vector $\mathbf{r}$ and a distribution with covariance $\mathbf{V}$. Here, $\mathbf{P}^{C}_{ij}$ encodes the directional confidence from the LTI model, while $w_{ij}$ adaptively rescales the precision based on the observed residual $\mathbf{r}_{ij}$, down-weighting measurements with larger than expected deviations.
The metric $1/(1+\mathbf{r}^{\top}\mathbf{P}\mathbf{r}/\nu)$ behaves similarly to the unnormalized softmax: if one residual is much smaller than the others, it dominates, while if all residuals are similar and not too large, the output approximates a uniform weighting. The key difference arises when all residuals are small relative to $\nu$: unlike softmax, which picks a dominant term whenever relative differences are large, this similarity score reflects uncertainty and maintains a roughly uniform attention distribution. In this sense, $\nu$ acts as a kind of noise floor.
Using the fact that $\log(1+x)\approx x,\,\,\text{for }x\ll 1$, we can expand the rational form of $w_{ij}$:
$w_{ij}=\frac{1}{1+\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}/\nu}\approx\exp(-\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}/\nu).$
which results in a more aggressive down-weighting of large residuals (and $\nu$ becomes a temperature parameter).
Appendix B Additional Material
B.1 Separable Approximation of the Precision Tensor
This section explores an alternative approach to address the memory bottleneck encountered in the diagonalizable (full tensor) case of Adaptive Filter Attention, which required $\mathcal{O}(m^{2}d)$ memory.
Recall that:
$\mathbf{D}[i,j]=\mathcal{P}^{C}[k,i,j]\cdot\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}+\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\cdot\Big[\mathcal{P}^{C}[k,i,j]\cdot\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}$
$-2\mathrm{Re}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\cdot\mathcal{P}^{C}[k,i,j]\cdot\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
Our previous approach — imposing a model with isotropic decay $\alpha$ and noise $\sigma,\eta$ — achieved $\mathcal{O}(m^{2})$ memory, thereby allowing the precision tensor $\mathcal{P}^{C}$ to be factored out of the summation over $k$. Here, we attempt to maintain the full $k$-dependence of the time-dependent precision tensor $\mathcal{P}^{C}[k,i,j]$ while still achieving $\mathcal{O}(m^{2}+md)$ memory. This is achieved by seeking a low-rank (separable) approximation of the precision kernel.
In particular, suppose we could approximate the precision tensor as a product of separable $i$ and $j$ terms:
$\mathcal{P}^{C}[k,i,j]\approx\mathbf{c}[k]+\mathbf{f}[k,i]\cdot\mathbf{g}[k,j].$
Then, we could write:
$\mathbf{D}[i,j]=\mathbf{c}[k]\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\Big[\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}+\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}-2\mathrm{Re}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\cdot\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
$+\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\Big[\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}+\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}$
$-2\mathrm{Re}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\cdot\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]\cdot\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
This can now be simplified into a sum of matrix products, which can be cheaply computed without storing a $d\times m\times m$ dimensional tensor:
$\mathbf{D}=\mathbf{c}[k]\Big[\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\big)^{\top}\big|\mathbf{\hat{Z}}_{\mathbf{k}}\big|^{2}+\sum_{k}\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\big|\mathbf{\hat{Z}}_{\mathbf{q}}\big|^{2}\big)-2\mathrm{Re}\Big(\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big)^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\Big)\Big]$
$+\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\mathbf{f}\big)^{\top}\big(\big|\mathbf{\hat{Z}}_{\mathbf{k}}\big|^{2}\odot\mathbf{g}\big)+\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\big|\mathbf{\hat{Z}}_{\mathbf{q}}\big|^{2}\odot\mathbf{f}\big)^{\top}\mathbf{g}$
$-2\mathrm{Re}\Big(\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\odot\mathbf{f}\big)^{\dagger}\big(\mathbf{\hat{Z}}_{\mathbf{k}}\odot\mathbf{g}\big)\Big)$
Recall that the precision tensor is defined by:
$\mathcal{P}^{C}[k,i,j]=\left(\lambda_{C,k}^{2}\odot\lambda_{Q,k}\odot\frac{1-e^{2\mathrm{Re}(\lambda_{k})(t_{i}-t_{j})}}{-2\mathrm{Re}(\lambda_{k})}+\lambda_{R,k}\odot e^{2\mathrm{Re}(\lambda_{k})(t_{i}-t_{j})}\right)^{-1}$
Let:
$\lambda_{\text{eff},k}=\frac{\lambda_{C,k}^{2}\,\lambda_{Q,k}}{-2\,\mathrm{Re}(\lambda_{k})}.$
Note that $\lambda_{\text{eff},k}>0$ for $\mathrm{Re}(\lambda_{k})<0$. The precision kernel is then,
$\mathcal{P}^{C}[k,i,j]=\left(\lambda_{\text{eff},k}+\big(-\lambda_{\text{eff},k}+\lambda_{R,k}\big)\,e^{2\,\mathrm{Re}(\lambda_{k})\,(t_{i}-t_{j})}\right)^{-1}$
$\frac{1}{\lambda_{\text{eff},k}}\left(1+\lambda_{r,k}\,e^{2\,\mathrm{Re}(\lambda_{k})\,(t_{i}-t_{j})}\right)^{-1},\quad\text{where }\lambda_{r,k}=-1+\frac{\lambda_{R,k}}{\lambda_{\text{eff},k}}>-1$
We use the separable approximation:
$(1+ae^{bx})^{-1}\approx 1-ce^{-dx},$
with constants chosen to match the value and slope at $x=0$.
We obtain the following separable approximation:
$\mathcal{P}^{C}_{k}(\Delta t)\approx\frac{1}{\lambda_{\text{eff},k}}\left(1-\frac{\lambda_{r,k}}{1+\lambda_{r,k}}\,e^{\tfrac{2\,\mathrm{Re}(\lambda_{k})\,\Delta t}{1+\lambda_{r,k}}}\right)=\frac{1}{\lambda_{\text{eff},k}}\left[1-\left(1-\frac{\lambda_{\text{eff},k}}{\lambda_{R,k}}\right)e^{-\tfrac{|\lambda_{C,k}|^{2}\,\lambda_{Q,k}}{\lambda_{R,k}}\,\Delta t}\right]$
This approximation has several useful properties:
  1. 1.
    It matches the true kernel at $\Delta t=0$ and converges to the same fixed value as $\Delta t\to\infty$.
  2. 2.
    It preserves the monotonic ordering of the true kernel.
  3. 3.
    It is always below the true kernel as a function of $\Delta t$, which means the approximation biases the estimate toward more local (recent) history.
Although the approximation introduces a finite-sample bias toward recent history and increases the effective variance, the resulting estimator is asymptotically consistent due to the kernel’s preservation of the boundary conditions. Furthermore, the $k$-dependence of $\mathbf{P}^{C}$ restores the ability to learn a dimension-specific memory profile, making the resulting mechanism far more expressive than scalar-precision alternatives.
Recall that the estimate is:
$\mathbf{\bar{Z}_{v}}[k,i]=\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\sum_{j\leq i}\frac{\mathcal{P}^{C}[k,i,j]}{1+\mathbf{D}[i,j]}\cdot\mathbf{\hat{Z}}_{\mathbf{v}}[k,j]$
The normalization term can be pulled outside the sum:
$\mathbf{\bar{Z}_{v}}[k,i]=\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\mathbf{N}[k,i]^{-1}\sum_{j\leq i}\frac{\mathbf{c}[k]+\mathbf{f}[k,i]\mathbf{g}[k,j]}{1+\mathbf{D}[i,j]}\cdot\mathbf{\hat{Z}}_{\mathbf{v}}[k,j]$
where:
$\mathbf{N}[k,i]=\sum_{j\leq i}\frac{\mathbf{c}[k]+\mathbf{f}[k,i]\mathbf{g}[k,j]}{1+\mathbf{D}[i,j]}$
We then get two sums:
$\mathbf{\bar{Z}_{v}}[k,i]$ $=\mathbf{c}[k]\,\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\mathbf{N}[k,i]^{-1}\cdot\sum_{j\leq i}\frac{1}{1+\mathbf{D}[i,j]}\cdot\mathbf{\hat{Z}}_{\mathbf{v}}[k,j]$
$+\big(\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\mathbf{f}[k,i]\big)\cdot\mathbf{N}[k,i]^{-1}\cdot\sum_{j\leq i}\frac{1}{1+\mathbf{D}[i,j]}\cdot\big(\mathbf{g}[k,j]\cdot\mathbf{\hat{Z}}_{\mathbf{v}}[k,j]\big)$
The normalizer is:
$\mathbf{N}[k,i]=\Big(\mathbf{c}[k]\sum_{j\leq i}\frac{1}{1+\mathbf{D}[i,j]}+\mathbf{f}[k,i]\cdot\sum_{j\leq i}\frac{1}{1+\mathbf{D}[i,j]}\cdot\mathbf{g}[k,j]\Big)$
Define $\mathbf{W}[i,j]:=\frac{1}{1+\mathbf{D}[i,j]}$ and $\mathbf{W}_{\text{causal}}:=\mathbf{W}\odot\mathbf{M}_{\text{causal}}$, where $\mathbf{M}_{\text{causal}}$ is a causal mask. The estimate can be written in matrix form as:
$\mathbf{\bar{Z}}_{\mathbf{v}}=\frac{1}{\mathbf{N}}\odot\bm{\Phi}_{\mathbf{v}}\odot\mathbf{U}$
$\mathbf{N}:=\mathbf{c}\odot\,\big(\mathbf{1}_{d}\odot(\mathbf{1}_{m}^{\top}\mathbf{W}_{\text{causal}}^{\top}\,)\big)+\mathbf{f}\odot(\mathbf{g}\,\mathbf{W}_{\text{causal}}^{\top})$
$\mathbf{U}:=\mathbf{c}\odot\,\left(\mathbf{\hat{Z}}_{\mathbf{v}}\,\mathbf{W}_{\text{causal}}^{\top}\right)+\mathbf{f}\odot\left((\mathbf{g}\odot\mathbf{\hat{Z}}_{\mathbf{v}})\,\mathbf{W}_{\text{causal}}^{\top}\right)$
(and where $1/\mathbf{N}$ denotes element-wise division).
While this formulation reduces the memory complexity from $\mathcal{O}(m^{2}d)$ to $\mathcal{O}(m^{2}+md)$, the reliance on numerous element-wise operations introduces a high constant factor to the $\mathcal{O}(m^{2}d)$ time complexity, making it less scalable than the isotropic decay model.
B.2 Recurrent Formulation
We now describe a way to incorporate state-dependent weighting from the full batch MLE while retaining the recursive structure of the Kalman Filter. The key idea is to use the innovation residual at each step to reweight the update, and iterate within each time step to approximate the effect of all pairwise comparisons. This iteration again takes place in the eigenbasis.
Prediction Step (Include Process Noise).
At time $t_{i}$, suppose we have the posterior from the previous time step $\mathbf{\bar{z}}_{\mathbf{s},i-1}^{+}$ with covariance $\bm{\Lambda}_{\mathbf{V},i-1}^{+}$ and precision $\bm{\Lambda}_{\mathbf{P},i-1}^{+}=(\bm{\Lambda}_{\mathbf{V},i-1}^{+})^{-1}$. Propagating through the linear dynamics with process noise $\bm{\Lambda}_{Q}$, the prior for the current time step becomes:
$\mathbf{\bar{z}}_{\mathbf{s}i}^{-}=\bm{\Lambda}\mathbf{\bar{z}}_{\mathbf{s},i-1}^{+},\quad\bm{\Lambda}_{\mathbf{V},i}^{-}=\bm{\Lambda}\bm{\Lambda}_{\mathbf{V},i-1}^{+}\bm{\Lambda}^{\dagger}+\bm{\Lambda}_{Q}=\big|\bm{\Lambda}\big|^{2}\bm{\Lambda}_{\mathbf{V},i-1}^{+}+\bm{\Lambda}_{Q},\quad\bm{\Lambda}_{\mathbf{P},i}^{-}=(\bm{\Lambda}_{\mathbf{V},i}^{-})^{-1}.$
The propagated precision for a single time step is constant: $\bm{\Lambda}_{\mathbf{P}}^{C}:=\bm{\Lambda}_{\mathbf{P}}^{C}(\Delta t_{i,i-1})$.
Weighted Update (Measurement + IRKF Iteration).
Within each iteration $k$ at time $t_{i}$, compute the innovation:
$\mathbf{y}_{\mathbf{s}i}^{(k)}=\mathbf{z}_{\mathbf{s}i}-\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}.$
Define the nominal posterior precision from the measurement model:
$\bm{\Lambda}_{\mathbf{P},i}^{+(k)}=\bm{\Lambda}_{\mathbf{P},i}^{-}+\bm{\Lambda}_{\mathbf{P}}^{C},$
and compute a residual-based robust weight:
$w_{i}^{(k)}=\frac{1}{1+(\mathbf{y}_{\mathbf{s}i}^{(k)})^{\dagger}\bm{\Lambda}_{\mathbf{P},i}^{+(k)}\mathbf{y}_{\mathbf{s}i}^{(k)}/\nu}.$
The weighted posterior precision is then:
$\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)}=\bm{\Lambda}_{\mathbf{P},i}^{-}+w_{i}^{(k)}\,\bm{\Lambda}_{\mathbf{P}}^{C},$
Notice that the posterior information vector can be written as:
$\bm{\tilde{\eta}}_{\mathbf{s}i}^{+}=\bm{\eta}_{\mathbf{s}i}^{-}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{z}_{\mathbf{s}i}=\bm{\Lambda}_{\mathbf{P},i}^{-}\mathbf{\bar{z}}_{\mathbf{s}i}^{-}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{z}_{\mathbf{s}i}$
$\bm{\Lambda}_{\mathbf{P},i}^{-}\mathbf{\bar{z}}_{\mathbf{s}i}^{-}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{y}_{\mathbf{s}i}^{(k)}.$
So the total posterior mean in the eigenbasis becomes:
$\mathbf{\bar{z}}_{\mathbf{s}i}^{+(k)}=(\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)})^{-1}\bm{\tilde{\eta}}_{\mathbf{s}i}^{+}$
$(\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)})^{-1}\Big[\bm{\Lambda}_{\mathbf{P},i}^{+(k)}\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{y}_{\mathbf{s}i}^{(k)}\Big].$
$(\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)})^{-1}\Big[\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)}\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}+w_{i}^{(k)}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{y}_{\mathbf{s}i}^{(k)}\Big]=\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}+w_{i}^{(k)}(\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)})^{-1}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{y}_{\mathbf{s}i}^{(k)}.$
(the robust weight only scales the measurement contribution, not the prior).
Alternatively, we can use a damped update:
$\mathbf{\bar{z}}_{\mathbf{s}i}^{+(k)}=\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k)}+\delta\,w_{i}^{(k)}(\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+(k)})^{-1}\,\bm{\Lambda}^{C}_{\mathbf{P}}\mathbf{y}_{\mathbf{s}i}^{(k)}.$
where $\delta\in(0,1]$. Set the prior for the next inner iteration as: $\mathbf{\bar{z}}_{\mathbf{s}i}^{-(k+1)}=\mathbf{\bar{z}}_{\mathbf{s}i}^{+(k)}$.
Iterate until convergence to produce the final posterior $\mathbf{\bar{z}}_{\mathbf{s}i}^{+}$ and $\bm{\tilde{\Lambda}}_{\mathbf{P},i}^{+}$ for the current time step.
While the batch MLE uses the global information provided by pair-wise residuals, the innovation-based weight $w_{i}$ in the recursive update depends only on the current residual $r_{i}=\mathbf{z}_{i}-\mathbf{\bar{z}}_{i}^{-}$. However, when the dynamics are slow (small eigenvalues), $\Delta t$ is sufficiently small, and the residuals are not too large, $\mathbf{\bar{z}}_{\mathbf{s}i}^{-}$ effectively encodes past measurements. In this case, with enough inner iterations, IRKF should approximate the batch IRLS solution, while retaining the recursive structure of the Kalman Filter. The number of pair-wise comparisons for a sequence of length $m$ is then reduced from $\mathcal{O}(m^{2})$ to $\mathcal{O}(Km)$, where $K\ll m$ is the number of inner iterations.
At inference time, we can approximately unroll our learned model into a reweighted Kalman Filter. The outer iteration propagates latent states through the sequence for fast, online inference, while the inner iterations at each step approximate the effect of full pairwise attention without incurring $\mathcal{O}(m^{2})$ computations. The recursive precision update acts as an implicit normalization of the "attention" weights.
Input: Prior state $\mathbf{\bar{z}}_{0}^{-}$, measurements $\{\mathbf{\hat{z}}_{\mathbf{s}ii}\}_{i=1}^{m}$ Parameters: Complex linear layers $\mathcal{L}_{Q},\mathcal{L}_{K},\mathcal{L}_{V},\mathcal{L}_{P}$; dynamics/noise parameters: $\alpha,\bm{\omega},\sigma,\eta$; timestep $\Delta t$; softmax temperature $\tau_{s}$, residual weight $\delta$. Constants: Causal mask $\mathbf{M}_{\text{causal}}\in\{0,1\}^{m\times m}$, stability constant $\epsilon$.
Complex-valued Projections:
Lift real inputs to complex by concatenating zeros: $\mathbf{z}=\left[\mathbf{z}_{r},\mathbf{0}\right]^{\top}$
$\big(\operatorname{Re}(\mathbf{z_{v}}),\operatorname{Im}(\mathbf{z_{v}})\big)\leftarrow\mathcal{L}_{V}\big(\operatorname{Re}(\mathbf{z}),\operatorname{Im}(\mathbf{z})\big).$
1ex
State transition matrix:
$\Phi_{\alpha}:=e^{-\alpha^{2}\Delta t},\quad\bm{\Phi}_{\bm{\omega}}:=e^{\bm{\omega}\Delta t},\quad\bm{\Phi}:=\Phi_{\alpha}\bm{\Phi}_{\bm{\omega}}$
Innovation covariance: $\sigma^{2}_{V}=\sigma^{2}\kappa+\eta^{2}\Phi^{2}_{\alpha},\quad\kappa=\frac{1-\Phi^{2}_{\alpha}+\epsilon}{2\alpha^{2}+\epsilon}\mathbf{1}_{\{\alpha^{2}>\epsilon\}}+\Delta t\mathbf{1}_{\{\alpha^{2}\leq\epsilon\}}$
Initial conditions: $\mathbf{\bar{z}}^{+}_{\mathbf{v},0}=\mathbf{z_{v}},\quad p_{0}^{-(0)}=1/\sigma_{V}^{2}$
1ex
for $i\leftarrow 1$ to $m$ do
   Initialize prior mean: $\mathbf{\bar{z}}_{\mathbf{v}i}^{-(0)}=\bm{\Phi}\mathbf{\bar{z}}^{+}_{\mathbf{v},i-1}$
 $p_{i}^{-(0)}=1/(\Phi_{\alpha}^{2}/p_{i-1}^{+}+\sigma^{2})$
 // Propagated precision
 for $k\leftarrow 0$ to K1K-1
  // Inner iteration within time step
 do
    Recompute queries and keys:
    $\big(\operatorname{Re}(\mathbf{\bar{z}}_{\mathbf{q}}^{-(k)}),\operatorname{Im}(\mathbf{\bar{z}}_{\mathbf{q}}^{-(k)})\big)\leftarrow\mathcal{L}_{Q}\big(\operatorname{Re}(\mathbf{\bar{z}}_{\mathbf{v}}^{-(k)}),\operatorname{Im}(\mathbf{\bar{z}}_{\mathbf{v}}^{-(k)})\big)$
    $\big(\operatorname{Re}(\mathbf{\bar{z}}_{\mathbf{k}}^{-(k)}),\operatorname{Im}(\mathbf{\bar{z}}_{\mathbf{k}}^{-(k)})\big)\leftarrow\mathcal{L}_{K}\big(\operatorname{Re}(\mathbf{\bar{z}}_{\mathbf{v}}^{-(k)}),\operatorname{Im}(\mathbf{\bar{z}}_{\mathbf{v}}^{-(k)})\big)$
     1exCompute innovations: $\mathbf{y}_{\mathbf{qk}i}^{(k)}=\mathbf{z}_{\mathbf{q}i}-\mathbf{\bar{z}}_{\mathbf{k}i}^{-(k)},\quad\mathbf{y}_{\mathbf{v}i}^{(k)}=\mathbf{z}_{\mathbf{v}i}-\mathbf{\bar{z}}_{\mathbf{v}i}^{-(k)}$
    Nominal posterior precision (before robust weighting): $p_{i}^{+(k)}=p_{i}^{-(k)}+1/\sigma^{2}_{V}$
    "Attention" weight: $w_{i}^{(k)}=-\log\left(1+p_{i}^{+(k)}\mathbf{y}_{\mathbf{v}i}^{(k)\dagger}\mathbf{y}_{\mathbf{v}i}^{(k)}+\epsilon\right)$
    Weighted posterior precision: $\tilde{p}_{i}^{+(k)}=p_{i}^{-(k)}+\exp(\tau_{s}w_{i}^{(k)})/\sigma_{V}^{2}$
    Posterior mean update:
$\mathbf{\bar{z}}_{\mathbf{v}i}^{+(k)}=\mathbf{\bar{z}}_{\mathbf{v}i}^{-(k)}+\text{sigmoid}(\delta)\,w_{i}^{(k)}\mathbf{y}_{\mathbf{v}i}^{(k)}/(\tilde{p}_{i}^{+(k)}\sigma_{V}^{2})$
Set prior for next inner iteration: $\mathbf{\bar{z}}_{\mathbf{v}i}^{-(k+1)}=\mathbf{\bar{z}}_{\mathbf{v}i}^{+(k)},\quad p_{i}^{-(k+1)}=\tilde{p}_{i}^{+(k)}$
 Final posterior for this step: $\mathbf{\bar{z}}_{\mathbf{v}i}^{+}=\mathbf{\bar{z}}_{\mathbf{v}i}^{+(K)}$, $p_{i}^{+}=p_{i}^{+(K)}$
 Propagate mean and precision to next time step:
 $\mathbf{\bar{z}}_{\mathbf{v},i+1}^{-}=\bm{\Phi}\mathbf{\bar{z}}_{\mathbf{v}i}^{+}$,  $p_{i+1}^{-}=1/(\Phi_{\alpha}^{2}/p_{i}^{+}+\sigma^{2})$
Map back to original basis: $\big(\operatorname{Re}(\mathbf{z_{p}}),\operatorname{Im}(\mathbf{z_{p}})\big)\leftarrow\mathcal{L}_{P}\big(\operatorname{Re}(\mathbf{z}_{\mathbf{v}i}^{+}),\operatorname{Im}(\mathbf{z}_{\mathbf{v}i}^{+})\big)$
Return: $\operatorname{Re}(\mathbf{z_{p}})$
Algorithm 4 Unrolled AFA
B.3 Generative Modeling
B.3.1 Propagating the Covariance through the Prediction Step
In order to use AFA as a generative model, we need to propagate the covariance through our one-step prediction.
In particular, given our current estimate for the $i$th measurement, $\mathbf{\bar{z}}_{i}(t_{j})$, the one-step forward prediction is:
$\mathbf{z}_{i}^{+}(t_{j}+\Delta t)=\mathbf{C}e^{\mathbf{A}\Delta t}\mathbf{C}^{-1}\mathbf{\bar{z}}_{i}(t_{j})$
In the eigenbasis, this simplifies to:
$\mathbf{z}_{\mathbf{s}i}^{+}(t_{j}+\Delta t)=e^{\bm{\Lambda}\Delta t}\mathbf{\bar{z}}_{\mathbf{s}i}(t_{j})$
Consider the innovation of the prediction at the next time step: $\mathbf{y}_{\mathbf{s}i}(t_{j}+\Delta t)=\mathbf{\bar{z}}_{\mathbf{s}i}(t_{j}+\Delta t)-\mathbf{z}^{C}_{\mathbf{s}i}(t_{j}+\Delta t)$, i.e. the difference between our predicted measurement and the true measurement. We want to compute $\text{Cov}\big(\mathbf{y}_{\mathbf{s}i}(t_{j}+\Delta t)\big)$.
The innovation covariance at the next step has two contributions:
  1. 1.
    Uncertainty due to our current state estimate (propagated through the dynamics).
  2. 2.
    Uncertainty due to the process and measurement noise that will occur between $t_{j}$ and $t_{j}+\Delta t$.
Suppose the current innovation covariance for the $i$th estimate is:
$\mathrm{Cov}\big(\mathbf{y}_{\mathbf{s}i}\big)=\bm{\Lambda}_{\mathrm{innov},i}^{-}.$
In particular,
$\mathrm{Cov}\big(\mathbf{y}_{\mathbf{s}i}\big)=\left(\sum_{j\leq i}w_{ij}\mathbf{P}^{C}_{ij}\right)^{-1},\quad\text{where:}\quad w_{ij}=\frac{1}{1+\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}/\nu}.$
(We already computed this during the estimation step.)
Propagating this through the deterministic dynamics gives:
$\bm{\Lambda}_{\mathrm{innov},i}^{+}=e^{2\mathrm{Re}(\bm{\Lambda})\Delta t}\,\bm{\Lambda}_{\mathrm{innov},i}^{-}.$
On the other hand, the contribution from the process noise and measurement noise for one time step is:
$\bm{\Lambda}_{\text{proc}}=\|\bm{\Lambda}_{C}\|^{2}\bm{\Lambda}_{Q}\frac{\mathbf{I}-e^{2\mathrm{Re}(\bm{\Lambda})\Delta t}}{-2\mathrm{Re}(\bm{\Lambda})}+\bm{\Lambda}_{R}$
Altogether, the total one-step propagated covariance for prediction $i$ is thus:
$\bm{\Lambda}_{i}^{+}=\bm{\Lambda}_{\mathrm{innov},i}^{+}+\bm{\Lambda}_{\text{proc}}$
We now have the marginal distribution of the mean and covariance of the prediction for every point in the sequence, i.e. $\mathbf{z}_{\mathbf{s}i}^{+}\sim\mathcal{N}(\mathbf{\bar{z}}_{\mathbf{s}i}^{+},\bm{\Lambda}_{i}^{+})$, and $\mathbf{z}_{i}^{+}=\mathbf{S}\mathbf{z}_{\mathbf{s}i}^{+}\mathbf{S}^{-1}$.
Hence, we can sample the final vector embedding from this distribution at the last layer of our network, before passing the embedding through the final linear layer. This is more principled than only using softmax-based sampling, because it properly accounts for the effects of the linear SDE model.
To sample a temporally consistent trajectory, we can use a form of backward recursive sampling based on the Rauch-Tung-Striebel (RTS) smoother: Starting at the final time step $i=m$ and moving backward to $i=1$, we sample the state: $\tilde{\mathbf{z}}_{i}$ conditioned on the previously sampled state $\tilde{\mathbf{z}}_{i+1}$ and the filtered estimate $\mathbf{z}_{i|i}$.
B.4 Model Predictive Control
B.4.1 Open loop waypoint control
Consider a linear system:
$\mathbf{\dot{x}}(\tau)=\mathbf{A}\mathbf{x}(\tau)+\mathbf{B}\mathbf{u}(\tau),\quad\mathbf{x}(0)=\mathbf{x}_{0}$
and we want to reach a waypoint $\mathbf{x}_{\mathrm{ref}}$ at time $t_{f}$ with a quadratic terminal cost:
$J=(\mathbf{x}(t_{f})-\mathbf{x}_{\mathrm{ref}})^{\top}\mathbf{Q}_{f}(\mathbf{x}(t_{f})-\mathbf{x}_{\mathrm{ref}})+\int_{0}^{t_{f}}\mathbf{u}(\tau)^{\top}\mathbf{R}_{c}\mathbf{u}(\tau)\,d\tau.$
The solution of the linear system is:
$\mathbf{x}(t_{f})=e^{\mathbf{A}t_{f}}\mathbf{x}_{0}+\int_{0}^{t_{f}}e^{\mathbf{A}(t_{f}-\tau)}\mathbf{B}\mathbf{u}(\tau)\,d\tau.$
Letting $\mathbf{e}_{0}:=\mathbf{x}_{\mathrm{ref}}-e^{\mathbf{A}t_{f}}\mathbf{x}_{0}$ and $\mathbf{M}(\tau):=e^{\mathbf{A}(t_{f}-\tau)}\mathbf{B}$, the cost becomes:
$J=\int_{0}^{t_{f}}\mathbf{u}^{\top}\mathbf{R}_{c}\mathbf{u}\,d+\left(\int_{0}^{t_{f}}\mathbf{M}(\tau)\mathbf{u}(\tau)\,d\tau-\mathbf{e}_{0}\right)^{\top}\mathbf{Q}_{f}\left(\int_{0}^{t_{f}}\mathbf{M}(\tau)\mathbf{u}(\tau)\,d\tau-\mathbf{e}_{0}\right).$
Taking the variational derivative:
$0=\frac{\delta J}{\delta\mathbf{u}(\tau)}=2\mathbf{R}_{c}\mathbf{u}(\tau)+2\mathbf{M}(\tau)^{\top}\mathbf{Q}_{f}\left(\int_{0}^{t_{f}}\mathbf{M}(\tau)\mathbf{u}(\tau)\,d\tau-\mathbf{e}_{0}\right).$
Defining $\mathbf{U}:=\int_{0}^{t_{f}}\mathbf{M}(\tau)\mathbf{u}(\tau)\,d\tau$ and solving for $\mathbf{u}(\tau)$, we obtain the open-loop feedforward control:
$\mathbf{u}_{\text{ref}}(\tau)=-\mathbf{R}_{c}^{-1}\mathbf{M}(\tau)^{\top}\mathbf{Q}_{f}(\mathbf{U}-\mathbf{e}_{0}).$
Substituting this into $\mathbf{U}$:
$\mathbf{U}=-\mathbf{G}(\mathbf{U}-\mathbf{e}_{0}),\quad\mathbf{G}:=\int_{0}^{t_{f}}e^{\mathbf{A}(t_{f}-\tau)}\mathbf{B}\mathbf{R}_{c}^{-1}\mathbf{B}^{\top}e^{\mathbf{A}^{\top}(t_{f}-\tau)}\mathbf{Q}_{f}\,d\tau,$
$(\mathbf{I}+\mathbf{G})\mathbf{U}=\mathbf{G}\mathbf{e}_{0}\quad\Rightarrow\quad\mathbf{U}=(\mathbf{I}+\mathbf{G})^{-1}\mathbf{G}\mathbf{e}_{0}.$
Hence,
$\mathbf{u}_{\text{ref}}(\tau)=-\mathbf{R}_{c}^{-1}\mathbf{M}(\tau)^{\top}\mathbf{Q}_{f}\Big((\mathbf{I}+\mathbf{G})^{-1}\mathbf{G}-\mathbf{I}\Big)\mathbf{e}_{0}.$
$\mathbf{R}_{c}^{-1}\mathbf{M}(\tau)^{\top}\mathbf{Q}_{f}(\mathbf{I}+\mathbf{G})^{-1}\mathbf{e}_{0}.$
B.4.2 Closed-form expressions for control and state evolution
If $\mathbf{A},\mathbf{B},\mathbf{Q}_{f},\mathbf{R}_{c}$ are diagonalizable, each mode $k$ decouples in the eigenbasis (recall that the $s$ subscript denotes quantities in the eigenbasis):
$\dot{x}_{s,k}=\lambda_{k}x_{s,k}+\lambda_{B,k}u_{s,k}.$
Let $c_{qr,k}:=\frac{q_{f,k}\big|\lambda_{B,k}\big|^{2}}{r_{k}}$, and define:
$G_{k}=c_{qr,k}\int_{0}^{t_{f}}e^{2\mathrm{Re}(\lambda_{k})(t_{f}-\tau)}\,d\tau=-c_{qr,k}\frac{1-e^{2\mathrm{Re}(\lambda_{k})t_{f}}}{2\mathrm{Re}(\lambda_{k})}$
Then, the feedforward control in mode $k$ is:
$u_{\text{ref},k}(\tau)=\frac{q_{f,k}\,\lambda_{B,k}^{*}}{r_{k}}\frac{1}{1+G_{k}}\big(x_{\mathrm{ref},k}-x_{s,k}(0)e^{\lambda_{k}t_{f}}\big)e^{\lambda_{k}^{*}(t_{f}-\tau)}.$
Plugging the control back into the dynamics:
$\dot{x}_{s,k}(\tau)=\lambda_{k}x_{s,k}(\tau)+\lambda_{B,k}u_{\text{ref},k}(\tau)$
Define:
$C_{k}:=c_{qr,k}\frac{1}{1+G_{k}}\big(x_{\mathrm{ref},k}-x_{s,k}(0)e^{\lambda_{k}t_{f}}\big)$
Then the dynamics become:
$\dot{x}_{s,k}(\tau)=\lambda_{k}x_{s,k}(\tau)+C_{k}e^{\lambda_{k}^{*}(t_{f}-\tau)}.$
The solution is:
$x_{s,k}(t_{f})=x_{s,k}(0)e^{\lambda_{k}t_{f}}-\frac{1-e^{2\mathrm{Re}(\lambda_{k})t_{f}}}{2\mathrm{Re}(\lambda_{k})}C_{k}$
Let $g_{k}:=-\frac{1-e^{2\mathrm{Re}(\lambda_{k})t_{f}}}{2\mathrm{Re}(\lambda_{k})}$ and $\beta_{k}:=\frac{c_{qr,k}\,g_{k}}{1+c_{qr,k}\,g_{k}}$.
The state under the control becomes:
$x_{s,k}(t_{f})=(1-\beta_{k})x_{s,k}(0)e^{\lambda_{k}t_{f}}+\beta_{k}x_{\mathrm{ref},k}$
This is a convex combination of the initial state, mapped through the dynamics, and the reference state. Substituting the current estimate $\bar{x}_{s,k}$ for $x_{s,k}(0)$ and one time step $\Delta t$ for $t_{f}$ gives:
$\tilde{x}_{s,k}^{+}:=(1-\beta_{k})\bar{x}_{s,k}e^{\lambda_{k}\Delta t}+\beta_{k}x_{\mathrm{ref},k}$
Or, in terms of the measurements:
$\tilde{z}_{s,k}^{+}:=(1-\beta_{k})\bar{z}_{s,k}e^{\lambda_{k}\Delta t}+\beta_{k}z_{\mathrm{ref},k}$
Covariance propagation is independent of the control for a linear system with additive noise, as long as the control is deterministic. So the full distribution under the control is simply: $\mathcal{N}\big(\tilde{z}_{s,k}^{+},\bm{\Lambda}_{i}^{+}\big)$.
This architecture separates the model into three parts: a world-model/filter (AFA), a low-level model predictive controller; and a next-step way-point planner. By embedding the output of the analytic feedforward control law in each layer, we are implementing a self-consistent open-loop controller that respects the model’s own learned world model.
B.5 Extension to 2D Data
Previously, we modeled sequence data as noisy observations of a linear SDE. For images, the analogous object is a random field or SPDE, but solving an SPDE exactly is intractable. Instead, we can locally approximate the field using separable 1D SDEs along each axis, yielding a precision-weighted estimator that fuses information from all four directions around a central pixel. This produces a learnable, adaptive 2D filter that generalizes convolution.
In 2D, the state $\mathbf{x}(x,y)$ is a function of spatial coordinates $(x,y)$ and evolves according to the linear stochastic partial differential equation:
$d\mathbf{x}(x,y)=\big(\mathbf{A}_{1}\mathbf{x}\,dx+\mathbf{A}_{2}\mathbf{x}\,dy\big)+\mathbf{G}\,d^{2}\mathbf{w}(x,y),$
where $d^{2}\mathbf{w}(x,y)$ is a 2D Wiener sheet with covariance:
$\mathbb{E}\!\big[d^{2}\mathbf{w}(x,y)\,d^{2}\mathbf{w}(x^{\prime},y^{\prime})^{\top}\big]=\mathbf{I}\,\delta(x-x^{\prime},y-y^{\prime})\,dxdy.$
Directly solving this SPDE for a precision-weighted estimator across a local field is computationally intractable. Therefore, we introduce a local approximation by modeling the relationships between pixels using decoupled, axial 1D SDEs along the $x$ and $y$ axes.
Given a state $\mathbf{x}_{ij}$ located at the $ij$th position on a grid, we define four SDEs radiating into position $ij$:
$\text{L$\to$R}:\quad d\mathbf{x}_{ij}^{l}$ $=\mathbf{A}_{l}\mathbf{x}_{ij}^{l}\,di+\mathbf{G}_{l}\,d\mathbf{w}^{(l)}$
$\text{R$\to$L}:\quad d\mathbf{x}_{ij}^{r}$ $=\mathbf{A}_{r}\mathbf{x}_{ij}^{r}\,di+\mathbf{G}_{r}\,d\mathbf{w}^{(r)}$
$\text{B$\to$T}:\quad d\mathbf{x}_{ij}^{b}$ $=\mathbf{A}_{b}\mathbf{x}_{ij}^{b}\,dj+\mathbf{G}_{b}\,d\mathbf{w}^{(b)}$
$\text{T$\to$B}:\quad d\mathbf{x}_{ij}^{t}$ $=\mathbf{A}_{t}\mathbf{x}_{ij}^{t}\,dj+\mathbf{G}_{t}\,d\mathbf{w}^{(t)}$
where $d\mathbf{w}^{(\ell)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}|di|)$, $\ell\in\{l,r\}$, $d\mathbf{w}^{(\ell)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}|dj|)$, $\ell\in\{b,t\}$ are Wiener increments. Define $\mathbf{Q}_{\ell}:=\mathbf{G}_{\ell}\mathbf{G}_{\ell}^{\top}$, for $\ell\in\{l,r,b,t\}$.
We define a single measurement model:
$\mathbf{z}_{ij}=\mathbf{x}_{ij}+\mathbf{v}_{ij},\quad\mathbf{v}_{ij}\sim\mathcal{N}(\mathbf{0},\mathbf{R}).$
We again assume that all system and noise matrices are simultaneously diagonalizable: $\mathbf{A}_{\ell}=\mathbf{S}\bm{\Lambda}_{\ell}\mathbf{S}^{-1}$, $\mathbf{Q}_{\ell}=\mathbf{S}\bm{\Lambda}_{Q,\ell}\mathbf{S}^{-1}$, $\mathbf{R}=\mathbf{S}\bm{\Lambda}_{R}\mathbf{S}^{-1}$, for some invertible $\mathbf{S}\in\mathbb{C}^{d\times d}$, and diagonal $\bm{\Lambda}_{\ell},\bm{\Lambda}_{Q,\ell},\bm{\Lambda}_{R}$.
To allow the SDE dynamics to operate along arbitrary orientations, we define a coordinate transformation. Given two vectors $\mathbf{z}_{a},\,\mathbf{z}_{b}\in\mathbb{R}^{d}$, located at positions $i_{a}j_{a}$ and $i_{b}j_{b}$, let $\Delta x=i_{a}-i_{b}$, $\Delta y=j_{a}-j_{b}$. Let $\mathbf{T}\in\mathbb{R}^{2\times 2}$, $\mathbf{b}\in\mathbb{R}^{2\times 1}$ and compute:
$\begin{pmatrix}\Delta x^{\prime}\\ \Delta y^{\prime}\end{pmatrix}=\mathbf{T}\\ \begin{pmatrix}\Delta x\\ \Delta y\end{pmatrix}+\mathbf{b}$
We will assume the same form of isotropic plus skew-symmetric dynamics as before:
$\bm{\Lambda}_{\ell}=\alpha_{\ell}\mathbf{I}+i\bm{\Lambda}_{\Omega,\ell},$
Define $\bm{\omega}_{l}:=\text{diag}\big(\bm{\Lambda}_{\Omega,\ell}\big)$ and:
$\alpha_{x}:=\alpha_{l}\,\mathbbm{1}(\Delta x^{\prime}>0)+\alpha_{r}\,\mathbbm{1}(\Delta x^{\prime}<0)$
$\alpha_{y}:=\alpha_{b}\,\mathbbm{1}(\Delta y^{\prime}>0)+\alpha_{t}\,\mathbbm{1}(\Delta y^{\prime}<0)$
$\bm{\omega}_{x}:=\bm{\omega}_{l}\,\mathbbm{1}(\Delta x^{\prime}>0)+\bm{\omega}_{r}\,\mathbbm{1}(\Delta x^{\prime}<0)$
$\bm{\omega}_{y}:=\bm{\omega}_{b}\,\mathbbm{1}(\Delta y^{\prime}>0)+\bm{\omega}_{t}\,\mathbbm{1}(\Delta y^{\prime}<0)$
Define state transition matrices:
$\bm{\Phi}_{x}(\Delta x^{\prime})=e^{-\alpha_{x}^{2}|\Delta x^{\prime}|}\,e^{-i\bm{\omega}_{x}\Delta x^{\prime}}$
$\bm{\Phi}_{y}(\Delta y^{\prime})=e^{-\alpha_{y}^{2}|\Delta y^{\prime}|}\,e^{-i\bm{\omega}_{y}\Delta y^{\prime}}$
$\bm{\Phi}_{xy}(\Delta x^{\prime},\Delta y^{\prime}):=\bm{\Phi}_{x}(\Delta x^{\prime})\odot\bm{\Phi}_{y}(\Delta y^{\prime})$
Mapping each measurement to the eigenbasis, $\mathbf{z}_{\mathbf{s},a}:=\mathbf{S}^{-1}\mathbf{z}_{a},\,\mathbf{z}_{\mathbf{s},b}:=\mathbf{S}^{-1}\mathbf{z}_{b}$, we can form an estimate of $\mathbf{z}_{\mathbf{s},a}$ from $\mathbf{z}_{\mathbf{s},b}$:
$\mathbf{\hat{z}}_{\mathbf{s},ab}:=\bm{\Phi}_{xy}(\Delta x^{\prime},\Delta y^{\prime})\odot\mathbf{z}_{\mathbf{s},b}$
Let us again assume isotropic process and measurement noise:
$\mathbf{Q}_{\ell}=\sigma_{\ell}^{2}\mathbf{I},\quad\mathbf{R}=\eta^{2}\mathbf{I}$
$\sigma_{x}:=\sigma_{l}\,\mathbbm{1}(\Delta x^{\prime}>0)+\sigma_{r}\,\mathbbm{1}(\Delta x^{\prime}<0)$
$\sigma_{y}:=\sigma_{t}\,\mathbbm{1}(\Delta y^{\prime}>0)+\sigma_{b}\,\mathbbm{1}(\Delta y^{\prime}<0)$
For a decoupled random field in 2D, the state covariance is a product of the 1D covariances propagated in each direction:
$\sigma_{V}^{2}=\Big(\sigma_{x}^{2}\frac{e^{-2\alpha_{x}^{2}|\Delta x^{\prime}|}-1}{-2\alpha_{x}^{2}}\Big)\Big(\sigma_{y}^{2}\frac{e^{-2\alpha_{y}^{2}|\Delta y^{\prime}|}-1}{-2\alpha_{y}^{2}}\Big)+\eta^{2}e^{-2\alpha_{x}^{2}|\Delta x^{\prime}|-2\alpha_{y}^{2}|\Delta y^{\prime}|}$
And the precision is then:
$\bm{\Lambda}^{C}_{\mathbf{P}(i_{a},j_{a}),(i_{b},j_{b})}=\frac{1}{\sigma_{V}^{2}}\mathbf{I}$
Given some neighborhood $\mathcal{N}$ (i.e. a convolutional kernel), we can now compute a precision-weighted average:
$\mathbf{\bar{z}}_{\mathbf{s},ij}=\frac{\displaystyle\sum_{u,v\in\mathcal{N}}\bm{\Lambda}^{C}_{\mathbf{P}(i,j),(i+u,j+v)}\,\mathbf{\hat{z}}_{\mathbf{s},(i,j),(i+u,j+v)}}{\displaystyle\sum_{u,v\in\mathcal{N}}\bm{\Lambda}^{C}_{\mathbf{P}(i,j),(i+u,j+v)}}$
We then map back to the original basis $\mathbf{\bar{z}}_{ij}=\mathbf{S}\mathbf{\bar{z}}_{\mathbf{s},ij}$. If the dynamics and process noise disappear, then the above reduces to ordinary convolution with fixed kernel weights.
The above is also very similar to the Gabor Filter, which is defined by a sinusoidal plane wave multiplied by a Gaussian envelope:
$g(\Delta x,\Delta y;\lambda,\theta,\psi,\sigma,\gamma)=\exp\!\left(-\frac{\Delta x^{\prime 2}+\gamma^{2}\Delta y^{\prime 2}}{2\sigma^{2}}\right)\cos\!\left(\frac{2\pi\Delta x^{\prime}}{\lambda}+\psi\right)$
where $\begin{pmatrix}\Delta x^{\prime}\\ \Delta y^{\prime}\end{pmatrix}=\mathbf{R}\begin{pmatrix}\Delta x\\ \Delta y\end{pmatrix}$, and where $\mathbf{R}=\begin{pmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{pmatrix}$ is a rotation matrix.
The differences are that we have a Laplacian envelope rather than Gaussian envelope, two oscillation terms rather than one, and a vector $\omega$ rather than scalar.
If $\alpha_{x}:=\alpha_{l}=\alpha_{r}$, $\alpha_{y}:=\alpha_{b}=\alpha_{t}$, $\bm{\omega}_{x}:=\bm{\omega}_{l}=\bm{\omega}_{r}$, $\bm{\omega}_{y}:=\bm{\omega}_{b}=\bm{\omega}_{t}$, then our kernel is:
[…]
B.6 Further Justification of the Radial-Tangential Structure
B.6.1 Normal plus Rank-1 Noise Covariance
In this section, we explore a generalization of the RT-SDE model by considering a normal-plus-rank-1 structure for $\mathbf{Q}$ and $\mathbf{R}$. Our goal is to determine whether relaxing the strict radial-tangential decomposition introduces additional expressivity in the propagated covariance or the resulting MLE.
Recall that the propagated covariance in the eigenbasis is:
$\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j}):=\int_{0}^{|\Delta t_{ij}|}e^{-\bm{\Lambda}\tau}(\mathbf{S}^{-1}\mathbf{Q}(t_{j}+\tau)\mathbf{S}^{-\dagger})e^{-\bm{\Lambda}^{\dagger}\tau}d\tau$
Let us now assume that $\mathbf{Q}$ is normal plus rank-1, i.e. $\mathbf{Q}=\mathbf{S}\left(\bm{\Lambda}_{Q}+\mathbf{q}_{ij}(\tau)\mathbf{q}_{ij}^{\dagger}(\tau)\right)\mathbf{S}^{\dagger}$, where $\mathbf{q}_{ij}(\tau)$ may depend on $\Delta t_{ij}$.
We already derived the diagonal part. The rank-1 part is:
$\int_{0}^{|\Delta t_{ij}|}e^{-\bm{\Lambda}\tau}\big(\mathbf{q}_{ij}(\tau)\mathbf{q}_{ij}^{\dagger}(\tau)\big)e^{-\bm{\Lambda}^{\dagger}\tau}d\tau$
$\int_{0}^{|\Delta t_{ij}|}\big(e^{-\bm{\Lambda}\tau}\mathbf{q}_{ij}(\tau)\big)\big(e^{-\bm{\Lambda}\tau}\mathbf{q}_{ij}(\tau)\big)^{\dagger}d\tau$
If we take: $\mathbf{q}(t_{j}+\tau)=e^{\bm{\Lambda}(\Delta t_{ij}+\tau)}\mathbf{p}_{i}$, so that $\mathbf{q}_{ij}$ moves with the LTI’s flow, this becomes:
$e^{\bm{\Lambda}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}^{\dagger}\Delta t_{ij}}\int_{0}^{|\Delta t_{ij}|}d\tau=e^{\bm{\Lambda}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}^{\dagger}\Delta t_{ij}}|\Delta t_{ij}|$
If we assume that $\bm{\Lambda}=\alpha\mathbf{I}+\bm{\Omega}$, where $\bm{\Omega}^{\dagger}=-\bm{\Omega}$, then we have:
$\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}e^{-\bm{\Lambda}_{\Omega}\tau}\big(\mathbf{q}_{ij}\mathbf{q}_{ij}^{\dagger}\big)e^{-\bm{\Lambda}_{\Omega}^{\dagger}\tau}d\tau$
Letting $\mathbf{q}(t_{j}+\tau)=ae^{\bm{\Lambda}_{\Omega}(\Delta t_{ij}+\tau)}\mathbf{p}$ for some constant $a$, we have:
$ae^{\bm{\Lambda}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}^{\dagger}\Delta t_{ij}}\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}d\tau=a\frac{e^{-2\alpha\Delta t_{ij}}-1}{-2\alpha}e^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}$
Altogether, we have:
$\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})=\sigma^{2}\kappa_{B}(\Delta t_{ij})\Big(\mathbf{I}+ae^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}\Big)$
Assuming measurement noise with a rank-1 component aligned with the rank-1 component of the process noise, the propagated measurement covariance is:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=e^{-\bm{\Lambda}\Delta t_{ij}}\Big[\big|\lambda_{C}\big|^{2}\sigma^{2}\kappa_{B}(|\Delta t_{ij}|)\Big(\mathbf{I}+ae^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}\Big)$
$+\eta^{2}\Big(\mathbf{I}+be^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}\Big)\Big]e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\big|\lambda_{C}\big|^{2}\sigma^{2}\kappa_{F}(|\Delta t_{ij}|)\Big(\mathbf{I}+a\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}\Big)+\eta^{2}e^{-2\alpha\Delta t_{ij}}\Big(\mathbf{I}+b\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}\Big)$
$\Big(\big|\lambda_{C}\big|^{2}\sigma^{2}\kappa_{F}(|\Delta t_{ij}|)+\eta^{2}e^{-2\alpha\Delta t_{ij}}\Big)\mathbf{I}+\Big(a\big|\lambda_{C}\big|^{2}\sigma^{2}\kappa_{F}(|\Delta t_{ij}|)+b\eta^{2}e^{-2\alpha\Delta t_{ij}}\Big)\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}$
Letting $\mathbf{p}_{i}=\mathbf{u}_{\mathbf{s}i}$, and setting $\sigma=\sigma_{t}$, $\eta=\eta_{t}$, $a=\frac{\sigma_{r}^{2}-\sigma_{t}^{2}}{\sigma_{t}^{2}}$, $b=\frac{\eta_{r}^{2}-\eta_{t}^{2}}{\eta_{t}^{2}}$, this reduces to the RT-SDE case:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})$ $=\Big(\big|\lambda_{C}\big|^{2}\sigma_{t}^{2}\kappa_{F}(|\Delta t_{ij}|)+\eta_{t}^{2}e^{-2\alpha\Delta t_{ij}}\Big)\mathbf{I}$
$+\Big(\big|\lambda_{C}\big|^{2}(\sigma_{r}^{2}-\eta_{t}^{2})\kappa_{F}(|\Delta t_{ij}|)+(\eta_{r}^{2}-\eta_{t}^{2})e^{-2\alpha\Delta t_{ij}}\Big)\mathbf{u}_{\mathbf{s}i}\mathbf{u}_{\mathbf{s}i}^{\dagger}$
$\sigma_{Vt}^{2}(|\Delta t_{ij}|)(\mathbf{I}-\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger})+\sigma_{Vr}^{2}(|\Delta t_{ij}|)\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$
In summary, the co-rotating normal-plus-rank-1 noise model provides a minor generalization of the RT-SDE structure by permitting anisotropy along directions not tied to the state. However, this comes at the cost of losing the clean hyper-spherical MLE formulation.
B.6.2 Expressiveness of Radial-Tangential SDEs
The efficiency of the memory-optimized AFA comes from restricting the dynamics to isotropic decay and noise. We now examine the expressive power of these constraints. Proposition 1 shows that a sequence of linear systems with isotropic decay can approximate the flow of any diagonalizable LTI system to arbitrary accuracy. Proposition 2 shows that a multi-head Radial-Tangential SDE can approximate any target covariance at a fixed time, and thus represent arbitrary state-dependent uncertainty structures.
Proposition 1: Expressiveness of Deterministic RT-SDEs.
Let $\mathbf{A}\in\mathbb{R}^{d\times d}$ be a diagonalizable matrix, with eigendecomposition $\mathbf{A}=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1}$, where $\mathbf{S}\in\mathbb{R}^{d\times d}$, $\bm{\Lambda}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{d})\in\mathbb{C}^{d\times d}$. For any $\epsilon>0$, there exists a sequence of matrices $\mathbf{A}_{\ell}=\mathbf{S}(\alpha_{\ell}\mathbf{I}+\bm{\Lambda}_{\Omega\ell})\mathbf{S}^{-1}$, where $\alpha_{\ell}\in\mathbb{R}$, $\bm{\Lambda}_{\Omega\ell}=\mathrm{diag}(i\omega_{\ell,1},\dots,i\omega_{\ell,d})\in i\mathbb{R}^{d\times d}$, and a choice of block durations $\Delta t>0$, such that
$\Big|e^{\mathbf{A}t_{k}}-\prod_{\ell=1}^{L}e^{\mathbf{A}_{\ell}\Delta t}\Big|<\epsilon,\quad\forall k=1,\dots,m.$
Proof Sketch:
Each $\mathbf{A}_{\ell}$ generates dynamics:
$e^{\mathbf{A}_{\ell}\Delta t}=\mathbf{S}\,\mathrm{diag}\big(e^{(\alpha_{\ell}+i\omega_{\ell,1})\Delta t},\dots,e^{(\alpha_{\ell}+i\omega_{\ell,d})\Delta t}\big)\mathbf{S}^{-1}.$
Each diagonal entry corresponds to a 1D complex exponential in the eigenbasis. By choosing sufficiently many blocks $L$ and appropriately selecting $\alpha_{\ell}$ and $\omega_{\ell,k}$, one can approximate:
$e^{\lambda_{k}t_{j}}\approx\prod_{\ell=1}^{L}e^{(\alpha_{\ell}+i\omega_{\ell,k})\Delta t},\quad\forall k,j,$
to arbitrary precision. Mapping back to the original basis with $\mathbf{S}$ yields the desired approximation. Hence, for any finite set of times, a product of radial-tangential blocks can approximate the target flow arbitrarily closely.
Remark: If we have $n$ heads, each producing its own low-dimensional output, stacking them allows combinations of rotations and scalings from different subspaces to mix. The resulting flow in the original $d$-dimensional space can approximate much richer dynamics than a single high-dimensional LTI with the same number of blocks. The effective number of distinct transformations is roughly the number of heads times the number of layers.
Proposition 2: Expressiveness of RT-SDE Covariance.
Let $\bm{\Sigma}^{\star}\in\mathbb{R}^{d\times d}$ be a symmetric positive definite (SPD) matrix and fix a time $t_{i}$. For any $\varepsilon>0$, there exists a number of heads $H$ and associated RT-SDE parameters:
$\{\mathbf{u}_{h}(t_{i}),\sigma_{Vr,h}^{2},\sigma_{Vt,h}^{2}\}_{h=1}^{H}$
such that the RT-SDE covariance satisfies:
$\big|\bm{\Sigma}(t_{i})-\bm{\Sigma}^{\star}\big|<\varepsilon,$
where
$\bm{\Sigma}(t_{i})=\sum_{h=1}^{H}\sigma_{Vt,h}^{2}\mathbf{I}+\sum_{h=1}^{H}\beta_{h}\,\mathbf{u}_{h}(t_{i})\mathbf{u}_{h}(t_{i})^{\top},\quad\beta_{h}=\sigma_{Vr,h}^{2}-\sigma_{Vt,h}^{2}.$
Proof Sketch:
Each head $h$ contributes an SPD term of the form:
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\sigma_{Vt}^{2}(|\Delta t_{ij}|)\mathbf{I}+\big(\sigma_{Vr}^{2}(|\Delta t_{ij}|)-\sigma_{Vt}^{2}(|\Delta t_{ij}|)\big)\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$
i.e. a rank-1 correction of identity aligned with the current state direction $\mathbf{u}(t_{i})$.
The total propagated covariance is the sum of the individual head covariances, assuming the uncertainty sources for each head are uncorrelated:
$\bm{\Sigma}(t_{i})\;=\;\Big(\sum_{h}\sigma_{Vt}^{2}(|\Delta t_{ij}|)\Big)\mathbf{I}+\sum_{h}\beta_{h}\,\mathbf{u}_{h}(t_{i})\mathbf{u}_{h}(t_{i})^{\top},\quad\beta_{h}=\sigma_{Vr}^{2}-\sigma_{Vt}^{2}.$
At a fixed time $t_{i}$, we can construct any target SPD matrix $\bm{\Sigma}^{\star}$ using the RT-SDE heads as follows. Let $\sigma_{\min}=\lambda_{\min}(\bm{\Sigma}^{\star}),\quad\sum_{h}\sigma_{Vt}^{2}=\sigma_{\min}$, so that the isotropic contributions sum to the minimal eigenvalue. Consider the spectral decomposition of the shifted matrix:
$\bm{\Sigma}^{\star}-\sigma_{\min}\mathbf{I}=\sum_{h}(\lambda_{h}-\sigma_{\min})\,\mathbf{v}_{h}\mathbf{v}_{h}^{\top},$
where $\{\mathbf{v}_{h}\}$ are the eigenvectors of $\bm{\Sigma}^{\star}$. Setting $\mathbf{u}_{h}(t_{i})=\mathbf{v}_{h},\,\beta_{h}=\lambda_{h}-\sigma_{\min}$, gives:
$\sum_{h}\sigma_{Vt}^{2}\mathbf{I}+\sum_{h}\beta_{h}\,\mathbf{u}_{h}(t_{i})\mathbf{u}_{h}(t_{i})^{\top}=\sigma_{\min}\mathbf{I}+(\bm{\Sigma}^{\star}-\sigma_{\min}\mathbf{I})=\bm{\Sigma}^{\star}.$
Remark: Note that this only holds at a fixed time, since $\mathbf{u}_{h}(t)$ evolves according to the underlying LTI dynamics, so we cannot freely choose the propagated covariance at all times independently. Nevertheless, this shows that the RT-SDE model is substantially more flexible than a diagonalizable LTI-SDE with isotropic noise, for which the propagated measurement covariance remains isotropic in the eigenbasis.